A Complete Bibliography of Publications in Linear Algebra and its Applications: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 2.11

Title word cross-reference

#3162 [Sta81].

(0, 1) [BFP88, BH85, CS89b, LN80, MSS89, Min85, Min87b, PS88, SS86]. (2^{m}, 2^{m} − 4) [KL88a]. (A_{1}) [Ber87]. (F, G) [Emr83]. (H, F) [Özg86]. (p, q) [LTT83, LTT88]. (S) [Wan86]. (s, t) [JP82]. (T) [Wan86]. (trA^{p})^{1/p} [Mag87]. (x − α)(x − α^{2})(x − α^{3}) [KL88a]. (Y A − DZ, Y C − BZ) = (E, F) [wEC87b]. + [Sil87b]. −2 [KRS82, Tor86a]. 0 [Fri85a]. {0, ±1} [Fuj84]. 1 [BP84a, CR86, Fri85a, RB89]. 2 [BN89, ELS82a, RW89, Ver84, HLG85]. 2 + 5 [BN89]. 27 [RS86a]. 2^{n} [Giv84, YH82]. 2 × 2 [Tho82b, Uhl81a, Uhl83]. 3 [AL80, Bix82]. 3 × 3 [Piz83, AO82]. ^{∗} [dOdS83]. ^{2}A_{2d−1}(p^{f}) [IS89]. A [BG82, Web86]. A(λ) = A_{0}λ^{t} + A_{1}λ^{t−1} + ⋯ + A_{t−1}λ + A_{t} [BZ89]. A > 0 [Sha86a]. A^{k}b [CLS87]. A^{m} = dl + λJ [Ma84]. A^{m} = λJ [MW87a]. AB [Sha86a]. A ∘ A^{−1} ≥ I [BK87]. AX + Y B = C [Zie85]. AX − XB = C [LLT84, dSB81b]. AX = B [Uhl87, JUW82, dSB81a]. AX = C [Mit84]. Ax = λBx [BG82, LS87, LM80]. AXB + CY D = E [BK80b].

AXB − CXD = E [wEC87b, HG89]. B [BG82]. B > 0 [Sha86a]. c [Beb86, LT83b, LT88b, BP85, GS82, GS83a, Tsi84]. C^{∗} [Gro82, LP81]. c^{2} [vB85]. C^{n} [Hol85]. H [Yan84]. d [DG87, BH84, Cai84, Har80, Tog80]. D_{1}AD_{2} [Bap82]. ΔH [Gab87]. d ≥ 3 [IS89]. dn [KP80]. e [Fri88, Sta87]. E(t) [Eva87]. ℓ_{1} [dC87]. ℓ_{∞} [Flo82]. ℓ_{p} [Gol86a, GS83b, Gol84]. F (x) + M(x)G(1/x) = 0 [Ng87]. F (X) + M(X)G(X^{−1}) = 0 [Eba89]. G [NT80, MW87a, Che83d, GW85, Mar81a]. Γ [OH85]. GF(2^{m}) [KL88a]. H [ABO88, FH88a, HS85, LP82a, NV84, Neu86, Pol87, Tad81]. H^{∞} [LA88, LST88]. HR [BG82, BG81b]. H^{∞} [JJS89]. K [Yua86, Bea83, Bea88, BL89a, Pif87]. K^{n} [Eba89]. K_{1} [VM87, Bro89b]. l^{p} [Dru87]. l_{1}(Γ) [Pol86]. L_{2} [Tad81]. Λ^{k}R^{n} [Mor85]. ≤ n [HHM89]. LP [Naz84]. LU [ABO88, Meh84c, Mur83, OS89, FP81, GMSW87]. M [Joh82a, Lew89a, Ste81b, Zha87b, Wol89a, AS82, BS87a, EP88, FMN83, FJMN85, FJM87, FM88b, FM88a, FHV85, FP81, HRS88, HS88a, Ima83, Ima84, JORvdD87, KR83, KB85, LN80, Li89, LP80, Meh84b, MNR84, Neu86, Neu81b, NP87, Sch84a, Smi87, ST87a, Szu89a, dG82]. M_{n}(Z/h) [Har81b]. M ≥ B^{∗}MB [BKK83]. MMA [Fie88b]. m × n × (mn − q) [Bin86]. n [AN83, Bou88, KP80, KS84, Tak85, TST86, Wer85a, Joh82b]. N_{0} [Joh85, Smi86]. N × N [GT89a]. O(m^{2}) [Kle87]. O(N^{2}) [GT89a]. O^{+}(V) [KN87b]. ω [HB84, HM86]. P [BG84a, Fan89, HJ86d, Tho85, VNC84, WN87, HB83, HJ86c]. P (t) [Eva87]. P_{k}^{3} [Bre84]. P_{0} [Fan89, RM85, HB83]. Per(J_{2} ⊗ M) ≥ 2^{n}[Per(M)]^{2} [Pat88a]. Ψ [OH85]. p × p [Tau85]. Q [JP84, RCS89, CvRS81, Mor88]. QD [Ker83]. QL [Zha89, JZ85]. QR [Cha87a, HS86b, BG86, Dem89, Faĭ89]. r [GB87, Gov89b, Wer87]. R^{3} [PL88]. R^{n} [MLS83, KL87, MLS87, MLS88]. S [Kle87, DSD88, Wat87a, Fie84c]. Σ [GVK^{+}83]. Sx^{+} − T x^{−} = y [KL87]. t [BF82, DS89a, Eva87, RCS89, Wil82b, Hal86b]. T (H^{1/n}T)^{n} = K [Fur88]. {t^{i}e^{λ_{k}t}} [Har82]. τ [HB84, HM86]. V [Meh84c]. W [JP88a, TW80]. W^{∗} [Had86, Brü84d]. x [GR89a]. X − AXB = C [Wim88, LLT84]. XA + BX + C = 0 [Mil88]. XA − BX = R [Dat88]. XA = A^{T} X [DD87b]. XAX^{−1} + B [Sil86]. XB = D [Mit84]. y [GR89a]. Z [HS88b, Smi88, HS88a, Ram86]. Z/pZ [Wat88b]. Z_{p} [Str87]. Z_{q} [Jun89]. ZME [Stu88a]. |D(X^{k})| = |kX^{k−1}| [Sun82b].

-adic [Tho85]. -algebra [Had86]. -algebras [Brü84d, Gro82]. -analogues [RCS89]. -circulant [MW87a]. -circulants [Min85, Min87b]. -classes [Yan84]. -connectivity [Bix82]. -convexity [LP81]. -cyclic [ELS82a, VNC84]. -D [RW89]. -decomposition [Mur83]. -designs [DS89a, RCS89, Wil82b]. -dimensional [DG87]. -frames [KS84]. -functions [Che83d, NT80]. -Hessenberg [Pif87]. -ideals [Hal86b]. -invariant [Özg86, Emr83]. -inverse [HLG85]. -irreducible [TW80]. -like [BG86]. -linear [Bou88]. -lossless [GVK^{+}83]. -majorization [GW85]. -matrices [AS82, BP85, BS87a, BH85, CvRS81, EP88, FMN83, Fie84c, FJMN85, FJM87, FM88b, FHV85, FP81, Gab87, HB83, HB84, HS85, HJ86c, HM86, HRS88, HS88a, Ima83, Joh82b, KR83, KB85, LN80, LP80, Meh84b, MNR84, NV84, Neu86, Neu81b, NP87, Pol87, RM85, Ram86, SS86, Smi86, ST87a, Stu88a, Szu89a, ABO88, BG84a, Fan89, Fie88b, JP84, Joh82a, Joh85, Kle87, Lew89a, Meh84c, Smi88, Ste81b, Zha87b]. -matrix [FM88a, Ima84, JORvdD87, LN80, Li89, Mor88, Sch84a, Smi87, dG82, dOdS83, HJ86d, HS88b]. -monotonicity [Wer87]. -networks [vB85]. -norm [dC87]. -numerical [Beb86, GS82, GS83a, LTT83, LTT88, Tsi84]. -potent [JP82]. -preservers [Bea83]. -proper [Web86]. -reguli [BF82]. -selfadjoint [LP82a]. -special [LT83b]. -spectral [LT88b]. -splittings [Sch84a]. -stability [Tad81, Tog80]. -stable [BH84, Cai84, Har80]. -structure [RB89]. -symmetric [Wat87a]. -tensors [AL80]. -Toeplitz [GB87, Gov89b]. -type [Faĭ89]. -unitary [FH88a]. -way [Tak85].

100 [Ano88a, BS89]. 1968/1988 [Ano88a, BS89]. 1978 [Ple81]. 1980 [BP81]. 1981-1984 [Ano84i]. 1983 [BU84]. 1984 [GL85a, Hol85, Lew85b]. 1984-1986 [Ano86n]. 1987 [BU88].

20 [Ano88a, Sta81]. 23 [BKR80]. 2D [BFM89b, BFM89a].

30 [Cai82]. 30-41 [Ano88-28]. 3rd [Gov89a].

41-56 [AB85]. 41-60 [Ano84i]. 42-54 [Ano88-28].

57 [Sta81].

61-80 [Ano86n].

70 [Sha89]. 70th [FL89b]. 75th [Ano85s, EH85].

80b [BKR80]. 82c [Cai82]. 86m [Sha89]. 89h [BS89].

Abelian [Bai85, BM82, EKC85, Kob85, Wim87]. abéliens [EKC85]. Abraham [Bar81a]. absolute [Kit89]. accelerate [Hug85]. accelerated [HY80]. Accelerating [EN89]. Acceleration [CM82, YJ80]. acceptance [Bon85]. accounting [Zen89]. Accuracy [Pai80, Sco85]. accurate [dG82]. acyclic [BH84, Dua89, Her86]. Adaptive [Pol89]. Addendum [Ano86a]. addition [Ma85a, Ma85b]. additional [ST88b]. additive [Rom86, SK84]. adic [Tho85]. adjacency [Bak87]. adjoint [Cro81, Hoc87, Mar84a]. adjoints [YB81]. Adjustment [GP80]. admissibility [BM89a]. Admissible [BM85]. Advanced [Cam84b]. affine [FH80, KL87, LP82b, Tch89]. Aggregation [CM82, MP80, MS83, Man84]. agree [Haa84]. Albert [Tho83a]. Alefeld [Ral86]. Alexander [Bru84c, Car85, Gau83b, SDGT81]. Alfred [HM84b]. Algebra [AB85, Ano81u, Ano81v, Ano83l, Ano84q, BV89, BS89, BKR80, Bun87, Cai82, Car85, For86, Gov89a, NdOH89, Pul87, Sha89, Sta81, dO83, LAA68, Ano80n, Ano80o, Ano80p, Ano81p, Ano81q, Ano81r, Ano81s, Ano81t, Ano82n, Ano88-28, BBV88b, BF83, Bur86, BCS87, CW85, CMR81, CZ84, Dad87, DALSP89, Had86, MR84, Mor85, Neu82b, OR88, RWZ88, Tre85a, Zas83, Gov88a, Gov89a]. Algebraic [DM82, Jód88a, Kra81, VV87b, Adk89, Ano80q, Ant83, AZ83, Bye87a, Chu88, Chu89, Cop81, Joh87, Ker83, KÖ85b, Lin87b, LPC89, RR84b, RV88, Sch83b, Sha86b, Tau80a, Tau80b, Uch86, VB88, Vas85, Wim84, Wim85a, Zim80]. Algebras [Laf81a, AL89, Bar89b, Brü84d, BC85, Cos89, Far84, Fei84, Gam88b, Gro82, Gur85a, HG88, Hel87, Hig87b, Hol87, Isa80, Laf81b, Laf83b, Laf85a, Lei89b, Lew82, Mac89, MM89, MCCF85, MCCF89, Per86, Per88b, Per88a, Que87a, Ros84b, Sha87, SM88, Tan86, Tom85, Tow80, VM87, WB81, dGH83, Wat80]. algèbre [Tre85a]. algèbres [Hel87, MM89, MCCF85, MCCF89]. Algorithm [APP88, Bai88, BV88, BG84b, CV87, O’L80a, Pai80, PV86, Sim84, dH87, ACHS89, Ant88, AS80, BN83, Boj86, BG81b, Dan89, DD87b, Giv82, Gow80, HM89b, Hu87, JZ85, JM89b, Kle87, Lin87a, O’L80b, Par80, Ram86, Saa86, She87b, Zha89]. Algorithms [Bun87, CS86b, Gam88b, JY86, SHW86, Woź80, AL86b, AFW81, BLAK89, DL89, EHL81, ER87, GKKL87, GM88b, Hen83, JY87, Ker83, Naz84, NM85, OS89, Pow87, Mon84]. Allen [Mon84]. Allowable [SS88]. Almost [TvdW88, Väl89a, Adi82, CH83a, Rho89, RS86b, Sha83]. Alston [Ano80a]. alternate [CL82b]. alternating [GL84]. alternative [FH80, Man81]. always [Ogr88]. among [Vei86]. analogies [Beb86]. analogue [GH87, Sun82b, Wim88]. analogues [AP83, AP88, RCS89]. Analyses [GLN86b]. Analysis [APP88, Ano88-32, EL88, Kie87, OvdD88, Sim84, Woź80, AL86b, ACC88, Bjö87, Bro85b, BG81b, CD85, CLP82, Dry85, FHV85, Joh88, Ker87, Man84, May88, May85, Neu81a, Neu84, Ric83, SS85, Sha85f, Sha89, ST88b, TvdW88, And87a]. analytic [Gor89]. angles [Cve88]. angularity [RC81]. anisotropic [Jam88, Jam83]. annihilated [How80]. Announcement [Ano81a, Ano82a]. ANOVA [Tak85]. Anti [GS84, Sou86a, Wie85a]. Anti-Hadamard [GS84]. Anti-invariant [Sou86a]. anti-Morishima [Wie85a]. Antieigenvalues [Mir83]. Antiinvariant [Sou86b]. Antipodal [Gar82a]. antiprisms [Bro85a]. Any [Nak84]. anyway [Car86]. Apollonian [Wil82a]. appear [Wie85c]. Appl [AB85, BS89, BKR80, Cai82, Sha89, Sta81]. Application [MS80a, Mon84, BR89a, Chu84, Dio89, FT89b, FHV85, GR87, HF85, Hun82, SS86, Spi82, Ste85, TG87, Usm87]. Applications [Bap86a, BV89, Har82, Jin88, Mar84b, NdOH89, Oko86a, Oko86b, RR84b, Rot81, AT80, LAA68, Ano82n, Ber86c, BF83, Bru88b, Cla88, Dah83, Dat84, Dat88, EHL81, Far87, FH88a, Gad88, GW85, GT82, HHM89, HLG85, Kön86, Lin89, MP82, Mne89, MG85, MM88, Özg86, Ran89, RY85b, Rei86, SH88, TC86, Uhl87, Vei82, VV86, Wer83, Wol81, Ano81u, Ano81v, Ano83l, Ano84q, Gai89, dO83, Bru84c, Gov89a, Ple83, Pul87, Tho83a, Bar87]. Applied [NdPV84, Sch84c, Hun82, Neu81a, NV84, Cam84c, Car85, Gov88a, Gov89a].

Approach [Chu86b, VV87c, BFM89b, BFM89a, Dio89, GT89a, KGÖ83, LA88, Man82, MW84, Neu80, OM89, RY85a, Sha86b, Shr86, Wim89c]. approaches [Mal89, Mur89]. Approximate [Lip86, AP86, BS83, HW89, Zie87]. Approximating [Van84, HRR87]. Approximation [And85b, CM82, GHS87, Gro80b, HW88, Chu86a, Que83, Wei83, dBP82, dB84]. Approximations [Flo82, LR81, LM87, NS89]. approximators [Far85a]. arbitrary [BS82a, DGK83, Fia83, Har84b, Lei89a, Lin88, VD83, Won82b]. architecture [HTRG88]. area [Mor85]. arising [Acz85, GM89, MS86]. arithmetic [And83, HTRG88, Vas85]. arithmetic-geometric-harmonic-mean [And83]. arithmetical [Hel87]. arithmétiques [Hel87]. Arnoldi [Saa80]. Aronszajn [BF88]. Array [ES86, Luk86, Urs82]. Arrays [SHW86, DR80a, HT84]. Aspects [Ruh83, Cop81, KÖ85b, Tau82, VD83]. Assignment [KNC89, OS86, BR86, DM83b, KG89, Rys85, ST88a, SY89, Zab89]. associated [BJL89, BR86, Che83d, DD87b, Dad87, FM81, Hun88, Ish87, McD80, Mer80b, Sai88, ST88a, Stu87, TW89, Tur85, Uch86, WE89, Wie85a, Wim83a, vB85]. association [Ban82, Bie84, IS89]. assumptions [BN88]. Asymmetric [BJ86, Sch86c]. Asymptotic [Lin89, RS86b, Sza85, BS86b, May87, Nel85]. Asymptotically [BA80]. Asymptotics [Bas85]. asynchronous [Lei89a]. Athens [AL86a]. Auflösbarkeit [Nic82]. August [AL86a]. Author [Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano81b, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81c, Ano81i, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano82g, Ano82h, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano83a, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84i, Ano84f, Ano84g, Ano84h, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano85i, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86n, Ano86j, Ano86k, Ano86l, Ano86m, Ano87j, Ano87a, Ano87k, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano87i, Ano88b, Ano88m, Ano88n, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano88h, Ano88i, Ano88j, Ano88k, Ano88l, Ano89a, Ano89b]. Author [Ano89c, Ano89d, Ano89e, Ano89f, Ano89j, Ano89g, Ano89h, Ano89i]. automata [Cob86, Rob80a]. automates [Rob80a]. automatically [Hug85]. Automorphism [Bar89b]. Automorphisms [Isa80, CS89a, Dam80, Per86, Per88b, SM88]. Auxiliary [DS82]. Average [WWWS86, ER89]. average-cost [ER89]. Avoiding [Sco86]. award [Ano80a]. Axelsson [Bea85]. Axis [RO89, BS83].

B [Bas85, Cam81, Cam84d, Gov89a, Sil87b]. Balanced [Gre88, OM89, DR80a]. balancing [Zen89]. Banach [Bur86, Ste86, Tre85a, VM87]. band [BAEGL87, BK81b, BC83, Ded88, Dub85, DG81, FG80, Hal80, MH81, Mee83, Tis87, Tre85b, Usm87, dBP82]. Banded [SS87a, BF81a, Bar84b, CDMS81, Che83a, EP88, Pif87, Rom86, dBP82]. bandstructure [Lás80]. bandstructure-invariant [Lás80]. baric [Per86, Per88b]. Barnes [SW85]. Barnett [Cam84e]. Bart [Lan81]. Bartels [Pow87]. based [BFM89a, CW80, DGK80, Kam89, LA89, LVWC88, Naz84]. bases [BI88, BH81b, LP82a, Lem88, Mer81b, Whi80]. Basilevsky [Car85]. basis [Bar80b]. basis-exchange [Bar80b]. Bauer [Li88b]. Baxter [Faĭ89]. Bayesian [CD85]. BC [OvdD88]. be [Gam88a, MS85]. behaved [Ogr88]. Behavior [Gre89, AZ83, Naz84, Rob83, vV87a]. behind [Wat88b]. belonging [Rot89a]. Bergman [Gha87]. Berlekamp [JM89b]. Berman [Bar81a]. Bernard [For86]. Bernhard [Sch84c]. Bernstein [Cos89, MM89]. best [LR81, Mat85]. Between [VV87b, Bak86, Beb86, Bol85, BN89, Cri88, Djo81, Djo82, Gar82c, GH83, GR86, GA85, HR84, Her89, JE87, KS84, Krä87, LT83b, LT88b, MZ89, MFPW85, OP82, Pol89, Ran89, Ste81a, Tan86, Tau85, Uhl81b, Uhl83, WB81]. Bézier [GM88b]. Bézout [BS88a, Fie88a, FP88, Hel89, Ptá84c]. Bézoutian [LT88a, SA89, PW85]. Bézoutians [FP87a, FD89, HF89, Ptá84a, Ptá84b, Wim89a, Wim89b]. bi [dBJP82]. biadditive [Ng87]. Bidiagonal [Chu86b, HI86]. bidirected [Hur89]. biflag [HS87]. bigger [Jam81]. biinfinite [SW84]. bijections [KL87]. bilinear [AFW81, Fan85, FH80, Gow81, LN89, Moo82, Tsi84, Wat87b]. bimatrix [CMO89]. binary [Cun82, Eli88, GT89b, KK89, KL88a]. Binet [HCR88, Nis88]. binomial [Str87, Wat88b]. Binyamin [FL89b]. Biorthogonal [Ptá83a]. Bipartite [GLN84, Ito82, Rot89a]. bireflectionality [Nie87]. Birkhoff [BS82a, FSZ85, SSZ83]. birthday [Ano85s, EH85, FL89b, Gau83b]. bit [Eli88]. Biunitarily [GS89]. bivariate [GM89]. Block [AC89, BCR88, CS86b, Els84a, Fre87, NdPV84, O’L80a, SHW86, Vit82, Wat80, ABI84, AP86, Axe86, Bhi88, BF89, ELS82a, Eli88, EGL85, FH88b, Gor89, GP89, HPS83, HPS88, HI86, HRS89, KK86, Laf81c, Pol87, RCA88, SH88, Tis87, Wei89b, Wim89b, Woe89]. Block-iterative [AC89]. block-matrix [ABI84, Axe86]. block-Toeplitz [Tis87]. blocks [Cai81, dO82]. blockwise [Pol87]. BLU [Wer85b]. board [Ano80k, Ano81o, Ano81n, Ano86s, Ano88-30, Ano89s, Ano80h, Ano80i, Ano80j, Ano80l, Ano80m, Ano81j, Ano81k, Ano81l, Ano81m, Ano82i, Ano82j, Ano82k, Ano82l, Ano82m, Ano83g, Ano83h, Ano83i, Ano83j, Ano83k, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano84p, Ano85k, Ano85l, Ano85m, Ano85n, Ano85o, Ano85p, Ano85q, Ano85r, Ano86o, Ano86p, Ano86q, Ano86r, Ano86t, Ano86u, Ano87l, Ano87m, Ano87n, Ano87o, Ano87p, Ano87q, Ano87r, Ano88q, Ano88r, Ano88s, Ano88t, Ano88u, Ano88v, Ano88w, Ano88x, Ano88y, Ano88z, Ano89k, Ano89l, Ano89m, Ano89n, Ano89o, Ano89p, Ano89q, Ano89r]. Boas [Bea87a]. Book [Ano88o]. Boolean [BP84a, BN83, CZ84, MR84, RR82, SL87, dCG81, Ple83]. Boolean-rank- [BP84a]. Boolean-rank-preserving [BP84a]. Border [Bin86]. bordered [Nom80]. Bose [Cam84c, Son87]. Böttcher [Bas85]. bottleneck [BCS87, Sco86]. Bound [Els85, PNO85, Bap86b, DEHS84a, DEHS84b, Joh89, Jon84, Neu81b, Ron82, Sta87, Szu88, Szu89b, Tao84, You81a, Yua88, dG82]. Boundary [Jód86, Jód87, Jód88b, MS86]. bounded [Dan88, ORV88]. bounding [Sco85]. Bounds [AL80, Fri88, GS87, HR84, LO81, Lew89b, MSW83, O’L90, Pow89, Rot80, WS80a, BH81a, Bea84, Bea87b, Ben86, Cao87b, Che83b, Con85, Ded88, Dem86a, DW83, JKW85, Lás82, Leó88, Lew89c, LA88, MŽ83, MW85a, Pat88b, RT85, Sch86c, WS80b, Yua89]. Brauer [HM84b]. Brenner [Bea87a]. Brewer [Nau87]. bridge [AB87]. Brownian [Gov88b]. Brualdi [MW87b, Min87a]. Bunce [Nau87]. Bunch [Cam81]. Burnside [Ros84b]. burst [Was84].

C [Cam81, Cam84c]. cône [Yca82]. Calculation [KG83, Mir83, Ste81a]. calculations [Fos86]. Calculus [Bru84c, Ber86a, Pol85]. Callier [Cam84c]. Campbell [Poo81, Mar84b]. Canada [OvdD88]. cancellation [GL82]. Canonical [BV88, BI88, PW83, Wei84, BGoMSTAUTA^{+}86, Har81b, HPW89, HH88, Lin88, ML85, OM89, VB89]. Cao [Pul87]. Capelli [And85a]. caps [Kes82]. Carl [Poo81]. Carmeli [Nor84]. Carolina [Neu82b]. Cartan [Zas83]. Cartesian [HT84]. cascade [GS87]. case [BGoMSTAUTA^{+}86, BNV82, DEHS84b, Hil82, PY82, Uhl81a, Uhl81b, Uhl83, Zie85]. cases [Tho88]. Castle [BP81]. Categories [RPV87]. category [SK82]. Cauchy [HM81, HCR88, Jód87, MM81]. Cayley [Bus86, OR88]. cell [HHM89, Tak85]. cellulaires [Rob80a]. cellular [Rob80a]. central [LM87]. Centro [Lee80]. Centro-Hermitian [Lee80]. Certain [BB86, Gro80a, BdP86, Beb86, BHJ85, Cri88, Ded88, Deu80, FM86a, FG80, HPS83, HPS88, HW81, HMS87, Jef86, Joh84, Joh87, KB89b, KR81a, Kwo88, LO81, Loe88, Loe89b, Lon81, NS85, Pat82, Pat84, ST88a, Son88, Vei86, Wie82, Wie85b]. Cesàro [Spr81]. CG [EL89]. Chain [FM86b]. chains [Fry85, GM85, Har87b, HS84, Hun86, IS82, Koš88, Sen80, Wim89d, Mey84]. change [Bha81, MP83]. chaotic [BEN88, BF89]. Characteristic [Wan83, Bry82, DvdP81, MS80a, Sha82c, Sha82a, Tau83, TC86, Uhl81a, dO82, vB85]. characteristics [HS89, Her89, LT83a]. Characterization [BH84, KS84, LR81, Lem88, Bea86, FM88b, Ger89, HM84a, HCR88, MSS89, Nom80, Zie88b]. Characterizations [FG81, Fie88a, Fie88b, HRS88, Hun88, PZ88, RS82, Won82a, Des88, Gow85a, HS86a, Özg86, Wer85a, Wer84]. characterized [IS89]. characterizing [Gov89b]. Charakterisierungen [Wer84]. Charles [And87a, Ste87b]. cheap [Gee89b]. cheap-control [Gee89b]. Chebyshev [Kam89]. chief [Ano88-27, Ano89t]. Chipart [FD89, Jur82]. Choleski [Sco86]. Cholesky [GHL86, Kie87, Olk85]. chordal [BJL89]. circle [EG89, Rot80, Sco85, Wat88a]. circuit [Cla88, DSD88]. circulant [Bel81, CD83, Cha88, Dub85, Lon81, MW87a, Mad83, RCA88, ST88a, ZM81, Hay82]. Circulants [Fei84, CZ84, Chi87, Chi88, Ma84, Min85, Min87b, Par82, Tau85]. Circular [Bar89c]. circulations [RS82]. Class [Mar81a, Woź80, Bir81, BM89c, BC85, But88, Chi87, Chi88, EG89, Har81b, Kay85a, MV88a, O’L80b, RSR89, RU87, San87, Wan89, Wie84, Zie87].

Classes [BR86, FMN83, BHJ85, BH82, BL82b, Chi86, DR83, DR88, Gow81, HM86, HMS87, Joh87, KH84, Kit89, Meh84b, Mer80b, MW81, Mer81b, Pat82, Piz83, Tam87a, Wie82, Yan84]. Classical [VV87b, Cal82, Ish83, Lei89b]. classification [ALZ81, Lew82, Wat82b]. classifications [HRS88, KS86a]. Clebsch [ML83]. close [BAS86, vV87a]. closed [Lin87b]. closed-loop [Lin87b]. closest [Kas89]. Closure [Joh87, FM88b]. clustering [TG87]. Clusters [CV87]. cocircuits [Cun82]. code [KL88a]. Codes [Cal80, BPB88, Eli88, PR88]. codimension [FS82, Wat84]. coefficient [Str87, Wat88b]. coefficients [BS86b, ML83, ML85, Nel85, RP85a, Sen84, Yca82]. Coherent [Hig87b]. Coimbra [GL85a, dO83]. cokernel [MR83, MR88b]. Collatz [Ano85s, EH85, FN89, TW89]. Collinearity [VV87c]. color [Jef88]. colorability [DF87]. colorations [PS82]. column [BP88b, Bru80, BM89c, Fos86, PS88, Roh81, RS89, YL85]. column-structured [YL85]. Com [Con81]. Combination [CM83, Mar84a, Nak84, Nis85]. combinations [Cha87b, LA89]. Combinatorial [Ano88-32, BH87b, Bru85, Joh88, BG81a, Car88, CCH89, HRS89, May88, Mur89, Rys80, Shr86, OvdD88]. combinatorially [CH89]. Combinatorics [BB82a, Ano80q, BB82b]. combinatorics/graph [Ano80q]. combined [JM89a, Neu80]. Combining [Cal80]. Comment [Ano85j]. Comments [Jur82]. Common [CP85, Koš88, Lan86a, She84, BL89b, DW89, She87a]. Communication [Saa86]. community [Luk81]. commutant [Hol82]. Commutation [GS81, DvdP81]. Commutative [For86, Nau87, dGH83, BFM89a, BK88c, HH85, Isa80, Laf85a]. Commutativity [Rad84, CL82c, CJR87]. Commutativity-preserving [Rad84]. commutator [ET81, Laf80]. commutators [Fon86, Tau80a]. commute [Eva85, Gof81, Wea83, Wea88]. commuting [Grü81, LA89]. compact [BPR87, GS89, Kön86, Lei89b, ST87b]. companion [Ell87, Gał81, Gov89b, Har87b, HJ86b, HI86, LLT84, Ptá83a, Ptá86b, Wer83]. Comparative [Gov89a]. Comparison [And89, Mee87, GA81, Neu86, RV88, Ste85]. compatible [Zha87b]. compensation [BFM89a]. Complement [Tru80, Bre88, Fie81b, Ima84, JMRT87, Mad88]. Complementable [CH83a]. complementarity [Aga81, Ber81, CPV89, JP87, KP80]. Complementary [BH88, CM81a, Lem88, Mur82]. Complementation [LT86]. complements [BF88, BM88, Car86, Oue81]. Complete [BCS83, BCS88, Ber88, Vin89, BH88, Bea86, Sil88]. Completely [PH81, BHH84, BH87b, Coh86, CMO89, HL83, Loe80]. Completing [FM86a, DM86]. completion [ACC88]. completions [BJL89, FM87, GJMW84, Woe89]. Complex [Dam80, PS87, Tad86, BHH87b, Djo81, Gro80a, Had84, Lon84, PL86, Str85, Stu88c, TST86, Uhl83, Wat82b]. Complexity [ISS86, GKKL87, Saa86, Vav89]. component [Des88, TC86]. components [HK83, Rod80]. comportement [Rob83]. composite [TG87]. composition [Sis86]. compound [Li86a, Vit82]. compounds [Min87b]. Computation [BV88, FG86, Ruh84, AC89, Eli88, GM89, KMK85, Lin88, May87, RP83, San87, Str83, Gro85]. computational [BC83, BK80c, DR86, VD83]. computations [Dat89, Gad88, Val88a, Vav89, Ral86, Ste87b]. compute [AR87]. computed [BG82]. computer [DM86, Sco86]. Computers [DKH86]. Computing [Bai88, CP81, Dem86b, DK87, Hig87a, Hig88, Luk86, PV86, Saa80, Sen80, AL86b, BN83, BG81b, GSV86b, GT89a, HS86b, Lin87a, Lin87b, Tao84, TST86, Usm87]. comrade [Bar84a]. concept [Beb86, MP82]. Concerning [Har80, Chu89, ET81, Kha83, Tau80b, Tau85, Tho83c, Wie82, Wie85a, Wie85b]. Condition [Dem86a, Gau83a, Van87, WWWS86, Zie88a, CR86, Els84a, FHP80, FT84, Sch86b, Tad81, TC86, Wol89b, Zie83]. conditional [Mar81c]. conditionally [MPJ81]. Conditioned [ES86]. Conditions [Eva87, Gam88a, Hal86a, HT84, JY86, LVWC88, MS85, Bec80, BK81c, BS86b, GM80, Jód88b, JY87, KG89, KB89b, Kom83, Sch81a, Sch89a, Syl86, Wim84]. Cone [BNS86, Gow85a, Har89, BT81b, HM84a, HW87, LT86, Mar84a, Pol86, RV89, Ste87a, ST87b, Tam81a, Tam81c, Tam85, TW89, Yca82]. cone-preserving [TW89]. Cones [Pol86, Adi82, Bar81b, BHH84, CP89, Eav84, NS89, PL88, Ste82, Väl88c]. Conference [Ano88p, Ano88-32, BU84, BU88, BCS86, BHL89a, BHL89c, BV89, Bru88a, GL85a, Gai89, Hol85, KLM86, NdOH89, OvdD88, dO83, BF85, Gro85, Laf85b, Lew85b, Neu82b, O’C85, SW85, Tim85, Wim85a]. configurations [Bry82, McD80]. conflict [Pol89]. confluent [HG80]. conformal [Hen83]. Confounding [Kob85, EKC85]. confusions [EKC85]. Congenial [Bar81c]. Congruence [BY80, YB81, BPR87, HHJ86, Tau83]. conjecture [BdP88, BWT89, BKR80, CL81, CL89, FW84, Hof85, Hwa86b, Laf83a, Lon82, MS80b, McD80, MV89, Min87a, PN85, RP85a, Sha85b, Sha82c, Sha82a, Zha87a, dPB86, vL81, Hwa87]. conjectured [Wat83]. conjectures [Hwa86a, Mor88]. Conjugate [Axe80, DT87, Gre89, HM84a, JY83, JY86, Mon84, O’L80a, Vei89a, Woź80, YJ80, CW80, GB84, JY87, MC81, NM85, O’L80b, Sch89a]. Conjugate- [JY83]. Conjugate-Gradient [Gre89, JY87]. conjugate-Toeplitz [GB84]. conjugation [Dam80, GS81]. Conjunctivity [BY80, YB81]. connectant [Tre85a]. connected [Neu85, Rod80, dBS86]. connecting [Tre85a]. connection [Ber86b, Gab87, HY80]. connections [MZ89, Tau85]. connectivity [Bix82, LSS89]. conquer [GT89a]. Considerations [AGG88, Car88]. consimilarity [HH88]. consistent [BI82, Neu80, Wei89a]. constant [BPR87, FHP80, Gen86, KG89, LN89, PS88, SL82]. constants [FT88, RWZ88]. Constrained [GGvM89, Tsi84]. Constraints [HS89, Jai84, MJ81, Poh89, RS82, ST88b]. constructed [MV88a]. Constructing [Cve88, CM84]. Construction [BK81b, Dua89, EKC85, VV86, ACHS89, MH81, dG82, EKC85]. constructions [Pan89c]. constructive [CK89]. containing [Bar89c, Laf81c, Meh84b, Son88]. context [Uch86]. contiguous [DGK80]. contingency [Joe85]. continued [VB89]. Continuity [Gee89b, Bur86, CH80, Con81, Kit89, Pok85]. Continuous [Sak85, Bro89b, Chu86a, ER89, RV88]. continuous- [RV88]. continuous-time [ER89]. continuously [Fer81]. contractants [Rob80a]. contractibility [BCS83, BCS88]. Contraction [SZ86, CZ86, Ver84]. contractions [Fan88, SZ89]. contractors [Rob80a]. contribution [Szu89a]. contributions [Bea87a]. Control [Bar80a, Cam84e, KNC89, Ano81w, BF83, Cla88, Dat84, Dat89, Faĭ89, Gee89b, GdHZ89, HS89, Har84a, Har89, HS83, Meh89, Pol89, Tch89, TvdW88, Zab88]. Controllability [Car84, Pan89b, dSB81a, dSB81b, BCS83, BCS88, CS87c, HG80, Uhl87, WH85a]. controllable [CS87a, KP89, MK89]. controller [LA88]. controllers [BKH88]. controls [Gee89a]. conventional [JE87]. Convergence [Bai88, BF89, FM89, Lei89a, MNP85, NP87, NdPV84, PV86, Tao84, Tho89b, Art86, Hac84, HPY86, HN88, Har86a, Hug85, Kri84, MS83, Man84, Mar86b, MK88, May84, May85, May87, NV84, Neu81b, SL87, You81b, Zha89, vV87a]. converse [Li88b, Tho83b]. Convex [CL86, AC89, BPR87, Cha87b, CvRS81, Els82b, HK83, Mar84a, Ste87a, SK82, SK84, Sun82c, ST87b, Tam85, Vei89a, Zăl85]. Convexity [Bon85, Nus86, Els84b, LP81, Mar88]. Copositive [MJ81, Gow89, Had83, Väl86, Väl89a, Väl89b]. coprime [OM89, SO89]. Corner [GM88b]. Correction [Sta81]. Corrections [Cai82]. correlation [BS86b, ML85, Nel85, Sza85, Wat88c]. correspondence [Hel89]. corresponding [FP87b, HL84]. cosine [CKN85]. cosine-sine [CKN85]. cost [ER89, Gee89a]. cotriangular [Hal83]. countable [IS82]. counterexample [Laf83a]. Counterexamples [Mor88, Tho83c]. Counterparts [Hol85]. Counting [Mer80a]. coupled [JAK89]. Covariance [Rob85, BS86b, BN88, KR83, Nel85, Rob84, Sch83c]. covariances [Wer85b]. cover [Ant83]. Cramer [Ver82, BI82, Wan86, Wan89, OR88]. Criteria [Gar82b, Väl86, Car88, CH89, Fer81, Mar81c, Väl89b]. criterion [BM89a, FD89, Jur82, Laf81b]. crossings [Sch89b]. cubes [Rei87]. cubic [MP89a, Rei87, Sha82b]. curve [Adk89]. curves [Bre84, BF82, GM88b, MP89a, MP89b, Sch89b, Vin89]. cut [BR89b]. cutting [CCH89, GM88b]. cutting-plane [CCH89]. Cvetković [Bru89]. cycle [CP89, Kun81]. Cycles [MV88b, GT89b, Tho86]. Cyclic [Mad83, Tho89c, ELS82a, Har86a, Stu88b, VNC84, WN87].

D [AB85, Bru89, Gre83a, Poo81, RW89]. D. [MS80b]. Daniel [Gov89a]. dans [EKC85, MCCF85, MCCF89, Tre85a]. Darroch [BR89a]. Data [FG86, BK81b, Dua89, MH81, Sza85]. David [Bar80a]. Davis [Hay82]. Decay [EP88]. Decentralized [KW89a, VV86, Zhe87]. Decomposability [Pat85]. Decomposable [Moo82, For80, Gam88a, Li86a, Tam87c, Tam87b, Tam87d, dOdS83]. Decomposition [Bai88, Chu86b, CD85, GL87, GP80, Liu88, PV81, Pai85, PV86, Chu86a, DGK80, Fos86, FP81, Gol86b, GB87, Gur81a, HH89, HHM89, MPJ81, Meh84c, Mur83, Nar86, Naz84, Olk85, RCA88, Stu87, Uhl81a, WH85b, You83]. decompositions [CL82a, wEC87a, wEC89, Fie84b, Hos86, Ler80, LVWC88, RWZ88, Tim85]. deconvexification [ER89]. decoupling [Dio89]. decrease [Loe89a]. dedicated [DHLW85]. Dedication [GDM87]. deduced [Sai87]. d’effets [EKC85]. deficient [DS80, MN87]. Definite [CM83, DI86, And83, CH83a, CHP88, CF86, GJMW84, HM84a, Hu87, JE87, MM88, Mur82, Rob82, Sai82, Sha85e, Ste85, Ste87a, Sun82a, WH85c, Wu80]. Definiteness [CC84, Kwo88, MJ81, Mar81c, Väl88c]. deflation [Cao87a, Che83a, Lás80]. Degeneracy [Fle88]. degenerate [FS82]. degré [Tre85a]. degree [BS88a, Ber86c, BN88, Dow85, Far85b, LA88, PS89, Rei86, Tre85a, Web86]. Deif [Cam84b]. Dense [ISS86]. Dense-Linear-System [ISS86]. density [SD80]. Departamento [GL85a]. Department [BU84, BU88]. depend [Eva87, Sch84a]. dependence [Fia89, GM85, Joe85, Wol89b, Wol89a]. dependent [Avr89, Kam89]. derivation [NM85, Nel85]. Derivations [MCCF89, Hol87, LT89, Per88b, MCCF85, MCCF85, MCCF89]. derivative [Eva85, Gof81, Rob83, Sti87]. derivatives [DN85]. derived [RSR89]. Dérivée [Rob83]. description [RU85, Vin89]. descriptor [MK89]. design [Bai85, Tor86b, Bar80a]. designs [DS89a, EKC85, Hos86, MFPW85, RCS89, Shr86, Wil82b]. Desoer [Cam84c]. detectability [Hau83]. detection [ACHS89]. determinacy [JORvdD87]. Determinant [Tis87, BS86a, BQ85, BdP86, HG80, MM81, Mne89, vzG87, Mne89]. Determinantal [And85a, BJ84a, BJ86, BJL89, Lav80, Lew85a, MV89, Sty81, Tho83b, Vin89, Wat88c]. Determinants [Loe83, Loe84, Chu89, GM89, GA85, KS80, MG85, Vei82, VJ88, dPB86]. Determination [AN83, DN84, Gov89b, LP82a]. determines [Tho89a]. Determining [CS87b, Väl88b, PL86]. deux [Mne89, Tre85a]. development [Vei89a]. diagonability [Bec80]. diagonal [Aga81, BR80a, Cai81, Che83d, ELS82a, FJMN85, FM87, Gar82c, Gen86, HS85, HHJ86, KW89a, KH84, Loe86, Lor81, SL82, SH88, Tho83c, Vav88]. diagonalizable [Fia83, GS89]. Diagonalization [Brü84d, PV82, RCA88, Adk89, BC84, Mit82]. diagonally [ES80]. diameter [BL82b, Gar82a]. Dichotomy [MC81, BAG89]. difference [BC87, LM87, Lip86, MV88a]. different [DALSP89, Mur83]. differentiable [Fer81]. Differential [Chu86b, Ros84a, Acz85, BNS86, Dah83, Jód86, Jód87, Jód88a, Jód88b, JAK89, LS85, NS89, Pol85, Spi82]. differentials [Won85]. differentiation [Ste86]. digraph [Sha85c, YB88]. digraphs [DSD88, God82, Ma85b, Sme81, Wie85a]. dihedral [Wan81b]. Dimension [Loe88, Laf85a, Sou86a, Sou86b, Syl86]. Dimensional [Bar80a, CLP82, Gre83a, BG81a, CH80, Con81, DG87, FS85, Gam88b, GG82, Gro82, Hil82, LR81, Mac88, Neu81a, Ong82, Ste81a, ST87b, TW80, Wil83]. Dimensionality [SW84]. dimensions [JS89, Mat85]. Direct [AFW81, BD80, MNP85, RWZ88, SS87a, CH83b, DM86, FW84, Neu80, Pat84]. Direct-iterative [MNP85, Neu80]. directed [Tho86]. direction [NM85]. Directions [DT87, May88]. disc [Nic82]. discovering [Wat87c]. Discrete [Che87, NS89, Rob83, AN83, BAG89, BH89, BK86, Caf81, Gad88, HS89, Ish87, KBM86, KB89a, Lin87b, LPC89, Piz83, RV88, Rob83]. discrete-time [BH89, HS89, KB89a, Lin87b, RV88]. discretized [Sch86b]. discriminant [GD84, Men86]. discriminant-like [Men86]. discriminants [Bap89]. discs [Bar89c]. disk [SZ86, SZ89]. dispersion [Des88, Neu85, OP82]. Displacement [GB87, SSS82]. displacement-rank [SSS82]. Dissection [GP80]. dissipative [BdP86, Fan88]. Distance [Wei89b, BAEGL87, Cri88, Far87, GH83, GR86, Gow85b, Ito82, Krä87, Mat85, Neu89, OP82, SPZ82]. Distance-regular [Wei89b, Ito82]. distances [GH83, HR84]. Distributed [NW87, Pan89b]. Distribution [FM86b, AO85, Bie84, BH81b, KL88a, McK81, Neu85, Par86, PL86, Sen80, Sen88, She89, TC86]. distributions [AN83, Chi86, Hun86, KR83, NS85, Sza85, Was84]. disturbances [Far85a]. Dittert [Hwa87]. divergence [NV84]. divide [GT89a]. divide-and-conquer [GT89a]. Division [Bar84a, GS81, HP81, Lew82, Mac89, SM88]. divisor [BL89b]. Divisors [MO87, Bru85, Lan86a, McF80, MW81, Rod80, Wat87a]. Djoković [MS80b]. do [Haa84]. Doković [Hwa86a]. domain [BLAK89, HN88, TvdW88]. domains [BB89a, CL86, NV84, PV82, Rao83]. dominance [Aga81, Che83d]. dominant [FZ84]. Dongarra [Cam81]. Doob [Bru89]. d’orthogonaux [EKC85]. Double [PL82, CH83b]. Doubly [And89, Bap83, BS82a, Mar84c, BR80a, Bru84a, Bru88b, CF86, For80, Fry85, Gib80a, Hwa86a, Lon81, SO89, Gib80b]. Douglas [CD83]. down [KG89]. Drazin [Wer85a, AR87, BHH87a, CG80, GH84, Har81b, JG80, Pye80, Ron82]. Drazin-monotonicity [Wer85a]. Dual [FH80, Zen84, Eav84, Haa84, IS89, LP82b, Sch81b, Spi85]. Duality [LT83b, LT88b, Wim87, Zim80, Ban82, Dix81, Kob85, MFPW85, Sch89a, Tam85, Vei89a]. Dublin [Gai89, BF85, GL85a, Gro85, Hol85, Laf85b, Lew85b, O’C85, SW85, Tim85, Wim85a]. due [HN88, Sak85]. Duff [Ple81]. d’une [Rob83]. duplication [Per88a]. during [Dan89]. Dynamic [BFM89b, BK88c, DM83b, KW89a]. dynamical [Mur89].

E. [Hwa87]. Eckart [GHS87]. econometric [SS86]. ed [Gov89a]. edges [Fri88, Row83, Row88, Sta87]. edited [Bru88a, Cam84d, Koh84, Mar84b, Ple81, Sch84c, Son87]. edition [Har88]. Editor [Ano89t]. Editor-in-Chief [Ano89t]. Editorial [Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano80m, Ano81j, Ano81k, Ano81l, Ano81m, Ano81o, Ano81n, Ano82i, Ano82j, Ano82k, Ano82l, Ano82m, Ano83g, Ano83h, Ano83i, Ano83j, Ano83k, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano84p, Ano85k, Ano85l, Ano85m, Ano85n, Ano85o, Ano85p, Ano85q, Ano85r, Ano86o, Ano86p, Ano86q, Ano86r, Ano86s, Ano86t, Ano86u, Ano87l, Ano87m, Ano87n, Ano87o, Ano87p, Ano87q, Ano87r, Ano88q, Ano88r, Ano88s, Ano88t, Ano88u, Ano88v, Ano88w, Ano88x, Ano88y, Ano88z, Ano89k, Ano89l, Ano89m, Ano89n, Ano89o, Ano89p, Ano89q, Ano89r, Ano89s, Ano88-30]. Editors [Ano88-27]. Editors-in-chief [Ano88-27]. Effectiveness [Pai80]. effects [EKC85, KL88b]. Efficient [GKK86, PW89]. Egoritsjev [vL81]. Egorychev [Bap84a]. Eigendecompositions [Dem86b, DK87]. Eigenelements [Saa80]. eigenfunctions [FB81b]. Eigenproblem [Pai80, DM86]. eigenstructure [Lin87a, VD83]. eigensystem [GT89a]. Eigenvalue [AGG88, BG84b, CFKK82, DKH86, ES82, GGvM89, HMS87, JR81, Ruh84, SHW86, Ash87, BF88, Bea84, Bea87b, BK81c, BB89b, Cao87a, Cao87b, Chu84, Chu88, Dia86, DGKW84, Doo82, DL89, Faĭ89, FG80, Fri85a, HRR87, Ker83, LS87, MS87, McK81, MW85a, Neu82a, Roe86, Roe87, Rot89a, SD80, She89, Tor86a, Tre85b, Uhl81b, WE89, Yua86, vB85]. Eigenvalues [And89, BG82, God82, JvdD88, Jef88, Kön86, PNO85, SWR87, Van84, Van87, AB84, AB85, BH87a, BH81a, Bar89c, BQ85, BG81b, But88, CH83b, Cve88, DD87a, Doo89, EP87b, Fri86, Gar82c, GS87, Gen86, Gow80, HP82, Han86, HR84, How83, KRS82, Lan86a, LO81, Leó88, Lew89b, Lin86, Lin87b, MSW83, Pow89, Que87b, Rot80, RT85, Sch81a, SS85, Sha86a, Sha83, SS86, Sil87a, TZ87, Usm87, Wea83, Wea88, Wim85b, WS80a, WS80b, Yua89, dG82, vV87a, Vit82]. Eigenvector [Mey89, Ash87]. Eigenvectors [Van87, CP85, CS85, Giv84, Har84a, Rot89a, She84, YH82, Zij84, Pow88]. electrical [BBV88b]. element [HP86, Leu86]. Elementary [MW81, Wat87a, Bru85, DSD88, McF80, TZ87]. elements [AZ83, BT81a, Dad87, Gar82c, Hea83, KMV87, LR81, Mer84, Tho83c, Bru84b]. Elimination [GN85, AR87, FP81, GN82, Saa86]. ellipsoids [GH83]. elliptic [Hes86, Ish87, LM87, Lip86]. embeddability [Bus80]. embedding [DGK83]. embrace [dP80]. Emilie [CMU86]. endomorphisms [Bye87b]. engineering [BBV88b]. Engineers [Cam84b]. ensembles [Rob80a]. entries [Con85, FM86a, Har84a, Mur83, Zab86]. Entropy [Elf80, Sch89a]. enumeration [Kov87, MW84]. EP [RP85b]. equal [Bap83, Bea81b, HL84]. Equality [dOdS83, CL82b, Lew89c, Rob80b]. Equation [Chu86b, AS80, BK80b, BS88a, BC87, Bye87a, CKN85, DD87b, Dat88, Don87, Eba89, Faĭ89, Fur88, HG89, Jód87, Jon84, KH84, KL87, Lin87b, Mil88, Ng87, Ptá81, Uhl87, Ver84, Wim84, Wim85a, Wim88, Zie85, dSB81a, vB85]. Equations [Axe80, BDM83, BA80, Roh89, WWWS86, Zie88a, dH87, Acz85, Adk85, Ama86, AS80, BI82, Bjö87, BKR80, wEC87b, wEC87a, wEC89, Cla88, Cro81, Dah83, DM82, GG82, GKK86, GKKL87, GR87, Hal80, Har89, Har81b, HJ86b, HS88b, Jód86, Jód88a, Jód88b, Kwo88, LM87, LT88a, Lip86, Lor81, LS85, MR84, MV88a, Mit84, Neu84, Neu87, Neu80, Nic82, Nor89, Par80, Ptá83b, RR84b, RV88, Ric83, Rot89c, Spi82, Str85, Uch86, Ver82, Wan86, Wan89, Wim80, Wim89c, Zie83, Ros84a]. equiangular [Neu89]. equiconvergence [SSZ83]. Equivalence [Ran89, Bak87, Gow81, Gur82, KK89, MM82, Pok86, Sch88, Tad81, Tak85]. equivalent [Fia83, Sha85a, WE89, Wol89b]. Ergodic [FM86b]. Ergodicity [IS82, FK85, SS81, Sen84]. Ernst [Gol85]. Errata [BS89]. Erratum [AB85, Sha89]. Error [AL86b, KMK85, Kie87, PNO85, Woź80, Zie88a, BM89a, Cao87b, Des88, Eli88, Pow87, Ron82, Ver82, Zie83]. error-component [Des88]. Error-free [KMK85]. Errors [VV87c, CM85, Fle88]. errors-in-variables [CM85]. essentially [Rot89a]. Estimates [AB84, MS86, Qi84, Zie88a, Fon86, Zie83, AB85]. Estimating [ML85, Olk85, Van87]. Estimation [CM85, Sha86a, VV87b, AO85, Kam89, Mir83, NM85, ST88b, Sza85, Wer85b, Zen89]. estimators [BM85, BM89a, RP85a]. Euclidean [Ant88, Bar88a, DS89a, Els84a, Far87, Gow85b, PW89]. Euclidian [Mat85]. Euler [BS83, Jód87]. evaluation [FT88, Nis88]. even [Mad83, Tho86]. eventually [Fei89]. everywhere [Sha83]. evolution [Ama86]. Exact [NV84]. excess [Big80]. exchange [Bar80b, Whi80]. Existence [BB89a, Meh89, RV88, Sch89a, Zab86, BT81a, GM80, Hur89, RCS89, Sch81a, Wim84]. Expansion [Ste84]. Expansions [Rot81]. expected [McK81]. experience [AG89]. experimental [MFPW85]. Explicit [HG89, Jód88b, Ptá84b, Sen84, Uhl81a, Usm87, Cyb87, dB84]. exponent [Sha85b, Sha85c, YB88, Zha87a]. Exponential [PM80, Tho88, Tho89b, Tho89c]. exponentials [CFKK82, Coh88, Wer89]. exponents [DSD88]. exposed [PL88]. expressing [Rei87]. Expressions [KR81a, Ptá84b]. Extended [ST87a, KL88a, MP82]. Extending [Dav80]. extension [BR89a, Cli81, FH88a, Man81]. Extensions [CM84, DG81, Cor87, JB85, Rao85, Sid89]. exterior [Mor85]. external [Sch88]. extra [MW85a]. extraction [HK83]. extrapolated [BF89, EL88]. extrapolation [HY80, Had84, Yey84]. Extremal [Sha85e, BS86a, DEHS84b, Li88a, Mes89, Pin85, Ptá86a, Zie88b]. Extrémales [Yca82]. Extremals [Yca82]. Extreme [Adi82, Dru87, AT88, CF86, FS85, RS82].

F [Cam84c, Nau87, Pul87, Ste87b]. face [LT86]. faces [Loe88, Loe89b, Son88, TW80]. factor [Bro85b, DM83b, May87, OM89, RCA88, Sha85f, Sha89, Zab89]. Factored [LLT84]. Factorial [Bai85, EKC85]. factoriels [EKC85]. Factorization [BZ89, EI87, EGL85, GR80, GHL86, HH83, Kie87, Lan81, Mar85, OS86, ABO88, ABI84, AP86, Axe86, BGoMSTAUTA^{+}86, BH88, BK88b, BK86, CDMS81, CG89, Dem89, GLR83, GW80b, HL83, HS86b, Kay87, Ker87, Lip86, Mar86a, PR81, Sco86, Wu89]. Factorizations [Cha87a, Laf89, BG86, CP81, Che84, Faĭ89, Gad88, GG82, JP81, KBM86, OS89, Pol87, Pow87, Rob82, SF85, SO89]. factors [GMSW87, GS82, GS83a, Gol87, Mar88, Que83, SB87, SH88, SB88, Sil87b, Tho82b, Uhl83, Zab87]. facts [Tau80b]. Faddeev [Giv82, Gow80]. Falikman [Bap84a]. families [GS89]. family [Lei81, Stu88a, Tur85]. Fan [Meh84a, Wat87b]. Fast [BA80, Dem89, KBM86, LQ87, AS80, BLAK89, NM85, TST86, Ros80]. faster [dP80]. Feasibility [GM80, AC89]. Feedback [Cam84c, BS87a, BFM89a, Bon83, BK88c, Bro83, Dio89, DM83b, LM88, LN89, Ran89, Tch89]. feedbacks [SY89, Zhe87]. Feingold [Wil87]. Fejér [ORV88]. Fermat [Tau85]. few [Ong87]. Fiedler [CL89]. field [BH81b, BS87b, Bry82, ER87, Gow81, HTRG88, KP89, Koh85, Kov87, Kun81, Lor81, MV88b, Tau80b, Uhl81b, Won82b]. fields [Bea86, Bre86, GL85a, Har84b, Ish83, Lem88, Pet82, Tau80a, Uhl83]. Fike [Li88b]. Finding [CG86, CM83, Jam81, Wan89]. finis [EKC85, Rob80a]. Finite [Bar80a, Bry82, CH80, Con81, Ger84, GG82, Kes82, Mac88, Mar81c, Neu81a, Pow82, dBS86, BG81a, BH81b, Bre86, Bye87b, CS89a, Chi87, Chi88, EKC85, FRS82, GL85a, Gam88b, Gow81, GR89b, Gro82, Gur85a, Hil82, HS84, HTRG88, Ish83, Kob85, KP89, Koh85, Kov87, Kun81, LR81, Lem88, Leu86, Lin87c, MV88b, Ong82, Pet82, Rob80a, RT83, ST87b, TW80, Wil83, Wim87, Gre83a]. Finite-Dimensional [Bar80a, CH80, Con81, GG82, Mac88, Neu81a, BG81a, Gam88b, Gro82, Hil82, LR81, Ong82, ST87b, TW80, Wil83]. finitely [Tho82b]. first [DN85, Had84, Hug85]. first-order [Had84]. Fischer [GM87, JB85, Pat87]. Fisher [DS89a]. fit [Kas89, Mat85]. Fitting [Mon84]. five [JS89]. fixed [Bea81a, BKH88, Bru80, HS87, LT83b, LT88b, NS85]. flag [HS87]. Fletcher [Pow87]. flexible [ER89]. Flow [FG86, BR89b]. flows [FT89b, Kar89, WE83, WE88, WE89]. FMDR [DL89]. Form [BV88, PL82, Bap84a, BC84, BN83, BG82, Cla88, FH88b, GM88b, HI86, HPW89, HHJ86, HH88, Lei89a, Lin88, Neu85, Rei87, Sch86a, Str87, Tho82b, Tsi84, Wei89b, Wil82a, Wim87]. formal [LS85, Wim80]. former [Ano88-30, Ano88-31]. Forms [Koh84, AT80, BH88, Bec80, Bou88, CC84, Cha85, DR80b, Dru87, Fan85, FH80, GH89, Gow81, GL84, Har81b, Hel89, LLT84, Lei89b, Lew82, Lew85b, Loe80, Lut84, Mad88, MJ81, Mar81c, Ng87, OM89, Olk83, PL86, Pow82, PW83, PY82, Rei87, Sen84, Tch89, Tur85, Wat87b, Wei84]. formula [BR80b, BS83, CK89, Cyb87, Mne89, Nis88, Tho88, Tho89c, Wim85b]. formulae [BJ84a, BJ86, BJL89, dB84]. formulas [Min85]. formulation [CD83, Dan86, MM81]. Formule [Mne89]. four [Nak84]. Fourier [Ros80, Gad88, Ger84, Par82]. Fourth [BHL89c, NW87, Gai89]. Fourth-Order [NW87]. Fox [Ker87]. fraction [KP86, VB89]. fractional [You84]. frames [Kov80, KS84, KS88, Lee89]. framework [HH83]. Fréchet [Sti87]. free [Bre84, GM88b, Gur85a, KMK85, Lip86, Uch86]. French [EKC85, Hel87, MM89, MCCF85, MCCF89, Mne89, Rob80a, Rob83, Tre85a, Yca82]. frequencies [Tak85]. frequency [TvdW88]. frequency-domain [TvdW88]. Frobenius [Tho82a, Ash87, CS85, Rob80b, Sch82, Uch86]. Fuhrmann [Cop84]. Full [ST88b, Dia81, Laf81b, Laf81c, Laf85a, Lew85b]. fully [JP87]. function [AMT84, BFP88, Bye87a, Fuj84, Gof81, HCR88, Väl88b, Wol89a, WSCP89]. functional [Acz85, CKN85, HG80]. functional-differential [Acz85]. functionals [Nav89, Oko86a, Oko86b, RV89]. Functions [PY80, Sta81, BR87, Bap86c, Bar86, BH88, BP88a, Bro89b, Brü84d, CL81, Che83d, Chi87, Chi88, Ell87, Eva85, Fer81, Fuh83, Fuh84, GR89a, GVK^{+}83, Ger84, GM88a, Hea83, HJ86a, Hel89, KÖ85b, Kwo89, MW85b, ML85, NT80, Pat85, RR84b, RP83, SF85, Sha88, Sti87, Tre86, Wol89b, Won82b, You81a, Lan81]. functorial [Hel89]. fundamental [Bea86, Kem81]. Further [Neu87]. Fuzzy [BP86].

G [Cam81, Gol85, Gov89a, Ple81]. G. [dPB86]. Galerkin [Ric83]. Galois [Bro83]. games [Bir81, Coh86, CMO89]. gametic [MCCF85, MCCF89]. gamétiques [MCCF85, MCCF89]. Gastinel [DR86]. Gauss [AR87, Mil87]. Gaussian [Bre86, GN82, Saa86]. Gen [Con81]. Gene [Ste87b]. General [Ber86a, Pai85, Axe86, Bak85, BS86b, BN88, CH89, GMSW87, Har81a, Hen83, LM87, LST88, Uhl81b, Vei89b, Zăl85]. Generalization [GHS87, Kem81, BK87, Bre86, CKN85, Fri81, Joh82b, Kha83, Kit87, Li87, MG85, Nis88, VJ88, Wan81a, Wim83b]. Generalizations [Cha85, Har84b, Che83d, NPW82]. Generalized [Bel81, BKR80, BG84b, Chi87, Chi88, Dat84, DT87, DR80a, Eag88, ES82, Emr83, Gib80a, Hun82, HP86, JY83, JY86, Kau87, KL88b, LT88a, Mal89, Mar84b, Pai85, RY85a, RY85b, RP87, Rob87b, Rot89c, Sch84b, SZ89, Ste81b, Tam81a, WH85a, Wim89b, YJ80, Zij84, dC87, AR87, Bar84a, Bar88a, BJ88b, BHK81, Bhi88, BM88, Cao87a, Cao87b, CL81, CZ84, Chu86a, wEC87a, Dem86a, Deu80, Don85, DL89, FG80, Flo82, FT89a, HPY86, HK81, HH83, HRS88, HS88a, HRS89, Hun88, JY87, KMK85, Koe85, LS87, MS85, Meh84a, MW85b, MNR84, Nom80, O’L80b, RM85, RR82, Rao83, SS85, Sod87, Tam81b, Tam87a, Tha89, Wan81b, WE89, Wer83, Wim89a, Wol89a, Wu80, Zho88, Zie87, Poo81]. Generalizing [Pat87, Pat88a]. generated [HJ86a, Laf81a, Loe88, Loe89b, Poh89, Sha82c, Sha82a, Tho82b, Tre86]. Generating [GB84, KR81b, GM88a, Laf81b, Laf83b, Sal81]. Generation [Ish83]. generator [KL88a]. generators [AL89, Lei89b, Uhl87]. Generic [YL85, Adk85, DM83b, Str83]. genetic [Hol87, WB81]. Geodesics [KS88]. Geodetic [GP80]. Geometric [LPC89, Sch81a, And83, AB87, BTD84, Mal89, RY85a, Tam81b, Tog80, Vei89a, Hel87]. géométrico [Hel87]. géométrico-arithmétiques [Hel87]. Geometries [LP82b, FH80, Far81, Kes82, Tha89]. Geometry [Fie81a, NS85, Tom85]. German [HH85, Nic82, PN85, Wer84]. Gersgorin [New80]. Gershgorin [Joh89, Sco85]. Gershgorin-type [Joh89]. Girth [Big80]. given [AHnLR88, Cve88, Mat85, OP82, RWZ88]. Givens [GH80, GLN84, GLN86a, GLN86b, GL87, HS86b]. Gleichungssysteme [Nic82]. Global [BR87, Adk85, Gur80, OS89]. GM [MR88a]. GM-matrices [MR88a]. Gohberg [BKL85, Cop84, Dat86, Lan81, CK89]. Goldberg [Tho89b]. Golub [Ste87b, Pow87, She87b]. Good [Cal80]. Gordan [BTD84, Man81, ML83]. Götz [Ral86]. government [CMR81]. Graciano [GL85a]. graded [PV82]. Gradient [Axe80, Gre89, JY83, JY86, Mon84, O’L80a, Woź80, YJ80, Bar80b, CW80, GH84, JY87, O’L80b]. gradients [MC81]. gradings [PZ89]. Graham [Bru84c]. Gram [GA85, Har82, Pat88b, Ptá83b, Ruh83]. grammars [Uch86]. Graph [Bru89, Man82, Neu89, Pow88, Ano80q, BB82a, BB82b, DEHS84a, DEHS84b, Doo82, Far85b, GLN84, GLN86a, GJS89, Hur89, KB85, McK81, MW84, Moh82, Mül87, PS82, Pow89, Rot89a, Sch84a, Sch86a, Zha87b]. Graph-theoretical [Man82]. Graphs [BT81b, KRS82, BJL89, Bee86, Bie84, BM82, BN89, BL82b, Doo89, FRS82, Fri88, Gar82a, GM80, God82, GM88a, HH89, Ito82, JvdD88, LSS89, MRW81, MO87, Moh88, Row83, Row88, ST88a, She89, Sta87, Ter87, Tho86, Tor86a, Wan83, Wei87, Wei89b, Yua88]. Grassman [BF81b]. Grassmann [BF85]. greater [Tor86a]. Greatest [BL89b, Krä87]. Greece [AL86a]. Green [Rob86]. grid [HT84]. Group [AP83, AP88, HW81, MR83, MR88b, Bec89, BV80, Cha81, Dam80, Dia81, Dad87, Djo81, KN87a, Lip80, Mit87, Spi82, Wag89, Wan80, Wan81b]. Group-invariant [AP83, AP88]. groupes [EKC85]. groups [Bai85, Bar89b, BM82, Bro83, CL82a, Chi87, Chi88, CKN85, Dad87, EKC85, EN81, GHW82, Gur85a, Ish83, Jam83, Jam88, Kob85, ML83, Nie87, Sch85, Tak82, Vas85, Wim87, dOdS83]. growth [Moh88, Pow87]. Guide [Cam81]. Gundelfinger [Tho82a]. Gutman [Bru89].

H [Ano87s, Lan81, Pul87, Ste87b]. Hadamard [AP83, AP88, Bus80, Fie83, FM88a, Giv84, GS84, JB85, JE87, YH82]. Haifa [BHL89c, BCS86, BHL89a]. half [Djo82]. Hamilton [OR88]. Hamiltonian [Fuh84, PV81, Van84]. Hang [Ple83]. Hankel [Dat84, Fie84a, Fie84b, Fie85, Fie88a, Fuh86, GH89, GR89b, HJ86a, HR88, HR89, HHM89, Hol85, JM89b, Pow82, Ptá84c, Ptá86b, Wim89b, You83]. Hardy [Sis86]. harmonic [And83]. Hasse [Tho82a]. Hausdorff [GH83, NW84]. having [CM81a, Cai81, HL84, Her83, JG80, Pat88b]. Hayes [Ker87]. Haynsworth [CMU86]. height [Her89]. held [AL86a]. Helmut [DHL85, DHLW85]. hemispaces [MLS83, MLS88]. Hermite [Byr83, SSZ83]. Hermitian [AB85, AB84, BPR87, BC84, Bin84, BK88b, BLAK89, CM81a, Cai81, Cha85, CG89, CM83, Dan87, Djo81, DGKW84, Far81, FH88a, Gof81, GR87, GJMW84, Han86, HHJ86, Laf83a, LN85, Lee80, Lew82, Lew85b, Mar81b, Mat86, Meh84b, Nak84, PH81, PY82, Sai82, Sha82c, Sha82a, Sha82b, Wat83, Wim83a, dB80]. Hermitian-preserving [PH81]. Herzberger [Ral86]. Hessenberg [CS86a, Pif87, Zho88]. Hessenberg-like [CS86a]. High [DKH86]. High-Performance [DKH86]. higher [Cla88, Li86b, MS88, Rei86]. higher-index [Cla88]. Hilbert [Cop84, Bus86, Gha87, Gow85a, Grü82]. Hilbertian [AK89]. Hill [HG87, CH83b, Hoc84]. HMDR [DL89]. Hoerl [RP85a]. Hoffman [Sak85]. Hölder [Oku87]. Hölder-type [Oku87]. holomorphic [Brü84d, Gur88]. homogeneous [KK89, MJ81, Ng87]. homographic [Ded88]. Homotopy [LS87, Sch89b, Chu84, Chu88]. Honoring [SDGT81]. hook [JP88b]. Hopf [GG82, GLR83]. Horn [And87a]. Householder [Ano80a, GL87, Kau87]. hulls [Sun82c]. Hurwitz [Bos88, Byr83, Nav89]. Hyperbolic [APP88, DI86, Djo82]. hyperboloids [CW85]. hyperplanes [Jam81]. hyperquadrics [Djo81]. hypothesis [Roe86, Roe87].

I. [HH85]. Iain [Ple81]. ideal [PV82]. ideals [AK89, And85a, Hal86b, Pok85, Tau80b]. ideas [Dat89]. idempotency [Kha80]. idempotent [Kha83, Kha85]. idempotents [Laf81a, Laf89, SW85, Stu87]. identical [Tau83]. Identifiability [Sha85f, Sha89]. identification [LS85, Pol89, RWZ88]. Identities [Vei86, And85a, BPB88, BK86, Chu89, ET81, KW89b, MW85b, Ros80, Wat88a]. identity [Fan85, Lew85a, MG85, Son88, VJ84, Wat87b]. if [IS89]. II [BJ86, Bha81, Chi88, DEW80, DEHS84b, GLN86a, Gol84, Har84b, Hoc87, Jam83, Jam88, LT88b, MM89, MCCF85, Oko86b, RR84b, RR82, Roe87, SD80, Sha82a, Szu89b, Tam87d, Tom85, WH85a, dO82]. II. [LT83b]. III [BNV82, Dan86, GLN86b, MCCF89]. Ill [ES86, Sch86b]. Ill-Conditioned [ES86]. ill-posed [Sch86b]. image [Bro85b, EHL81]. images [Fia89]. Imaginary [Lás82]. immanant [GM87]. immanants [Gro85, JP88b]. Immittance [BLAK89]. Immittance- [BLAK89]. implementation [Boj86]. Implicit [GLN86a]. Improved [BV88]. Improving [KMV87, MW85a, Mil87]. inadmissible [PN85]. Incidence [Wil82b, BHK81, CS89a, Kes82]. Incline [Pul87]. Incomplete [Pol87, ABI84, Axe86, Rob82]. Inconsistent [Axe80, Ver82, Wan86, Zie85]. increasing [FK85]. Indefinite [Dat86, Wim83a, BK80c, GR87, Hes86, LN85, RU85]. independence [BCS87]. Index [Ano82b, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano80r, Ano81b, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81c, Ano81i, Ano82c, Ano82d, Ano82e, Ano82f, Ano82g, Ano82h, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano83a, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84i, Ano84f, Ano84g, Ano84h, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano85i, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86n, Ano86j, Ano86k, Ano86l, Ano86m, Ano87j, Ano87a, Ano87k, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano87i, Ano88b, Ano88m, Ano88n, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano88h, Ano88i, Ano88j, Ano88k, Ano88l, Ano88a]. index [Ano89a, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89j, Ano89g, Ano89h, Ano89i, AR87, BS89, Cla88, Lip80, Row83, Row88, SL87, WN87]. indices [Avr89, Rot80]. Induced [Mer81b, Bye87b, CL82b, FB81b, MW81]. industry [CMR81]. inefficiency [Rao85]. Inequalities [Bap86c, Fan88, JP88b, MW85b, And83, AB87, Bap87, Bar89a, BK88a, CE82, Cha85, Che87, CFKK82, Coh88, DI85, DS89a, EJ89, Fuj84, HW89, JR81, JB85, KO85a, Mar88, Moh88, NT80, Oku87, Pat82, Sai87, Sai88, Sod87, Spi85, Spi87, Stu88a, Tho85, Tor86b, Wat87c, Zab88, Tho83a]. inequality [AT80, AP83, AP88, BKK83, BM89a, Bap86a, Cha87b, Dav80, FJMN85, FM88a, FK87, Fur89, GL88, GP89, GH87, GM87, KMV87, Kha83, Lav80, MM80, MM81, MJ81, Meh84a, Olk83, Pat84, Pat87, Pat88a, Rao85, Sak85, Sty81, Tho83b, Wat88c]. Inertia [BAG89, BC87, Cai80, Cai82, LT83a, Uhl81b, Bin84, CM81a, Cai81, Car84, Dan86, DD87b, HF85, JMRT87, Mad88, Mar81b, RC81, Väl88b, WH85a, WH85c, Wim83b]. inertial [WH85b]. inertias [Dan87, Dan88]. Infimum [Mit88]. Infinite [Gre83a, Tor86a, Art86, Bol85, GM88a, HW89, Lin86, Moh82, MO87, Mül87, Ptá83a, Ptá86b, Sen80, SWR87, Stu88a, You83, dBJP82]. infinity [LPC89, Mee87]. inflators [Stu87]. influence [Sch86a]. informal [CMR81]. Information [Cha81, AN83, Bap86a, Bro85b]. Ingram [Tho83a]. inheritance [OS89]. inhomogeneous [CS82, IS82, MS87]. initial [MC81]. Inner [Wat88a, BWT89, Cri88, Dam80, RU85, Wim83a]. inner-product [Cri88, Wim83a]. Input [Wil83, CS87c, MK89]. Input-output [Wil83]. inscribed [MP89a]. instability [HG87, Tch89]. Integer [BHK81, AO82, Bak87, Cam89, DEHS84a, DEHS84b, Laf89, Wim87, Wim89d]. integral [Ash87, BH82, BV80, HK81, Har81b, Hur89, Koe85, Kön86, Rao83, Tau80b, Tau85]. integrality [Sch81b]. integrals [Poh89, Sai88]. integration [BS83, FHP80]. intensive [HM89b]. interaction [Tak85]. interchange [BL82b]. interchanges [Fos86]. Interesting [Tha89]. interior [Har80]. Interlacing [Bar89a, Sod87, Zab88, Zab89, Car83, DD87a, Fri85b, HMS87, LN85, Mar88, Que87b]. International [Bru88a, BV89]. Interpolating [Sch83a]. interpolation [Ant88, AA89, BGoMSTAUTA^{+}86, Dub85, FSZ85, FT89a, GM89, LA88, MR84, SSZ83]. interpretation [CW80, Gee89a, Hol87, JM89b, Wat87c]. intersection [DR80a]. Intersections [BF82, Rys81]. Intertwining [CS89a, FP87b, FP87a]. Interval [Ral86, Roh81, Roh89, Bos88, CG86, GA81, May84, May85, May87, Neu84, Neu87, Nic82]. Intervall [Nic82]. Intervall-Gleichungssysteme [Nic82]. Intervals [BG84a, HG87]. Introduction [Cam84a, O’C85, Ral86]. invariance [Bea82, Bea88, BL89a, BH87c, GH89, Tam87c, Tam87b, Tam87d, TvdW88]. Invariant [Hea83, Que83, AP83, AP88, BGoMSTAUTA^{+}86, Bie84, Che83a, Chi86, CD85, DM83b, Els82b, Emr83, Fei89, GR86, KS85, Lás80, Lon84, Mar88, MM82, MW84, Ong82, Ong87, Özg86, RR84a, RR84b, RU85, Sil86, Sil87b, Sou86a, SRPM82, Ste82, Tho82b, Wil83, Zab87, Zab89]. Invariants [Gur85a, Coh87, Dio89, LS85, Ray83]. inventories [Mer80b]. Inverse [Joh82a, Joh85, LY88, Urs82, AR87, BR87, Ber86b, BHH87a, BK81c, BK86, Caf81, CH83b, CG80, Cli81, DGK80, Dia86, Dic84, Dry85, FG81, FMN83, FM86a, FJM87, FM88b, FM88a, Flo82, GH84, Hoc84, HLG85, Ima83, Ima84, Joh83, KMK85, KB85, Koe85, Lav80, LN80, Lew89a, Lub87, MS86, PR81, RY85a, Rob84, Rob87a, Ron82, Sch83c, Sha83, Sme81, Tho89a, Usm87, Zho88]. inverse-partitioned-matrix [Dry85]. inverse-positive [FG81]. Inverses [BF81a, O’L90, Pif87, Poo81, Riz84, Ste89, Vav84, BJ84a, Bar84b, BJ86, Bel81, BF87, BJ88b, BHK81, Bhi88, BKR80, CS86a, Ded88, Dem86a, Don85, DG81, Eag88, EP88, GB84, Gre83b, HK81, Har81b, HW81, HR88, HR89, Hun82, Hun88, HP86, KK86, Kle82, MR88a, Mee83, Mee87, MR83, MR88b, Mit87, MNR84, Nom80, Pye80, RM85, RR82, Rao83, RY85b, Rob85, RPV87, RP87, Rob87b, Rom86, She86, Tam81a, Tam81b, VV86, Wie82, Wie85b, Wim89a, Won82b, Wu80, Zho88, Zie87, dC87, Mar84b]. Inversion [BK88b, BAS86, Fei84, Fuh86, Gus84, SSS82, Tho82b]. inversive [Wil82a]. Invert [PS87]. invertibility [Dem82, Bas85]. invertible [Rob85, Wie85a, dBJP82]. involution [Hel87, Que87a, Wea83, Wea88]. involutions [KN87b, KN87a, Liu88]. Involutory [Won82b, CM84]. Involving [Gau83a, Car88, CL81, Cha85, FHV85, Lav80, Pat84, PL86, Pon87, Stu88a, VJ84]. Iohvidov [Tho82a]. Irreducibility [FM81, FB81a, BS87a]. Irreducible [Dia81, JZ85, MŽ83, Neu81b, San87, Sha85d, SL87, TW80, Vin89]. Irredundant [Laf83b]. irregular [Bea86]. Ising [Ste81a]. Isometries [LT88b, Gro80a, LT83b]. isometry [Lew82]. isomorphic [Des88, Pol86]. isomorphism [Gur82]. Isomorphisms [Bol85]. isonemal [HT84]. Isoperimetric [Moh88]. Isospectral [Cal82]. issue [Ano80q, Ano85t, Ano86a, Ano86v, ORS85b]. issues [Ano88-34]. iterated [Poh89]. iterates [Tad86]. iteration [BEN88, Lei89a, NVD80, OS84, Rob83, Rob83]. Iterations [Rob80a, Ded88, Hug85, Mil87, Rob80a]. Iterative [CS86b, EHL81, HRR87, NdPV84, YJ80, Abe82, AC89, ABI84, EL88, GR87, HPS83, HPY86, HPS88, MS83, Man84, MNP85, Mil88, Mis85, NV84, Neu80, NP87, Sch84b, dP80]. IV [Ano80a].

J [Bar81a, Cam81, Gov89a, Nau87]. Jacobi [BF89, Gab87, Har86a, KG83, Mil83, Mil87, SD80, SHW86, Tho82a, WN87, dBS86]. Jacobi-like [Gab87, Har86a]. Jacobson [Har87b]. James [Ano87s, Bea87a]. January [BU88, BHL89c]. Järvinen [Gre83a]. Joel [Bea87a]. John [Bea87a]. Johnson [And87a, Lew89c]. joint [CS89b]. Jordan [AB85, AB84, AR87, HPW89, Koš88, Laf81c, LP82a, MP83, RB89, Sch86a, Str87, Wim87, Wim89d]. Jr. [Poo81]. June [BU84]. Jürgen [Ral86].

Kågström [Cam84d]. Kaashoek [Lan81]. Kahan [Neu81b]. Kantorovic [Dav80]. Kantorovich [Cha87b, Rao85, Wat87c]. Kantorovich-type [Wat87c]. Kato [Cam84a]. Kennard [RP85a]. kernel [MR83, MR88b]. kernels [Hun88, RP87, Sai82, Sai87, Sai88]. Ki [Ple83]. Kim [Pul87, Ple83]. Kind [Ong87]. kinds [Mur83]. Kippenhahn [Laf83a, Sha82c, Sha82a]. Kogbetliantz [Bai88, CV87, PV86]. Kojima [Hof85]. Korte [Sch84c]. Korzhik [Was84]. Krause [Wil82a]. Kreın [Ban82]. Kreisscheiben [Nic82]. Kreisscheiben- [Nic82]. Kronecker [BV88, Coh87, Hal86a, JAK89, Lin88, MV88a, Oko86a, Oko86b, Ste81a, Bru84c]. Kruskal [BWT89]. Krylov [Ruh84].

L [Bar80a, BKL85, Bru84b, Cop84, Dat86, Mar84b, Poo81]. L. [Ker87]. Lancaster [BKL85, Cop84, Dat86, Har88]. Lanczos [CW80, Gre89, Pai80, Par80, Sim84, Cyb87, ER87]. Laplacian [Bap86b]. Large [GP80, MNP85, Saa80, Cob86, Dat89, EHL81, EL89, HPS83, HPS88, McK81, Neu80]. Large-Scale [GP80, MNP85, Dat89]. largest [Ash87, Neu82a, Yua86]. last [Tau85]. Lat [CH80]. Latin [BF82]. lattice [Boj86, LT86, Lon84]. lattices [DS89b, GR86, MM82]. Lavoie [Kha83, Sty81]. law [Dan86, Har86b, Wim83b]. Lax [FT84]. Least [APP88, BD80, BB86, Bro89a, ES86, Fre87, GP80, MNP85, NdPV84, VV87b, VV87c, Bak85, BT81a, BI82, Bjö87, Boj86, Dax88, Doo82, EL89, GH80, Kam89, KRS82, MN87, RP85a, Rao85, Spi87, Tor86a, Ver82, Wan86, Wei89a]. least-norm [BI82, Ver82]. least-squared-error [Ver82]. Least-Squares [BB86, GP80, MNP85, NdPV84, VV87b, Spi87, Wan86]. leave [MK89]. leaving [Els82b]. Ledermann [GL85b]. Lee [Bea87a]. lemma [Hal86a, MS80a, Mar88]. length [DSD88]. Leon [BP84b]. Leontief [FHV85, SS87b]. less [JS89]. Letter [Ano88-28]. Letters [Ano80n, Ano80o, Ano80p, Ano81p, Ano81q, Ano81r, Ano81s, Ano81t]. level [Her89]. Leverrier [Giv82, Gow80]. Levi [RWZ88]. Levine [CM84]. Lewin [Zha87a]. Lexicographical [MLS87]. Lie [Tow80, AL89, Bro83, Far84, Gam88b, Lei89b, PZ89, RWZ88, Zas83]. Liénard [FD89, Jur82]. light [Har87a]. like [BG86, CS86a, Gab87, Gau83a, Har86a, Men86]. likelihood [AO85, Don85]. Limit [Fri86, HS84, Doo89, Spr81]. limiting [Tor86a]. limits [Sch83a]. Line [Sme81, BS81, FS85]. Linear [AB85, LAA68, Ano81w, Ano82n, Axe80, Bea81a, Bea82, Bea88, BL89a, BHJ85, BS87a, BCR88, BDM83, BA80, BD80, BBV88b, BF83, Bro83, BV89, BS89, BKR80, Bun87, Cai82, Cam84e, Cam84a, CL82c, CLT87, CJR87, Cop81, CM81b, CM83, Dax88, Dad87, Far85a, Fia89, GKKL87, Gov88a, Gov89a, Gre83a, Hal83, HJ86c, HM86, How80, ISS86, JY83, LTT83, LTT88, Loe89a, MK89, MR84, Mon84, NdOH89, Pai85, Poo81, Roh89, SS87a, Sha89, Sta81, Tam87c, Tam87b, Tam87d, Wat82a, Wat86, WWWS86, Wim89c, Wol89a, You84, Zie88a, dO83, Adk85, Aga81, Ano80n, Ano80o, Ano80p, Ano80q, Ano81p, Ano81q, Ano81r, Ano81s, Ano81t, Ano88-28, Ant83, AL86b, AS80, BM85, Bak85, BM89a, Bar80b, BI82, BI88, BB89a, Ber81, BNS86, BKH88, BFM89a, Bjö87]. linear [Boj86, Bou88, BH89, BPB88, BCS87, CMR81, CE82, wEC87a, Chu88, wEC89, CD85, CPV89, Cri88, DI85, DW89, DS89b, Dia81, Dio89, DM82, Don87, Dry85, EHL81, Emr83, EL88, FHV85, Fuj84, GS89, Gee89a, GH80, GL88, GKK86, GdHZ89, GR87, GA81, Gur85a, HPS83, HPS88, HH83, HS89, Hia87, HPW87, HPW89, HW89, Jai84, JKvdD87, JP87, JJS89, Kam88, KP80, KW89a, Kay85a, Kay85b, KP86, KR81b, KP89, Kra81, Kun81, LY88, LR81, LT83b, LT88b, Lip80, Loe88, Loe89b, LPC89, Mal89, Man82, MŽ83, MJ81, Meh89, Mil83, MN87, MV88b, Nak84, Neu81a, Neu84, Neu87, Neu80, Neu82b, NS89, Nic82, Nie89, Nis85, Oko86a, Oko86b, Par80, Pet82, PH81, PW83, RY85a, RV89, Rei87, RO89, Roh81, RS82, Rot89c, Sch84b]. linear [Sch88, Sin88, Spi85, Spi87, ST88b, Ste85, SY89, Syl86, Sza85, Tan86, TW80, Val88a, Ver82, Wan86, Wei83, WN87, Wil83, Wim80, Wol89b, Zie83, Zie85, dSB81a, Ano81u, Ano81v, Ano83l, Ano84q, Zim80, Cop84, For86, Gov89a, Nau87]. linear-quadratic [Gee89a, JJS89, SY89]. linearer [Nic82]. lines [Caf81, KBM86, Neu89]. linked [DALSP89]. Linpack [Cam81]. Liouville [Usm87]. little [Jam81]. Loan [Ste87b]. Local [AZ83, Adk85, EI87, FN89, Gur80, Gur81b, LP82a, MS83, Rob83, Zhe87]. local-global [Gur80]. Localization [HB83, Sen84]. locally [MFPW85, Pol86]. location [Gen86, PY82]. loci [DALSP89]. Loewner [Fie84a, FP88, Vav88]. log [Nus86]. Logan [BU84, BU88]. loglinear [BR89a]. Longstaff [Stu88c]. look [Par80, Rys85]. lookup [HTRG88]. loop [Lin87b]. lossless [GVK^{+}83]. Lothar [Ano85s, EH85]. low [Cob86]. Lower [Bea87b, Ben86, Con85, JKW85, Pat88b, Bar83, BJ88a, Bar88b, Che83a, Che83b, Joh89, Lew89c, Szu88, Szu89b, Yua89, dG82]. Löwner [Vav84]. LQ [Pol89]. Lyapunov [BC87, HJ86b, HS85, Jon84, KH84, New83, Nor89, Ptá81, Ptá83b, Ptá84c, SB87, SH88, SB88]. Lyndon [Bea87a].

M [Bru84b, Bru89, Cam84c, Gau83b, Lan81, SDGT81]. M. [She86]. MacLaurin [BS83]. MacWilliams [BPB88]. main [BH88]. Maintaining [GMSW87]. majorants [Dah83]. Majorization [And89, Gar82c, Tho83a, BS85a, Bru84a, Fis87, GW85, Her89, Kom83, Pan89a, Zab89]. majorized [BT81a]. Making [Dan89]. Manifestations [Mer80a, Bre88]. manifold [AL86b, HS87]. Manifolds [Kov80, KS88]. manufacturing [ER89]. many [Tau82]. map [HJ86c, HM86, Kit89, Mar84a, TW89, Wat88b]. mapping [Hen83]. mappings [Ber86c, Cri88, Tur85]. maps [BT81b, CL82b, CJR87, FB81a, FB81b, Hia87, How80, MŽ83, Tam81a, Tam81c, Tan86, Tom85, Wat82a]. March [GL85a, Hol85, BF85, Gro85, Laf85b, Lew85b, O’C85, SW85, Tim85, Wim85a]. marginal [NS85]. marked [HK83, Sch86a]. Markham [CL89, Smi87]. Markov [Mey84, Dat84, FB81a, FB81b, Fry85, FM86b, GM85, HJ86a, HS84, Hun86, IS82, JM89a, Sen80]. Markovian [Hun88]. Marshall [Tho83a]. Massey [JM89b]. Master [GT89b]. matching [Bro89a, Emr83]. Matemática [GL85a]. Mathematical [Bar81a, Bea87a, GL85b]. mathematician [HM84b]. Mathematics [BU84, BU88, Sch84c, DR86, Bar80a]. Matrices [And89, BY80, Bar81a, Ber81, BS81, Bru80, Bye87b, Chu86b, CS86b, CM83, Dem86b, DM80, Els82c, Els85, ES80, FB81b, GL85a, Gau83a, GR80, KK86, Kau87, KG83, KB85, Li88a, LQ87, MT85, MSS84, Mey84, NW87, PV81, PS87, Rot81, Rys81, Saa80, Sha85a, Sil87a, Zab87, dPB86, dO82, Adk89, AHnLR88, ABO88, AS82, AB84, AB85, And83, And85b, And87b, AO82, Art83, Art86, Ash87, Ata89, AP86, BF85, Bak87, BK81a, BP85, Bak85, BH87a, BGoMSTAUTA^{+}86, BPR87, BR80a, Bap82, Bap83, Bap85, Bap86a, BR89a, Bap89, BT81a, Bar89b, BH81a, Bar89c, Bar81c, BF81a, BJ84a, BJ84b, Bar84b, BJ86, Bar83, BH88, BJ88a, Bar88b, Bea81b, Bea82, Bea88, BL89a, Bea84, Bea87b, BQ85, BdP86, Bee80, Bek88, Bel81, BAS86, Ben86, BS82a, Ber86b]. matrices [BF87, BH84, BHJ85, BH87b, BS87a, BJ88b, BL89b, BHK81, BH82, BHH87b, BM89b, BG84a, BK81b, BB89b, BC83, BK88b, BLAK89, BH81b, BS87b, BG82, BL82a, BL82b, Bru84a, BH85, BR86, BS87c, Bru88b, BM89c, BS82b, But88, Caf81, Cai84, Cal82, Cam89, CS86a, CH83a, Car83, Car84, CP85, Car88, CH89, CDMS81, Cha81, CL82c, CR86, CLT87, Cha82, Cha88, Che83a, Che83c, Che84, CS89b, CL86, CLLW87, CHP88, CF86, CG80, CM84, CS82, Con85, CvRS81, CPV89, Cri88, Dan88, Dat84, Dav80, Ded88, Dem89, Dem82, Den81, Deu81, DS89b, Dic84, DR83, DR88, DR80b, DGKW84, DG87, Dow85, DS82, Dua89, Dub85, DG81, EP88, ELS82a, Ell87, EGL85, Els82b, Els84b, EP87a, EJ89, EP87b, Fan89, Far87, Fer80, FG81, Fie81a, FM81, FMN83]. matrices [Fie83, Fie84a, Fie84c, Fie84b, Fie85, FJMN85, FP87a, FP87b, FJM87, Fie88a, FM88b, Fie88b, FP88, FS85, Fis87, FG80, FL89a, FH88a, Fri85a, Fri85b, Fri86, FH88b, Fuh86, FD89, FHV85, FP81, Gab87, Gad88, GLS83, Gał81, Gar82b, Gen86, Ger89, Gib80a, GT89a, Giv84, Gof81, GG82, GR86, GS83b, Gol84, Gol86a, Gol86b, Gor89, GP89, GB84, GB87, Gov88b, Gov89b, GL84, Gow89, Gow80, Gow85b, GR87, GS84, GW80b, Gre83b, Gro80a, GJMW84, GJMW87, Grü81, GJS89, GN82, Gur80, Gur81b, Gur85b, Gur88, Had83, HM84a, HF85, Han86, HL83, Har80, Har81a, Har84a, HL84, HP81, Har82, Har84b, HS86a, Har87a, Har87b, HW88, HJ86a, HJ86b, HR88, HHM89, HI86, HB83, Her83, HB84, HS85, HJ86c, HM86, Her86]. matrices [HJ86d, HMS87, HRS88, HS88a, HH85, Hia87, HW87, HG87, HHJ86, HH88, How80, Hwa86a, Ima83, Ish87, JG80, JL87, JvdD88, JP82, JP84, JP87, JP88a, JZ85, Joh80, Joh82b, Joh82a, Joh83, Joh84, Joh85, Joh87, JE87, KR83, KO85a, KH84, Kay87, KGÖ83, Kha80, Kha83, Kha85, Kle87, Kle82, KL86, Koe85, KS80, Koš88, Kov87, Laf80, Laf85b, Laf86, Laf89, LLT84, LO81, LN85, Lás80, LR81, Lee80, Lei81, Leó88, LN80, Lew89a, Lew85b, LT83b, LT88b, Lin89, Liu88, Loe83, Loe84, Loe86, Lon81, Lon88, LP80, Luk81, Mac88, Mad83, MM80, Mar81b, Mar84c, MJ81, MPJ81, MSS89, MS86, MH81, MR88a, May84, McF80, Mee83, Meh84b, Meh84c, MV88a, Mit82, MP82, MO86, Mit88, MNR84, MM88, nM87, NV84, Neu86, Neu81b, NP87, Nom80]. matrices [O’C85, Oku87, OvdD82, OP82, Ong84, OS84, Pan89c, Par86, Pat88b, PL86, Pif87, Pin85, Piz83, Pol87, PK87, Ptá83b, PW85, Ptá86b, PS88, PV82, Rad84, Rag84, RM85, Ram86, RR82, Rao83, RY85b, RC81, Ray83, Rho89, Riz84, Rob82, Rob86, RPV87, Rob87b, Rom86, RT85, RS89, RCA88, RU87, SD80, Sai82, Sch82, Sch81a, SS88, Sch86c, SL82, Sha87, Sha85b, Sha85d, Sha82c, Sha82a, Sha85e, SH88, She84, SWR87, SS86, SS87b, Sil88, SPZ82, Sin81, Smi86, Smi88, Son88, Spi82, SRPM82, Ste85, Ste87a, Ste81b, ST87a, Sto86, Stu88a, Stu88c, Stu88b, Szu89a, Tad86, Tam81b, Tam87a, Tam87c, Tam87b, Tam87d, Tan83, Tao84, Tau80a, Tau83, Tho86, Tho82b, Tho83c, Tis87, Tre85b, Tre86, TST86, Uhl81b, Väl86, Väl88c, Väl89a, Väl89b, VNC84]. matrices [Vav84, Vav88, Vei86, Vit82, Wan80, Wan81b, Wat86, Wat87a, WH85c, Wat88c, Wea83, Wea88, Wer84, Wer85a, Wie82, Wie84, Wie85a, Wie85b, Wie85c, WN87, Wil82b, Wim83a, Wim87, Wim89a, Wim89b, Wim89d, Woe89, Won82a, Won82b, WB81, Wu80, Wu87, Wu88, YL85, Yan84, YH82, Yca82, YB81, YB84, You81a, Zab86, ZM81, Zen84, Zen89, Zha87b, Zha87a, Zho88, Zie87, Zie88b, Zij84, dBJP82, dCG81, Gib80b, Wat80, dB80, Bas85, Har88, Hay82, Nor84, Dat86]. Matrix [Ano88-32, BP81, Bea86, BCS86, BHL89a, BHL89c, BH83, Bru84b, Bru84c, BG86, Cam84b, Car85, Ded88, DK87, ES86, Elf80, Fer81, FG86, GL85a, Gai89, GHS87, Gur82, Har84b, HR89, Hig87a, Hig88, Hol85, Jin88, Lan81, Lás81, Ler80, OS86, OvdD88, Ple83, Qi84, Ros80, Sch89a, Val88a, Van84, Zie88a, ABI84, Axe86, BF85, BK80b, BKK83, BP85, BM89a, BPS89, BR87, Bap84b, Bap86b, Bar84a, BFP88, BJL89, BH88, BP88a, BF88, BZ89, BHH87a, Bhi88, Bol85, Bro85b, BS86b, BN88, BK86, BG81b, Bye87a, BV80, CM81a, Cai81, CG86, CL81, CH83b, Che83b, wEC87b, wEC87a, wEC89, CFKK82, Coh86, Coh88, Cop84, CM81b, Cro81, Dah83, Dan87, DD87b, Dat88, Dat89, DN84, DGK80, DGK83, Des88, Deu80, DW83, Dio89, DM82, DvdP81, Don87]. matrix [DEHS84a, DEHS84b, Dry85, ES80, ER87, EP87b, EG89, Eva85, Fei84, Fia83, FJMN85, FM86a, FM87, FM88a, For80, Fos86, FK87, Gar82c, Gek81, GS87, Gha87, GMSW87, GW85, GH83, GLR83, GHW82, Gro85, Grü82, Gus84, HG80, Hal80, Hal86a, HRR87, Hea83, HS86b, HG89, HJ86d, HS88b, HRS89, Hoc84, Hol82, How83, Hu87, Ima84, Isa80, Jia82, JAK89, JKW85, JORvdD87, Joh88, KBM86, KMV87, Kem81, KP86, KB89b, Kit87, Koh85, Kra83, Krä87, Kwo88, Kwo89, Laf81b, Laf81c, Laf83b, Laf85a, Laf85b, LT83a, Lan86a, LY88, LT88a, LN80, Lew85b, Lew89b, Lew89c, Li86a, Li89, Lin87a, Lon84, Lub87, MP83, Mar85, Mar86a, Mat85, May88, Mee87, Mer84, MW85b, Mes89, Mil88, Mit84, Mit86, Mor88, Mur83, Mur82, Nak84, Nel85, Neu85, NPW82]. matrix [Neu82b, Nor89, O’C85, OH85, Ple81, Pok86, Pol85, Pon87, Pow87, Ptá83a, RR84b, Rob85, RB89, Rod80, Ros84b, Rot89a, Rot80, RP83, Rys80, SA89, San87, Sch85, Sch84a, Sch86a, SZ86, SZ89, Sco85, Sco86, SS81, SF85, Sha88, SL87, Shr86, SS86, Sil86, Sil87b, Smi87, SW85, Spr81, Sta81, Sti87, Str87, Stu88a, Sun82a, Sza85, Szu88, Szu89b, Tan86, TZ87, Tau80a, Tho89a, Tho85, Tho88, Tim85, Tom85, Uhl83, Uhl87, Urs82, Usm87, Väl88b, VB89, VD83, Vav89, Wat84, Wat88b, Wea83, Wea88, Wer89, Wer83, Wim85b, Wim85a, Wim88, Wim89c, Wol81, You81b, You83, Zen89, Zie83, Zie85, dBP82, dBS86, dG82, dOdS83, And87a, Bar87, BKL85, Cam84d, Ste87b]. matrix-fraction [KP86]. Matrix-theoretic [Fer81]. Matrizen [HH85, Wer84]. matroid [CS87b, Wag89]. matroids [BR89b, Cor87, Cun82, GT89b, LW81, Law82]. Matthews [Pow87]. max [BR89b, OR88]. max-flow [BR89b]. Maximal [SH88, BS81, Fri85a, KS86b, Laf85a, RR84a, RR84b, RU85, Row83, Row88]. Maximization [Elf80]. maximizing [WT85]. Maximum [AO85, BAEGL87, BR80b, Don85, Ish87, Leó88, RT85, She89, Sou86a, Sou86b, Sto86]. Maximum-likelihood [AO85]. May [Ano88-32, OvdD88, Bon83]. McDonald [For86]. McIntosh [Mon84]. mean [And83, AB87, BM89a, Gek81, Sza85]. means [AMT84, Fur89, Lut84]. measures [AN83, Ata89, EP87a, GM88a, Har81a]. Mehta [Bru84b].

Meixner [Chi86]. members [Ano88-30]. Memory [GHL86, Ano87s, HM89b]. memory-intensive [HM89b]. Merten [BFP88]. mesh [CZ86, Ver84]. metacyclic [ML83]. Method [BD80, BH83, Cal80, MNP85, Saa80, Van84, Axe86, Bar80b, Bjö87, BF89, BK80c, CZ86, Cao87a, Chu84, Chu88, CM84, DM86, DI85, DGK80, Dry85, EL88, GT89a, GA81, HY80, Har81b, HS86b, Ker87, Kri84, LS87, Lin87b, MS83, Man84, Mar86b, May87, Mir83, NM85, NV84, Ogr88, Pan89a, Ric83, SSS82, Sid89, Spi85, Spi87, Tao84, TST86, VB88, VNC84, Ver84, Wat87c, Yey84]. Methods [Axe80, CM82, Elf80, Fre87, JY83, MP80, NdPV84, O’L80a, PL82, Ruh84, SS87a, Sim84, YJ80, AC89, Ano80q, ABI84, AP86, BEN88, BG86, CE82, DM82, EL89, FM89, Gab87, HPY86, HRR87, Kra81, Lás81, Ler80, LM80, MK88, MN87, NP87, Ric83, RTW83, San87, Sch84b]. metric [Bus86, Sch82]. Metrics [Had86]. Meyer [Poo81]. MGR [RTW83]. min [BR89b]. min-cut [BR89b]. Minc [Nis88]. Mini [dO83, Neu82b]. Mini-Conference [dO83, Neu82b]. Minimal [Bre84, SF85, Woe89, Adi82, BC85, Ele87, GLS83, Laf85a, OM89, Sal81, Sha82b, VB88, VB89, WS83, dGH83, Lan81]. minimally [YB88]. Minimax [Car83, ST88b]. minimization [Kab85, LM80, LST88, Mor85, Pon87, Sch89a]. minimizing [PW89]. Minimum [BK80a, For87, Son88, BS87c, For80, MS80a, Rot89a, Tre85a, Wan86, Tre85a]. minimum-norm [Wan86]. Minkowski [NVD80]. minor [SS87b]. minorants [Dah83]. minors [Eag88, HL84, Her83, HMS87, JORvdD87, KR81a, Loe86, MNR84, Mur82, RM85, Ram86, VV86]. minus [Bak86, Mit86]. Miron [Har88]. Mirsky [BP84b, GHS87]. missing [Sza85]. missing-data [Sza85]. Missized [FG86]. Mixed [Bap89, Gol86a, Bap86c, Coh86, CMO89, Gol87]. Model [Pai85, Bak85, CZ86, ER89, Emr83, Far85a, FHV85, GLN84, GLN86a, GLN86b, KGÖ83, SS87b, Ste81a]. Models [BEN88, Mon84, BM85, BR89a, CM85, CD85, Des88, SF85, SS86, Ste85, Wer85b, Zen89, Zhe87]. Modern [Nie89, Sch84c]. Modification [AGG88]. modified [BF88, Giv82, Gow80, Mar86b, Mil83]. modular [Bec89]. module [Wim89c]. modules [Gur81a, Gur82, GL82, Ler80, Oko86a, Oko86b, Tho82b, You84]. moduli [Wim85b]. modulus [HRR87, Leó88, RT85]. Moler [Cam81]. Moment [Tor86b]. moments [JM89a]. Monge [ACHS89]. monic [Rod80]. Monomial [CLS87, Bre84]. monomorphisms [SK82]. Monotone [OvdD82, Ber86c, KS86b, Kwo89, Sto86, Wer84]. monotonen [Wer84]. monotonic [MS84]. Monotonicity [Lon81, Hwa86a, KO85a, NPW82, Wer85a, Wer87]. Moore [Cli81, Lav80, MR83, MR88b, PR81, Rob84, Rob85, RPV87, Rob87a, Sch83c, Sme81, Won82b]. Morishima [Wie85a, Wie85b]. morphism [PR81]. morphisms [MR83, MR88b, Rob80b, RP85b, RP87]. Moshe [Nor84]. Most [DKH86]. motions [RS86b]. Motzkin [Fri81]. MR [BS89, BKR80, Cai82, Sha89, Sta81]. multi [MK89]. multi-input [MK89]. multicommodity [Kar89]. Multidimensional [Cam84c, Bap82, BR89a, FL89a, Rag84, Son87]. Multigrid [Hac84, CZ86, Ver84]. multilinear [Pat85]. multimodal [BCS83, BCS88, CS87a]. multinormal [KR83]. multipass [RO89]. multiple [Gow80, Mar85]. multiples [Lan86a]. multiplexes [For87]. Multiplication [Vav88, JE87]. Multiplicative [RT83, AN83, Dia86, Lin87c]. Multiplicativity [GS83a, GS83b, Gol84, Gol87, GS82, Gol86a]. multiplicity [Wie84]. multiplier [Ong84]. multipliers [Ma85a]. multiply [Law82]. Multiprocessor [GHL86, ISS86]. multiprocessors [Saa86]. multisplitting [FM89, NP87]. multivariable [HS89, KW89a, Pow82, Cam84c]. Multivariate [Chi86, AO85, Bon85, Bos88, CM85, DM83a, Kab85, Kom83, Sza85]. multiway [KS86a]. mutation [DALSP89].

N [Cam84c, Son87, dPB86]. N. [BKR80]. Nahikian [CM84]. Napoleon [CD83]. Nashed [She86]. Near [Pet82, AZ83, EP88, KS88, Uhl81a]. near-characteristic [Uhl81a]. Near-fields [Pet82]. near-geodesics [KS88]. Nearest [Hig88]. Nearly [MSS84, CR86, For80, Tre87]. Necessary [Bec80, JY86, JY87, KG89, Kom83]. négatif [Yca82]. Negative [Mey84]. Nested [BT81a, DW83]. network [Zen89]. networks [vB85]. Neumaier [HN88]. Neumann [CD83, CL89, Fis87, Lew85b, MS80a, May85]. Newman [MW87b]. Nilpotent [DM80, Lon84, Tow80, Wu87]. nilpotents [Laf89]. nilradicals [RWZ88]. ninetieth [Gau83b]. no [Sta81]. Noble [Gov89a]. Noël [DR86]. Noisy [HM89b]. Non [YJ80, BK88b, BLAK89, CS87b, Gow85b, GR87, SPZ82, Yca82, Mey84]. Non- [YJ80]. non-Euclidean [Gow85b]. non-Hermitian [BK88b, BLAK89, GR87]. Non-Negative [Mey84]. non-Radon [CS87b]. non-tree-realizable [SPZ82]. noncentral [Neu85]. noncommutative [Sun82b]. noncooperative [Bir81]. nondegenerate [Gow81]. nondiagonable [BNS86]. nonequivalence [Lin86]. nonidentity [Mac88]. noninteger [Har81b]. noninteracting [TvdW88]. Noninvertible [GN85]. Nonlinear [Zen89, Ama86, FK85, HF85, Jef86, Pan89a]. Nonnegative [BGP85, BJ88b, Deu81, EJ89, GW80b, HK81, HL83, JG80, JP82, Lin87c, Pye80, SRPM82, Ash87, Bak87, BWT89, BL82a, CP81, Che84, CS82, Con85, DW83, DEHS84a, DEHS84b, Eag88, Els84b, FN89, Fri86, GJS89, Har81a, Har89, HRR87, Her83, JP81, Joh83, Kay87, Leó88, Lew89b, Lin89, nM87, RU85, Rot89a, Rot80, RT85, San87, Sch86a, Sch86c, SS81, Sha85e, Ste85, Ste87a, Szu88, Szu89b, Tam81b, Wea83, Wea88, Wu80, Yan84, Yca82, Zij84, Bar81a]. Nonnegativity [MNR84, RM85, Jai84, Mat86, Ram86]. nonnormal [Jia82]. nonnormality [EP87a]. nonpositive [Pan89c]. nonprincipal [Loe83, Loe84]. nonselfadjoint [LA89]. nonsingular [MP89a, Neu81b, Tho86]. Nonsingularity [Car88, CH89, JUW82]. nonstationary [BAG89]. nonsymmetric [HPS83, HPS88, OS84]. nonsymmetrical [BB89a]. Nonsymmetrizable [JY83]. nontrivial [Els82b]. Norm [Fon86, And85b, BK80a, BI82, Flo82, Li86b, Li87, Oku87, Ong84, Pok85, Tao84, Ver82, Wal86, Wan86, dC87]. Normal [Fia83, GJMW87, AO85, Bar89c, BQ85, CD85, Dav80, DR80b, Gab87, Hos86, Kab85, Lem88, Loe83, MM80, PL86, RC81, Sid89, Sun82c, Sza85, Tch89, Tho83c, Wim87]. normality [BN88]. Normalizable [Fan83, PK87]. normalized [OM89, RS82]. Normally [NW87]. normed [LR81]. Norms [Bar87, Zie88a, Deu80, FZ84, GS82, GS83b, Gol84, Gol86a, LT83b, LT88b, MS84, Mee87, Mer81a, PW89, Sta81, Tau80a, Wat82a, You81a, Zen84, Zie83, Koh84]. Note [Ano89t, Bai88, Ber87, Che83b, Cli81, Fre87, Hof85, Kie87, BP85, CS86a, Chu88, Cos89, Des88, DvdP81, Dry85, Eli88, Els88, EL89, Fie83, GH87, Gur80, Gus84, HG80, HN88, Har87a, Hwa86b, JP81, Kha85, Laf81c, Leó88, Mur82, Per88a, RU85, RTW83, ST88a, Sin88, SK84, Sun82a, Tre87, Wat87b, YH82, YB84]. Notes [FJM87, HB84, vL81, MV89, Smi88]. null [Fos86, Joh84]. Nullities [Rob87a]. nullspace [HS88a, HRS89]. nullspaces [AR87, HRS88]. nuls [Yca82]. Number [WWWS86, CZ86, Fri85a, Gow81, KP89, MT85, Row83, Row88, Sil86, Tau80a, Tau80b, Ver84]. Numbers [Zie88a, Dem86a, FN89, Que83, Sch86b, TC86, Zie83]. Numerical [AG89, Bun87, FHP80, Joh81, Naz84, Ruh83, Beb86, DEW80, Fie81a, FT88, GT82, GS82, GS83a, LTT83, Li86a, Li86b, Li87, LTT88, LT89, MS85, MS88, Moo82, Neu81a, Tam87c, Tam87b, Tam87d, Tsi84, Wei83].

Obituaries [Ano88-29]. observability [dSB81b]. observations [Giv84, Gro82]. observers [Hau83]. obvious [RPV87]. occasion [Ano85s, EH85, FL89b]. octets [Wil82a]. October [BP81]. odd [Mad83]. odd-even [Mad83]. off [Gen86]. off-diagonal [Gen86]. Olga [Koh84]. Oliveira [GL85a, BdP88, MV89, dPB86]. Olkin [Tho83a]. Omissions [BS89]. one [BC84, EN89, Kam89, Man84, Pat88b, Ste81a, Urs82, Wat86, dBP82]. one-dependent [Kam89]. one-dimensional [Ste81a]. ones [Bru80, BL82b, BS87c, BM89c, EP87b, Fri85a, Jam81]. only [RPV87]. onto [AL86b, HM86]. open [Sha85f, Sha89, SS87b]. operation [Hoc87]. operational [Ber86a]. Operations [Sch84c, Joh87]. Operator [AL86a, BDM83, GS82, Lan81, AB87, Bar86, BKR80, ER89, FN89, Fur88, Fur89, GW80a, Jód86, Jód87, Jód88a, Jód88b, Mer81a, Pat82, Tak83, Tam85, Tim85, Wei83, Wu89, Zen84]. Operators [BP88a, Cop84, Jin88, Adi82, AMT84, AT88, BP84a, BGP85, BP86, Bea88, BL89a, Bha81, BK88a, Bus86, CHP88, DD87a, Dan88, Fan83, Fan88, FK85, GS89, Gol86b, Gow85a, GR89b, Gro82, HW81, Hes86, Hol85, Kön86, KS86b, LP82a, LTT83, LTT88, LA89, Loe88, Mil83, Ong82, Ong87, PY80, Ptá86a, Rad84, Rei86, RS86b, Sid89, Sun82c, Tam87c, Tam87b, Tam87d, TW80, Web86, Cam84a]. Oppenheim [GH87]. Optimal [ELS82a, Els85, HPS83, HPS88, Wei83, BR86, Els84a, Gee89a, Had84, Hil82, Meh89, MFPW85, Rys85, Str83, SY89, Tor86b, WT85]. Optimality [FT84, KB89b]. Optimization [Nav89, Sch84c, Ste86, Acz85, CCH89, CW80, JMRT87, Kra83, RS89, Wol81, Yey84]. orbit [Bee86, Pok86].

Orbits [DM80, CL82a]. Order [NW87, Ste84, BKH88, Cro81, DN85, Fur89, Giv84, Had84, Har81b, Har86b, Hug85, Jód86, Jód87, Jód88b, LM87, Mit86, Mit87, Ran89, RS86a, RSR89, RV89, SM88, YH82]. ordered [Kle88, Law82]. Ordering [Ste85, BP85, GLN84, GLN86a, GLN86b, HS86a, Joe85, OS89, Wu80]. orderings [Bak86, BH87a, BPS89, GW85, Ste87a]. orders [Bec89]. ordinary [Spi82]. orientable [Wag89]. orientations [HM89a]. Oriented [Law82, Cor87, LW81]. orthant [MS84]. orthants [CLLW87]. Orthogonal [Gau83a, GL87, GP80, LSS89, LQ87, BF82, EN81, EG89, GR87, KN87a, Nie87, SL82, dBS86]. Orthogonality [DEW80, KS86a, Tak85]. Orthogonalization [Ruh83]. orthogonals [EKC85]. oscillating [Gar82c]. osculation [Hen83]. Ostrowski [Gau83b, SDGT81, Tho83b, Wim83b]. Other [Bre88, Sch82]. outer [She86]. output [DM83b, Wil83]. Over- [WN87]. Overdetermined [Jai84, Spi87, Str85]. overrelaxation [HY80, Kri84, Mar86b, MN87, NV84, VNC84]. overview [Joh88].

P [BKL85, Bea87a, Cop84, Dat86]. Padé [dB84]. pair [Laf80, Mit88, Tan83]. Pairs [GLS83, GL84, Kle88, Lei89b, BH88, HHJ86, Rag84, Sil88, Sun82a, Wat84]. papers [Ano86v]. Parallel [BCR88, BB86, Dat89, DI86, FG86, GHL86, GSV86b, Leu86, OS86, SS87a, SHW86, AC89, AP86, BEN88, BF89, DM86, FM89, Gad88, GKKL87, Mis85, MO86, NP87, PW89, Sco86]. parallelepiped [BV80]. parameter [BB89b, DS82, Man84, Pan89b, Väl88b]. parameters [AO85, Dat84, HJ86a, IS89, JM89a, KL88b]. parametric [CG86]. parametrization [FT89a]. parametrizations [DW89]. parametrized [Adk89]. part [Lás82]. Partial [BH87a, Bec89, Jun89, Sha88, BPS89, GL83, GJMW84, HS86a, JM89a, Mit86, ORV88, Ran89, Sha86b, ST88b, Tha89, VB88, VB89]. partially [Kle88, Pan89c]. particular [MSS89, Zie85]. partition [PS82]. Partitioned [GR80, Joh84, Bek88, Bhi88, Dry85, EHL81, RY85b]. partitioning [Pow88]. partitions [CS87b, PM80]. parts [dG82]. path [Bee84]. paths [HM89a]. Pattern [Mer80b, AHnLR88, Tho89a, Wie85c]. pattern-reduction [Wie85c]. patterned [Sza85]. patterns [BJ86, BF87, CLS87, DR80a, EJ89, FG81, JUW82, Joh83, JS89, SRPM82]. Paul [Cop84]. penalty [Bar80b]. Pencil [BV88, Bin84, Lin88, Pok86, Sha82c, Sha82a, Tau83, Väl88b]. Pencils [DK87, Bar89a, Bea84, Bea87b, CG89, ER87, FS82, Laf83a, LY88, Lin86, Sha82b, Wat83, Cam84d]. Penrose [Cli81, Lav80, MR83, MR88b, PR81, Rob84, Rob85, RPV87, Rob87a, Sch83c, Sme81, Won82b]. Performance [DKH86, Ric83]. performing [HTRG88]. periodic [BC87, BH89, Dub85, LPC89]. periodicity [GJS89, RS86b]. Permanent [vzG87, Bap86b, BS86a, BP87, DEHS84a, DEHS84b, For80, HCR88]. Permanental [Far85b, Gib80b, MRW81, Min87b, Krä87]. Permanents [DG87, Bap84b, Bap84a, Bap85, Bha84, Chu89, Els88, For87, GH87, Hwa86a, Hwa86b, KR81a, Lon81, Min85, Min87b, Nis88, Pat84, Pat88b, Son88, VJ84, VJ88]. Permutation [GHW82, JL87, Pat82, Cha82, Gib80a, Sha85a]. Permutation-matrix [GHW82]. Permutation-operator [Pat82]. Permutational [ML83]. permutations [Bar83, Bar88b, MV88b, Sun82c]. permute [BS87b]. Perron [Ash87, Bap87, CS85, DW83, DN85, Mar84a, Mey89, San87, Szu88, Szu89b, Uch86]. Perturbation [BDM83, Cam84a, Coh86, CMO89, ES82, GdHZ89, HP82, KS85, Ste84, BK88a, Hac84, Kra81, LM88, Ron82, Sen88, Wei89a]. Perturbations [KS86b, Bea85, CR86, KL88b, MP83, Pok86, SS88, TZ87]. Perturbed [Gre89, Hun86, LVWC88, Pan89b]. Peter [Har88]. Phadke [BKR80]. phenomenon [CW80]. Philip [Hay82]. pi [CLP82]. Picturing [Lon84]. Piecewise [KL87, GLR83]. Pillis [Bea87a]. pipelined [HS86b]. pivoting [FP81, GN82]. Pivots [GN85]. planar [Bry82]. plane [BN83, CCH89, Djo82, RS86a]. planes [LP82b, RSR89]. plans [EKC85]. plants [LST88]. Plemmons [Bar81a]. plus [HR88, HR89, Wat86]. Point [AL86b, Cor87, Sch83b, ST87b]. points [AT88, BV80, CF86, CH80, Con81, Doo89, Dru87, FS85, HS87, Str85, Zie88b]. pointsets [Tha89]. pointwise [Adk85, Gek81]. pointwise-local-global [Adk85]. Poisson [Ver84]. polar [IS89, Uhl81a, Uhl83]. Polarity [Cor87]. Pole [KNC89, SY89]. poles [Kam88, WSCP89]. Politecnica [BV89]. Polya [Mer80a]. polygons [MP89a]. Polyhedra [Kar89, BN83, CP89, CS85]. polyhedral [Väl88c, Vei89b]. Polynômes [Tre85a]. Polynomial [Coh87, KP86, Ma85b, Özg86, BP87, Bee84, Bee86, BZ89, BS88a, BN88, DvdP81, Ele87, FP87b, How80, KL88a, Ker87, KGÖ83, Lás82, Lin87a, Ma85a, MZ89, PW85, Ram86, Rod80, Sal81, SF85, Sha82c, Sha82a, Sha82b, Tau83, Uhl81a, VD83, Vav89, Wer83, Wim83a, Wim89a, Zhe87, dO82]. polynomial-model [KGÖ83]. Polynomials [BS87b, Fie85, Gau83a, Hol82, Wei87, Bar84a, Bar88a, Bee80, Bos88, Cal82, Cop84, Cyb87, EG89, GLS83, Gał81, GLR83, GB84, GR87, HM81, Har84b, JL87, Kam89, KK89, LT83a, Lan86a, Lás81, Lon82, MS80a, Mar85, Mar86a, Men86, MRW81, Nav89, Sch83a, She87a, Sil86, Tak83, TG87, Tre85a, Wan83, dBS86, BKL85, Cam84e]. polytope [CS87b, PS89, ST88a]. polytopes [BK80a, Eav84, GT89b, Stu88b, Gib80b]. Portugal [GL85a]. posed [Sch86b]. positifs [Yca82]. position [Zăl85]. Positive [AHnLR88, Bon83, CF86, CS87c, CM83, DI86, Fer80, GM85, GJMW84, Hig88, Kay85a, Kay85b, RV89, Sai82, AN83, Adi82, AMT84, AT88, And83, And85b, And87b, Bap85, Bap89, BHH84, BJ84b, Bek88, BH87b, BH89, Bus86, CG86, CHP88, CS82, Dem82, Dru87, FG81, FJMN85, Fri85b, Gow85a, GW80b, Gro82, GHW82, HM84a, HL83, HW88, HW81, HW87, Hu87, JE87, KH84, KS80, Lei81, Loe84, Loe88, Loe89b, Lut84, Mar84c, MPJ81, Meh84b, MM88, Mur82, OvdD82, Pin85, PH81, Pol86, Rob82, RU87, Sch82, Sch81a, Stu88b, Tan86, Tom85, WH85c, Wu88, Yca82, dBP82, dBJP82]. positive-semidefinite [MPJ81]. positive-semidefinite-preserving [BHH84]. positively [Ste82]. positivity [BHJ85, Ber88, Joh87, RO89].

Possible [BJ84b, Dan87, May88, MM88, Tau85]. possibly [Wer85b]. potent [JP82]. potential [Hur89, Neu81a]. potentially [JS89]. pour [Mne89]. Power [Sin81, Lut84, Sid89, SRPM82, Wim80, You81b]. Powers [Kha80, Wim85b, BJ88b, BF81b, Bha84, BL82a, FH88b, Hea83, HJ86d, KMV87, KO85a, Lin89, May84, nM87, PL86]. PPS [WT85]. Practical [RP83]. prediction [Boj86]. Preface [Ano86v, BHL89b, BPS80, BBV88a, BS88b, BS88c, ENV84, GSV86a, Lan86b, ORS85a, Rot89b, dO89]. prescribed [BH88, BQ85, BM89c, CM81a, Cai81, Fri85a, Mar84c, Mat85, Per86, Per88b, Roh81, Row83, Row88, Sil87a, VV86, Zab86, Zab87, dO82]. Presence [CV87, Fle88, vV87a]. Present [Ano88-30, Ano88-31]. preservation [Wat86]. preserve [Bea81a, BP88a, BHJ85, CL82c, How80, Loe89a]. preserver [LT83b, LT88b]. preservers [Bea83, BP87, CLT87]. preserving [BT81b, BHH84, BP84a, BGP85, BP86, CJR87, FM87, Fur89, Hia87, LTT83, LTT88, Mar84a, Mat86, PH81, Rad84, Tam81a, Tam81c, TW89, Wat82a]. prespecified [RS89]. Prestabilization [VM87]. primal [Eav84, Spi85]. primal-dual [Spi85]. primary [BC85]. prime [Bry82, KGÖ83]. prime-field [Bry82]. primeness [Tak83]. Primes [dCG81]. primitive [DSD88, Sha85b, Sha85c, Sha85a, YB88, Zha87a]. primitivity [FB81a]. Principal [Loe86, CM81a, Dan87, DGK80, Eag88, HL84, Her83, HMS87, JR81, Joh81, MNR84, Mur82, Pat88b, PV82, RM85, Ram86, Sil87a, TC86, TST86, dO82]. principle [Adk85, Gur80, HY80, Naz84]. principles [Ish87, Sto86]. probabilistic [Wat87c]. probability [AN83, AT80, Ata89, Hun82, Kab85, NS85, WT85]. Problem [AGG88, BG84b, ES82, GGvM89, Mey89, Acz85, AA89, AO82, BAEGL87, Ber81, BR86, BK80c, Caf81, Cam89, Cao87a, CH83b, CPV89, Dia86, Dic84, DGKW84, DL89, Faĭ89, FG80, FH88a, Gee89a, Gee89b, GLN86b, GL83, Hac84, Had84, Ima84, JP87, Jód87, Kab85, Kra83, LN80, Luk81, Mar81a, MN87, MC81, Pan89a, Pon87, Ptá86a, Roe86, Roe87, Rys85, Sch83b, SS85, Sha86b, Stu88c, Tog80, Tre85b, VB88, VB89, WE89, vB85]. Problems [BD80, BB86, ES86, Fre87, FG86, MNP85, NdPV84, VV87c, AC89, Ant83, ACC88, BK80a, BR87, BF88, BK81c, BB89b, Bjö87, Boj86, Bro89a, CZ86, Chu84, Chu88, Dat84, DD87b, EL89, FSZ85, GM89, GH80, Hun82, Ish87, Jód86, Jód87, JJS89, KP80, Ker83, Kov87, LY88, Leu86, LT83b, LS87, LT88b, MS86, MS87, Meh89, Mes89, O’L80b, Pin85, Prz83, Sch86b, Sha83, Sha85e, Sha85f, Sha89, She86, Shr86, Wei89a, Wu89, Zab89]. Proceedings [AL86a, Ano88-32, BHL89c, NdOH89, Ple81, Bru88a]. process [BG82]. processes [Abe82, Har86a, Lin87c, RO89, RT83, SK82, SK84]. Processing [APP88]. Processor [Luk86]. processors [AP86]. product [AB84, AB85, Cha87b, Cri88, Fie83, FM88a, Gek81, McF80, MV88a, Mur82, RU85, Ste81a, Vei89b, Wal86, Wat88a, Wim83a]. Products [Bru84c, KN87b, Sha85d, Wu87, Wu88, Art83, Art86, BR80a, BS85a, BWT89, Bru85, CL82b, CFKK82, CS82, FMN83, FR82, GW80a, GR89a, GL84, JAK89, KB89b, KN87a, Laf89, Oku87, Pat84, Pol85, SS81, Won82a, Laf85b, Dat86]. Profiles [Ano88-33]. programming [Bar80b, BI88, BK80c, CM81b, Dax88, GA81, HF85, LP80, O’L80b, Pan89a, Väl89b, Zim80]. programs [Vei89a]. progress [Dan89]. projection [AC89, AL86b, FT88, Hal80, Har81b, Spi85, Spi87]. projection-recurrence [Hal80]. Projectionally [PL88]. Projections [O’L90, Ste89, ALZ81, DI85, MZ89, Nak84, Nis85, Ptá84a, Tre85a]. Projective [KK89, Adk89, CL82a, Djo81, Far81, Kes82, MM82]. projector [Eva87, KS88]. projectors [Kov80, KS84]. promote [Bon83]. proof [BR89b, CK89, Fur89, Ger89, MS83, ORV88, Tho89b, Was84, vL81]. proofs [CCH89]. proper [Fan88, KÖ85b, LT86, Sch86b, Tow80, Web86]. Properties [Gow85b, NW87, She87b, Zie87, BPS89, BC83, BH89, Che83c, Dem82, Els84b, Fri85b, Gee89b, Hel87, JP84, JP88a, Joh84, Joh87, KO85a, LM88, Li88a, Lin89, MS88, MR88a, MM88, Pan89b, Sch86a, Smi87, Spr81, Wu80, YL85, Zhe87]. property [BS86a, Ber87, Doo82, Dow85, GH89, KS80, Que87b, RP85a, Wat83, Wer85a, Whi80]. property- [Wer85a]. proportional [Tak85]. Propriétés [Hel87]. Pseudo [O’L90, Ste89]. Pseudo-Inverses [O’L90, Ste89]. Pseudoconsimilarity [BHH87b]. pseudoinverses [DS80, GW80a, JG80]. Pseudomonotone [Gow89]. Pseudosimilarity [GL82]. pseudosymmetric [BG82]. Pukkila [Bru88a]. Puntanen [Bru88a]. Pythagorean [Tau82].

Q [Pul87]. quadrangles [Tha89]. Quadratic [Bai88, Koh84, Loe80, LP80, PV86, Sai87, Sai88, Väl89b, AT80, Bec80, BK80c, CC84, Che87, GLS83, Gee89a, JJS89, LY88, Lut84, Mad88, MJ81, Mar81c, Meh89, Neu85, O’L80b, PL86, SY89, Tur85]. Quadratic-programming [Väl89b]. quadrics [LP82b, Wat82b]. Qualitative [Jef86, JKvdD87, KB89a, Man82, Wie82, Wie84]. quantitative [Dan86]. Quasi [CHP88, BK88b, BLAK89]. Quasi-positive [CHP88]. quasi-Toeplitz [BK88b, BLAK89]. quasicommutativity [CP85]. quasiconvexity [Fer81]. Quasidirect [Fie84b]. quasitriangular [GB87]. quasitridiagonal [Riz84]. quaternion [ET81]. Quaternionic [Far84]. questions [BG81a, GVK^{+}83]. quiver [HPW87]. quotient [LM80].

R [And87a, Cam81, For86]. R. [Bea87a, Min87a]. radial [Tam87a]. Radical [EN81, HP86]. radii [GS82, GS83a, Moo82]. radius [BFP88, BN89, BH85, BS87c, Bur86, Els84b, FN89, Fri88, GT82, Li86a, Li86b, Li87, Nus86, PY80, San87, Sch86c, Sta87, Tam87b, Yua88]. Radon [CS87b]. Ramsey [FRS82]. Random [NW87, Nor84, Tak82, Kam89, KS86b, MT85, OP82]. range [Beb86, DEW80, Fie81a, LTT83, LTT88, LT89, MS85, Tam87c, Tam87d, Tsi84, WH85b, WH85c]. ranges [Joh81, MS88]. Rank [AGG88, Bea83, Cha87a, FM87, Fos86, Str83, BPR87, Bea81a, Bea81b, BP84a, BGP85, BP86, BP87, Bea88, BP88b, BL89a, BS81, Bin86, BH81b, BR80b, BC85, CP81, CR86, Che83b, Che84, Cob86, DS80, DS89b, EN89, FH88b, Gol86b, GB87, GR89b, HS86a, HHM89, HS84, JP81, Laf80, Lic85, Loe89a, MN87, Pat88b, Pow82, Rob80b, SSS82, SS86, Syl86, Wat86, Woe89, dGH83]. Rank- [Bea83, Bea88, BL89a, CR86]. rank-deficient [DS80]. rank-one [EN89]. rank-one-plus-scalar [Wat86]. Rank-preserving [FM87, BGP85, BP86]. ranks [AL80, Bea82, Bea83, Pon87]. Ratcliff [BR89a]. rate [Bha81, Eli88, You81b, Zha89]. rates [DALSP89, EP88]. ratio [CZ86, Ver84]. Rational [Ant88, Hel89, Ruh84, Sch85, AA89, BR87, BH88, BM82, BF82, Dow85, Fuh83, Fuh84, GM89, HJ86a, KÖ85b, LST88, Ma84, RR84b, Ray83, SD80, SF85, Sha88, Tau80a, Tre86]. ratios [MSW83, TC86]. Rayleigh [LM80, MSS84]. Rayleigh-quotient [LM80]. reachability [BNS86, CS87c, NS89]. reachable [HPW89]. reaching [Lin87c, RT83]. Real [Hig87a, PS87, Wat82b, Wea83, Wea88, AG89, BH81a, BG82, Bro89b, But88, BS83, Den81, FJMN85, GS87, HHM89, HS85, HMS87, Lás80, Lei89b, MS85, Nis85, Rob82, Sco85, Tau83, Uhl81b, Uhl83, dG82]. realizable [SPZ82]. realization [CS89b, GL83, JM89a, Sha86b, Sha88, VB88, VB89, WS83]. realizations [BH88, Gre88, Ran89]. Recognition [OS89]. recognizing [Kle87]. recombination [DALSP89]. reconstruction [EHL81]. rectangular [BS82b, CG80, Dem86a, Gro80a, Mit82, Neu80, PM80]. Recurrence [Min85, Hal80]. Recurrences [Gre89]. recurring [Sin88]. Recursive [JM89a, Kle87, AN83, APP88, GKK86, GKKL87]. Redheffer [RB89]. reduced [Bar80b, HTRG88, Ran89, Sch86a, Wat82a]. reduced-gradient [Bar80b]. reducible [nM87]. reducing [Lin86]. reduction [BG82, HHJ86, Laf86, Mad83, SA89, Wie85c]. Reed [KL88a]. Referees [Ano89u, Ano88-31, Ano81u, Ano81v, Ano83l, Ano84q]. referring [BH87a]. Refined [PNO85]. Refinement [MSS84, MP89b]. refinements [Sen88]. Reflections [GL87, Djo81]. Reflexive [HG88, KMK85]. regions [Bon85]. Regression [VV87b, Dry85, RP85a, ST88b, TC86, Wer85b]. Regular [CLLW87, Rob82, Bar89a, Bea86, Bie84, CL86, FRS82, GM80, HH89, Ito82, Lew85b, Li89, Lin86, Lin87a, McK81, MR83, MR88b, Wei89b]. regularity [Gar82b]. Regularization [ES86, Hil82, Mac88]. regulation [BFM89b]. regulator [Sch83b]. reguli [BF82]. Reinhardt [Tim85]. Related [BA80, O’L80a, BFP88, BG84a, Bur86, Fei84, Fon86, GVK^{+}83, KO85a, Kar89, Kit87, RC81, Sch86a, WH85b]. relation [GA85, Her89, Kob85, Ste81a]. Relations [Djo81, Djo82, Uhl83, EN81, Rob86, Uhl81b]. Relationship [WB81, Bak86, GH83, JE87]. Relationships [VV87b]. Relative [Han86, Tak83]. relatively [Gur85a]. relaxed [FM89]. remainders [Bar88a]. Remark [BS85b, Cha88, DR80b, Grü82, Ros84b]. Remarks [ET81, Fie81b, Fuh86, Lew89c, Ong82, SSS82, Tau85, Vav89, Zie88a, dB80, BdP88, LP81, Sch82, Wer89, Zie83]. removal [Har83]. reorderings [OH85]. Reorthogonalization [Sim84]. Report [BF85, GL85a, Gro85, Hol85, Laf85b, Lew85b, O’C85, SW85, Tim85, Wim85a, KLM86]. Reports [Ano82o, Ano82p, Bar82, Ano88p, Ano88-28]. Representation [Vei89b, BHH87a, GM88b, HK83, Kle88, Lip80, Mag87, Que87a, Urs82]. Representations [Jam83, Jam88, Dam80, Dia81, Hal83, HR89, KP86, LSS89, Neu89, Tau80b, Vin89, Wil83]. represented [BN83]. reproducing [Sai82, Sai87, Sai88]. rescaling [Hu87]. Research [Sch84c]. Residual [Cao87b]. resolutions [Bre84]. Resolvent [Rot81, Ele87, Tad81]. Resonance [Lor81]. respect [BM89a, Hes86, LT83b, LT88b, Rob85]. Restricted [AGG88, Mad88, BM85, Don85, Ger84]. result [ORV88, Was84]. resultant [KW89b, Lub87]. Resultants [Har87b, Nor89, Wan80]. Results [Bru89, CMR81, Ant83, Ber86b, BH87b, DM83b, Kit87, Kwo89, LN85, MK88, Mur89, Neu87, Sha85f, Sha89, WH85b, Wie82, Wie84, Wie85a, dP80]. retention [WT85]. retrocirculant [Wan81a]. Revealing [Cha87a]. reversal [DGKW84]. reverse [Har86b]. Review [And87a, Bar81a, Bar87, Bar80a, BKL85, Bas85, Bru84b, Bru84c, Bru88a, Bru89, Cam81, Cam84e, Cam84b, Cam84c, Cam84d, Cam84a, Car85, Cop84, Dat86, For86, Gov88a, Gre83a, Har88, Hay82, Koh84, Lan81, Mar84b, Mey84, Mon84, Nau87, Nor84, Ple81, Ple83, Poo81, Pul87, Ral86, Ros84a, Sch84c, Son87, Ste87b, Tho83a, Gov89a, Lan86a]. reviews [Ano88o]. revisited [CC84, Gre83b, Har83, Har86b, KLM84, New80, New83]. revolution [CW85]. Riaz [Gov88a]. Riccati [BC87, Bye87a, JAK89, Lin87b, RR84b, RV88, Wim84, Wim85a]. Richard [Gre83a]. Richardson [OS84]. ridge [TC86]. Riemann [Roe86, Roe87]. Ring [ISS86, EI87, Gib80a, Gov88b, HP81, Koh85, Piz83]. rings [Bol85, BK88c, Bro89b, Emr83, GS81, Gur81b, Isa80, KÖ85b, Lew85b, You84, For86, Nau87]. Ritz [MSS84, vV87a]. Robert [Bar81a]. Robust [Bos88, KNC89]. Rodman [BKL85, Cop84, Dat86]. Roger [And87a, Bea87a]. role [PL86]. rook [BP87]. rook-polynomial [BP87]. Root [BH83, Ter87, Wim89d, DW83, Lip86, Sco86, Szu88, Szu89b, TG87, Bap87]. root-free [Lip86]. Roots [Den81, Hig87a, TG87, AS82, FSZ85, Far85b, HM81, Krä87, Mar84a, Mar85, SSZ83, She87a, TC86, TST86]. Rosenberg [BS82b, BNV82]. Rotation [APP88, Kit87]. Rotations [DI86, GL87, ER87, GH80]. Roth [Gur81a, Gur85b, Har83, Wim88]. Rounding [Kie87]. Roundoff [Woź80, Fle88]. Roush [Pul87]. Row [GLN84, GLN86a, GLN86b, Bru80, BM89c, PS88, RS89]. Row-ordering [GLN84, GLN86a, GLN86b]. rows [Zab87]. Ruhe [Cam84d]. rule [BI82, Har83, Tho82a, Tho82b, Ver82, Wan86, Wan89]. Russell [Bar80a]. Ryser [BR80b].

S [Ano80a, Cam84b, Mar84b, Nau87, Ple81]. S. [BKR80]. S.O.R. [Neu81b]. Saddle [Str85]. Saigal [Hof85]. sample [BS86b, Nel85, SS85]. sampling [BS83, WT85]. satisfying [Jai84, Joh84, Ma84, RS82, Rot89c, Syl86, dPB86]. Saturated [Ata89]. Satz [HH85]. Scalar [Dat86, VB88, Wat86]. Scale [GP80, Dat89, LVWC88, MNP85, Neu80]. scale-based [LVWC88]. Scaled [O’L90, Ste89, Lee89]. Scaling [Elf80, PL82, BR89a, ELS82a, Els84a, FL89a, Sch89a, SB87, SH88, SB88].

Scalings [RS89, Rot89c]. scattering [BLAK89, Bos88, BK86]. scattering-domain [BLAK89]. Schatten [Kit89]. Scheduling [OS86, ER89]. Schein [DS89b]. scheme [Had84, PW89]. schemes [Ban82, Bie84, GLN84, GLN86a, GLN86b, HPS83, HPS88, IS89]. Schmidt [Ruh83]. Schur [HH85, ACC88, BS85a, Bap86c, BF88, BH83, Bre88, BM88, Car86, Chu86a, Fie81b, HH85, Ima84, JMRT87, KO85a, KL88b, Mad88, Oku87, Ong84, Oue81, PV81, Wal86]. Schwarz [FL89b, MM81]. Sciences [Bar81a, Car85]. Scientists [Cam84b]. scrambling [Rho89]. Second [Bru88a, Cro81, Ste84, Ano86a, BN88, DN85, GM87, Har88, Jód86, Jód87, Jód88b, LM87, Neu82a]. second-order [Jód86, Jód87]. Seidel [Mil87]. Selection [VV87c]. Selective [LM88]. Self [MP89a, WE83, WE88, WE89, Fon86, Lem88]. self-commutators [Fon86]. self-complementary [Lem88]. Self-equivalent [WE89]. Self-inscribed [MP89a]. Self-similar [WE83, WE88]. selfadjoint [DD87a, Dan88, LP82a, LT83a, RR84b, Wat86]. Semencul [CK89]. semicommutative [WB81]. semiconsimilarity [BHH87b]. Semidefinite [Hig88, AHnLR88, AT88, And85b, Bap85, Bap89, BHH84, CG86, Gow85a, GW80b, HM84a, HW88, HW87, Loe84, MPJ81, Meh84b, RR84a, RR84b, RU87, Sch84b, Wu88]. semidefiniteness [Bek88, CC84]. semigroup [Yan84, dCG81]. Semigroups [Ama86, RS86b, Sis86, Tam81b]. semilinear [BHH87a]. semimonotone [JP87]. seminormal [Bjö87]. Semiring [BP88b, BP88a, Tam81c]. Semisimilarity [HP81]. semisimple [AL89]. semistability [HS85, SH88]. semistate [Cla88]. semisymmetric [AN83]. Seneta [Mey84]. Sensitivity [FM86b, Sen88, Fei89, FHV85, HS89, Jon84, LST88, TZ87]. Separating [Cun82]. separation [How83, MLS87, SS85]. September [BV89, NdOH89]. Sequence [Ruh84, CLS87, JM89a]. sequences [ACHS89, BAEGL87, DS82, MT85, PS89, Sin88, VB88, dBS86]. serial [HG88]. series [May85, Tho89b, Wim80, You81b]. sesquilinear [Gow81]. set [AT88, CS87a, Els82b, Had86, Ray83, Rod80, Rys81, Sha85b, Sha85e, ST87b, Wol89a, Zha87a]. Sets [CS82, Ata89, BPR87, Bea82, Bea83, Bry82, Bye87b, CP85, Cha82, CvRS81, DS82, Gar82b, Gur85b, HK83, Kle88, Laf83b, Laf86, Law82, Ma85a, Ma85b, NVD80, Neu89, Rob80a, TW89, Zăl85, Zij84]. several [She87a]. sexually [DALSP89]. Shared [GHL86]. Shared-Memory [GHL86]. Sharp [Yua89, Mit87]. sharpenings [ORV88]. shear [Bak87]. Sherk [RS86a]. Shift [BGoMSTAUTA^{+}86, PS87, Bak87, JZ85, Zha89]. shifts [Tho89c]. Short [Cam84a, BR89b, Ger89, Vei82]. Shorted [Gol86b, MP82, Mit86]. shortest [DSD88, HM89a]. sided [MŽ83]. Sign [BF87, JUW82, Joh83, JORvdD87, Tho86, Bye87a, FG81, Gar82b, How83, JS89, Kle87, Tho89a, Wie85b]. Sign-nonsingular [Tho86]. Signal [APP88]. Signsolvability [KLM84]. Silbermann [Bas85]. similar [ES80, WE83, WE88]. similarities [BJ88a, KS84]. Similarity [Gur81b, Gur88, Hia87, Joh80, Piz83, AO82, Bar83, Bar88b, BM89b, DR83, DR88, Gur80, KL86, Laf86, Loe86]. Simo [Bru88a]. Simple [Qi84, Bix82, Chu84, JM89b, Vas85]. simplex [Dan89, GA81, Ogr88]. Simplification [JY83, JY86, JY87]. Simultaneous [Adk89, BPR87, Laf80, Laf86, LM80, Mit82, Bec80, BC84, DI85, dB84]. Simultaneously [PK87, CP85]. sine [CKN85]. Singer [BM82]. single [JP88b]. single-hook [JP88b]. Singular [Bai88, BV88, Chu86b, wEC87a, KNC89, Luk86, Pai85, PV86, PS88, Qi84, Ste84, BH87a, BNV82, CR86, Che83c, CLLW87, DR83, DR88, Gee89a, GP89, GR89b, Hac84, Har81b, HS83, Joh89, Kra81, Lew89b, Lew89c, LT83b, LT88b, LS85, Meh89, Neu80, Par86, PY82, Que87b, Sch84a, Sod87, Wan89, Wat84, Wer85b, You83, Ros84a]. singular-value [You83]. singularly [LVWC88]. Six [BM88]. Skew [KGÖ83, GL84, Laf85b, Lee80]. skew-centro-Hermitian [Lee80]. Skew-prime [KGÖ83]. skew-symmetric [GL84, Laf85b]. skewfields [Lew85a]. Slightly [Gre89]. Small [BS88a, BL82b, Ste84, MP83]. smallest [Joh89, Lew89c]. Smith [Tho82b]. smooth [Vin89]. smoothing [Mac88]. social [Zen89]. solitonic [LA89]. Solomon [KL88a]. solubility [BK81c]. soluble [Gam88b]. Solution [Acz85, BCR88, BDM83, BA80, BD80, BB86, DI86, GH80, Hal80, ISS86, Poh89, Tre86, Ant83, wEC87b, wEC87a, GKK86, GR87, Had84, Har81b, HG89, Jon84, Kra81, Mil88, VB89, Ver82, Wan86, Wan89, WT85, dSB81b]. Solutions [HS88b, Adk85, AC89, BI82, BS88a, wEC89, DM86, Don87, FT89a, HW89, Jód88b, KH84, Kay85a, Kay85b, Kwo88, LLT84, MW87a, Meh89, Spi87, Stu88a, Wim84, Wim85a]. solvability [JUW82, Kle87, Man82, Nic82]. solve [Ric83, VB88]. solved [ER89]. solver [AG89, Mis85]. Solvers [DKH86]. Solving [Bye87a, SS87a, WWWS86, dH87, AS80, CZ86, DD87b, EL88, Lip86, MN87, Neu80, O’L80b, Pan89a, Par80, Spi85, Tre87, Ver84]. Some [BPS89, Beb86, BdP88, BH82, Bru88b, BK86, Che83d, Chu89, DM83b, Elf80, FSZ85, Giv84, Gro82, JP84, KP80, KO85a, Kov87, Kwo89, LN85, LP81, MK88, MR88a, May88, MV89, Mur89, Pin85, Qi84, RC81, Smi88, Tau80b, Wie82, Wie85a, Wol81, Zie83, Zie88a, dB84, ACC88, AL80, ABI84, BG81a, BK88a, Cha82, Che83c, Cop81, Dat89, DM82, Fei84, Gab87, GM89, HPY86, HS86a, HHM89, HG87, KW89b, LT83b, LT88b, Li88a, McD80, MP82, MG85, MM88, Prz83, Sch82, Sen88, Sha85f, Sha89, She86, Shr86, Spr81, Uch86, Vei82, Wer83, Wie85c, Wu80, ZM81]. sometimes [Bon83]. SOR [EL89, Fre87]. sort [HM89b]. sorting [HM89b]. South [Neu82b]. space [BFM89b, Djo81, Far81, Fos86, Gow85a, HHM89, Joh84, KN87a, Lip86, RU85, Sal81, Sch83b, Wil83, Cop84]. Spaces [Bea81b, Kas89, BG81a, BP84a, BS82a, BH81b, Bre86, Cob86, GS81, Hal83, IS89, LR81, Loe80, Nar86, Sis86, Ste86, Syl86, Tak85, Tim85, TW80, Vei89b, Wim83a, Gre83a]. Spain [BV89]. Spanning [JB85]. Spanning-tree [JB85]. Sparse [BD80, Fre87, GR80, GL87, Kau87, NdPV84, BJ84a, BJ86, GH80, GLN84, GLN86a, GLN86b, GMSW87, HPS83, HPS88, HS86b, O’C85, Ple81]. sparsity [AHnLR88]. Special [Ano80q, Ano85t, Ano87s, Ano88-34, BB82b, BF83, DHLW85, ORS85b, Tho88, Ano86a, Ano86v, BP85, GSV86b, LT83b, MV88a, Sch85, Vav89]. specified [FM86a]. Spectra [Els82c, HJ86d, BJ84b, Bha81, BF81b, BM82, Con85, Fan89, Gro82, HB83, Her83, Lei81, Pok85, Zen84, Bru89]. Spectral [ALZ81, BDM83, BC83, CG89, Coh88, Els85, JJS89, NT80, Wie84, BR87, BFP88, BH87c, BK81b, BN89, BH85, BS87c, Bur86, Dua89, Els84b, EJ89, FN89, FZ84, Fri88, GM88a, Har84b, Hoc84, Jia82, LY88, LM88, Li86b, Li87, LT88b, MŽ83, MH81, Nus86, Prz83, PY80, Rot80, RCA88, San87, SS88, Sch86c, Sta87, Tam87a, Yua88]. spectral-circle [Rot80]. Spectrally [Sil88]. spectraloid [Deu81]. spectrum [AZ83, Bur86, DN84, Gha87, Hes86, Mar84c, Moh82, MO87, Moh88, Mül87, Sen84, Smi86, dP80]. sphere [Zie88b]. spheres [Tur85]. spherically [CvRS81]. Spline [Gro80b, Dub85]. splines [DM83a, Sch83a]. splitting [Lei89a, Wie84]. splittings [Li89, Sch84a, Zha87b]. spread [Bee80, JKW85, Sco85]. spreads [Jun89]. Square [BH83, Hig87a, AS82, Bak87, BM89a, Hea83, RSR89, Sco86, Tho89a]. square-root [Sco86]. squared [Cri88, Ver82]. squared-distance [Cri88]. Squares [APP88, BD80, BB86, ES86, Fre87, GP80, VV87c, Bak85, Bjö87, Boj86, Bro89a, BF82, Dax88, ET81, EL89, GH80, Kam89, MNP85, MN87, NdPV84, RP85a, Rao85, Spi87, VV87b, Wan86, Wei89a]. squaring [KG89]. Squeezing [DKH86]. SSOR [HN88, MK88, NV84]. Stability [Bjö87, Bun87, GN82, Her86, RR84a, RR84b, BB89a, Bir81, Bon83, Car84, DD87b, FD89, Gee89a, Gee89b, Jef86, JKvdD87, Jef88, Jur82, KB89a, LS85, Meh89, Ran89, RO89, Tad81, Tog80]. Stabilization [BKH88, Fei89, Abe82, KW89a, LN89, Ran89, VV86]. stabilizer [Tan83]. Stable [ABO88, Dem86b, DK87, AA89, BH84, Cai84, FZ84, Har80, JS89, KÖ85b, Man81, OvdD82, SS87b, Tad86, YB84]. stall [Dan89]. standard [Far85a]. star [Bak86, Far85b, Gow89, HS86a]. State [BU84, BU88, RW89, BFM89b, Cob86, Sch83b, Tch89, Wil83]. state-space [BFM89b, Sch83b, Wil83]. states [Had86]. static [Tch89]. stationarity [Lin87c, RT83]. Stationary [FM86b, Hun86, Sen80, Sen88]. Statistical [Car85, RY85b, Nor84]. statistics [Ano85t, Ano86a, Ano86v, HLG85, ORS85b, Oue81, Wie85c, Won85, Bru88a]. Stein [BS82b, BNV82]. step [DS82, May87]. Stephen [Poo81]. Stewart [Cam81, Ple81, Sun82a]. Stieltjes [Luk81]. Stiemke [BTD84]. stiff [MC81]. Stochastic [And89, Cob86, CS86b, Joh80, MSS84, PL82, Ata89, BR80a, Bap83, BS82a, Bru84a, Bru88b, CF86, FS85, For80, Fry85, Gib80a, Gre88, Har84a, Hwa86a, LO81, Lon81, Mar84c, nM87, Rho89, Sin81, Spr81, Gib80b]. strain [FHP80]. Strang [Gov89a]. Strategies [PS87]. Straus [Gol85]. strengthenings [PN85]. strict [Tad81]. strictly [BdP86, Fan88, FK85, Pin85, Ste82, dBP82]. Strong [Bak87, Bak85, Bun87, BCS87, FK85, Hau83, Ish87, SS81, Kle87, YB88]. stronger [Bap84a]. Strongly [FRS82, HH89, Bie84, Tad86]. Structural [BH89, Cla88, Mal89, Tch89, Wie82]. Structure [JP88a, Yan84, dBJP82, CS87a, DW89, EP87b, Fry85, GKK86, GKKL87, HS83, HRS89, Kle87, Kun81, LPC89, Lub87, MP83, MT85, MLS83, MLS88, Nis85, RWZ88, RB89, Rom86, Sch84a, SS86]. structured [MV88a, YL85]. structures [CS89a, Kes82]. study [DM82, LA89, SH88, Tog80]. Sturm [Bro89b, Usm87]. subalgebra [Zas83]. subalgebras [Gam88b, Laf81c, Laf85a, Tow80]. subclass [JP87]. subdefinite [Fer80]. subdeterminant [MM80]. Subdivisions [Eav84]. subdominant [HRR87, Leó88, RT85]. subgradient [Tao84]. subgroup [Tan83]. subgroups [Vas85]. Subject [Ano80r, MJ81, Tch89]. sublattices [Vei89b]. Submatrices [SPZ82, CM81a, Dan87, DGK80, JR81, Joh81, Loe83, Loe84, Mer81a, Rob87a, Sil87a]. submatrix [DN84, Pat88b, Son88]. submodular [FT89b, Fuj84]. submodules [Emr83]. suboptimal [FT89a]. subpermanents [Bap83]. subpolytopes [ST88a]. subsemilattices [Vei89b]. Subset [VV87c]. subsets [Bur86]. subspace [Eva87, HW88, MM82]. Subspaces [BDM83, Lee89, BGoMSTAUTA^{+}86, GR86, KS85, LR81, Lon84, Mat86, Ong82, Ong87, Özg86, RR84a, RR84b, RU85, She87a, Sou86a, Sou86b]. Substochastic [Fis87, Rob86, Rob87b]. substructured [Leu86]. subtangentiality [ST87a]. subtractivity [HS86a]. Successive [CM82, MN87, Kri84, NV84, SY89, VNC84]. Sufficient [BK81c, CPV89, JY86, Bec80, JY87, KG89, Kom83]. suitable [AP86]. sum [AT80, BQ85, Bru80, FW84, GHW82, HP86, Kha83, Mer84, Olk83, PW89, Rei87, ST81]. summability [MO86]. summation [BS83]. Sums [GR89a, AFW81, Avr89, BS81, Bre86, BM89c, ET81, Her83, KMV87, Mar81b, ORV88, PM80, PS88, Roh81, RS89]. superalgebra [Val88a]. superfast [AG89]. support [Ger84, SW85]. surface [Sch89b]. Surjectivity [Dem82]. surprising [Doo82]. survey [CMR81, Sch86a, Shr86, Vei82]. surveys [SS85]. SVD [CV87]. sweep [BN83]. sweep-plane [BN83]. Sweeping [KL86]. Sweeping-similarity [KL86]. Sylvester [Dan86, Lub87, MG85, Wim83b]. symbol [DR80a, Gor89, Har81b, Tho82a]. Symbolic [GR80]. Symmetric [AMT84, AGG88, BG84b, wEC89, DI86, GR80, Hal86b, Hig88, Pai80, Sim84, BH81a, BC84, Bha84, BC83, Bou88, BG82, BN88, Cao87a, CH89, CL82c, Cha88, Chu84, Con85, Dan88, Don87, ES80, ER87, EL88, Fan85, Fuh83, GT89a, GL84, HF85, HHJ86, JZ85, Kri84, Laf85b, Lás80, Lin88, Loe80, Mad83, MSS89, NV84, Par80, Pat85, Rad84, Ray83, Rei86, Rot89a, Sco85, Sha88, Sin81, Tad86, Tau83, VNC84, Wan83, Wat87a, Wat87b, Wea83, Wea88, Wei84, Wie85b, Won82a]. Symmetries [FR82]. symmetrizability [FJMN85]. symmetrizable [YJ80]. symmetrized [Gam88a, Uhl81b, dOdS83]. Symmetrizers [BK81a]. symmetrizing [Uhl81b]. Symmetry [MS88, DGKW84, Mer80b, MW81, Mer81b, ML83, Pat82, Tam87a]. Symplectic [Van84, BG86, LP82b]. Symposium [AL86a]. synthesis [Caf81, Lem88, Prz83]. System [HS83, ISS86, KÖ85b, Acz85, Fuj84, Kam88, Mis85, Ter87, Usm87]. System-theoretic [KÖ85b]. Systems [Axe80, Bar80a, BCR88, BA80, Cam84c, Cop84, DI86, JY83, KNC89, Nau87, Roh89, Ros84a, SS87a, Son87, dH87, Ant83, AS80, BCS83, BCS88, BAG89, BB89a, BNS86, BKH88, BFM89b, BFM89a, BF83, Bro83, BH89, Cao87b, Coh87, CS87a, Cop81, CS87c, DW89, Dem86a, Dio89, EHL81, Emr83, EL88, Fei89, FHV85, GL88, GKK86, GKKL87, GdHZ89, GR87, HPS83, HPS88, HH83, HS89, HPW87, HPW89, HW89, Jai84, Jef86, JKvdD87, JAK89, KW89a, KB89a, Kay85a, Kay85b, KP86, KP89, Kra81, LM88, LPC89, LVWC88, LN89, Mal89, Man82, MK89, Mur89, NW84, Neu81a, Neu80, NS89, Nic82, Nie89, OM89, Pan89b, Par80, PW83, Ptá83a, RV88, RW89, Roh81, Sch84b, Sch89b, Sch88, SSS82, SW84, Spi85, Spi87, Tch89, Tre86, Tre87]. systems [WN87, Wil83, Zhe87, Cam84e, Cam84c]. Systolic [Boj86, ES86, SHW86]. Szego [Avr89].

T [HM84b]. table [HTRG88]. tables [Joe85]. Tactical [CL82a, Hos86]. tame [Kle88]. Tampere [Bru88a]. Tannaka [Ban82]. Tannaka-Kreın [Ban82]. Tarmo [Bru88a]. Taussky [Fri81, Koh84, Tho83b]. Taylor [Mne89]. teacher [HM84b]. Technique [LQ87, MSS84]. Techniques [Hug85, Bar80b, Lip86, Neu84]. Tensor [Gro80b, Pol85, Bha84, Bru85, FR82, Gam88a, McF80]. tensorial [Lic85]. tensors [AL80, Bin86, Mer80b, MW81, Mer81b, Pat82, Poh89, Str83, Tam87a, Wei84, dOdS83]. Term [BP87, BR80b]. Term-rank [BP87]. terms [SF85]. Ternary [Koh84]. Testing [Väl88c, FP87b, Ram86, Sza85]. tests [Bon85, Jef88, MFPW85]. th [TST86, Yua86]. Thakare [BKR80]. Their [Bar87, BF87, Dat84, Eva85, Gof81, Hun82, IS89, Kle82, MR88a, MM88, RCS89, Rei86, SH88, Sta81, TC86, Wie82, Wim89a, YB81]. themselves [HJ86c, HM86]. Theorem [GHS87, Art83, Avr89, Bap84a, BR89a, BTD84, BS82a, Bix82, BR89b, BS85b, Bou88, Bro85a, Bro89b, BS82b, BS83, Byr83, CDMS81, CD83, CLP82, DD87a, Fri81, Gek81, HF85, Har84b, HH85, HCR88, Hoc84, Li88b, Mag87, Man81, Mar86a, MW87b, Que87a, Ros84b, Rys80, Sch82, Sco85, SS85, Sun82a, Tau85, Wie85b, Wil82a, Wil87, Wim88, Mer80a]. Theorems [BNV82, ES82, Sch84a, Bap82, BAG89, BC87, Car83, Dix81, Far87, Gur81a, Gur85b, HS84, Mad88, RV88, RC81, ZM81]. theoretic [DEHS84a, DEHS84b, Fer81, KÖ85b, Wim89c]. Theoretical [AGG88, Man82]. Theory [AL86a, Bar81b, Bar80a, BHL89c, Bru89, Cam84b, Cam84a, Cla88, GL85a, Gai89, Har88, Hol85, Nor84, Ple83, Tho83a, AT80, Ano80q, Ano81w, Ant83, BF85, Bap86a, Bar86, BF88, Ber86c, BB82a, BB82b, Cai80, Cai82, Coh86, CMO89, Dub85, Faĭ89, FRS82, GL88, Gro85, Har81a, Har84a, Hen83, Hil82, HPW87, JJS89, Laf85b, Lew85b, LA88, LA89, Mac89, MW84, Neu81a, Neu82b, Nie89, O’C85, Rys80, Sai87, SW85, Spi82, Szu89a, Tim85, Tor86b, Uch86, WH85a, Wim85a, Wol81, Zab88, BCS86, BHL89a, Bru84b, Cam84c, Son87]. third [Ano86v]. Three [Ber86b, Far87, FS85, Gał81, ET81, Gar82a, Ito82, Lip86, Liu88, SM88]. Three-dimensional [FS85]. three-variable [Gał81]. Three-way [Gał81]. Tight [Sch86c]. time [BH89, DGKW84, ER89, Fei89, HS89, Kam88, KB89a, KP86, Lin87b, Lin87c, LPC89, RV88, RT83, Wil83]. time-invariant [Fei89, Wil83]. time-varying [Kam88, KP86]. times [GR89a]. Tismenetsky [Har88]. Toeplitz [Bas85, GB84, GB87, Gov89b, AG89, Bar84b, BAS86, BC83, BK88b, BLAK89, BA80, BB86, Caf81, DGK83, Dic84, EP88, ES86, Fie84b, FH88a, Gor89, GB84, Gov89b, Gre83b, Grü81, GN82, HR88, HR89, Hol85, KK86, LQ87, Mee83, MM88, NW84, Par86, SSS82, Sha87, Tis87, Tre85b, Tre86, Tre87, dH87]. Toeplitz-plus-Hankel [HR89, HR88]. Topological [Web86]. Torgasev [Bru89]. Torus [FG86]. Tosio [Cam84a]. Total [VV87b, VV87c, May87, Sch81b, Tru80]. Totally [And87b, Stu88b, BJ84b, Dem82, Fri85b, Ger89, Pin85, dBP82, dBJP82]. trace [And85b, FJMN85, KB89b, Mer84, PZ88, SS86, Ste81a]. traces [MSW83, WS80a, WS80b]. Transfer [Bar86, Dio89, Fuh83, Fuh84, GVK^{+}83, WSCP89]. transform [Ger84, Par82, Tau80a, Ros80]. Transformation [Kau87, BHH87a, HS87, Kun81, Lin86, Tau83]. Transformations [Sch88, Bar83, Bar88b, Bea81a, Bea82, BHJ85, BFM89a, CL82c, EI87, Fia89, GLN84, GLN86a, GLN86b, HJ86c, HM86, KR81b, Loe89b, Loe89a, MK89, MW81, Pet82, PH81, RY85a, Syl86, Wat86, WH85c, You84, Poo81]. transforms [Gad88, Joh80, Vav88]. transitive [KB85]. Translates [DM83a]. translation [RSR89]. transmission [Caf81, KBM86]. transportation [BK80a]. transpose [Haa84, Tau80a]. treatment [Tam81b]. tree [JvdD88, JB85, JS89, Neu82a, SPZ82, Yua86]. Treediagonal [Kle82]. trees [Cve88, Wie85a, Yua89]. Triadic [PR88]. triangles [Tau82]. Triangular [ES86, Luk86, Bar89b, Bar83, BH88, BJ88a, Bar88b, DGK80, FH88b, HRS89, Ima83, SH88, Tre87, Urs82, Vei86]. triangularizable [CP85]. Triangularization [LQ87, Bar83, BJ88a, Bar88b, Laf80, Leu86, Wat80]. Tridiagonal [BCR88, Gad88, Ima83, Car84, CL86, CLLW87, DN84, GT89a, Grü81, GN82, JZ85, MS86]. Tridiagonalization [Stu88c, Lon88]. Tridiagonals [PNO85]. trigonometrical [ST81]. trinomial [GD84]. trireflectionality [Nie87]. Trivalent [Bee86]. Truemper [BR89b]. Tseng [BR89b]. turns [Djo82]. Tutte [Ger89, Wag89]. twice [Fer81]. Two [Bap87, BB89b, CM83, Els85, Fre87, MŽ83, SS87b, Ste87a, Wer89, Wim80, AS80, BQ85, Bec80, Bur86, CZ86, Cha85, DR80a, DALSP89, GL84, Jód88b, KH84, KN87a, KS84, Krä87, Laf80, Laf81a, McF80, Mes89, Mne89, Neu89, Nis85, OP82, Sha82c, Sha82a, She84, Tre85a, PN85]. two-distance [Neu89]. two-linked [DALSP89]. Two-minor [SS87b]. Two-sided [MŽ83]. two-symbol [DR80a]. twos [BM89c]. Type [Axe80, AN83, Ber86c, BNV82, DS89a, Faĭ89, IS89, Joh89, Kle88, Li88b, Oku87, Ptá81, Sch85, Wat87c, Yca82]. Typical [Lic85].

unbiased [Far85a]. Uncoupled [MSS84]. Uncoupling [Mey89]. underrelaxation [WN87]. undirected [Kar89]. unification [Mar81a]. unified [Bak85, NM85]. unified-least-squares [Bak85]. Uniform [MP89b, Hos86]. unimodular [Ger89]. unimodularity [Tru80]. Unions [LW81]. unipotent [HS87]. Unique [HK83, Whi80]. uniquely [Tho89a]. uniqueness [Hoc87, Meh89, SB87, SB88, Tho82b, Wim84]. Unit [BWT89, EG89, SZ86, SZ89, Wat88a, Zie88b]. Unitary [BH87c, BM89b, Djo81, FH88a, Jam83, Jam88]. unity [FSZ85, SSZ83]. Universidad [BV89, GL85a]. University [BU84, BU88, OvdD88, Neu82b, dO83]. Unmixed [Wim85a]. Unobservable [She87a]. unsolvability [Sha83]. Unstable [LQ87]. Unsymmetric [Axe80, Saa80, FG80]. unzulässige [PN85]. updates [EN89]. Updating [PNO85, Pow87]. upon [CW80, Naz84]. Upper [Bar83, BJ88a, Bar88b, Bea84, RT85, DEHS84a, DEHS84b, Ima83, Leó88]. Use [PNO85, CMR81, Don85, MV88a, Won85]. useful [RY85b]. User [Cam81]. Using [AR87, SHW86, VV87c, Bar84a, Bar80b, DM83b, Fos86, FB81a, GH80, HRS88, LM88, MSW83, MW85a, WS80a, WS80b]. Usmani [Gov88a]. USSOR [MK88]. Utah [BU84, BU88, BU84, BU88].

V [BKR80]. Valencia [BV89, NdOH89]. valency [Big80, Ito82]. valuation [Piz83]. Value [Bai88, Chu86b, Pai85, PV86, wEC87a, ER87, For80, Gek81, GP89, Jód86, Jód87, Jód88b, Joh89, Kit89, Lew89c, MS86, MC81, You83]. valued [Tho85]. Values [Luk86, Qi84, Ste84, BH87a, Brü84d, DR83, DR88, GR89b, Lew89b, LT83b, LT88b, Par86, Que87b, Sod87, Uhl83, vV87a]. Vandermonde [Dem89, FD89, Gau83a, HG80, SA89]. Vandermonde-like [Gau83a]. vanishing [HG87, Wol89b, Wol89a]. Varga [HN88]. variable [Gał81]. Variables [VV87c, AS80, CM85, CM81b, Kam89, Lip86, Wol89a]. variant [Sch82]. Variation [BF81b, Bha84, Els82c, Els85, Beb86, BH87c, Els88, Jia82]. Variations [Saa80, Har84a]. various [Joh87]. varying [Kam88, KP86]. vec [Jin88]. vech [Jin88]. Vector [DF87, FM86b, AP86, BS82a, BH81b, Bre86, Bru84a, DN85, Eva87, GS81, KN87a, Nar86, Pan89a, San87]. Vectors [Ruh83, BWT89, Bru80, Fuj84, OP82, Wim89d]. verification [Bru85]. Vermutung [PN85]. versions [ABI84, Sak85]. Verstärkungen [PN85]. versus [BP88b, BLAK89, GL87]. vertauschbare [HH85]. vertices [MP89a]. via [CG89, Dax88, Fur89, Joh80, KBM86, LN89, Par82, RS89, Tor86b, Zhe87]. vibrating [LM88]. Victoria [Ano88-32, OvdD88]. view [Sch83b]. visibility [BS85b]. Vitek [Zha87a]. Vleck [Nau87]. VLSI [HTRG88]. Vol [Ano88a, BS89]. Volume [Ano82b, BS89, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano82c, Ano82d, Ano82e, Ano82f, Ano82g, Ano82h, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano83a, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano84g, Ano84h, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano85i, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano86m, Ano86a, Ano87j, Ano87a, Ano87k, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano87i, Ano87s, Ano88m, Ano88n, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano88h, Ano88i, Ano88j, Ano88k, Ano88l, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e]. volume [Ano89f, Ano89g, Ano89h, Ano89i, BN83, BB82b, BF83, DHLW85, GSV86b]. volumes [Ano81c, Ano84i, Ano86n, Ano88-28, Ano89u, Ano89j].

W [AB85, Cam81, DHLW85, Gov89a, Nau87, Ple81, Pul87, Tho83a]. Waerden [Lon82, McD80, vL81]. Walk [GM88a, GM80, PS82]. walk-regular [GM80]. walks [Tak82]. Walsh [SSZ83]. Walter [GL85b]. way [Gał81, Tak85]. Weak [Bun87, Fri85b, Ish87, SK82, Wer87]. weakly [Wag89, WN87]. weight [KL88a]. Weighted [LST88, Sis86, Hur89]. weights [Hes86, LST88]. Weinstein [BF88]. well [Lee89, Ogr88]. well-scaled [Lee89]. Wendroff [FT84]. Wheatstone [AB87]. which [Bea81a, BS87b, Eva85, Gof81, HM86, Hug85, Loe89a, MK89, RS89, Sch84a, Tau80a, Wea83, Wea88, Wie85c]. whose [Ash87, Har81b, Laf80, Tow80, Wie85a, Wie85b]. Wiegold [Bea87a]. Wielandt [BS85b, DHL85, DHLW85, FN89, Sak85, TW89]. Wiener [GG82, GLR83]. wild [HPW87]. Wilkinson [PN85, Ano87s, PN85, She87b, Zha89]. Winograd [Par82]. Wirtinger [Che87]. Wishart [Neu85]. without [Fos86, FP81, Gee89a, Gee89b, GN82]. work [Bea87a, GL85b]. Wronskian [Wol89b, Har82]. Wronskians [Wol89a]. Wronsky [GA85].

XAY [Sil87b].

Yang [Faĭ89]. year [Ano88a, BS89]. Young [GHS87].

Z [Pul87, She86, MS80b]. Z-Q [Pul87]. Zero [PY82, BJ86, BS81, EJ89, Gam88a, KG89, McD80, SRPM82, VV86, Yca82]. zeros [Bru80, BL82b, BS87c, BM89c, EP87b, Har81b, Kam88, Lás82, Lon82, WS83, WSCP89]. ZME [SS88]. ZME-matrices [SS88]. Zwei [PN85]. Zygotic [DALSP89].

References

Antoulas:1989:PSR

[AA89] A. C. Antoulas and B. D. O. Anderson. On the problem of stable rational interpolation. Linear Algebra and its Applications, 122/123/124(??):301–329, 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alikakos:1984:EEJ

[AB84] N. Alikakos and P. W. Bates. Estimates for the eigenvalues of the Jordan product of Hermitian matrices. Linear Algebra and its Applications, 57(??):41–56, February 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901757. See er-

ratum [AB85].

Alikakos:1985:EEE

[AB85] Nicholas D. Alikakos and Peter W. Bates. Erratum: Estimates for the eigenvalues of the Jordan product of Hermitian matrices: Nicholas D. Alikakos and Peter W. Bates, Linear Algebra Appl. 57:41-56 (1984). Linear Algebra and its Applications, 65(??):282, February 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585901041. See [AB84].

Ando:1987:GMO

[AB87] T. Ando and John W. Bunce. The geometric mean, operator inequalities, and the Wheatstone bridge. Linear Algebra and its Applications, 97(??):77–91, December 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587901418.

Abel:1982:SIP

[Abe82] Ulrich Abel. On the stabilization of iterative processes. Linear Algebra and its Applications, 43(??):1–15, March 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582902415.

Axelsson:1984:SVI

[ABI84] O. Axelsson, S. Brinkkemper, and V. P. Ilín. On some versions of incomplete block-matrix factorization iterative methods. Linear Algebra and its Applications, 58 (??):3–15, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584902003.

Ahac:1988:SFM

[ABO88] Alan A. Ahac, John J. Buoni, and D. D. Olesky. Stable LU factorization of H-matrices. Linear Algebra and its Applications, 99(??):97–110, February 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901279.

Aharoni:1989:BIP

[AC89] Ron Aharoni and Yair Censor. Block-iterative projection methods for parallel computation of solutions to convex feasibility problems. Linear Algebra and its Applications, 120(??):165–175, August 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589903753. Proceedings

of the Fourth Haifa Matrix Theory Conference (Haifa, 1988).

Arsene:1988:SAS

[ACC88] Gr. Arsene, Zoia Ceauşescu, and T. Constantinescu. Schur analysis of some completion problems. Linear Algebra and its Applications, 109(??):1–35, October 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901954.

Alon:1989:ADC

[ACHS89] Noga Alon, Steven Cosares, Dorit S. Hochbaum, and Ron Shamir. An algorithm for the detection and construction of Monge sequences. Linear Algebra and its Applications, 114/115(??):669–680, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Aczel:1985:SSF

[Acz85] J. Aczél. Solution of a system of functional-differential equations arising from an optimization problem. Linear Algebra and its Applications, 71(??):9–13, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902290.

Adin:1982:EPO

[Adi82]  Ron Adin.  Extreme positive operators on min-

imal and almost minimal cones. Linear Algebra and its Applications, 44(??):61–86, April 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582900052.

Adkins:1985:PLG

[Adk85] William A. Adkins. The pointwise-local-global principle for solutions of generic linear equations. Linear Algebra and its Applications, 66(??):29–43, April 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585901247.

Adkins:1989:SDM

[Adk89] William A. Adkins. Simultaneous diagonalization of matrices parametrized by a projective algebraic curve. Linear Algebra and its Applications, 116(??):101–108, April 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958990400X.

Auslander:1981:DSB

[AFW81] L. Auslander, E. Feig, and S. Winograd. Direct sums of bilinear algorithms. Linear Algebra and its Applications, 38(??):175–192, June 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581900197.

Ammar:1989:NES

[AG89] Gregory S. Ammar and William B. Gragg. Numerical experience with a superfast real Toeplitz solver. Linear Algebra and its Applications, 121(??):185–206, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589907015. Linear

algebra and applications (Valencia, 1987).

Aganagic:1981:DDL

[Aga81] Muhamed Aganagic. On diagonal dominance in linear complementarity. Linear Algebra and its Applications, 39 (??):41–49, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902883.

Arbenz:1988:RRM

[AGG88] Peter Arbenz, Walter Gander, and Gene H. Golub. Restricted rank modification of the symmetric eigenvalue problem: Theoretical considerations. Linear Algebra and its Applications, 104(??):75–95, June 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588903096.

Agler:1988:PSM

[AHnLR88] Jim Agler, J. William Helton, and Scott McCullough nd Leiba Rodman. Positive semidefinite matrices with a given sparsity pattern. Linear Algebra and its Applications, 107(??):101–149, August 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902406. Proceedings

of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Abhyankar:1989:HI

[AK89] Shreeram Abhyankar and Devadatta M. Kulkarni. On Hilbertian ideals. Linear Algebra and its Applications, 116 (??):53–79, April 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589903984.

Atkinson:1980:BRS

[AL80] M. D. Atkinson and S. Lloyd. Bounds on the ranks of some 3-tensors. Linear Algebra and its Applications, 31 (??):19–31, June 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580902025.

Amann:1986:PSO

[AL86a] H. (Herbert) Amann and Peter Lancaster, editors. Proceedings of the Symposium on Operator Theory: held in Athens, Greece, August 1985, volume 84 of Linear Algebra and its Applications. North-Holland Publishing Co., Amsterdam, The Netherlands, 1986. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arioli:1986:EAA

[AL86b] M. Arioli and A. Laratta. Error analysis of algorithms for computing the projection of a point onto a linear manifold. Linear Algebra and its Applications, 82(??): 1–26, October 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901412.

Albuquerque:1989:GSL

[AL89] H. Albuquerque and F. Silva Leite. On the generators of semisimple Lie algebras. Linear Algebra and its Applications, 119(??):51–56, July 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589900682.

Aupetit:1981:SCP

[ALZ81] Bernard Aupetit, Thomas J. Laffey, and Jaroslav Zemánek. Spectral classification of projections. Linear Algebra and its Applications, 41(??):131–135, December 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900938.

Amann:1986:SNE

[Ama86] Herbert Amann. Semigroups and nonlinear evolution equations. Linear Algebra and its Applications, 84(??): 3–32, December 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586903058.

Proceedings of the symposium on operator theory (Athens, 1985).

Anderson:1984:SFM

[AMT84] W. N. Anderson, Jr., T. D. Morley, and G. E. Trapp. Symmetric function means of positive operators. Linear Algebra and its Applications, 60(??):129–143, August 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584900752.

Aczel:1983:DAS

[AN83] J. Aczél and C. T. Ng. Determination of all semisymmetric recursive information measures of multiplicative type on n positive discrete probability distributions. Linear Algebra and its Applications, 52/53(??):1–30, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ando:1983:AGH

[And83] T. Ando. On the arithmetic-geometric-harmonic-mean inequalities for positive definite matrices. Linear Algebra and its Applications, 52/53(??):31–37, 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anderson:1985:DIC

[And85a] Ian M. Anderson. Determinantal ideals and the Capelli identities. Linear Algebra and its Applications, 69(??): 269–277, August 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585900813.

Ando:1985:ATN

[And85b] T. Ando. Approximation in trace norm by positive semidefinite matrices. Linear Algebra and its Applications, 71(??):15–21, November 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902307.

Ando:1987:RBA

[And87a] T. Ando. Review of Matrix Analysis, by Roger A. Horn and Charles R. Johnson. Linear Algebra and its Applications, 90(??):245–247, May 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587903181.

Ando:1987:TPM

[And87b] T. Ando. Totally positive matrices. Linear Algebra and its Applications, 90(??):165–219, May 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587903132.

Ando:1989:MDS

[And89] T. Ando. Majorization, doubly stochastic matrices, and comparison of eigenvalues. Linear Algebra and its Applications, 118(??):163–248, June 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589905806.

Anonymous:1980:AHA

[Ano80a] Anonymous. Alston S. Householder award IV (1981). Linear Algebra and its Applications, 32(??):1, August 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580900026.

Anonymous:1980:AI

[Ano80b] Anonymous. Author index. Linear Algebra and its Applications, 34(??):401, December 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901743.

Anonymous:1980:AIVa

[Ano80c] Anonymous. Author index to volume 29. Linear Algebra and its Applications, 29(??):531–532, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580902608.

Anonymous:1980:AIVb

[Ano80d] Anonymous. Author index to volume 30. Linear Algebra and its Applications, 30(??):251, April 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901998.

Anonymous:1980:AIVc

[Ano80e] Anonymous. Author index to volume 31. Linear Algebra and its Applications, 31(??):249, June 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580902244.

Anonymous:1980:AIVd

[Ano80f] Anonymous. Author index to volume 32. Linear Algebra and its Applications, 32(??):233, August 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580900191.

Anonymous:1980:AIVe

[Ano80g] Anonymous. Author index to volume 33. Linear Algebra and its Applications, 33(??):233, October 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580901068.

Anonymous:1980:EBa

[Ano80h] Anonymous. Editorial Board. Linear Algebra and its Applications, 29(??):??, February 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580902256.

Anonymous:1980:EBb

[Ano80i] Anonymous. Editorial Board. Linear Algebra and its Applications, 30(??):??, April 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379580901767.

Anonymous:1980:EBc

[Ano80j] Anonymous. Editorial Board. Linear Algebra and its Applications, 31(??):??, June 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379580902001.

Anonymous:1980:EBd

[Ano80k] Anonymous. Editorial board. Linear Algebra and its Applications, 32(??):??, August 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379580900014.

Anonymous:1980:EBe

[Ano80l] Anonymous. Editorial Board. Linear Algebra and its Applications, 33(??):??, October 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379580900919.

Anonymous:1980:EBf

[Ano80m] Anonymous. Editorial Board. Linear Algebra and its Applications, 34(??):??, December 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901548.

Anonymous:1980:LLAa

[Ano80n] Anonymous. Letters in linear algebra. Linear Algebra and its Applications, 31(??):213–214, June 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580902207.

Anonymous:1980:LLAb

[Ano80o] Anonymous. Letters in linear algebra. Linear Algebra and its Applications, 32(??):199–200, August 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580900166.

Anonymous:1980:LLAc

[Ano80p] Anonymous. Letters in linear algebra. Linear Algebra and its Applications, 33(??):153, October 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901032.

Anonymous:1980:SIL

[Ano80q] Anonymous. Special issue linear algebraic methods in combinatorics/graph theory. Linear Algebra and its Applications, 33(??):iv, October 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580900920.

Anonymous:1980:SI

[Ano80r] Anonymous. Subject index. Linear Algebra and its Applications, 34(??):403–404, December 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901755.

Anonymous:1981:A

[Ano81a] Anonymous. Announcement. Linear Algebra and its Applications, 35(??):v, February 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379581902603.

Anonymous:1981:AI

[Ano81b] Anonymous. Author index. Linear Algebra and its Applications, 35(??):298, February 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581902834.

Anonymous:1981:AIVf

[Ano81c] Anonymous. Author index for volumes 21–40 (1978– 1981). Linear Algebra and its Applications, 40(??):283– 296, October 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581901609.

Anonymous:1981:AIVa

[Ano81d] Anonymous. Author index to volume 36. Linear Algebra and its Applications, 36(??):297, March 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581902330.

Anonymous:1981:AIVb

[Ano81e] Anonymous. Author index to volume 37. Linear Algebra and its Applications, 37(??):297, April 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581901890.

Anonymous:1981:AIVc

[Ano81f] Anonymous. Author index to volume 38, 1981. Linear Algebra and its Applications, 38(??):295, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900288.

Anonymous:1981:AIVd

[Ano81g] Anonymous. Author index to volume 39, 1981. Linear Algebra and its Applications, 39(??):297, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958190313X.

Anonymous:1981:AIVe

[Ano81h] Anonymous. Author index to volume 40. Linear Algebra and its Applications, 40(??):279, October 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581901592.

Anonymous:1981:AIVg

[Ano81i] Anonymous. Author index to volume 41. Linear Algebra and its Applications, 41(??):299, December 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581901051.

Anonymous:1981:EBa

[Ano81j] Anonymous. Editorial Board. Linear Algebra and its Applications, 35(??):??, February 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581902597.

Anonymous:1981:EBb

[Ano81k] Anonymous. Editorial Board. Linear Algebra and its Applications, 36(??):??, March 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379581902147.

Anonymous:1981:EBc

[Ano81l] Anonymous. Editorial Board. Linear Algebra and its Applications, 37(??):??, April 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379581901610.

Anonymous:1981:EBd

[Ano81m] Anonymous. Editorial Board. Linear Algebra and its Applications, 39(??):??, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379581902846.

Anonymous:1981:EBf

[Ano81n] Anonymous. Editorial board. Linear Algebra and its Applications, 40(??):247–248, October 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581901555.

Anonymous:1981:EBe

[Ano81o] Anonymous. Editorial board. Linear Algebra and its Applications, 40(??):??, October 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
002437958190135X.

Anonymous:1981:LLAa

[Ano81p] Anonymous. Letters in linear algebra. Linear Algebra and its Applications, 35(??):259–260, February 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581902779.

Anonymous:1981:LLAb

[Ano81q] Anonymous. Letters in linear algebra. Linear Algebra and its Applications, 36(??):185, March 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581902317.

Anonymous:1981:LLAc

[Ano81r] Anonymous. Letters in linear algebra. Linear Algebra and its Applications, 38(??):251–252, June 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581900240.

Anonymous:1981:LLAd

[Ano81s] Anonymous. Letters in linear algebra. Linear Algebra and its Applications, 39(??):261–262, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581903098.

Anonymous:1981:LLAe

[Ano81t] Anonymous. Letters in linear algebra. Linear Algebra and its Applications, 41(??):273–274, December 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581901026.

Anonymous:1981:LAAa

[Ano81u] Anonymous. Linear Algebra and Its Applications: Referees for 1980. Linear Algebra and its Applications, 35 (??):297, February 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902822.

Anonymous:1981:LAAb

[Ano81v] Anonymous. Linear Algebra and Its Applications: Referees for 1981. Linear Algebra and its Applications, 41(??): 300, December 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581901063.

Anonymous:1981:LCT

[Ano81w] Anonymous. Linear control theory. Linear Algebra and its Applications, 38(??):1, June 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958190001X.

Anonymous:1982:A

[Ano82a] Anonymous. Announcement. Linear Algebra and its Applications, 42(??):1–2, February 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582901331.

Anonymous:1982:AIVa

[Ano82b] Anonymous. Author index to volume 42. Linear Algebra and its Applications, 42(??):299, February 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582901586.

Anonymous:1982:AIVb

[Ano82c] Anonymous. Author index to volume 43. Linear Algebra and its Applications, 43(??):297, March 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582902610.

Anonymous:1982:AIVc

[Ano82d] Anonymous. Author index to volume 44. Linear Algebra and its Applications, 44(??):203, April 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582900180.

Anonymous:1982:AIVd

[Ano82e] Anonymous. Author index to volume 45. Linear Algebra and its Applications, 45(??):199, June 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582902191.

Anonymous:1982:AIVe

[Ano82f] Anonymous. Author index to volume 46. Linear Algebra and its Applications, 46(??):243, August 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582900386.

Anonymous:1982:AIVf

[Ano82g] Anonymous. Author index to volume 47. Linear Algebra and its Applications, 47(??):217, October 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582902397.

Anonymous:1982:AIVg

[Ano82h] Anonymous. Author index to volume 48. Linear Algebra and its Applications, 48(??):485–486, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582901306.

Anonymous:1982:EBa

[Ano82i] Anonymous. Editorial Board. Linear Algebra and its Applications, 42(??):??, February 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958290132X.

Anonymous:1982:EBb

[Ano82j] Anonymous. Editorial Board. Linear Algebra and its Applications, 43(??):??, March 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379582902403.

Anonymous:1982:EBc

[Ano82k] Anonymous. Editorial Board. Linear Algebra and its Applications, 45(??):??, June 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
002437958290204X.

Anonymous:1982:EBd

[Ano82l] Anonymous. Editorial Board. Linear Algebra and its Applications, 46(??):??, August 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379582900192.

Anonymous:1982:EBe

[Ano82m] Anonymous. Editorial Board. Linear Algebra and its Applications, 47(??):??, October 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379582902208.

Anonymous:1982:LAA

[Ano82n] Anonymous. Linear algebra and its applications. Linear Algebra and its Applications, 48(??):487, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582901318.

Anonymous:1982:Ra

[Ano82o] Anonymous. Reports. Linear Algebra and its Applications, 42(??):261, February 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379582901549.

Anonymous:1982:Rb

[Ano82p] Anonymous. Reports. Linear Algebra and its Applications, 44(??):187, April 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379582900155.

Anonymous:1983:AIVf

[Ano83a] Anonymous. Author index for volume 110. Linear Algebra and its Applications, 110(??):319, November 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958390143X.

Anonymous:1983:AIVa

[Ano83b] Anonymous. Author index to volume 49. Linear Algebra and its Applications, 49(??):293, February 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583901118.

Anonymous:1983:AIVb

[Ano83c] Anonymous. Author index to volume 50. Linear Algebra and its Applications, 50(??):657, April 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583900745.

Anonymous:1983:AIVc

[Ano83d] Anonymous. Author index to volume 51. Linear Algebra and its Applications, 51(??):183, June 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958390160X.

Anonymous:1983:AIVd

[Ano83e] Anonymous. Author index to volume 54. Linear Algebra and its Applications, 54(??):183, October 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958390215X.

Anonymous:1983:AIVe

[Ano83f] Anonymous. Author index to volume 55. Linear Algebra and its Applications, 55(??):181, December 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583901763.

Anonymous:1983:EBa

[Ano83g] Anonymous. Editorial Board. Linear Algebra and its Applications, 49(??):ii–iii, February 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583900903.

Anonymous:1983:EBb

[Ano83h] Anonymous. Editorial Board. Linear Algebra and its Applications, 51(??):ii–iii, June 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379583901441.

Anonymous:1983:EBc

[Ano83i] Anonymous. Editorial Board. Linear Algebra and its Applications, 54(??):ii–iii, October 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958390201X.

Anonymous:1983:EBd

[Ano83j] Anonymous. Editorial Board. Linear Algebra and its Applications, 55(??):ii–iii, December 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583901611.

Anonymous:1983:EBe

[Ano83k] Anonymous. Editorial Board. Linear Algebra and its Applications, 110(??):ii–iii, November 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583901283.

Anonymous:1983:LAA

[Ano83l] Anonymous. Linear Algebra and Its Applications: Referees for 1983. Linear Algebra and its Applications, 55(??): 183, December 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901775.

Anonymous:1984:AIVa

[Ano84a] Anonymous. Author index to volume 56. Linear Algebra and its Applications, 56(??):295, January 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901344.

Anonymous:1984:AIVb

[Ano84b] Anonymous. Author index to volume 57. Linear Algebra and its Applications, 57(??):297, February 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901976.

Anonymous:1984:AIVc

[Ano84c] Anonymous. Author index to volume 58. Linear Algebra and its Applications, 58(??):461, April 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958490226X.

Anonymous:1984:AIVd

[Ano84d] Anonymous. Author index to volume 59. Linear Algebra and its Applications, 59(??):215, June 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901691.

Anonymous:1984:AIVe

[Ano84e] Anonymous. Author index to volume 60. Linear Algebra and its Applications, 60(??):201, August 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900818.

Anonymous:1984:AIVg

[Ano84f] Anonymous. Author index to volume 61. Linear Algebra and its Applications, 61(??):295, September 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584900417.

Anonymous:1984:AIVh

[Ano84g] Anonymous. Author index to volume 62. Linear Algebra and its Applications, 62(??):295, November 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901071.

Anonymous:1984:AIVi

[Ano84h] Anonymous. Author index to volume 63. Linear Algebra and its Applications, 63(??):293, December 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901514.

Anonymous:1984:AIVf

[Ano84i] Anonymous. Author index to volumes 41-60 (1981-1984). Linear Algebra and its Applications, 60(??):203– 216, August 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958490082X.

Anonymous:1984:EBa

[Ano84j] Anonymous. Editorial Board. Linear Algebra and its Applications, 56(??):ii–iii, January 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901083.

Anonymous:1984:EBb

[Ano84k] Anonymous. Editorial Board. Linear Algebra and its Applications, 57(??):ii–iii, February 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901708.

Anonymous:1984:EBc

[Ano84l] Anonymous. Editorial Board. Linear Algebra and its Applications, 58(??):ii–iii, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379584901988.

Anonymous:1984:EBd

[Ano84m] Anonymous. Editorial Board. Linear Algebra and its Applications, 59(??):i–ii, June 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379584901538.

Anonymous:1984:EBe

[Ano84n] Anonymous. Editorial Board. Linear Algebra and its Applications, 60(??):ii–iii, August 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900661.

Anonymous:1984:EBf

[Ano84o] Anonymous. Editorial Board. Linear Algebra and its Applications, 62(??):ii–iii, November 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900831.

Anonymous:1984:EBg

[Ano84p] Anonymous. Editorial Board. Linear Algebra and its Applications, 63(??):ii–iii, December 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901356.

Anonymous:1984:LAA

[Ano84q] Anonymous. Linear Algebra and Its Applications: Referees for 1984. Linear Algebra and its Applications, 63(??): 295, December 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901526.

Anonymous:1985:AIVa

[Ano85a] Anonymous. Author index to volume 64. Linear Algebra and its Applications, 64(??):279, January 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902848.

Anonymous:1985:AIVb

[Ano85b] Anonymous. Author index to volume 65. Linear Algebra and its Applications, 65(??):281, February 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958590103X.

Anonymous:1985:AIVc

[Ano85c] Anonymous. Author index to volume 66. Linear Algebra and its Applications, 66(??):279, April 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585901387.

Anonymous:1985:AIVd

[Ano85d] Anonymous. Author index to volume 67. Linear Algebra and its Applications, 67(??):279, June 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902022.

Anonymous:1985:AIVe

[Ano85e] Anonymous. Author index to volume 68. Linear Algebra and its Applications, 68(??):279, July 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902265.

Anonymous:1985:AIVf

[Ano85f] Anonymous. Author index to volume 69. Linear Algebra and its Applications, 69(??):279, August 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900825.

Anonymous:1985:AIVg

[Ano85g] Anonymous. Author index to volume 70. Linear Algebra and its Applications, 70(??):369, October 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585900655.

Anonymous:1985:AIVh

[Ano85h] Anonymous. Author index to volume 71. Linear Algebra and its Applications, 71(??):341–342, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902617.

Anonymous:1985:AIVi

[Ano85i] Anonymous. Author index to volume 72. Linear Algebra and its Applications, 72(??):183, December 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585901545.

Anonymous:1985:C

[Ano85j] Anonymous. Comment. Linear Algebra and its Applications, 70(??):301, October 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379585900618.

Anonymous:1985:EBa

[Ano85k] Anonymous. Editorial Board. Linear Algebra and its Applications, 64(??):ii–iii, January 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902629.

Anonymous:1985:EBb

[Ano85l] Anonymous. Editorial Board. Linear Algebra and its Applications, 65(??):ii–iii, February 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900837.

Anonymous:1985:EBc

[Ano85m] Anonymous. Editorial Board. Linear Algebra and its Applications, 66(??):ii–iii, April 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379585901223.

Anonymous:1985:EBd

[Ano85n] Anonymous. Editorial Board. Linear Algebra and its Applications, 67(??):ii–iii, June 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
002437958590179X.

Anonymous:1985:EBe

[Ano85o] Anonymous. Editorial Board. Linear Algebra and its Applications, 68(??):ii–iii, July 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379585902034.

Anonymous:1985:EBf

[Ano85p] Anonymous. Editorial Board. Linear Algebra and its Applications, 70(??):ii–iii, October 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900370.

Anonymous:1985:EBg

[Ano85q] Anonymous. Editorial Board. Linear Algebra and its Applications, 71(??):ii–iii, November 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902277.

Anonymous:1985:EBh

[Ano85r] Anonymous. Editorial Board. Linear Algebra and its Applications, 72(??):i–iii, December 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585901399.

Anonymous:1985:LCO

[Ano85s] Anonymous. Lothar Collatz — on the occasion of his 75th birthday. Linear Algebra and its Applications, 68 (??):??, July 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585902046.

Anonymous:1985:SIS

[Ano85t] Anonymous, editor. Special issue on statistics, volume 70 of Linear Algebra and its Applications. Elsevier Science Publishers, Amsterdam, The Netherlands, 1985. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:1986:AVS

[Ano86a] Anonymous. Addendum to volume 70: the second volume of the special issue on statistics. Linear Algebra and its Applications, 75(??):277–279, March 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586901953.

Anonymous:1986:AIVa

[Ano86b] Anonymous. Author index to volume 73. Linear Algebra and its Applications, 73(??):281, January 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902430.

Anonymous:1986:AIVb

[Ano86c] Anonymous. Author index to volume 74. Linear Algebra and its Applications, 74(??):281, February 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586901291.

Anonymous:1986:AIVc

[Ano86d] Anonymous. Author index to volume 75. Linear Algebra and its Applications, 75(??):281, March 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586901965.

Anonymous:1986:AIVd

[Ano86e] Anonymous. Author index to volume 76. Linear Algebra and its Applications, 76(??):281, April 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902302.

Anonymous:1986:AIVe

[Ano86f] Anonymous. Author index to volume 77. Linear Algebra and its Applications, 77(??):357, May 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586901771.

Anonymous:1986:AIVf

[Ano86g] Anonymous. Author index to volume 78. Linear Algebra and its Applications, 78(??):249, June 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586900285.

Anonymous:1986:AIVg

[Ano86h] Anonymous. Author index to volume 79. Linear Algebra and its Applications, 79(??):249, July 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586903022.

Anonymous:1986:AIVh

[Ano86i] Anonymous. Author index to volume 80. Linear Algebra and its Applications, 80(??):261, August 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902855.

Anonymous:1986:AIVj

[Ano86j] Anonymous. Author index to volume 81. Linear Algebra and its Applications, 81(??):281, September 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902600.

Anonymous:1986:AIVk

[Ano86k] Anonymous. Author index to volume 82. Linear Algebra and its Applications, 82(??):281, October 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586901576.

Anonymous:1986:AIVl

[Ano86l] Anonymous. Author index to volume 83. Linear Algebra and its Applications, 83(??):167, November 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902739.

Anonymous:1986:AIVm

[Ano86m] Anonymous. Author index to volume 84. Linear Algebra and its Applications, 84(??):421, December 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586903290.

Anonymous:1986:AIVi

[Ano86n] Anonymous. Author index to volumes 61-80 (1984-1986). Linear Algebra and its Applications, 80(??):263– 275, August 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586902867.

Anonymous:1986:EBa

[Ano86o] Anonymous. Editorial Board. Linear Algebra and its Applications, 76(??):ii–iii, April 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379586902090.

Anonymous:1986:EBb

[Ano86p] Anonymous. Editorial Board. Linear Algebra and its Applications, 79(??):ii–iii, July 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379586902879.

Anonymous:1986:EBc

[Ano86q] Anonymous. Editorial Board. Linear Algebra and its Applications, 80(??):ii–iii, August 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902740.

Anonymous:1986:EBd

[Ano86r] Anonymous. Editorial Board. Linear Algebra and its Applications, 81(??):ii–iii, September 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902442.

Anonymous:1986:EBe

[Ano86s] Anonymous. Editorial board. Linear Algebra and its Applications, 82(??):ii–iii, October 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586901400.

Anonymous:1986:EBf

[Ano86t] Anonymous. Editorial Board. Linear Algebra and its Applications, 83(??):ii–iii, November 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902612.

Anonymous:1986:EBg

[Ano86u] Anonymous. Editorial Board. Linear Algebra and its Applications, 84(??):ii–iii, December 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586903034.

Anonymous:1986:PTS

[Ano86v] Anonymous. Preface to the third special issue of papers on statistics. Linear Algebra and its Applications, 82 (??):143, October 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901473.

Anonymous:1987:AIVb

[Ano87a] Anonymous. Author index for volume 86. Linear Algebra and its Applications, 86(??):283, February 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587902989.

Anonymous:1987:AIVd

[Ano87b] Anonymous. Author index for volume 90. Linear Algebra and its Applications, 90(??):249, May 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587903193.

Anonymous:1987:AIVe

[Ano87c] Anonymous. Author index for volume 91. Linear Algebra and its Applications, 91(??):217, June 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587900735.

Anonymous:1987:AIVf

[Ano87d] Anonymous. Author index for volume 92. Linear Algebra and its Applications, 92(??):273, July 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587902643.

Anonymous:1987:AIVg

[Ano87e] Anonymous. Author index for volume 93. Linear Algebra and its Applications, 93(??):281, August 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/S0024379587903302.

Anonymous:1987:AIVh

[Ano87f] Anonymous. Author index for volume 94. Linear Algebra and its Applications, 94(??):273, September 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587900966.

Anonymous:1987:AIVi

[Ano87g] Anonymous. Author index for volume 95. Linear Algebra and its Applications, 95(??):217, October 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587900371.

Anonymous:1987:AIVj

[Ano87h] Anonymous. Author index for volume 96. Linear Algebra and its Applications, 96(??):249, November 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958790348X.

Anonymous:1987:AIVk

[Ano87i] Anonymous. Author index for volume 97. Linear Algebra and its Applications, 97(??):249, December 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587901546.

Anonymous:1987:AIVa

[Ano87j] Anonymous. Author index to volume 85. Linear Algebra and its Applications, 85(??):281, January 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587902230.

Anonymous:1987:AIVc

[Ano87k] Anonymous. Author index to volume 87. Linear Algebra and its Applications, 87(??):279, March 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587901753.

Anonymous:1987:EBa

[Ano87l] Anonymous. Editorial Board. Linear Algebra and its Applications, 86(??):ii–iii, February 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587902837.

Anonymous:1987:EBb

[Ano87m] Anonymous. Editorial Board. Linear Algebra and its Applications, 87(??):ii–iii, March 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379587901558.

Anonymous:1987:EBc

[Ano87n] Anonymous. Editorial Board. Linear Algebra and its Applications, 90(??):ii–iii, May 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379587902990.

Anonymous:1987:EBd

[Ano87o] Anonymous. Editorial Board. Linear Algebra and its Applications, 93(??):ii–iii, August 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/S0024379587903077.

Anonymous:1987:EBe

[Ano87p] Anonymous. Editorial Board. Linear Algebra and its Applications, 94(??):??, September 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587900747.

Anonymous:1987:EBf

[Ano87q] Anonymous. Editorial Board. Linear Algebra and its Applications, 95(??):ii–iii, October 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958790022X.

Anonymous:1987:EBg

[Ano87r] Anonymous. Editorial Board. Linear Algebra and its Applications, 96(??):ii–iii, November 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587903314.

Anonymous:1987:SVM

[Ano87s] Anonymous, editor. Special volume in memory of James H. Wilkinson, volume 88/89 of Linear Algebra and its Applications. Elsevier, Amsterdam, The Netherlands, 1987. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:1988:YIV

[Ano88a] Anonymous. 1968/1988. 20 year index: Vol. 1–100. Linear Algebra and its Applications, 100(??):x + 86, 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:1988:AI

[Ano88b] Anonymous. Author index. Linear Algebra and its Applications, 100(??):17–86, March 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588903862.

Anonymous:1988:AIVc

[Ano88c] Anonymous. Author index for volume 101. Linear Algebra and its Applications, 101(??):281, April 1988.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901553.

Anonymous:1988:AIVd

[Ano88d] Anonymous. Author index for volume 102. Linear Algebra and its Applications, 102(??):247, April 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958890328X.

Anonymous:1988:AIVe

[Ano88e] Anonymous. Author index for volume 103. Linear Algebra and its Applications, 103(??):249, May 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902303.

Anonymous:1988:AIVf

[Ano88f] Anonymous. Author index for volume 104. Linear Algebra and its Applications, 104(??):250, June 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588903163.

Anonymous:1988:AIVg

[Ano88g] Anonymous. Author index for volume 105. Linear Algebra and its Applications, 105(??):249, July 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588900171.

Anonymous:1988:AIVh

[Ano88h] Anonymous. Author index for volume 106. Linear Algebra and its Applications, 106(??):280, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588900353.

Anonymous:1988:AIVi

[Ano88i] Anonymous. Author index for volume 107. Linear Algebra and its Applications, 107(??):359, August 1988.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588902546.

Anonymous:1988:AIVj

[Ano88j] Anonymous. Author index for volume 108. Linear Algebra and its Applications, 108(??):281, September 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901930.

Anonymous:1988:AIVk

[Ano88k] Anonymous. Author index for volume 109. Linear Algebra and its Applications, 109(??):279, October 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588902121.

Anonymous:1988:AIVl

[Ano88l] Anonymous. Author index for volume 111. Linear Algebra and its Applications, 111(??):327, December 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588900699.

Anonymous:1988:AIVa

[Ano88m] Anonymous. Author index for volume 98. Linear Algebra and its Applications, 98(??):434, January 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901759.

Anonymous:1988:AIVb

[Ano88n] Anonymous. Author index for volume 99. Linear Algebra and its Applications, 99(??):278, February 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901371.

Anonymous:1988:BR

[Ano88o] Anonymous. Book reviews. Linear Algebra and its Applications, 100(??):4–5, March 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588903825.

Anonymous:1988:CR

[Ano88p] Anonymous. Conference reports. Linear Algebra and its Applications, 100(??):3, March 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588903357.

Anonymous:1988:EBa

[Ano88q] Anonymous. Editorial Board. Linear Algebra and its Applications, 98(??):ii–iii, January 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901565.

Anonymous:1988:EBb

[Ano88r] Anonymous. Editorial Board. Linear Algebra and its Applications, 100(??):ii–iii, March 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588903291.

Anonymous:1988:EBc

[Ano88s] Anonymous. Editorial Board. Linear Algebra and its Applications, 101(??):i–ii, April 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379588901383.

Anonymous:1988:EBd

[Ano88t] Anonymous. Editorial Board. Linear Algebra and its Applications, 102(??):ii–iii, April 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379588903175.

Anonymous:1988:EBe

[Ano88u] Anonymous. Editorial Board. Linear Algebra and its Applications, 103(??):ii–iii, May 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379588902169.

Anonymous:1988:EBf

[Ano88v] Anonymous. Editorial Board. Linear Algebra and its Applications, 104(??):ii–iii, June 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379588903023.

Anonymous:1988:EBg

[Ano88w] Anonymous. Editorial Board. Linear Algebra and its Applications, 105(??):ii–iii, July 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379588900018.

Anonymous:1988:EBh

[Ano88x] Anonymous. Editorial Board. Linear Algebra and its Applications, 107(??):i–ii, August 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902315.

Anonymous:1988:EBi

[Ano88y] Anonymous. Editorial Board. Linear Algebra and its Applications, 108(??):ii–iii, September 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901760.

Anonymous:1988:EBj

[Ano88z] Anonymous. Editorial Board. Linear Algebra and its Applications, 109(??):ii–iii, October 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901942.

Anonymous:1988:EC

[Ano88-27] Anonymous. Editors-in-chief. Linear Algebra and its Applications, 100(??):viii, March 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
002437958890331X.

Anonymous:1988:LLA

[Ano88-28] Anonymous. Letter in linear algebra, volumes 30-41, and reports, volumes 42-54. Linear Algebra and its Applications, 100(??):16, March 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588903758.

Anonymous:1988:O

[Ano88-29] Anonymous. Obituaries. Linear Algebra and its Applications, 100(??):6, March 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379588903734.

Anonymous:1988:PFM

[Ano88-30] Anonymous. Present and former members of the editorial board. Linear Algebra and its Applications, 100 (??):ix, March 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588903321.

Anonymous:1988:PFR

[Ano88-31] Anonymous. Present and former referees. Linear Algebra and its Applications, 100(??):7–15, March 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588903746.

Anonymous:1988:PVC

[Ano88-32] Anonymous, editor. Proceedings of the Victoria Conference on Combinatorial Matrix Analysis: May 20–23, 1987, volume 107 of Linear Algebra and its Applications. Elsevier, Amsterdam, The Netherlands, 1988. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:1988:P

[Ano88-33] Anonymous. Profiles. Linear Algebra and its Applications, 100(??):2, March 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379588903345.

Anonymous:1988:SI

[Ano88-34] Anonymous. Special issues. Linear Algebra and its Applications, 100(??):1, March 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379588903333.

Anonymous:1989:AIVa

[Ano89a] Anonymous. Author index for volume 112. Linear Algebra and its Applications, 112(??):247, January 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589905995.

Anonymous:1989:AIVb

[Ano89b] Anonymous. Author index for volume 113. Linear Algebra and its Applications, 113(??):183, February 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589902966.

Anonymous:1989:AIVc

[Ano89c] Anonymous. Author index for volume 116. Linear Algebra and its Applications, 116(??):183, April 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589904047.

Anonymous:1989:AIVd

[Ano89d] Anonymous. Author index for volume 117. Linear Algebra and its Applications, 117(??):185, May 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589905569.

Anonymous:1989:AIVe

[Ano89e] Anonymous. Author index for volume 118. Linear Algebra and its Applications, 118(??):249, June 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589905818.

Anonymous:1989:AIVf

[Ano89f] Anonymous. Author index for volume 119. Linear Algebra and its Applications, 119(??):183, July 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589900773.

Anonymous:1989:AIVh

[Ano89g] Anonymous. Author index for volume 121. Linear Algebra and its Applications, 121(??):711–712, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589907222.

Anonymous:1989:AIVi

[Ano89h] Anonymous. Author index for volume 125. Linear Algebra and its Applications, 125(??):151, December 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589900396.

Anonymous:1989:AIVj

[Ano89i] Anonymous. Author index for volume 126. Linear Algebra and its Applications, 126(??):151, December 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958990013X.

Anonymous:1989:AIVg

[Ano89j] Anonymous. Author index for Volumes 101–120. Linear Algebra and its Applications, 120(??):279–289, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589903911.

Anonymous:1989:EBa

[Ano89k] Anonymous. Editorial Board. Linear Algebra and its Applications, 112(??):ii–iii, January 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958990582X.

Anonymous:1989:EBb

[Ano89l] Anonymous. Editorial Board. Linear Algebra and its Applications, 113(??):ii–iii, February 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589902838.

Anonymous:1989:EBc

[Ano89m] Anonymous. Editorial Board. Linear Algebra and its Applications, 116(??):ii–iii, April 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379589903923.

Anonymous:1989:EBd

[Ano89n] Anonymous. Editorial Board. Linear Algebra and its Applications, 117(??):ii–iii, May 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379589905405.

Anonymous:1989:EBe

[Ano89o] Anonymous. Editorial Board. Linear Algebra and its Applications, 118(??):ii–iii, June 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379589905685.

Anonymous:1989:EBf

[Ano89p] Anonymous. Editorial Board. Linear Algebra and its Applications, 120(??):ii–iii, August 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589903637.

Anonymous:1989:EBg

[Ano89q] Anonymous. Editorial Board. Linear Algebra and its Applications, 121(??):ii–iii, August 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589906861.

Anonymous:1989:EBh

[Ano89r] Anonymous. Editorial Board. Linear Algebra and its Applications, 125(??):ii–iii, December 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589900293.

Anonymous:1989:EBi

[Ano89s] Anonymous. Editorial board. Linear Algebra and its Applications, 126(??):ii–iii, December 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589900013.

Anonymous:1989:NEC

[Ano89t] Anonymous. Note from the Editor-in-Chief. Linear Algebra and its Applications, 125(??):149, December 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589900372.

Anonymous:1989:RV

[Ano89u] Anonymous. Referees for volumes 101–120. Linear Algebra and its Applications, 120(??):273–277, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958990390X.

Antoulas:1983:NRA

[Ant83] A. C. Antoulas. New results on the algebraic theory of linear systems: the solution of the cover problems. Linear Algebra and its Applications, 50(??):1–43, April 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583900526.

Antoulas:1988:RIE

[Ant88] A. C. Antoulas. Rational interpolation and the Euclidean algorithm. Linear Algebra and its Applications, 108(??):157–171, September 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901851.

Appelgate:1982:SPI

[AO82] Harry Appelgate and Hironori Onishi. The similarity problem for 3 × 3 integer matrices. Linear Algebra and its Applications, 42(??):159–174, February 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958290146X.

Anderson:1985:MLE

[AO85] T. W. Anderson and I. Olkin. Maximum-likelihood estimation of the parameters of a multivariate normal distribution. Linear Algebra and its Applications, 70(??): 147–171, October 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585900497.

Anderson:1983:GIA

[AP83] Steen A. Anderson and Michael D. Perlman. Groupinvariant analogues of Hadamard’s inequality. Linear Algebra and its Applications, 110(??):91–116, November 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583901349.

Axelsson:1986:AFM

[AP86] Owe Axelsson and Ben Polman. On approximate factorization methods for block matrices suitable for vector and parallel processors. Linear Algebra and its Applications, 77(??):3–26, May 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958690159X.

Special volume on parallel computing.

Andersson:1988:GIA

[AP88] Steen A. Andersson and Michael D. Perlman. Groupinvariant analogues of Hadamard’s inequality. Linear Algebra and its Applications, 110(??):91–116, 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alexander:1988:ARL

[APP88] S. T. Alexander, C.-T. Pan, and R. J. Plemmons. Analysis of a recursive least squares hyperbolic rotation algorithm for signal processing. Linear Algebra and its Applications, 98(??):3–40, January 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901589.

Anstreicher:1987:UGJ

[AR87] Kurt M. Anstreicher and Uriel G. Rothblum. Using Gauss–Jordan elimination to compute the index, generalized nullspaces, and Drazin inverse. Linear Algebra and its Applications, 85(??):221–239, January 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587902199.

Artzrouni:1983:TPM

[Art83] Marc Artzrouni. A theorem on products of matrices. Linear Algebra and its Applications, 49(??):153– 159, February 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901003.

Artzrouni:1986:CIP

[Art86] Marc Artzrouni. On the convergence of infinite products of matrices. Linear Algebra and its Applications, 74(??): 11–21, February 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901126.

Aspvall:1980:FAS

[AS80] Bengt Aspvall and Yossi Shiloach. A fast algorithm for solving systems of linear equations with two variables per equation. Linear Algebra and its Applications, 34(??):117–124, December 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901627.

Alefeld:1982:SRM

[AS82] G. Alefeld and N. Schneider. On square roots of Mmatrices. Linear Algebra and its Applications, 42(??): 119–132, February 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582901434.

Ashley:1987:PFE

[Ash87] Jonathan Ashley. On the Perron–Frobenius eigenvector for nonnegative integral matrices whose largest eigenvalue is integral. Linear Algebra and its Applications, 94(??):103–108, September 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587900814.

Anderson:1980:ISQ

[AT80] T. W. Anderson and John B. Taylor. An inequality for a sum of quadratic forms with applications to probability theory. Linear Algebra and its Applications, 30 (??):93–99, April 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580901846.

Anderson:1988:EPS

[AT88] William N. Anderson, Jr. and George E. Trapp. The extreme points of a set of positive semidefinite operators. Linear Algebra and its Applications, 106(??):209– 217, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588900298.

Atalla:1989:SSP

[Ata89] Robert E. Atalla. Saturated sets of probability measures and stochastic matrices. Linear Algebra and its Applications, 112(??):207–216, January 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958990596X.

Avram:1989:STS

[Avr89] Florin Avram. Szegő’s theorem for sums with dependent indices. Linear Algebra and its Applications, 120(??): 243–249, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589903820.

Axelsson:1980:CGT

[Axe80] Owe Axelsson. Conjugate gradient type methods for unsymmetric and inconsistent systems of linear equations. Linear Algebra and its Applications, 29(??):1– 16, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580902268.

Axelsson:1986:GIB

[Axe86] O. Axelsson. A general incomplete block-matrix factorization method. Linear Algebra and its Applications, 74(??): 179–190, February 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901217.

Aupetit:1983:LBS

[AZ83] Bernard Aupetit and Jaroslav Zemánek. Local behavior of the spectrum near algebraic elements. Linear Algebra and its Applications, 52/53(??):39–44, 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bitmead:1980:AFS

[BA80] Robert R. Bitmead and Brian D. O. Anderson. Asymptotically fast solution of Toeplitz and related systems of linear equations. Linear Algebra and its Applications, 34(??):103–116, December 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901615.

Ben-Artzi:1987:MDP

[BAEGL87] A. Ben-Artzi, R. L. Ellis, I. Gohberg, and D. C. Lay. The maximum distance problem and band sequences. Linear Algebra and its Applications, 87(??):93–112, March 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587901613.

Ben-Artzi:1989:ITN

[BAG89] A. Ben-Artzi and I. Gohberg. Inertia theorems for nonstationary discrete systems and dichotomy. Linear Algebra and its Applications, 120(??):95–138, August 1989.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589903728. Proceedings

of the Fourth Haifa Matrix Theory Conference (Haifa, 1988).

Bailey:1985:FDA

[Bai85] R. A. Bailey. Factorial design and Abelian groups. Linear Algebra and its Applications, 70(??):349–368, October 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585900643.

Bai:1988:NQC

[Bai88] Zhao Jun Bai. Note on the quadratic convergence of Kogbetliantz’s algorithm for computing the singular value decomposition. Linear Algebra and its Applications, 104 (??):131–140, June 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588903138.

Baksalary:1985:SUL

[Bak85] Jerzy K. Baksalary. Strong unified-least-squares matrices for a general linear model. Linear Algebra and its Applications, 70(??):61–65, October 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900424.

Baksalary:1986:RBS

[Bak86] Jerzy K. Baksalary. A relationship between the star and minus orderings. Linear Algebra and its Applications, 82(??):163–167, October 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586901497.

Baker:1987:SSE

[Bak87] Kirby A. Baker. Strong shift equivalence and shear adjacency of nonnegative square integer matrices. Linear Algebra and its Applications, 93(??):131–147, August 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0024379587903193.

Bannai:1982:TKD

[Ban82] Eiichi Bannai. Tannaka-kreĭn duality for association schemes. Linear Algebra and its Applications, 46(??): 83–90, August 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582900283.

Bapat:1982:TMM

[Bap82] Ravindra Bapat. D_{1}AD_{2} theorems for multidimensional matrices. Linear Algebra and its Applications, 48(??): 437–442, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582901252.

Bapat:1983:DSM

[Bap83] Ravindra Bapat. Doubly stochastic matrices with equal subpermanents. Linear Algebra and its Applications, 51 (??):1–8, June 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901453.

Bapat:1984:SFE

[Bap84a] R. B. Bapat. A stronger form of the Egorychev– Falikman theorem on permanents. Linear Algebra and its Applications, 63(??):95–100, December 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958490137X.

Bapat:1984:MP

[Bap84b] Ravindra Bapat. A matrix of permanents. Linear Algebra and its Applications, 56(??):57–62, January 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901137.

Bapat:1985:PPS

[Bap85] Ravindra Bapat. On permanents of positive semidefinite matrices. Linear Algebra and its Applications, 65(??):

83–90, February 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585900898.

Bapat:1986:AII

[Bap86a] R. B. Bapat. Applications of an inequality in information theory to matrices. Linear Algebra and its Applications, 78(??):107–117, June 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586900182.

Bapat:1986:BPL

[Bap86b] R. B. Bapat. A bound for the permanent of the Laplacian matrix. Linear Algebra and its Applications, 74(??): 219–223, February 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901242.

Bapat:1986:IMS

[Bap86c] R. B. Bapat. Inequalities for mixed Schur functions. Linear Algebra and its Applications, 83(??):143–149, November 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586902715.

Bapat:1987:TIP

[Bap87] R. B. Bapat. Two inequalities for the Perron Root. Linear Algebra and its Applications, 85(??):241–248, January 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587902205.

Bapat:1989:MDP

[Bap89] R. B. Bapat. Mixed discriminants of positive semidefinite matrices. Linear Algebra and its Applications, 126(??): 107–124, December 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589900098.

Barnett:1980:RBF

[Bar80a] S. Barnett. Review of Mathematics of Finite-Dimensional Control Systems: Theory and Design, by David L. Russell. Linear Algebra and its Applications, 33(??):155– 158, October 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580901044.

Bartels:1980:PLP

[Bar80b] Richard Bartels. A penalty linear programming method using reduced-gradient basis-exchange techniques. Linear Algebra and its Applications, 29(??):17–32, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958090227X.

Barker:1981:RBM

[Bar81a] G. P. Barker. Review of Nonnegative Matrices in the Mathematical Sciences, by Abraham Berman and Robert J. Plemmons. Linear Algebra and its Applications, 35(??): 293–295, February 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902810.

Barker:1981:TC

[Bar81b] George Phillip Barker. Theory of cones. Linear Algebra and its Applications, 39(??):263–291, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581903104.

Barnett:1981:CM

[Bar81c] Stephen Barnett. Congenial matrices. Linear Algebra and its Applications, 41(??):277–298, December 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958190104X.

Barker:1982:R

[Bar82] George Phillip Barker. Reports. Linear Algebra and its Applications, 48(??):471, December 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582901288.

Bart:1983:UTM

[Bar83] H. Bart. Upper triangularization of matrices by permutations and lower triangular similarity transformations. Linear Algebra and its Applications, 110(??):255– 261, November 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901398.

Barnett:1984:DGP

[Bar84a] Stephen Barnett. Division of generalized polynomials using the comrade matrix. Linear Algebra and its Applications, 60(??):159–175, August 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900776.

Barrett:1984:TMB

[Bar84b] Wayne W. Barrett. Toeplitz matrices with banded inverses. Linear Algebra and its Applications, 57(??):131– 145, February 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901824.

Bart:1986:TFO

[Bar86] H. Bart. Transfer functions and operator theory. Linear Algebra and its Applications, 84(??):33–61, December 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958690306X. Pro-

ceedings of the symposium on operator theory (Athens, 1985).

Barker:1987:RBN

[Bar87] George P. Barker. Review of Matrix Norms and Their Applications. Linear Algebra and its Applications, 87 (??):271, March 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958790173X.

Barnett:1988:ERG

[Bar88a] Stephen Barnett. Euclidean remainders for generalized polynomials. Linear Algebra and its Applications, 99(??):

111–122, February 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588901280.

Bart:1988:UTMb

[Bar88b] H. Bart. Upper triangularization of matrices by permutations and lower triangular similarity transformations. Linear Algebra and its Applications, 110(??):255– 261, 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baragana:1989:IIR

[Bar89a] Itziar Baragaña. Interlacing inequalities for regular pencils. Linear Algebra and its Applications, 121(??):521– 535, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589907209.

Linear algebra and applications (Valencia, 1987).

Barker:1989:AGA

[Bar89b] George Phillip Barker. Automorphism groups of algebras of triangular matrices. Linear Algebra and its Applications, 121(??):207–215, August 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589907027. Linear alge-

bra and applications (Valencia, 1987).

Barnes:1989:CDC

[Bar89c] Earl R. Barnes. Circular discs containing eigenvalues of normal matrices. Linear Algebra and its Applications, 114/115(??):501–521, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Basor:1985:RBA

[Bas85] Estelle Basor. Review of Invertibility and Asymptotics of Toeplitz Matrices, by A. Böttcher and B. Silbermann. Linear Algebra and its Applications, 68(??):275–278, July 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902253.

Ben-Artzi:1986:ITC

[BAS86] Asher Ben-Artzi and Tamir Shalom. On inversion of Toeplitz and close to Toeplitz matrices. Linear Algebra and its Applications, 75(??):173–192, March 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586901886.

Biggs:1982:CGT

[BB82a] N. L. Biggs and W. G. Bridges. Combinatorics and graph theory. Linear Algebra and its Applications, 46 (??):v, August 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582900209.

Biggs:1982:SVC

[BB82b] Norman L. Biggs and William G. Bridges, editors. Special volume: combinatorics and graph theory, volume 46 of Linear Algebra and its Applications. Elsevier Science Publishers, Amsterdam, The Netherlands, 1982. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bojanczyk:1986:PSC

[BB86] A. Bojańczyk and R. P. Brent. Parallel solution of certain Toeplitz least-squares problems. Linear Algebra and its Applications, 77(??):43–60, May 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586901618. Special vol-

ume on parallel computing.

Benzaouia:1989:ENS

[BB89a] Abdellah Benzaouia and Christian Burgat. Existence of nonsymmetrical stability domains for linear systems. Linear Algebra and its Applications, 121(??):217–231, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589907039. Linear

algebra and applications (Valencia, 1987).

Binding:1989:TPE

[BB89b] Paul Binding and Patrick J. Browne. Two parameter eigenvalue problems for matrices. Linear Algebra and its Applications, 113(??):139–157, February 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589902930.

Blake:1988:P

[BBV88a] I. F. Blake, N. K. Bose, and M. Vidyasagar. Preface. Linear Algebra and its Applications, 98(??):1–2, January 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901577.

Blake:1988:LAE

[BBV88b] Ian F. Blake, N. K. (Nirmal K.) Bose, and M. (Mathukumalli) Vidyasagar, editors. Linear algebra in electrical engineering, volume 96 of Linear Algebra and its Applications. North-Holland Publishing Co., Amsterdam, The Netherlands, 1988. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bini:1983:SCP

[BC83] Dario Bini and Milvio Capovani. Spectral and computational properties of band symmetric Toeplitz matrices. Linear Algebra and its Applications, 52/53(??):99–126, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Benedetti:1984:SDO

[BC84] R. Benedetti and P. Cragnolini. On simultaneous diagonalization of one Hermitian and one symmetric form. Linear Algebra and its Applications, 57(??):215– 226, February 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901897.

Buchi:1985:CPA

[BC85] Werner Büchi and Michael Clausen. On a class of primary algebras of minimal rank. Linear Algebra and its Applications, 69(??):249–268, August 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900801.

Bolzern:1987:ITP

[BC87] Paolo Bolzern and Patrizio Colaneri. Inertia theorems for the periodic Lyapunov difference equation and periodic Riccati difference equation. Linear Algebra and its Applications, 85(??):249–265, January 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587902217.

Bevilacqua:1988:PSB

[BCR88] Roberto Bevilacqua, Bruno Codenotti, and Francesco Romani. Parallel solution of block tridiagonal linear systems. Linear Algebra and its Applications, 104(??):39–57, June 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588903059.

Barker:1983:CCC

[BCS83] G. P. Barker, L. T. Conner, Jr., and D. P. Stanford. Complete controllability and contractibility in multimodal systems. Linear Algebra and its Applications, 110(??): 55–74, November 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901325.

Berman:1986:HCM

[BCS86] Abraham Berman, Yair Censor, and Hans Schneider. Haifa 1985 Conference on Matrix Theory. Linear Algebra and its Applications, 80(??):173–260, August 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902843.

Butkovic:1987:SLI

[BCS87] Peter Butkovič, Katarína Cechlárová, and Peter Szabó. Strong linear independence in bottleneck algebra. Linear Algebra and its Applications, 94(??):133–155, September 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587900851.

Barker:1988:CCC

[BCS88] G. P. Barker, L. T. Conner, Jr., and D. P. Stanford. Complete controllability and contractibility in multimodal systems. Linear Algebra and its Applications, 110(??):55–74, 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bjorck:1980:DMS

[BD80] Å. Björck and Iain S. Duff. A direct method for the solution of sparse linear least squares problems. Linear Algebra and its Applications, 34(??):43–67, December 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580901585.

Bhatia:1983:PSS

[BDM83] Rajendra Bhatia, Chandler Davis, and Alan McIntosh. Perturbation of spectral subspaces and solution of linear operator equations. Linear Algebra and its Applications, 52/53(??):45–67, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bebiano:1986:DCS

[BdP86] Natália Bebiano and J. da Providência. On the determinant of certain strictly dissipative matrices. Linear Algebra and its Applications, 83(??):117–128, November 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902685.

Bebiano:1988:SRC

[BdP88] Natália Bebiano and João da Providência. Some remarks on a conjecture of de Oliveira. Linear Algebra and its Applications, 102(??):241–246, April 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588903278.

Beasley:1981:LTW

[Bea81a] LeRoy B. Beasley. Linear transformations which preserve fixed rank. Linear Algebra and its Applications, 40(??):

183–187, October 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581901488.

Beasley:1981:SME

[Bea81b] LeRoy B. Beasley. Spaces of matrices of equal rank. Linear Algebra and its Applications, 38(??):227–237, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900227.

Beasley:1982:LTM

[Bea82] LeRoy B. Beasley. Linear transformations on matrices: the invariance of sets of ranks. Linear Algebra and its Applications, 48(??):25–35, December 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582900969.

Beasley:1983:RPP

[Bea83] LeRoy B. Beasley. Rank-k-preservers and preservers of sets of ranks. Linear Algebra and its Applications, 55(??): 11–17, December 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901635.

Beauwens:1984:UEB

[Bea84] Robert Beauwens. Upper eigenvalue bounds for pencils of matrices. Linear Algebra and its Applications, 62(??): 87–104, November 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584900880.

Beauwens:1985:AP

[Bea85] Robert Beauwens. On Axelsson’s perturbations. Linear Algebra and its Applications, 68(??):221–242, July 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902149.

Beard:1986:MFR

[Bea86] Jacob T. B. Beard, Jr. Matrix fields, regular and irregular: a complete fundamental characterization. Linear Algebra and its Applications, 81(??):137–152, September 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902545.

Beasley:1987:MWJ

[Bea87a] LeRoy B. Beasley. The mathematical work of Joel Lee Brenner. With contributions by R. P. Boas, John de Pillis, Roger Lyndon, and James Wiegold. Linear Algebra and its Applications, 90(??):1–13, May 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587903004.

Beauwens:1987:LEB

[Bea87b] Robert Beauwens. Lower eigenvalue bounds for pencils of matrices. Linear Algebra and its Applications, 85(??): 101–119, January 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587902114.

Beasley:1988:LOM

[Bea88] LeRoy B. Beasley. Linear operators on matrices: the invariance of rank-k matrices. Linear Algebra and its Applications, 107(??):161–167, August 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958890242X. Proceedings

of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Bebiano:1986:SAB

[Beb86] Natália Bebiano. Some analogies between the c-numerical range and a certain variation on this concept. Linear Algebra and its Applications, 81(??):47–54, September 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958690248X.

Becker:1980:NSC

[Bec80] Ronald I. Becker. Necessary and sufficient conditions for the simultaneous diagonability of two quadratic forms. Linear Algebra and its Applications, 30(??):

129–139, April 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580901883.

Beck:1989:POM

[Bec89] István Beck. Partial orders and the modular group. Linear Algebra and its Applications, 120(??):139–147, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958990373X. Pro-

ceedings of the Fourth Haifa Matrix Theory Conference (Haifa, 1988).

Beesack:1980:SMP

[Bee80] Paul R. Beesack. The spread of matrices and polynomials. Linear Algebra and its Applications, 31(??):145–149, June 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580902141.

Beezer:1984:PP

[Bee84] Robert A. Beezer. On the polynomial of a path. Linear Algebra and its Applications, 63(??):221–225, December 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901447.

Beezer:1986:TOP

[Bee86] Robert A. Beezer. Trivalent orbit polynomial graphs. Linear Algebra and its Applications, 73(??):133–146, January 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902351.

Bekker:1988:PSP

[Bek88] Paul A. Bekker. The positive semidefiniteness of partitioned matrices. Linear Algebra and its Applications, 111(??):261–278, December 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958890064X.

Bell:1981:GIC

[Bel81] C. Leary Bell. Generalized inverses of circulant and generalized circulant matrices. Linear Algebra and its Applications, 39(??):133–142, August 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581902974.

Bennett:1986:LBM

[Ben86] Grahame Bennett. Lower bounds for matrices. Linear Algebra and its Applications, 82(??):81–98, October 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586901436.

Bru:1988:MPC

[BEN88] Rafael Brú, Ludwig Elsner, and Michael Neumann. Models of parallel chaotic iteration methods. Linear Algebra and its Applications, 103(??):175–192, May 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588902273.

Berman:1981:MLC

[Ber81] Abraham Berman. Matrices and the linear complementarity problem. Linear Algebra and its Applications, 40 (??):249–256, October 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379581901567.

Berg:1986:GOC

[Ber86a] Lothar Berg. General operational calculus. Linear Algebra and its Applications, 84(??):79–97, December 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586903083. Proceedings

of the symposium on operator theory (Athens, 1985).

Berg:1986:TRC

[Ber86b] Lothar Berg. Three results in connection with inverse matrices. Linear Algebra and its Applications, 84(??):

63–77, December 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586903071.

Proceedings of the symposium on operator theory (Athens, 1985).

Berkovits:1986:DTM

[Ber86c] Juha Berkovits. On the degree theory for mappings of monotone type and applications. Linear Algebra and its Applications, 84(??):243–249, December 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586903174. Proceedings

of the symposium on operator theory (Athens, 1985).

Bercovici:1987:NP

[Ber87] Hari Bercovici. Note on property (A_{1}). Linear Algebra and its Applications, 91(??):213–216, June 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587900723.

Berman:1988:CP

[Ber88] Abraham Berman. Complete positivity. Linear Algebra and its Applications, 107(??):57–63, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902376. Proceedings

of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Barrett:1981:IBM

[BF81a] Wayne W. Barrett and Philip J. Feinsilver. Inverses of banded matrices. Linear Algebra and its Applications, 41(??):111–130, December 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581900926.

Bhatia:1981:VGP

[BF81b] Rajendra Bhatia and Shmuel Friedland. Variation of Grassman powers and spectra. Linear Algebra and its Applications, 40(??):1–18, October 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581901361.

Bruen:1982:IRR

[BF82] A. A. Bruen and J. W. Freeman. Intersections of treguli, rational curves, and orthogonal Latin squares. Linear Algebra and its Applications, 46(??):103–116, August 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582900301.

Brockett:1983:LAA

[BF83] Roger W. Brockett and Paul A. Furhmann, editors. Linear algebra and its applications. Special volume: linear systems and control, volume 50 of Linear Algebra and its Applications. Elsevier Science Publishers, Amsterdam, The Netherlands, April 1983. ISSN 0024-3795 (print), 1873-1856 (electronic).

Backhouse:1985:RDM

[BF85] N. B. Backhouse and A. G. Fellouris. Report on the Dublin matrix theory conference, March 1984: Grassmann matrices. Linear Algebra and its Applications, 68 (??):260–263, July 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585902216.

Berger:1987:SPM

[BF87] Marc A. Berger and Alexander Felzenbaum. Sign patterns of matrices and their inverses. Linear Algebra and its Applications, 86(??):161–177, February 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587902928.

Beattie:1988:SCW

[BF88] Christopher Beattie and David W. Fox. Schur complements and the Weinstein–Aronszajn theory for modified matrix eigenvalue problems. Linear Algebra and its Applications, 108(??):37–61, September 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901784.

Bru:1989:CPC

[BF89] Rafael Bru and Robert Fuster. Convergence of the parallel chaotic block extrapolated Jacobi method. Linear Algebra and its Applications, 120(??):252–257, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589903844.

Bisiacco:1989:SFC

[BFM89a] M. Bisiacco, E. Fornasini, and G. Marchesini. 2D systems feedback compensation: an approach based on commutative linear transformations. Linear Algebra and its Applications, 121(??):135–150, August 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589906976. Linear alge-

bra and applications (Valencia, 1987).

Bisiacco:1989:DRS

[BFM89b] Mauro Bisiacco, Ettore Fornasini, and Giovanni Marchesini. Dynamic regulation of 2D systems: a statespace approach. Linear Algebra and its Applications, 122/123/124(??):195–218, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barrett:1988:SRM

[BFP88] Wayne W. Barrett, Rodney W. Forcade, and Andrew D. Pollington. On the spectral radius of a (0, 1) matrix related to mertens’ function. Linear Algebra and its Applications, 107(??):151–159, August 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902418. Proceedings

of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Barany:1981:SCQ

[BG81a] I. Bárány and V. S. Grinberg. On some combinatorial questions in finite-dimensional spaces. Linear Algebra and its Applications, 41(??):1–9, December 1981.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900859.

Bunse-Gerstner:1981:AAC

[BG81b] A. Bunse-Gerstner. An analysis of the HR algorithm for computing the eigenvalues of a matrix. Linear Algebra and its Applications, 35(??):155–173, February 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581902718.

Brebner:1982:ERS

[BG82] M. A. Brebner and J. Grad. Eigenvalues of Ax = λBx for real symmetric matrices A and B computed by reduction to a pseudosymmetric form and the HR process. Linear Algebra and its Applications, 43(??):99–118, March 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582902464.

Bialas:1984:IMR

[BG84a] S. Białas and J. Garloff. Intervals of P -matrices and related matrices. Linear Algebra and its Applications, 58 (??):33–41, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584902027.

Bunse-Gerstner:1984:ASG

[BG84b] A. Bunse-Gerstner. An algorithm for the symmetric generalized eigenvalue problem. Linear Algebra and its Applications, 58(??):43–68, April 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584902039.

Bunse-Gerstner:1986:MFS

[BG86] Angelika Bunse-Gerstner. Matrix factorizations for symplectic QR-like methods. Linear Algebra and its Applications, 83(??):49–77, November 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958690265X.

Ball:1986:SIS

[BGoMSTAUTA^{+}86] Joseph A. Ball, Israel Gohberg, School of Mathematical Sciences Tel Aviv University Tel Aviv, Israel Raymond, and Beverly Sackler Faculty of Exact Sciences. Shift invariant subspaces, factorization, and interpolation for matrices. I. the canonical case. Linear Algebra and its Applications, 74(??):87–150, February 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586901175.

Beasley:1985:NRP

[BGP85] LeRoy B. Beasley, David A. Gregory, and Norman J. Pullman. Nonnegative rank-preserving operators. Linear Algebra and its Applications, 65(??):207–223, February 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585900989.

Barnes:1981:BER

[BH81a] E. R. Barnes and A. J. Hoffman. On bounds for eigenvalues of real symmetric matrices. Linear Algebra and its Applications, 40(??):217–223, October 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581901518.

Brawley:1981:DRB

[BH81b] J. V. Brawley and Max Hankins. On distribution by rank of bases for vector spaces of matrices over a finite field. Linear Algebra and its Applications, 39(??): 91–101, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902937.

Bevis:1982:SCI

[BH82] Jean H. Bevis and Frank J. Hall. Some classes of integral matrices. Linear Algebra and its Applications, 48(??): 473–483, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958290129X.

Bjorck:1983:SMS

[BH83] Å. Björck and S. Hammarling. A Schur method for the square root of a matrix. Linear Algebra and its Applications, 52/53(??):127–140, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Berman:1984:CAS

[BH84] Abraham Berman and Daniel Hershkowitz. Characterization of acyclic D-stable matrices. Linear Algebra and its Applications, 58(??):17–31, April 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584902015.

Brualdi:1985:SRM

[BH85] R. A. Brualdi and A. J. Hoffman. On the spectral radius of (0, 1)-matrices. Linear Algebra and its Applications, 65(??):133–146, February 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900928.

Baksalary:1987:POM

[BH87a] Jerzy K. Baksalary and Jan Hauke. Partial orderings of matrices referring to singular values or eigenvalues. Linear Algebra and its Applications, 96(??):17–26, November 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587903338.

Berman:1987:CRC

[BH87b] Abraham Berman and Daniel Hershkowitz. Combinatorial results on completely positive matrices. Linear Algebra and its Applications, 95(??):111–125, October 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587900292.

Bhatia:1987:UIS

[BH87c] Rajendra Bhatia and John A. R. Holbrook. Unitary invariance and spectral variation. Linear Algebra and its Applications, 95(??):43–68, October 1987.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587900267.

Bart:1988:CTF

[BH88] H. Bart and H. Hoogland. Complementary triangular forms of pairs of matrices, realizations with prescribed main matrices, and complete factorization of rational matrix functions. Linear Algebra and its Applications, 103(??):193–228, May 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902285.

Bru:1989:SPD

[BH89] Rafael Brú and Vicente Hernández. Structural properties of discrete-time linear positive periodic systems. Linear Algebra and its Applications, 121(??):171–183, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589907003. Linear

algebra and applications (Valencia, 1987).

Bhatia:1981:RCS

[Bha81] Rajendra Bhatia. On the rate of change of spectra of operators. II. Linear Algebra and its Applications, 36 (??):25–32, March 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902160.

Bhatia:1984:VST

[Bha84] Rajendra Bhatia. Variation of symmetric tensor powers and permanents. Linear Algebra and its Applications, 62(??):269–276, November 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901022.

Barker:1984:CPP

[BHH84] George Phillip Barker, Richard D. Hill, and Raymond D. Haertel. On the completely positive and positive-semidefinite-preserving cones. Linear Algebra and its Applications, 56(??):221–229, January 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901277.

Bevis:1987:DIS

[BHH87a] Jean H. Bevis, Frank J. Hall, and Robert E. Hartwig. The Drazin inverse of a semilinear transformation and its matrix representation. Linear Algebra and its Applications, 97(??):229–242, December 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587901522.

Bevis:1987:PSC

[BHH87b] Jean H. Bevis, Frank J. Hall, and Robert E. Hartwig. Pseudoconsimilarity and semiconsimilarity of complex matrices. Linear Algebra and its Applications, 90(??): 73–80, May 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587903065.

Bhimasankaram:1988:GIB

[Bhi88] P. Bhimasankaram. On generalized inverses of a block in a partitioned matrix. Linear Algebra and its Applications, 109(??):131–143, October 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902030.

Berman:1985:LTP

[BHJ85] Abraham Berman, Daniel Hershkowitz, and Charles R. Johnson. Linear transformations that preserve certain positivity classes of matrices. Linear Algebra and its Applications, 68(??):9–29, July 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902058.

Bevis:1981:IGI

[BHK81] Jean H. Bevis, Frank J. Hall, and Irving J. Katz. Integer generalized inverses of incidence matrices. Linear Algebra and its Applications, 39(??):247–258, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581903074.

Berman:1989:HCM

[BHL89a] Abraham Berman, Daniel Hershkowitz, and Leonid Lerer. Haifa 1988 Conference on Matrix Theory. Linear Algebra and its Applications, 120(??):237–243, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589903819.

Berman:1989:P

[BHL89b] Abraham Berman, Daniel Hershkowitz, and Leonid Lerer. Preface. Linear Algebra and its Applications, 120(??): 1, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589903649.

Berman:1989:PFH

[BHL89c] Abraham Berman, Daniel Hershkowitz, and Leonid Lerer, editors. Proceedings of the Fourth Haifa Matrix Theory Conference: January 3–5, 1988, volume 120 of Linear Algebra and its Applications. North-Holland Publishing Co., Amsterdam, The Netherlands, 1989. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ben-Israel:1982:CRL

[BI82] Adi Ben-Israel. A Cramer rule for least-norm solutions of consistent linear equations. Linear Algebra and its Applications, 43(??):223–226, March 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582902555.

Ben-Israel:1988:CBL

[BI88] Adi Ben-Israel. Canonical bases in linear programming. Linear Algebra and its Applications, 102(??): 95–119, April 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588903229.

Bier:1984:DIA

[Bie84] Thomas Bier. A distribution invariant for association schemes and strongly regular graphs. Linear Algebra and its Applications, 57(??):105–113, February 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901800.

Biggs:1980:GVE

[Big80] Norman Biggs. Girth, valency, and excess. Linear Algebra and its Applications, 31(??):55–59, June 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580902050.

Binding:1984:IHP

[Bin84] Paul Binding. The inertia of a Hermitian pencil. Linear Algebra and its Applications, 63(??):179–191, December 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901423.

Bini:1986:BRT

[Bin86] Dario Bini. Border rank of m × n × (mn − q) tensors. Linear Algebra and its Applications, 79(??):45–51, July 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902910.

Bird:1981:SCN

[Bir81] Charles G. Bird. On the stability of a class of noncooperative games. Linear Algebra and its Applications, 40(??): 119–127, October 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581901440.

Bixby:1982:STC

[Bix82] Robert E. Bixby. A simple theorem on 3-connectivity. Linear Algebra and its Applications, 45(??):123–126, June 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582902142.

Barrett:1984:DFM

[BJ84a] Wayne W. Barrett and Charles R. Johnson. Determinantal formulae for matrices with sparse inverses. Linear Algebra and its Applications, 56(??):73–88, January 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901150.

Barrett:1984:PST

[BJ84b] Wayne W. Barrett and Charles R. Johnson. Possible spectra of totally positive matrices. Linear Algebra and its Applications, 62(??):231–233, November 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584900983.

Barrett:1986:DFM

[BJ86] Wayne W. Barrett and Charles R. Johnson. Determinantal formulae for matrices with sparse inverses. II. asymmetric zero patterns. Linear Algebra and its Applications, 81(??):237–261, September 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902582.

Bart:1988:UTMa

[BJ88a] H. Bart and P. S. M. Kop Jansen. Upper triangularization of matrices by lower triangular similarities. Linear Algebra and its Applications, 103(??):229–248, May 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588902297.

Berman:1988:NGI

[BJ88b] A. Berman and S. K. Jain. Nonnegative generalized inverses of powers of nonnegative matrices. Linear Algebra and its Applications, 107(??):169–179, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902431. Proceedings

of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Barrett:1989:DFM

[BJL89] Wayne W. Barrett, Charles R. Johnson, and Michael Lundquist. Determinantal formulae for matrix completions associated with chordal graphs. Linear Algebra and its Applications, 121(??):265–289, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589907064. Linear

algebra and applications (Valencia, 1987).

Bjorck:1987:SAM

[Bjö87] Å. Björck. Stability analysis of the method of seminormal equations for linear least squares problems. Linear Algebra and its Applications, 88/89(??):31–48, 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bachem:1980:MNP

[BK80a] Achim Bachem and Bernhard Korte. Minimum norm problems over transportation polytopes. Linear Algebra and its Applications, 31(??):103–118, June 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580902116.

Baksalary:1980:ME

[BK80b] J. K. Baksalary and R. Kala. The matrix equation AXB + CY D = E. Linear Algebra and its Applications, 30(??):141–147, April 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901895.

Bunch:1980:CMI

[BK80c] James R. Bunch and Linda Kaufman. A computational method for the indefinite quadratic programming problem. Linear Algebra and its Applications, 34(??):341– 370, December 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958090172X.

Baksalary:1981:SM

[BK81a] J. K. Baksalary and R. Kala. Symmetrizers of matrices. Linear Algebra and its Applications, 35(??):51– 62, February 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902652.

Biegler-Konig:1981:CBM

[BK81b] Friedrich W. Biegler-König. Construction of band matrices from spectral data. Linear Algebra and its Applications, 40(??):79–87, October 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581901415.

Biegler-Konig:1981:SCS

[BK81c] Friedrich W. Biegler-König. Sufficient conditions for the solubility of inverse eigenvalue problems. Linear Algebra and its Applications, 40(??):89–100, October 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581901427.

Bruckstein:1986:SMF

[BK86] Alfred Bruckstein and Thomas Kailath. Some matrix factorization identities for discrete inverse scattering. Linear Algebra and its Applications, 74(??):157–172, February 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586901199.

Bapat:1987:G

[BK87] R. B. Bapat and Man Kam Kwong. A generalization of A ∘ A^{−1} ≥ I. Linear Algebra and its Applications, 93(??):107–112, August 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/S0024379587903156.

Bhatia:1988:SPI

[BK88a] Rajendra Bhatia and Fuad Kittaneh. On some perturbation inequalities for operators. Linear Algebra and its Applications, 106(??):271–279, August 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588900341.

Bistritz:1988:IFN

[BK88b] Yuval Bistritz and Thomas Kailath. Inversion and factorization of non-Hermitian quasi-Toeplitz matrices. Linear Algebra and its Applications, 98(??):77–121, January 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901619.

Brewer:1988:DFC

[BK88c] James Brewer and Lee Klingler. Dynamic feedback over commutative rings. Linear Algebra and its Applications, 98(??):137–168, January 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901632.

Bhattacharyya:1988:SLS

[BKH88] S. P. Bhattacharyya, L. H. Keel, and J. W. Howze. Stabilization of linear systems with fixed order controllers. Linear Algebra and its Applications, 98(??):57–76, January 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901607.

Baksalary:1983:MI

[BKK83] J. K. Baksalary, R. Kala, and K. Kłaczyński. The matrix inequality M ≥ B^{∗}MB. Linear Algebra and its Applications, 54(??):77–86, October 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583902069.

Bart:1985:RBP

[BKL85] H. Bart, M. A. Kaashoek, and L. Lerer. Review of Matrix polynomials, by I. Gohberg, P. Lancaster, and L. Rodman. Linear Algebra and its Applications, 64(??):267–272, January 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902824.

Bruning:1980:CVP

[BKR80] Jochen Brüning, Ki Hang Kim, and Fred N. Roush. On a conjecture of S. V. Phadke and N. K. Thakare: “Generalized inverses and operator equations” [Linear Algebra Appl. 23 (1979), 191–199, MR 80b:47017]. Linear Algebra and its Applications, 32(??):113–114, August 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580900105.

Brualdi:1982:PNM

[BL82a] Richard A. Brualdi and Mordechai Lewin. On powers of nonnegative matrices. Linear Algebra and its Applications, 43(??):87–97, March 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582902452.

Brualdi:1982:SDI

[BL82b] Richard A. Brualdi and Qiao Li. Small diameter interchange graphs of classes of matrices of zeros and ones. Linear Algebra and its Applications, 46(??):177– 194, August 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582900349.

Beasley:1989:LOM

[BL89a] LeRoy B. Beasley and Thomas J. Laffey. Linear operators on matrices: the invariance of rank-k matrices. Linear Algebra and its Applications, 120(??):249–251, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589903832.

Beslin:1989:GCD

[BL89b] Scott Beslin and Steve Ligh. Greatest common divisor matrices. Linear Algebra and its Applications, 118 (??):69–76, June 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589905727.

Bistritz:1989:IVS

[BLAK89] Yuval Bistritz, Hanoch Lev-Ari, and Thomas Kailath. Immittance- versus scattering-domain fast algorithms for non-Hermitian Toeplitz and quasi-Toeplitz matrices. Linear Algebra and its Applications, 122/123/124(??):847– 888, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bridges:1982:RSG

[BM82] W. G. Bridges and R. A. Mena. On the rational spectra of graphs with Abelian Singer groups. Linear Algebra and its Applications, 46(??):51–60, August 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582900258.

Baksalary:1985:ALE

[BM85] Jerzy K. Baksalary and Augustyn Markiewicz. Admissible linear estimators in restricted linear models. Linear Algebra and its Applications, 70(??):9–19, October 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585900394.

Butler:1988:SGS

[BM88] C. A. Butler and T. D. Morley. Six generalized Schur complements. Linear Algebra and its Applications, 106(??): 259–269, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958890033X.

Baksalary:1989:MIA

[BM89a] Jerzy K. Baksalary and Augustyn Markiewicz. A matrix inequality and admissibility of linear estimators with respect to the mean square error matrix criterion. Linear Algebra and its Applications, 112(??):9–18, January 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589905843.

Bhattacharya:1989:USM

[BM89b] Ratna Bhattacharya and Kalyan Mukherjea. On unitary similarity of matrices. Linear Algebra and its Applications, 126(??):95–105, December 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589900086.

Brualdi:1989:CMZ

[BM89c] Richard A. Brualdi and T. S. Michael. The class of matrices of zeros, ones, and twos with prescribed row and column sums. Linear Algebra and its Applications, 114/115 (??):181–198, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bieri:1983:SPA

[BN83] H. Bieri and W. Nef. A sweep-plane algorithm for computing the volume of polyhedra represented in Boolean form. Linear Algebra and its Applications, 52/53(??):69– 97, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Browne:1988:CMG

[BN88] M. W. Browne and H. Neudecker. The covariance matrix of a general symmetric second degree matrix polynomial under normality assumptions. Linear Algebra and its Applications, 103(??):13–20, May 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902182.

Brouwer:1989:GSR

[BN89] A. E. Brouwer and A. Neumaier. The graphs with spectral radius between 2 and \sqrt{2 + \sqrt{5}}. Linear Algebra and its Applications, 114/115(??):273–276, 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Berman:1986:CRN

[BNS86] Abraham Berman, Michael Neumann, and Ronald J. Stern. Cone reachability for nondiagonable linear differential systems. Linear Algebra and its Applications, 81(??):263–280, September 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902594.

Buoni:1982:TSR

[BNV82] John J. Buoni, Michael Neumann, and Richard S. Varga. Theorems of Stein–Rosenberg type. III. the singular case. Linear Algebra and its Applications, 42(??):183– 198, February 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582901483.

Bojanczyk:1986:SIL

[Boj86] A. W. Bojańczyk. Systolic implementation of the lattice algorithm for least squares linear prediction problems. Linear Algebra and its Applications, 77(??):27–42, May 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586901606. Special

volume on parallel computing.

Bolla:1985:IBI

[Bol85] Michael L. Bolla. Isomorphisms between infinite matrix rings. Linear Algebra and its Applications, 69(??): 239–247, August 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585900795.

Bone:1983:PFM

[Bon83] Terrence Bone. Positive feedback may sometimes promote stability. Linear Algebra and its Applications, 51 (??):143–151, June 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901556.

Bondar:1985:CAR

[Bon85] James V. Bondar. Convexity of acceptance regions of multivariate tests. Linear Algebra and its Applications, 67 (??):195–199, June 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585901958.

Bose:1988:RMS

[Bos88] N. K. Bose. Robust multivariate scattering Hurwitz interval polynomials. Linear Algebra and its Applications, 98(??):123–136, January 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901620.

Bousfield:1988:TSL

[Bou88] R. A. Bousfield. A theorem for symmetric n-linear forms. Linear Algebra and its Applications, 105(??): 183–193, July 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588900110.

Barnett:1981:CMO

[BP81] Sephen Barnett and Henry Power. Castle Matrix: 4 October 1980. Linear Algebra and its Applications, 41(??): 275–276, December 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581901038.

Beasley:1984:BRP

[BP84a] LeRoy B. Beasley and Norman J. Pullman. Boolean-rank-preserving operators and Boolean-rank-1 spaces. Linear Algebra and its Applications, 59(??):55–77, June 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901587.

Burkill:1984:LM

[BP84b] H. Burkill and Hazel Perfect. Leon Mirsky, 1918– 1983. Linear Algebra and its Applications, 61(??):1– 10, September 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958490017X.

Baksalary:1985:NMO

[BP85] Jerzy K. Baksalary and Friedrich Pukelsheim. A note on the matrix ordering of special C-matrices. Linear Algebra and its Applications, 70(??):263–267, October 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585900588.

Beasley:1986:FRP

[BP86] LeRoy B. Beasley and Norman J. Pullman. Fuzzy rank-preserving operators. Linear Algebra and its Applications, 73(??):197–211, January 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902405.

Beasley:1987:TRP

[BP87] LeRoy B. Beasley and Norman J. Pullman. Termrank, permanent, and rook-polynomial preservers. Linear Algebra and its Applications, 90(??):33–46, May 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587903028.

Beasley:1988:OPS

[BP88a] LeRoy B. Beasley and Norman J. Pullman. Operators that preserve semiring matrix functions. Linear Algebra and its Applications, 99(??):199–216, February 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901322.

Beasley:1988:SRV

[BP88b] LeRoy B. Beasley and Norman J. Pullman. Semiring rank versus column rank. Linear Algebra and its Applications, 101(??):33–48, April 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901413.

Brualdi:1988:MIL

[BPB88] Richard A. Brualdi, Vera S. Pless, and Janet S. Beissinger. On the MacWilliams identities for linear codes. Linear Algebra and its Applications, 107(??): 181–189, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588902443.

Proceedings of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Ballantine:1987:SCC

[BPR87] C. S. Ballantine, Stephen Pierce, and Leiba Rodman. Simultaneous congruence of convex compact sets of Hermitian matrices with constant rank. Linear Algebra and its Applications, 87(??):49–60, March 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587901571.

Bjorck:1980:P

[BPS80] Åke Björck, Robert J. Plemmons, and Hans Schneider. Preface. Linear Algebra and its Applications, 34(??): 1–2, December 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958090155X.

Baksalary:1989:SPM

[BPS89] Jerzy K. Baksalary, Friedrich Pukelsheim, and George P. H. Styan. Some properties of matrix partial orderings. Linear Algebra and its Applications, 119(??):57–85, July 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589900694.

Bebiano:1985:DST

[BQ85] Natália Bebiano and João Filipe Queiró. The determinant of the sum of two normal matrices with prescribed eigenvalues. Linear Algebra and its Applications, 71(??): 23–28, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585902319.

Bapat:1980:DPD

[BR80a] R. Bapat and T. E. S. Raghavan. On diagonal products of doubly stochastic matrices. Linear Algebra and its Applications, 31(??):71–75, June 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580902074.

Brualdi:1980:RMT

[BR80b] Richard A. Brualdi and Jeffrey A. Ross. On Ryser’s maximum term rank formula. Linear Algebra and its Applications, 29(??):33–38, February 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580902281.

Brualdi:1986:CMA

[BR86] Richard A. Brualdi and H. J. Ryser. Classes of matrices associated with the optimal assignment problem. Linear Algebra and its Applications, 81(??):1–17, September 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902454.

Ball:1987:GIS

[BR87] Joseph A. Ball and André C. M. Ran. Global inverse spectral problems for rational matrix functions. Linear Algebra and its Applications, 86(??):237–282, February 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587902977.

Bapat:1989:ETD

[BR89a] R. B. Bapat and T. E. S. Raghavan. An extension of a theorem of Darroch and Ratcliff in loglinear models and its application to scaling multidimensional matrices. Linear Algebra and its Applications, 114/115(??):705–715, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bixby:1989:SPT

[BR89b] Robert E. Bixby and Arvind Rajan. A short proof of the Truemper–Tseng theorem on max-flow min-cut matroids. Linear Algebra and its Applications, 114/115(??): 277–292, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bresinsky:1984:MFR

[Bre84] Henrik Bresinsky. Minimal free resolutions of monomial curves in p_{k}^{3}. Linear Algebra and its Applications, 59(??):

121–129, June 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901630.

Bremser:1986:GGS

[Bre86] Priscilla S. Bremser. A generalization of Gaussian sums to vector spaces over finite fields. Linear Algebra and its Applications, 81(??):35–45, September 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902478.

Brezinski:1988:OMS

[Bre88] Claude Brezinski. Other manifestations of the Schur complement. Linear Algebra and its Applications, 111(??): 231–247, December 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588900626.

Brockett:1983:LFS

[Bro83] R. W. Brockett. Linear feedback systems and the groups of Galois and Lie. Linear Algebra and its Applications, 50 (??):45–60, April 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583900538.

Broadie:1985:TAA

[Bro85a] Mark N. Broadie. A theorem about antiprisms. Linear Algebra and its Applications, 66(??):99–111, April 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585901272.

Browne:1985:IMI

[Bro85b] M. W. Browne. The information matrix for image factor analysis. Linear Algebra and its Applications, 70(??): 51–59, October 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585900412.

Brockett:1989:LSM

[Bro89a] R. W. Brockett. Least squares matching problems. Linear Algebra and its Applications, 122/123/124(??):761–777, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Browkin:1989:TSR

[Bro89b] Jerzy Browkin. A theorem of Sturm and K_{1} for rings of real continuous functions. Linear Algebra and its Applications, 113(??):1–6, February 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958990284X.

Brualdi:1980:MZO

[Bru80] Richard A. Brualdi. Matrices of zeros and ones with fixed row and column sum vectors. Linear Algebra and its Applications, 33(??):159–231, October 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901056.

Brualdi:1984:DSM

[Bru84a] Richard A. Brualdi. The doubly stochastic matrices of a vector majorization. Linear Algebra and its Applications, 61(??):141–154, September 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900272.

Brualdi:1984:RBM

[Bru84b] Richard A. Brualdi. Review of Elements of Matrix Theory, by M. L. Mehta. Linear Algebra and its Applications, 62(??):293–294, November 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958490106X.

Brualdi:1984:RBP

[Bru84c] Richard A. Brualdi. Review of Kronecker Products and Matrix Calculus with Applications, by Alexander Graham. Linear Algebra and its Applications, 60(??):199– 200, August 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584900806.

Bruning:1984:DHF

[Brü84d] Jochen Brüning. Diagonalization of holomorphic functions with values in W^{∗}-algebras. Linear Algebra and its Applications, 56(??):131–138, January 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901198.

Brualdi:1985:CVE

[Bru85] Richard A. Brualdi. Combinatorial verification of the elementary divisors of tensor products. Linear Algebra and its Applications, 71(??):31–47, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902332.

Brualdi:1988:RBS

[Bru88a] Richard A. Brualdi. Review of Proceedings of the Second International Tampere Conference in Statistics, edited by Tarmo Pukkila and Simo Puntanen. Linear Algebra and its Applications, 111(??):325–326, December 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588900687.

Brualdi:1988:SAD

[Bru88b] Richard A. Brualdi. Some applications of doubly stochastic matrices. Linear Algebra and its Applications, 107(??): 77–100, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958890239X.

Proceedings of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Brualdi:1989:RBR

[Bru89] Richard A. Brualdi. Review of Recent Results in the Theory of Graph Spectra by D. Cvetković, M. Doob, I. Gutman, and A. Torgašev. Linear Algebra and its Applications, 126(??):149–150, December 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589900128.

Brylawski:1982:FPF

[Bry82] Tom Brylawski. Finite prime-field characteristic sets for planar configurations. Linear Algebra and its Applications, 46(??):155–176, August 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582900337.

Berman:1981:MZL

[BS81] Abraham Berman and B. David Saunders. Matrices with zero line sums and maximal rank. Linear Algebra and its Applications, 40(??):229–235, October 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581901531.

Benson:1982:DSM

[BS82a] R. V. Benson and R. C. Shiflett. Doubly stochastic matrices over arbitrary vector spaces and the Birkhoff theorem. Linear Algebra and its Applications, 42(??):145– 158, February 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582901458.

Buoni:1982:SRT

[BS82b] J. J. Buoni and B. Subramanian. A Stein–Rosenberg theorem for rectangular matrices. Linear Algebra and its Applications, 48(??):177–190, December 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582901069.

Butzer:1983:EMS

[BS83] P. L. Butzer and R. L. Stens. The Euler–MacLaurin summation formula, the sampling theorem, and approximate integration over the real axis. Linear Algebra and its Applications, 52/53(??):141–155, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bapat:1985:MSP

[BS85a] R. B. Bapat and V. S. Sunder. On majorization and Schur products. Linear Algebra and its Applications, 72(??):107–117, December 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585901478.

Blokhuis:1985:RWV

[BS85b] A. Blokhuis and J. J. Seidel. Remark on Wielandt’s visibility theorem. Linear Algebra and its Applications, 71 (??):29–30, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379585902320.

Bapat:1986:EPP

[BS86a] R. B. Bapat and V. S. Sunder. An extremal property of the permanent and the determinant. Linear Algebra and its Applications, 76(??):153–163, April 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958690220X.

Browne:1986:ACM

[BS86b] M. W. Browne and A. Shapiro. The asymptotic covariance matrix of sample correlation coefficients under general conditions. Linear Algebra and its Applications, 82(??):169–176, October 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586901503.

Berman:1987:LFI

[BS87a] Abraham Berman and Ronald J. Stern. Linear feedback, irreducibility, and M-matrices. Linear Algebra and its Applications, 97(??):141–152, December 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587901467.

Brawley:1987:PWP

[BS87b] J. V. Brawley and George E. Schnibben. Polynomials which permute the matrices over a field. Linear Algebra and its Applications, 86(??):145–160, February 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587902916.

Brualdi:1987:MSR

[BS87c] Richard A. Brualdi and Ernie S. Solheid. On the minimum spectral radius of matrices of zeros and ones. Linear Algebra and its Applications, 85(??):81–100, January 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587902102.

Berenstein:1988:SDS

[BS88a] Carlos A. Berenstein and Daniele C. Struppa. Small degree solutions for the polynomial Bézout equation. Linear Algebra and its Applications, 98(??):41–55, January 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901590.

Brualdi:1988:Pa

[BS88b] Richard A. Brualdi and Hans Schneider. Preface. Linear Algebra and its Applications, 100(??):vii, March 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588903308.

Brualdi:1988:Pb

[BS88c] Richard A. Brualdi and Hans Schneider. Preface. Linear Algebra and its Applications, 107(??):1, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588902327.

Brualdi:1989:EOV

[BS89] Richard A. Brualdi and Hans Schneider. Errata: Omissions in Volume 100: “20 year index: Vol. 1–100, 1968/1988” [Linear Algebra Appl. 100 (1988), MR 89h:15001]. Linear Algebra and its Applications, 120 (??):271, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589903893.

See [?].

Barel:1981:NME

[BT81a] M. Barel and A. Tamir. Nested matrices and the existence of least majorized elements. Linear Algebra and its Applications, 38(??):65–72, June 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581900082.

Barker:1981:GCP

[BT81b] George Phillip Barker and Bit-Shun Tam. Graphs for cone preserving maps. Linear Algebra and its Applications, 37 (??):199–204, April 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581901804.

Barker:1984:GGS

[BTD84] G. P. Barker, B. S. Tam, and Norbol Davila. A geometric Gordan–Stiemke theorem. Linear Algebra and its Applications, 61(??):83–89, September 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900235.

Beasley:1984:USU

[BU84] LeRoy Beasley and E. E. Underwood. The 1983 Utah State University Department of Mathematics Conference, 9, 10, and 11 June 1983, Logan, Utah. Linear Algebra and its Applications, 59(??):183–211, June 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901678.

Beasley:1988:USU

[BU88] LeRoy Beasley and E. E. Underwood. The 1987 Utah State University Department of Mathematics Conference 29, 30, and 31 January 1987, Logan, Utah. Linear Algebra and its Applications, 104(??):165–249, June 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588903151.

Bunch:1987:WSS

[Bun87] James R. Bunch. The weak and strong stability of algorithms in numerical linear algebra. Linear Algebra and its Applications, 88/89(??):49–66, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Burlando:1986:TSB

[Bur86] Laura Burlando. On two subsets of a Banach algebra that are related to the continuity of spectrum and spectral radius. Linear Algebra and its Applications, 84(??):251–269, December 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586903186. Proceedings

of the symposium on operator theory (Athens, 1985).

Bush:1980:HE

[Bus80] K. A. Bush. On Hadamard embeddability. Linear Algebra and its Applications, 29(??):39–52, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580902293.

Bushell:1986:CHM

[Bus86] P. J. Bushell. The Cayley–Hilbert metric and positive operators. Linear Algebra and its Applications, 84(??): 271–280, December 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586903198.

Proceedings of the symposium on operator theory (Athens, 1985).

Butcher:1988:CMR

[But88] J. C. Butcher. On a class of matrices with real eigenvalues. Linear Algebra and its Applications, 103(??):1–12, May 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588902170.

Byrd:1980:GIP

[BV80] Ken Byrd and Theresa P. Vaughan. A group of integral points in a matrix parallelepiped. Linear Algebra and its Applications, 30(??):155–166, April 1980.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580901913.

Beelen:1988:IAC

[BV88] Th. Beelen and P. Van Dooren. An improved algorithm for the computation of Kronecker’s canonical form of a singular pencil. Linear Algebra and its Applications, 105 (??):9–65, July 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588900031.

Bru:1989:ICL

[BV89] R. Bru and J. Vitória. International Conference on Linear Algebra and Applications, 28–30 September 1987, Universidad Politecnica de Valencia, Valencia, Spain. Linear Algebra and its Applications, 121(??):537–710, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589907210.

Blokhuis:1989:UVN

[BWT89] Aart Blokhuis, Henny Wilbrink, and Gunter Teumer. Unit vectors with nonnegative inner products and Kruskal’s conjecture. Linear Algebra and its Applications, 117(??):7–9, May 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589905429.

Ballantine:1980:CCM

[BY80] C. S. Ballantine and E. L. Yip. Congruence and conjunctivity of matrices. Linear Algebra and its Applications, 32(??):159–198, August 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580900154.

Byers:1987:SAR

[Bye87a] Ralph Byers. Solving the algebraic Riccati equation with the matrix sign function. Linear Algebra and its Applications, 85(??):267–279, January 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587902229.

Byers:1987:MIE

[Bye87b] W. Byers. Matrices induced by endomorphisms of finite sets. Linear Algebra and its Applications, 93(??): 81–92, August 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0024379587903132.

Byrnes:1983:THH

[Byr83] Christopher I. Byrnes. On a theorem of Hermite and Hurwitz. Linear Algebra and its Applications, 50(??): 61–101, April 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958390054X.

Beitia:1989:FMP

[BZ89] Ma. Asunción Beitia and Ion Zaballa. Factorization of the matrix polynomial A(λ) = A_{0}λ^{t} + A_{1}λ^{t−1} + ⋯ + A_{t−1}λ+A_{t}. Linear Algebra and its Applications, 121(??): 423–432, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589907143.

Linear algebra and applications (Valencia, 1987).

Caflisch:1981:IPT

[Caf81] Russel E. Caflisch. An inverse problem for Toeplitz matrices and the synthesis of discrete transmission lines. Linear Algebra and its Applications, 38(??):207–225, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900215.

Cain:1980:IT

[Cai80] Bryan E. Cain. Inertia theory. Linear Algebra and its Applications, 30(??):211–240, April 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901962. See correc-

tions [Cai82].

Cain:1981:IHMb

[Cai81] Bryan E. Cain. The inertia of a Hermitian matrix having prescribed diagonal blocks. Linear Algebra and its Applications, 37(??):173–180, April 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581901762.

Cain:1982:CIT

[Cai82] B. E. Cain. Corrections to: “Inertia theory” [Linear Algebra Appl. 30 (1980), 211–240, MR 82c:15001]. Linear Algebra and its Applications, 42(??):285–286, February 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582901562.  See

[Cai80].

Cain:1984:ISM

[Cai84] Bryan E. Cain. Inside the D-stable matrices. Linear Algebra and its Applications, 56(??):237–243, January 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901290.

Calderbank:1980:GMC

[Cal80] Robert Calderbank. A good method of combining codes. Linear Algebra and its Applications, 32(??):115– 124, August 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580900117.

Calogero:1982:IMC

[Cal82] F. Calogero. Isospectral matrices and classical polynomials. Linear Algebra and its Applications, 44(??): 55–60, April 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582900040.

Campbell:1981:RBU

[Cam81] Stephen L. Campbell. Review of Linpack User’s Guide by J. J. Dongarra, C. B. Moler, J. R. Bunch, and G. W. Stewart. Linear Algebra and its Applications, 39(??):

293–294, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581903116.

Campbell:1984:RBS

[Cam84a] Stephen L. Campbell. Review of A Short Introduction to Perturbation Theory for Linear Operators, by Tosio Kato. Linear Algebra and its Applications, 56(??):293– 294, January 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901332.

Campbell:1984:RBMa

[Cam84b] Stephen L. Campbell. Review of Advanced Matrix Theory for Scientists and Engineers, by A. S. Deif. Linear Algebra and its Applications, 56(??):291–292, January 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901320.

Campbell:1984:RBMb

[Cam84c] Stephen L. Campbell. Review of Applied Multidimensional Systems Theory, by N. K. Bose, and Multivariable Feedback Systems, by F. M. Callier and C. A. Desoer. Linear Algebra and its Applications, 57(??):287– 289, February 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901940.

Campbell:1984:RBP

[Cam84d] Stephen L. Campbell. Review of Matrix Pencils: edited by B. Kågström and A. Ruhe. Linear Algebra and its Applications, 62(??):287–288, November 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901046.

Campbell:1984:RBL

[Cam84e] Stephen L. Campbell. Review of Polynomials and Linear Control Systems, by Stephen Barnett. Linear Algebra and its Applications, 61(??):289–290, September 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584900387.

Cameron:1989:PIM

[Cam89] Peter J. Cameron. A problem on integer matrices. Linear Algebra and its Applications, 114/115(??):199–206, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cao:1987:DMS

[Cao87a] Zhi Hao Cao. On a deflation method for the symmetric generalized eigenvalue problem. Linear Algebra and its Applications, 92(??):187–196, July 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587902564.

Cao:1987:REB

[Cao87b] Zhi Hao Cao. Residual error bounds of generalized eigenvalue systems. Linear Algebra and its Applications, 93(??):113–120, August 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/S0024379587903168.

Carlson:1983:MIT

[Car83] David Carlson. Minimax and interlacing theorems for matrices. Linear Algebra and its Applications, 54(??): 153–172, October 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583902112.

Carlson:1984:CIS

[Car84] David Carlson. Controllability, inertia, and stability for tridiagonal matrices. Linear Algebra and its Applications, 56(??):207–220, January 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901265.

Carlson:1985:RBM

[Car85] Roger Carlson. Review of Applied Matrix Algebra in the Statistical Sciences, by Alexander Basilevsky. Linear Algebra and its Applications, 64(??):273–277, January 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902836.

Carlson:1986:WSC

[Car86] David Carlson. What are Schur complements, anyway? Linear Algebra and its Applications, 74(??):257– 275, February 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901278.

Carlson:1988:NCM

[Car88] David Carlson. Nonsingularity criteria for matrices involving combinatorial considerations. Linear Algebra and its Applications, 107(??):41–56, August 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902364. Proceedings

of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Chabrillac:1984:DSQ

[CC84] Yves Chabrillac and J.-P. Crouzeix. Definiteness and semidefiniteness of quadratic forms revisited. Linear Algebra and its Applications, 63(??):283–292, December 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901502.

Chvatal:1989:CPP

[CCH89] V. Chvátal, W. Cook, and M. Hartmann. On cuttingplane proofs in combinatorial optimization. Linear Algebra and its Applications, 114/115(??):455–499, 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chang:1983:CFN

[CD83] Geng Zhe Chang and Philip J. Davis. A circulant formulation of the Napoleon–Douglas–Neumann theorem. Linear Algebra and its Applications, 54(??):87–95, October 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583902070.

Consonni:1985:DBA

[CD85] G. Consonni and A. P. Dawid. Decomposition and Bayesian analysis of invariant normal linear models. Linear Algebra and its Applications, 70(??):21–49, October 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585900400.

Cavaretta:1981:FTB

[CDMS81] A. S. Cavaretta, Jr., W. A. Dahmen, C. A. Micchelli, and P. W. Smith. A factorization theorem for banded matrices. Linear Algebra and its Applications, 39(??):229–245, August 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581903062.

Censor:1982:NML

[CE82] Yair Censor and Tommy Elfving. New methods for linear inequalities. Linear Algebra and its Applications, 42(??): 199–211, February 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582901495.

Christensen:1986:PDD

[CF86] Jens Peter Reus Christensen and Pal Fischer. Positive definite doubly stochastic matrices and extreme points. Linear Algebra and its Applications, 82(??):123– 132, October 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958690145X.

Cohen:1982:EIP

[CFKK82] Joel E. Cohen, Shmuel Friedland, Tosio Kato, and Frank P. Kelly. Eigenvalue inequalities for products of matrix exponentials. Linear Algebra and its Applications, 45(??):55–95, June 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582902117.

Cline:1980:DIR

[CG80] Randall E. Cline and T. N. E. Greville. A Drazin inverse for rectangular matrices. Linear Algebra and its Applications, 29(??):53–62, February 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958090230X.

Caron:1986:FPS

[CG86] R. J. Caron and N. I. M. Gould. Finding a positive semidefinite interval for a parametric matrix. Linear Algebra and its Applications, 76(??):19–29, April 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902119.

Clements:1989:SFH

[CG89] D. J. Clements and K. Glover. Spectral factorization via Hermitian pencils. Linear Algebra and its Applications, 122/123/124(??):797–846, 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Conway:1980:FDP

[CH80] J. B. Conway and P. R. Halmos. Finite-dimensional points of continuity of Lat. Linear Algebra and its Applications, 31(??):93–102, June 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580902104.

Carlson:1983:CAD

[CH83a] David Carlson and Emilie V. Haynsworth. Complementable and almost definite matrices. Linear Algebra and its Applications, 52/53(??):157–176, 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chari:1983:DIP

[CH83b] Bhagya Chari and Harry Hochstadt. A direct and inverse problem for a Hill’s matrix with double eigenvalues. Linear Algebra and its Applications, 52/53(??):177–197, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Carlson:1989:NCG

[CH89] David Carlson and Daniel Hershkowitz. Nonsingularity criteria for general combinatorially symmetric matrices. Linear Algebra and its Applications, 114/115(??):399– 416, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chalkley:1981:IAG

[Cha81] Roger Chalkley. Information about group matrices. Linear Algebra and its Applications, 38(??):121–133, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900136.

Chao:1982:SSP

[Cha82] Chong Yun Chao. On some sets of permutation matrices. Linear Algebra and its Applications, 45(??):131–134, June 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582902166.

Chang:1985:GTI

[Cha85] Sin Chung Chang. Generalizations of two inequalities involving Hermitian forms. Linear Algebra and its Applications, 65(??):179–187, February 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900965.

Chan:1987:RRF

[Cha87a] Tony F. Chan. Rank revealing QR factorizations. Linear Algebra and its Applications, 88/89(??):67–82, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chang:1987:PCC

[Cha87b] Sin Chung Chang. On the product of convex combinations and the Kantorovich inequality. Linear Algebra and its Applications, 93(??):9–38, August 1987.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0024379587903090.

Chao:1988:RSC

[Cha88] Chong-Yun Chao. A remark on symmetric circulant matrices. Linear Algebra and its Applications, 103(??): 133–148, May 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958890225X.

Chen:1983:IDL

[Che83a] Dao Qi Chen. An invariant deflation for lower banded matrices. Linear Algebra and its Applications, 49(??): 237–245, February 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901064.

Chen:1983:NLB

[Che83b] Gong Ning Chen. Note on lower bounds for the rank of a matrix. Linear Algebra and its Applications, 55(??): 125–132, December 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901714.

Chen:1983:SPS

[Che83c] Gong Ning Chen. On some properties of singular matrices. Linear Algebra and its Applications, 49(??):137– 151, February 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958390099X.

Chen:1983:SGD

[Che83d] Gong Ning Chen. Some generalizations of diagonal dominance associated with G-functions. Linear Algebra and its Applications, 55(??):169–180, December 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583901751.

Chen:1984:NRF

[Che84] Ji Cheng Chen. The nonnegative rank factorizations of nonnegative matrices. Linear Algebra and its Applications, 62(??):207–217, November 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958490096X.

Cheng:1987:DQW

[Che87] Sui-Sun Cheng. Discrete quadratic Wirtinger’s inequalities. Linear Algebra and its Applications, 85(??):57– 73, January 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587902084.

Chikuse:1986:MMC

[Chi86] Yasuko Chikuse. Multivariate Meixner classes of invariant distributions. Linear Algebra and its Applications, 82(??): 177–200, October 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901515.

Chillag:1987:GCC

[Chi87] David Chillag. Generalized circulants and class functions of finite groups. Linear Algebra and its Applications, 93 (??):191–208, August 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0024379587903259.

Chillag:1988:GCC

[Chi88] David Chillag. Generalized circulants and class functions of finite groups. II. Linear Algebra and its Applications, 108(??):199–212, September 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901887.

Choudhury:1988:QPD

[CHP88] Dipa Choudhury, Roger A. Horn, and Stephen J. Pierce. Quasi-positive definite operators and matrices. Linear Algebra and its Applications, 99(??):161–176, February 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901309.

Chu:1984:SAH

[Chu84] Moody T. Chu. A simple application of the homotopy method to symmetric eigenvalue problems. Linear Algebra and its Applications, 59(??):85–90, June 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901605.

Chu:1986:CAG

[Chu86a] Moody Chu. A continuous approximation to the generalized Schur decomposition. Linear Algebra and its Applications, 78(??):119–132, June 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586900194.

Chu:1986:DEA

[Chu86b] Moody T. Chu. A differential equation approach to the singular value decomposition of bidiagonal matrices. Linear Algebra and its Applications, 80(??):71–79, August 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902788.

Chu:1988:NHM

[Chu88] Moody T. Chu. A note on the homotopy method for linear algebraic eigenvalue problems. Linear Algebra and its Applications, 105(??):225–236, July 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588900158.

Chu:1989:SAI

[Chu89] W. C. Chu. Some algebraic identities concerning determinants and permanents. Linear Algebra and its Applications, 116(??):35–40, April 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589903960.

Choi:1987:LMP

[CJR87] M. D. Choi, A. A. Jafarian, and H. Radjavi. Linear maps preserving commutativity. Linear Algebra and its Applications, 87(??):227–241, March 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587901698.

Chun:1989:CPG

[CK89] J. Chun and T. Kailath. A constructive proof of the Gohberg–Semencul formula. Linear Algebra and its Applications, 121(??):475–489, August 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589907179. Linear alge-

bra and applications (Valencia, 1987).

Chung:1985:GCS

[CKN85] J. K. Chung, Pl. Kannappan, and C. T. Ng. A generalization of the cosine-sine functional equation on groups. Linear Algebra and its Applications, 66(??):259–277, April 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585901375.

Chan:1981:CIG

[CL81] G. H. Chan and M. H. Lim. A conjecture involving generalized matrix functions. Linear Algebra and its Applications, 36(??):157–163, March 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581902287.

Cameron:1982:TDO

[CL82a] P. J. Cameron and R. A. Liebler. Tactical decompositions and orbits of projective groups. Linear Algebra and its Applications, 46(??):91–102, August 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582900295.

Carlson:1982:EMI

[CL82b] David Carlson and Raphael Loewy. On equality of maps induced by alternate products. Linear Algebra and its Applications, 47(??):89–110, October 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582902270.

Chan:1982:LTS

[CL82c] G. H. Chan and M. H. Lim. Linear transformations on symmetric matrices that preserve commutativity. Linear Algebra and its Applications, 47(??):11–22, October 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582902221.

Cheng:1986:CRD

[CL86] Sui Sun Cheng and Tzon Tzer Lu. Convex regular domains of tridiagonal matrices. Linear Algebra and its Applications, 79(??):103–125, July 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902958.

Chen:1989:CFM

[CL89] J. C. Chen and M. X. Lai. On a conjecture by Fiedler, Markham, and Neumann. Linear Algebra and its Applications, 117(??):99–103, May 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958990551X.

Clark:1988:SFH

[Cla88] Kenneth D. Clark. A structural form for higherindex semistate equations. I. theory and applications to circuit and control theory. Linear Algebra and its Applications, 98(??):169–197, January 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901644.

Cline:1981:NEM

[Cli81] Randall E. Cline. Note on an extension of the Moore– Penrose inverse. Linear Algebra and its Applications, 40 (??):19–23, October 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379581901373.

Cheng:1987:RSO

[CLLW87] Sui-Sun Cheng, Horng Jaan Li, Tzon-Tzer Lu, and Shu Hei Wu. Regular and singular orthants of tridiagonal matrices. Linear Algebra and its Applications, 94(??):181–191, September 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587900887.

Curtis:1982:DAP

[CLP82] W. D. Curtis, J. David Logan, and W. A. Parker. Dimensional analysis and the pi theorem. Linear Algebra and its Applications, 47(??):117–126, October 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582902294.

Coxson:1987:MPS

[CLS87] Pamela G. Coxson, Loren C. Larson, and Hans Schneider. Monomial patterns in the sequence A^{k}b. Linear Algebra and its Applications, 94(??):89–101, September 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587900802.

Chan:1987:LPM

[CLT87] Gin Hor Chan, Ming Huat Lim, and Kok-Keong Tan. Linear preservers on matrices. Linear Algebra and its Applications, 93(??):67–80, August 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/S0024379587903120.

Cain:1981:IHMa

[CM81a] Bryan E. Cain and E. Marques de Sá. The inertia of a Hermitian matrix having prescribed complementary principal submatrices. Linear Algebra and its Applications, 37(??):161–171, April 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581901750.

Craven:1981:LPM

[CM81b] B. D. Craven and B. Mond. Linear programming with matrix variables. Linear Algebra and its Applications, 38 (??):73–80, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581900094.

Chatelin:1982:AAS

[CM82] Françoise Chatelin and Willard L. Miranker. Acceleration by aggregation of successive approximation methods. Linear Algebra and its Applications, 43(??):17–47, March 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582902427.

Crawford:1983:FPD

[CM83] C. R. Crawford and Yiu Sang Moon. Finding a positive definite linear combination of two Hermitian matrices. Linear Algebra and its Applications, 51(??):37–48, June 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583901489.

Cline:1984:ELN

[CM84] Randall E. Cline and Robert M. McConnel. Extensions of the Levine–Nahikian method for constructing involutory matrices. Linear Algebra and its Applications, 57(??):247–270, February 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901915.

Chan:1985:EME

[CM85] N. N. Chan and T. K. Mak. Estimation in multivariate errors-in-variables models. Linear Algebra and its Applications, 70(??):197–207, October 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900527.

Cohen:1989:PTC

[CMO89] Joel E. Cohen, Ezio Marchi, and Jorge A. Oviedo. Perturbation theory of completely mixed bimatrix games. Linear Algebra and its Applications, 114/115(??):169–180, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Campbell:1981:RIS

[CMR81] Stephen L. Campbell, Carl D. Meyer, Jr., and Nicholas J. Rose. Results of an informal survey on the use of linear algebra in industry and government. Linear Algebra and its Applications, 38(??):289–294, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900276.

Carlson:1986:EH

[CMU86] David Carlson, Thomas L. Markham, and Frank Uhlig. Emilie Haynsworth, 1916–1985. Linear Algebra and its Applications, 75(??):269–276, March 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586901941.

Cobham:1986:SAL

[Cob86] Alan Cobham. Stochastic automata with large state spaces and low rank. Linear Algebra and its Applications, 75(??):57–66, March 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379586901813.

Cohen:1986:PTC

[Coh86] Joel E. Cohen. Perturbation theory of completely mixed matrix games. Linear Algebra and its Applications, 79 (??):153–162, July 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586902971.

Cohen:1987:PSK

[Coh87] Nir Cohen. Polynomial systems and Kronecker invariants. Linear Algebra and its Applications, 87(??):257– 265, March 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587901716.

Cohen:1988:SIM

[Coh88] Joel E. Cohen. Spectral inequalities for matrix exponentials. Linear Algebra and its Applications, 111(??): 25–28, December 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588900481.

Conway:1981:FDP

[Con81] John B. Conway. Finite-dimensional points of continuity of Gen and Com. Linear Algebra and its Applications, 35(??):121–127, February 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958190269X.

Constantine:1985:LBS

[Con85] Gregory Constantine. Lower bounds on the spectra of symmetric matrices with nonnegative entries. Linear Algebra and its Applications, 65(??):171–178, February 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585900953.

Coppel:1981:LSS

[Cop81] W. A. Coppel. Linear systems: some algebraic aspects. Linear Algebra and its Applications, 40(??):257– 273, October 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581901579.

Coppel:1984:RBS

[Cop84] W. A. Coppel. Review of Linear Systems and Operators in Hilbert Space, by Paul A. Fuhrmann and matrix polynomials, by I. Gohberg, P. Lancaster, and L. Rodman. Linear Algebra and its Applications, 59(??):213–214, June 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958490168X.

Cordovil:1987:PPE

[Cor87] Raul Cordovil. Polarity and point extensions in oriented matroids. Linear Algebra and its Applications, 90 (??):15–31, May 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587903016.

Costa:1989:NBA

[Cos89] Roberto Costa. A note on Bernstein algebras. Linear Algebra and its Applications, 112(??):195–205, January 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589905958.

Campbell:1981:CNR

[CP81] Stephen L. Campbell and George D. Poole. Computing nonnegative rank factorizations. Linear Algebra and its Applications, 35(??):175–182, February 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958190272X.

Carlson:1985:CEQ

[CP85] David Carlson and Stephen Pierce. Common eigenvectors and quasicommutativity of sets of simultaneously triangularizable matrices. Linear Algebra and its Applications, 71(??):49–55, November 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902344.

Coullard:1989:CCP

[CP89] Collette R. Coullard and William R. Pulleyblank. On cycle cones and polyhedra. Linear Algebra and its Applications, 114/115(??):613–640, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cottle:1989:SML

[CPV89] R. W. Cottle, J.-S. Pang, and V. Venkateswaran. Sufficient matrices and the linear complementarity problem. Linear Algebra and its Applications, 114/115(??):231– 249, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chan:1986:CNS

[CR86] Tony F. Chan and Diana C. Resasco. On the condition of nearly singular matrices under rank-1 perturbations. Linear Algebra and its Applications, 76(??):223–232, April 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902247.

Critchley:1988:CLM

[Cri88] Frank Critchley. On certain linear mappings between inner-product and squared-distance matrices. Linear Algebra and its Applications, 105(??):91–107, July 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588900067.

Crone:1981:SOA

[Cro81] Lawrence Crone. Second order adjoint matrix equations. Linear Algebra and its Applications, 39(??):61–71, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581902901.

Cohen:1982:SNM

[CS82] Joel E. Cohen and Peter H. Sellers. Sets of nonnegative matrices with positive inhomogeneous products. Linear Algebra and its Applications, 47(??):185–192, October 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582902373.

Courtois:1985:PPF

[CS85]  P.-J. Courtois and P. Semal.  On polyhedra of

Perron–Frobenius eigenvectors. Linear Algebra and its Applications, 65(??):157–170, February 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900941.

Cao:1986:NIH

[CS86a] Wei Lu Cao and William J. Stewart. A note on inverses of Hessenberg-like matrices. Linear Algebra and its Applications, 76(??):233–240, April 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902259.

Courtois:1986:BIA

[CS86b] P. J. Courtois and P. Semal. Block iterative algorithms for stochastic matrices. Linear Algebra and its Applications, 76(??):59–70, April 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902144.

Conner:1987:SCS

[CS87a] Luther T. Conner, Jr. and David P. Stanford. The structure of the controllable set for multimodal systems. Linear Algebra and its Applications, 95(??):171–180, October 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587900334.

Cordovil:1987:DMP

[CS87b] Raul Cordovil and Ilda P. Silva. Determining a matroid polytope by non-radon partitions. Linear Algebra and its Applications, 94(??):55–60, September 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587900772.

Coxson:1987:PIR

[CS87c] Pamela G. Coxson and Helene Shapiro. Positive input reachability and controllability of positive systems. Linear Algebra and its Applications, 94(??):35–53, September 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587900760.

Camina:1989:IAF

[CS89a] Alan Camina and Johannes Siemons. Intertwining automorphisms in finite incidence structures. Linear Algebra and its Applications, 117(??):25–34, May 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589905454.

Chen:1989:JRM

[CS89b] Yong Chuan Chen and Aditya Shastri. On joint realization of (0, 1) matrices. Linear Algebra and its Applications, 112(??):75–85, January 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589905909.

Cunningham:1982:SCB

[Cun82] William H. Cunningham. Separating cocircuits in binary matroids. Linear Algebra and its Applications, 43 (??):69–86, March 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582902440.

Charlier:1987:KSA

[CV87] J.-P. Charlier and P. Van Dooren. On Kogbetliantz’s SVD algorithm in the presence of clusters. Linear Algebra and its Applications, 95(??):135–160, October 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587900310.

Cvetkovic:1988:CTG

[Cve88] Dragoš Cvetković. Constructing trees with given eigenvalues and angles. Linear Algebra and its Applications, 105(??):1–8, July 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958890002X.

Cottle:1981:SCS

[CvRS81] Richard W. Cottle, Rabe von Randow, and Richard E. Stone. On spherically convex sets and Q-matrices. Linear Algebra and its Applications, 41(??):73–80, December 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900884.

Cullum:1980:LPI

[CW80] Jane Cullum and Ralph A. Willoughby. The Lanczos phenomenon—an interpretation based upon conjugate gradient optimization. Linear Algebra and its Applications, 29(??):63–90, February 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580902311.

Cabrera:1985:AHR

[CW85] Javier F. Cabrera and Geoffrey S. Watson. The algebra of hyperboloids of revolution. Linear Algebra and its Applications, 70(??):115–130, October 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900473.

Cybenko:1987:EFL

[Cyb87] George Cybenko. An explicit formula for Lánczos polynomials. Linear Algebra and its Applications, 88/89(??):99– 115, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chao:1984:GCB

[CZ84] Chong Yun Chao and Mou Cheng Zhang. On generalized circulants over a Boolean algebra. Linear Algebra and its Applications, 62(??):195–206, November 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584900958.

Cao:1986:CNM

[CZ86] Zhi Hao Cao and Yin Zhang. The contraction number of a multigrid method with mesh ratio two for solving model problems. Linear Algebra and its Applications, 79 (??):23–32, July 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586902892.

DiasdaSilva:1987:LGA

[Dad87] J. A. Dias da Silva and M. a da Purificação Coelho. Linear groups associated with elements of a group algebra. Linear Algebra and its Applications, 94(??):165– 179, September 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587900875.

Dahlquist:1983:MMM

[Dah83] Germund Dahlquist. On matrix majorants and minorants, with applications to differential equations. Linear Algebra and its Applications, 52/53(??):199–216, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Durand-Alegria:1989:ZAT

[DALSP89] A. I. Durand-Alegría, J. López-Sánchez, and A. Pérez de Vargas. Zygotic algebra for two-linked loci with sexually different recombination and mutation rates. Linear Algebra and its Applications, 121(??):385–399, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589907118. Linear

algebra and applications (Valencia, 1987).

Damhus:1980:CCG

[Dam80] Ture Damhus. Complex conjugation of group representations by inner automorphisms. Linear Algebra and its Applications, 32(??):125–135, August 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580900129.

Dancis:1986:QFS

[Dan86] Jerome Dancis. A quantitative formulation of Sylvester’s law of inertia. III. Linear Algebra and its Applications, 80(??):141–158, August 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/002437958690282X.

Dancis:1987:PIH

[Dan87] Jerome Dancis. The possible inertias for a Hermitian matrix and its principal submatrices. Linear Algebra and its Applications, 85(??):121–151, January 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587902126.

Dancis:1988:ISM

[Dan88] Jerome Dancis. On the inertias of symmetric matrices and bounded selfadjoint operators. Linear Algebra and its Applications, 105(??):67–75, July 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588900043.

Dantzig:1989:MPD

[Dan89] George B. Dantzig. Making progress during a stall in the simplex algorithm. Linear Algebra and its Applications, 114/115(??):251–259, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Datta:1984:GHM

[Dat84] B. N. Datta. Generalized Hankel matrices of Markov parameters and their applications to control problems. Linear Algebra and its Applications, 62(??):139–154, November 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584900910.

Datta:1986:RBI

[Dat86] B. N. Datta. Review of Matrices and Indefinite Scalar Products, by I. Gohberg, P. Lancaster, and L. Rodman. Linear Algebra and its Applications, 74(??):277– 279, February 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958690128X.

Datta:1988:MEA

[Dat88] Karabi Datta. The matrix equation XA − BX = R and its applications. Linear Algebra and its Applications, 109(??):91–105, October 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902005.

Datta:1989:PLS

[Dat89] Biswa Nath Datta. Parallel and large-scale matrix computations in control: some ideas. Linear Algebra and its Applications, 121(??):243–264, August 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589907052. Linear alge-

bra and applications (Valencia, 1987).

Davis:1980:EKI

[Dav80] Chandler Davis. Extending the Kantorovič inequality to normal matrices. Linear Algebra and its Applications, 31 (??):173–177, June 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580902177.

Dax:1988:LPL

[Dax88] Achiya Dax. Linear programming via least squares. Linear Algebra and its Applications, 111(??):313–324, December 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588900675.

deBruijn:1980:RHM

[dB80] N. G. de Bruijn. Remarks on Hermitian Matrices. Linear Algebra and its Applications, 32(??):201–208, August 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580900178.

deBruin:1984:SEF

[dB84] Marcel G. de Bruin. Some explicit formulae in simultaneous Padé approximation. Linear Algebra and its Applications, 63(??):271–281, December 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901496.

deBoor:1982:SIB

[dBJP82] C. de Boor, Rong Qing Jia, and A. Pinkus. Structure of invertible (bi)infinite totally positive matrices. Linear Algebra and its Applications, 47(??):41–55, October 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582902257.

deBoor:1982:ATP

[dBP82] Carl de Boor and Allan Pinkus. The approximation of a totally positive band matrix by a strictly banded totally positive one. Linear Algebra and its Applications, 42(??):81–98, February 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582901392.

deBoor:1986:FSO

[dBS86] Carl de Boor and Edward B. Saff. Finite sequences of orthogonal polynomials connected by a Jacobi matrix. Linear Algebra and its Applications, 75(??):43–55, March 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586901801.

deChela:1987:GIU

[dC87] Delia Flores de Chela. Generalized inverses under the ℓ_{1}- norm. Linear Algebra and its Applications, 94(??):237– 261, September 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587900942.

deCaen:1981:PSB

[dCG81] D. de Caen and D. A. Gregory. Primes in the semigroup of Boolean matrices. Linear Algebra and its Applications, 37(??):119–134, April 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581901725.

Dancis:1987:ITE

[DD87a] Jerome Dancis and Chandler Davis. An interlacing theorem for eigenvalues of selfadjoint operators. Linear Algebra and its Applications, 88/89(??):117–122, 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Datta:1987:MEA

[DD87b] Biswa Nath Datta and Karabi Datta. The matrix equation XA = A^{T} X and an associated algorithm for solving the inertia and stability problems. Linear Algebra and its Applications, 97(??):103–119, December 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587901431.

Dedieu:1988:MHI

[Ded88] Jean-Pierre Dedieu. Matrix homographic iterations and bounds for the inverses of certain band matrices. Linear Algebra and its Applications, 111(??):29–42, December 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588900493.

Donald:1984:GTUa

[DEHS84a] John Donald, John Elwin, Richard Hager, and Peter Salamon. A graph theoretic upper bound on the permanent of a nonnegative integer matrix. I. Linear Algebra and its Applications, 61(??):187–198, September 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584900302.

Donald:1984:GTUb

[DEHS84b] John Donald, John Elwin, Richard Hager, and Peter Salamon. A graph theoretic upper bound on the permanent of a nonnegative integer matrix. II. the extremal case. Linear Algebra and its Applications, 61(??):199–218, September 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900314.

Demko:1982:SIP

[Dem82] Stephen Demko. Surjectivity and invertibility properties of totally positive matrices. Linear Algebra and its Applications, 45(??):13–20, June 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582902063.

Demko:1986:CNR

[Dem86a] Stephen Demko. Condition numbers of rectangular systems and bounds for generalized inverses. Linear Algebra and its Applications, 78(??):199–206, June 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586900248.

Demmel:1986:CSE

[Dem86b] James Weldon Demmel. Computing stable eigendecompositions of matrices. Linear Algebra and its Applications, 79(??):163–193, July 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379586902983.

Demeure:1989:FFV

[Dem89] Cédric J. Demeure. Fast QR factorization of Vandermonde matrices. Linear Algebra and its Applications, 122/123/124(??):165–194, 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Denman:1981:RRM

[Den81] Eugene D. Denman. Roots of real matrices. Linear Algebra and its Applications, 36(??):133–139, March 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581902263.

Deschamps:1988:NIC

[Des88] Philippe J. Deschamps. A note on isomorphic characterizations of the dispersion matrix in error-component models. Linear Algebra and its Applications, 111(??): 147–150, December 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588900560.

Deutsch:1980:CGM

[Deu80] Emeric Deutsch. On certain generalized matrix norms. Linear Algebra and its Applications, 29(??):113–129, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580902335.

Deutsch:1981:NSM

[Deu81] Emeric Deutsch. Nonnegative spectraloid matrices. Linear Algebra and its Applications, 39(??):149–154, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581902998.

Das:1980:ONR

[DEW80] Kiran Chandra Das and Mary Embry-Wardrop. Orthogonality and the numerical range. II. Linear Algebra and its Applications, 30(??):63–67, April 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901810.

DiasdaSilva:1987:VC

[DF87] J. A. Dias da Silva and Amélia Fonseca. Vector colorability. Linear Algebra and its Applications, 94(??): 67–88, September 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587900796.

Dym:1981:EBM

[DG81] Harry Dym and Israel Gohberg. Extensions of band matrices with band inverses. Linear Algebra and its Applications, 36(??):1–24, March 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581902159.

deGroen:1982:CAL

[dG82] Pieter P. N. de Groen. The construction of an accurate lower bound for the real parts of the eigenvalues of an M-matrix. Linear Algebra and its Applications, 43(??): 161–166, March 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582902506.

Dow:1987:PDM

[DG87] Stephen J. Dow and Peter M. Gibson. Permanents of d-dimensional matrices. Linear Algebra and its Applications, 90(??):133–145, May 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587903119.

deGroote:1983:CAM

[dGH83] Hans F. de Groote and Joos Heintz. Commutative algebras of minimal rank. Linear Algebra and its Applications, 55(??):37–68, December 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583901659.

Delsarte:1980:MMI

[DGK80] Ph. Delsarte, Y. Genin, and Y. Kamp. A method of matrix inverse triangular decomposition based on contiguous principal submatrices. Linear Algebra and its Applications, 31(??):199–212, June 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580902190.

Delsarte:1983:TEA

[DGK83] P. Delsarte, Y. Genin, and Y. Kamp. On the Toeplitz embedding of an arbitrary matrix. Linear Algebra and its Applications, 51(??):97–119, June 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583901519.

Dongarra:1984:EPH

[DGKW84] J. J. Dongarra, J. R. Gabriel, D. D. Koelling, and J. H. Wilkinson. The eigenvalue problem for Hermitian matrices with time reversal symmetry. Linear Algebra and its Applications, 60(??):27–42, August 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584900685.

deHoog:1987:NAS

[dH87] Frank de Hoog. A new algorithm for solving Toeplitz systems of equations. Linear Algebra and its Applications, 88/89(??):123–138, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Deutsch:1985:HW

[DHL85] E. Deutsch, K. P. Hadeler, and T. J. Laffey. Helmut Wielandt. Linear Algebra and its Applications, 71(??): 1–8, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585902289.

Deutsch:1985:SVD

[DHLW85] E. Deutsch, K. P. (Karl Peter) Hadeler, T. J. Laffey, and Helmut W. Wielandt, editors. Special volume dedicated to Helmut W. Wielandt, volume 71 of Linear Algebra and its Applications. Elsevier, Amsterdam, The Netherlands, 1985. ISSN 0024-3795 (print), 1873-1856 (electronic).

DePierro:1985:SPM

[DI85] Alvaro R. De Pierro and Alfredo N. Iusem. A simultaneous projections method for linear inequalities. Linear Algebra and its Applications, 64(??):243–253, January 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902800.

Delosme:1986:PSS

[DI86] Jean-Marc Delosme and Ilse C. F. Ipsen. Parallel solution of symmetric positive definite systems with hyperbolic rotations. Linear Algebra and its Applications, 77 (??):75–111, May 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901631.

Special volume on parallel computing.

DiasdaSilva:1981:IRF

[Dia81]  J. A. Dias da Silva.  Irreducible representations

of the full linear group. Linear Algebra and its Applications, 40(??):161–182, October 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581901476.

DiasdaSilva:1986:MIE

[Dia86] J. A. Dias da Silva. On the multiplicative inverse eigenvalue problem. Linear Algebra and its Applications, 78 (??):133–145, June 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586900200.

Dickinson:1984:IPT

[Dic84] Bradley W. Dickinson. An inverse problem for Toeplitz matrices. Linear Algebra and its Applications, 59(??): 79–83, June 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901599.

Dion:1989:TMA

[Dio89] Jean Michel Dion. A transfer matrix approach to feedback invariants of linear systems: application to decoupling. Linear Algebra and its Applications, 121(??): 491–510, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589907180.

Linear algebra and applications (Valencia, 1987).

Dixon:1981:DT

[Dix81] John D. Dixon. On duality theorems. Linear Algebra and its Applications, 39(??):223–227, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581903050.

Djokovic:1981:RBR

[Djo81] Dragomir Ž. Djoković. Relations between reflections in the unitary group and Hermitian hyperquadrics in the complex projective space. Linear Algebra and its Applications, 35(??):129–153, February 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581902706.

Djokovic:1982:RBH

[Djo82] Dragomir Ž. Djoković. Relations between the half turns of the hyperbolic plane. Linear Algebra and its Applications, 47(??):57–88, October 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582902269.

Demmel:1987:CSE

[DK87] James Weldon Demmel and Bo Kågström. Computing stable eigendecompositions of matrix pencils. Linear Algebra and its Applications, 88/89(??):139–186, April 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dongarra:1986:SME

[DKH86] Jack J. Dongarra, Linda Kaufman, and Sven Hammarling. Squeezing the most out of eigenvalue solvers on high-performance computers. Linear Algebra and its Applications, 77(??):113–136, May 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586901643. Special vol-

ume on parallel computing.

Dzeng:1989:HFA

[DL89] D. C. Dzeng and W. W. Lin. HMDR and FMDR algorithms for the generalized eigenvalue problem. Linear Algebra and its Applications, 112(??):169–187, January 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589905934.

Djokovic:1980:ONM

[DM80] Dragimir Ž. Djoković and J. Malzan. Orbits of nilpotent matrices. Linear Algebra and its Applications, 32(??): 157–158, August 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580900142.

Djaferis:1982:AMS

[DM82] T. E. Djaferis and S. K. Mitter. Algebraic methods for the study of some linear matrix equations. Linear Algebra and its Applications, 44(??):125–142, April 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582900088.

Dahmen:1983:TMS

[DM83a] Wolfgang Dahmen and Charles A. Micchelli. Translates of multivariate splines. Linear Algebra and its Applications, 52/53(??):217–234, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Djaferis:1983:SGI

[DM83b] T. E. Djaferis and S. K. Mitter. Some generic invariant factor assignment results using dynamic output feedback. Linear Algebra and its Applications, 50(??): 103–131, April 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583900551.

Davies:1986:DMC

[DM86] Roy O. Davies and J. J. Modi. A direct method for completing eigenproblem solutions on a parallel computer. Linear Algebra and its Applications, 77(??):61–74, May 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958690162X.

Deift:1984:DTM

[DN84] P. Deift and T. Nanda. On the determination of a tridiagonal matrix from its spectrum and a submatrix. Linear Algebra and its Applications, 60(??):43–55, August 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584900697.

Deutsch:1985:FSO

[DN85] Emeric Deutsch and Michael Neumann. On the first and second order derivatives of the Perron vector. Linear Algebra and its Applications, 71(??):57–76, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902356.

deOliveira:1982:MPC

[dO82] G. N. de Oliveira. Matrices with prescribed characteristic polynomial and principal blocks. II. Linear Algebra and its Applications, 47(??):35–40, October 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582902245.

deOliveira:1983:UCM

[dO83] G. N. de Oliveira. The University of Coimbra Mini-Conference on Linear Algebra and Applications. Linear Algebra and its Applications, 54(??):111–138, October 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583902094.

deOliveira:1989:P

[dO89] Graciano de Oliveira. Preface. Linear Algebra and its Applications, 121(??):1–2, August 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589906873.

deOliveira:1983:EDS

[dOdS83] G. N. de Oliveira and J. A. Dias da Silva. Equality of decomposable symmetrized tensors and ^{∗}-matrix groups. Linear Algebra and its Applications, 49(??):191– 219, February 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901039.

Don:1985:UGI

[Don85] F. J. Henk Don. The use of generalized inverses in restricted maximum likelihood. Linear Algebra and its Applications, 70(??):225–240, October 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900540.

Don:1987:SSL

[Don87] F. J. Henk Don. On the symmetric solutions of a linear matrix equation. Linear Algebra and its Applications, 93 (??):1–7, August 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0024379587903089.

Doob:1982:SPL

[Doo82] Michael Doob. A surprising property of the least eigenvalue of a graph. Linear Algebra and its Applications, 46 (??):1–7, August 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582900210.

Doob:1989:LPE

[Doo89] Michael Doob. The limit points of eigenvalues of graphs. Linear Algebra and its Applications, 114/115(??):659– 662, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Downs:1985:DPR

[Dow85] T. Downs. A degree property of rational matrices. Linear Algebra and its Applications, 72(??):73–84, December 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585901430.

dePillis:1980:HEY

[dP80] John de Pillis. How to embrace your spectrum for faster iterative results. Linear Algebra and its Applications, 34(??):125–143, December 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901639.

daProvidencia:1986:MSC

[dPB86] J. da Providência and N. Bebiano. Matrices satisfying a conjecture of G. N. de Oliveira on determinants. Linear Algebra and its Applications, 78(??):187–198, June 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586900236.

Deza:1980:GIP

[DR80a] M. Deza and I. G. Rosenberg. Generalized intersection patterns and two-symbol balanced arrays. Linear Algebra and its Applications, 30(??):9–40, April 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580901780.

Dlab:1980:RNF

[DR80b] Vlastimil Dlab and Claus Michael Ringel. A remark on normal forms of matrices. Linear Algebra and its Applications, 30(??):109–114, April 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958090186X.

Djokovic:1983:SVS

[DR83] Dragomir Ž. Djoković and Mark P. Richards. On singular values and similarity classes of matrices. Linear Algebra and its Applications, 110(??):1–28, November 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583901295.

DellaDora:1986:NGC

[DR86] Jean Della Dora and François Robert. Noël Gastinel and computational mathematics. Linear Algebra and its Applications, 78(??):237–247, June 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586900273.

Djokovic:1988:SVS

[DR88] Dragomir Ž. Djoković and Mark P. Richards. On singular values and similarity classes of matrices. Linear Algebra and its Applications, 110(??):1–28, 1988. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drury:1987:EPP

[Dru87] S. W. Drury. Extreme points for positive forms on l^{p}. Linear Algebra and its Applications, 97(??):219– 228, December 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587901510.

Drygas:1985:NIP

[Dry85] Hilmar Drygas. A note on the inverse-partitioned-matrix method in linear regression analysis. Linear Algebra and its Applications, 67(??):275–277, June 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902010.

Deuflhard:1980:RDP

[DS80] P. Deuflhard and W. Sautter. On rank-deficient pseudoinverses. Linear Algebra and its Applications, 29(??): 91–111, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580902323.

Drake:1982:ASM

[DS82] David A. Drake and Sharad S. Sane. Auxiliary sets of matrices with new step parameter sequences. Linear Algebra and its Applications, 46(??):131–153, August 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582900325.

Delsarte:1989:FTI

[DS89a] P. Delsarte and J. J. Seidel. Fisher type inequalities for Euclidean t-designs. Linear Algebra and its Applications, 114/115(??):213–230, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DiNola:1989:SRM

[DS89b] Antonio Di Nola and Salvatore Sessa. On the Schein rank of matrices over linear lattices. Linear Algebra and its Applications, 118(??):155–158, June 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589905788.

deSouza:1981:CLE

[dSB81a] Eurice de Souza and S. P. Bhattacharyya. Controllability and the linear equation Ax = b. Linear Algebra and its Applications, 36(??):97–101, March 1981.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581902226.

deSouza:1981:COS

[dSB81b] Eurice de Souza and S. P. Bhattacharyya. Controllability, observability and the solution of AX −XB = C. Linear Algebra and its Applications, 39(??):167–188, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581903013.

Dong:1988:EPD

[DSD88] Wei Quan Dong, Jia Yu Shao, and Chun Fei Dong. On the exponents of primitive digraphs with the shortest elementary circuit length s. Linear Algebra and its Applications, 104(??):1–27, June 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588903035.

Dennis:1987:GCD

[DT87] J. E. Dennis, Jr. and Kathryn Turner. Generalized conjugate directions. Linear Algebra and its Applications, 88/89(??):187–209, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Duarte:1989:CAM

[Dua89] António Leal Duarte. Construction of acyclic matrices from spectral data. Linear Algebra and its Applications, 113(??):173–182, February 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589902954.

Dubeau:1985:BCM

[Dub85] François Dubeau. On band circulant matrices in the periodic spline interpolation theory. Linear Algebra and its Applications, 72(??):177–182, December 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585901533.

Don:1981:NCP

[DvdP81] F. J. Henk Don and Adriaan P. van der Plas. A note on the characteristic polynomial of the commutation matrix. Linear Algebra and its Applications, 37(??): 135–142, April 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581901737.

Deutsch:1983:NBP

[DW83] Emeric Deutsch and Helmut Wielandt. Nested bounds for the Perron root of a nonnegative matrix. Linear Algebra and its Applications, 52/53(??):235–251, 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Deistler:1989:CSP

[DW89] M. Deistler and Liqun Wang. The common structure of parametrizations for linear systems. Linear Algebra and its Applications, 122/123/124(??):921–941, 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eagambaram:1988:GIN

[Eag88] N. Eagambaram. Generalized inverses with nonnegative principal minors. Linear Algebra and its Applications, 111(??):293–312, December 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588900663.

Eaves:1984:SPD

[Eav84]  B. Curtis Eaves.  Subdivisions from primal and

dual cones and polytopes. Linear Algebra and its Applications, 62(??):277–285, November 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901034.

Ebanks:1989:E

[Eba89] Bruce R. Ebanks. On the equation F (X)+M(X)G(X^{−1}) = 0 on K^{n}. Linear Algebra and its Applications, 125(??):

1–17, December 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958990030X.

Espanol:1989:CMO

[EG89] Francisco Marcellán Español and Isabel Rodríguez González. A class of matrix orthogonal polynomials on the unit circle. Linear Algebra and its Applications, 121(??):233–241, August 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589907040. Linear alge-

bra and applications (Valencia, 1987).

Ellis:1985:FBM

[EGL85] Robert L. Ellis, Israel Gohberg, and David Lay. Factorization of block matrices. Linear Algebra and its Applications, 69(??):71–93, August 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900692.

Elsner:1985:LCO

[EH85] L. Elsner and K. P. Hadeler. Lothar Collatz — on the occasion of his 75th birthday. Linear Algebra and its Applications, 68(??):1–8, 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eggermont:1981:IAL

[EHL81] P. P. B. Eggermont, G. T. Herman, and A. Lent. Iterative algorithms for large partitioned linear systems, with applications to image reconstruction. Linear Algebra and its Applications, 40(??):37–67, October 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581901397.

Ellers:1987:FTL

[EI87] Erich W. Ellers and Hiroyuki Ishibashi. Factorization of transformations over a local ring. Linear Algebra and its Applications, 85(??):17–27, January 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587902059.

Elsner:1989:NMZ

[EJ89] Ludwig Elsner and Charles R. Johnson. Nonnegative matrices, zero patterns, and spectral inequalities. Linear Algebra and its Applications, 120(??):225–236, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589903807. Proceedings

of the Fourth Haifa Matrix Theory Conference (Haifa, 1988).

ElMossadeq:1985:COG

[EKC85] A. El Mossadeq, A. Kobilinsky, and D. Collombier. Construction d’orthogonaux dans les groupes abéliens finis et confusions d’effets dans les plans factoriels. (French) [Construction of orthogonals in finite Abelian groups and confounding effects in factorial designs]. Linear Algebra and its Applications, 70(??):303–320, 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958590062X.

Evans:1988:ASE

[EL88] D. J. Evans and C. Li. Analysis of a symmetric extrapolated iterative method for solving linear systems. Linear Algebra and its Applications, 103(??):149–173, May 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588902261.

Evans:1989:NSC

[EL89] D. J. Evans and C. Li. A note on the SOR and CG methods for large least squares problems. Linear Algebra and its Applications, 121(??):163–170, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958990699X. Linear

algebra and applications (Valencia, 1987).

Elezovic:1987:MPR

[Ele87] Neven Elezović. On the minimal polynomial of a resolvent. Linear Algebra and its Applications, 86(??): 1–9, February 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587902849.

Elfving:1980:SME

[Elf80] Tommy Elfving. On some methods for entropy maximization and matrix scaling. Linear Algebra and its Applications, 34(??):321–339, December 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901718.

Elia:1988:NCB

[Eli88] Michele Elia. A note on the computation of the bit error rate for binary block codes. Linear Algebra and its Applications, 98(??):199–210, January 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901656.

Eller:1987:FCM

[Ell87] József Eller. On functions of companion matrices. Linear Algebra and its Applications, 96(??):191–210, November 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587903442.

Eisenstat:1982:OBD

[ELS82a] Stanley C. Eisenstat, John W. Lewis, and Martin H. Schultz. Optimal block diagonal scaling of block 2-cyclic matrices. Linear Algebra and its Applications, 44(??): 181–186, April 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582900143.

Elsner:1982:MLI

[Els82b] L. Elsner. On matrices leaving invariant a nontrivial convex set. Linear Algebra and its Applications, 42(??):

103–107, February 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582901410.

Elsner:1982:VSM

[Els82c] L. Elsner. On the variation of the spectra of matrices. Linear Algebra and its Applications, 47(??):127–138, October 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582902300.

Elsner:1984:BSO

[Els84a] L. Elsner. Block scaling with optimal Euclidean condition. Linear Algebra and its Applications, 58(??): 69–73, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584902040.

Elsner:1984:CPS

[Els84b] L. Elsner. On convexity properties of the spectral radius of nonnegative matrices. Linear Algebra and its Applications, 61(??):31–35, September 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958490020X.

Elsner:1985:OBS

[Els85] L. Elsner. An optimal bound for the spectral variation of two matrices. Linear Algebra and its Applications, 71(??): 77–80, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585902368.

Elsner:1988:NVP

[Els88] L. Elsner. A note on the variation of permanents. Linear Algebra and its Applications, 109(??):37–39, October 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901966.

Emre:1983:GMM

[Emr83] Erol Emre. Generalized model matching and (F, G)- invariant submodules for linear systems over rings. Linear Algebra and its Applications, 50(??):133–166, April 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583900563.

Ellers:1981:RRO

[EN81] Erich W. Ellers and Wolfgang Nolte. Radical relations in orthogonal groups. Linear Algebra and its Applications, 38(??):135–139, June 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581900148.

Eirola:1989:ARO

[EN89] Timo Eirola and Olavi Nevanlinna. Accelerating with rank-one updates. Linear Algebra and its Applications, 121(??):511–520, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379589907192.

Linear algebra and applications (Valencia, 1987).

Elsner:1984:P

[ENV84] L. Elsner, W. Niethammer, and R. S. Varga. Preface. Linear Algebra and its Applications, 58(??):1, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958490199X.

Elsner:1987:MNM

[EP87a] L. Elsner and M. H. C. Paardekooper. On measures of nonnormality of matrices. Linear Algebra and its Applications, 92(??):107–123, July 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587902539.

Ermers:1987:ESM

[EP87b] Ruud Ermers and Ben Polman. On the eigenvalues of the structure matrix of matrices of zeros and ones. Linear Algebra and its Applications, 95(??):17–41, October 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587900255.

Eijkhout:1988:DRI

[EP88] Victor Eijkhout and Ben Polman. Decay rates of inverses of banded M-matrices that are near to Toeplitz matrices. Linear Algebra and its Applications, 109(??): 247–277, October 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958890211X.

Ericsson:1987:LAF

[ER87] Thomas Ericsson and Axel Ruhe. Lánczos algorithms and field of value rotations for symmetric matrix pencils. Linear Algebra and its Applications, 88/89(??):733–746, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eaves:1989:CTA

[ER89] B. Curtis Eaves and Uriel G. Rothblum. A continuoustime average-cost flexible manufacturing and operator scheduling model solved by deconvexification over time. Linear Algebra and its Applications, 114/115(??):417– 428, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Engel:1980:MDS

[ES80] Gernot M. Engel and Hans Schneider. Matrices diagonally similar to a symmetric matrix. Linear Algebra and its Applications, 29(??):131–138, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580902347.

Elsner:1982:PTG

[ES82] Ludwig Elsner and Ji Guang Sun. Perturbation theorems for the generalized eigenvalue problem. Linear Algebra and its Applications, 48(??):341–357, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582901203.

Elden:1986:SAR

[ES86] Lars Eldén and Robert Schreiber. A systolic array for the regularization of ill-conditioned least squares problems with triangular Toeplitz matrix. Linear Algebra and its Applications, 77(??):137–147, May 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586901655. Special vol-

ume on parallel computing.

Estes:1981:RCS

[ET81] Dennis R. Estes and Olga Taussky. Remarks concerning sums of three squares and quaternion commutator identities. Linear Algebra and its Applications, 35(??): 279–285, February 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902792.

Evard:1985:MFW

[Eva85] Jean-Claude Evard. On matrix functions which commute with their derivative. Linear Algebra and its Applications, 68(??):145–178, July 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902125.

Evard:1987:CVS

[Eva87] Jean-Claude Evard. Conditions for a vector subspace E(t) and for a projector P (t) not to depend on t. Linear Algebra and its Applications, 91(??):121–131, June 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587900668.

Faibusovich:1989:TFY

[Faĭ89] L. E. Faĭbusovich. qr-type factorizations, the Yang– Baxter equation, and an eigenvalue problem of control theory. Linear Algebra and its Applications, 122/123/124 (??):943–971, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fan:1983:NO

[Fan83] Ky Fan. Normalizable operators. Linear Algebra and its Applications, 52/53(??):253–263, 1983. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fan:1985:ISB

[Fan85] Ky Fan. An identity for symmetric bilinear forms. Linear Algebra and its Applications, 65(??):273–279, February 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585901028.

Fan:1988:IPC

[Fan88] Ky Fan. Inequalities for proper contractions and strictly dissipative operators. Linear Algebra and its Applications, 105(??):237–248, July 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958890016X.

Fang:1989:SM

[Fan89] Li Fang. On the spectra of P - and P_{0}-matrices. Linear Algebra and its Applications, 119(??):1–25, July 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589900657.

Farmer:1981:HGP

[Far81] K. Bolling Farmer. Hermitian geometries in projective space. Linear Algebra and its Applications, 35(??):37– 50, February 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902640.

Farnsteiner:1984:QLA

[Far84] Rolf Farnsteiner. Quaternionic Lie algebras. Linear Algebra and its Applications, 61(??):225–231, September 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584900338.

Farebrother:1985:LUA

[Far85a] R. W. Farebrother. Linear unbiased approximators of the disturbances in the standard linear model. Linear Algebra and its Applications, 67(??):259–274, June 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902009.

Faria:1985:PRS

[Far85b] Isabel Faria. Permanental roots and the star degree of a graph. Linear Algebra and its Applications, 64(??): 255–265, January 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585902812.

Farebrother:1987:TTA

[Far87] R. W. Farebrother. Three theorems with applications to Euclidean distance matrices. Linear Algebra and its Applications, 95(??):11–16, October 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587900243.

Friedman:1981:IPU

[FB81a] Nathan Friedman and Abraham Boyarsky. Irreducibility and primitivity using Markov maps. Linear Algebra and its Applications, 37(??):103–117, April 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581901713.

Friedman:1981:MEI

[FB81b] Nathan Friedman and Abraham Boyarsky. Matrices and eigenfunctions induced by Markov maps. Linear Algebra and its Applications, 38(??):141–147, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958190015X.

Fuhrmann:1989:BVM

[FD89] P. A. Fuhrmann and B. N. Datta. On Bézoutians, Vandermonde matrices, and the Liénard–Chipart stability criterion. Linear Algebra and its Applications, 120(??): 23–37, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589903674.

Proceedings of the Fourth Haifa Matrix Theory Conference (Haifa, 1988).

Feinsilver:1984:CIC

[Fei84] Philip Feinsilver. Circulants, inversion of circulants, and some related matrix algebras. Linear Algebra and its Applications, 56(??):29–43, January 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901113.

Feintuch:1989:SSE

[Fei89] Avraham Feintuch. Stabilization and sensitivity for eventually time-invariant systems. Linear Algebra and its Applications, 122/123/124(??):105–114, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ferland:1980:PSM

[Fer80] Jacques A. Ferland. Positive subdefinite matrices. Linear Algebra and its Applications, 31(??):233–244, June 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580902220.

Ferland:1981:MTC

[Fer81] Jacques A. Ferland. Matrix-theoretic criteria for the quasiconvexity of twice continuously differentiable functions. Linear Algebra and its Applications, 38(??):51–63, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900070.

Fletcher:1980:GEP

[FG80] R. Fletcher and D. F. Griffiths. The generalized eigenvalue problem for certain unsymmetric band matrices. Linear Algebra and its Applications, 29(??):139– 149, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580902359.

Fiedler:1981:CSP

[FG81] Miroslav Fiedler and Robert Grone. Characterizations of sign patterns of inverse-positive matrices. Linear Algebra and its Applications, 40(??):237–245, October 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581901543.

Funderlic:1986:TDF

[FG86] R. E. Funderlic and A. Geist. Torus data flow for parallel computation of missized matrix problems. Linear Algebra and its Applications, 77(??):149–164, May 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586901667.

Farmer:1980:DAG

[FH80] K. Bolling Farmer and Mark P. Hale, Jr. Dual affine geometries and alternative bilinear forms. Linear Algebra and its Applications, 30(??):183–199, April 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580901949.

Freund:1988:EPU

[FH88a] Roland Freund and Thomas Huckle. An extension problem for H-unitary matrices with applications to Hermitian Toeplitz matrices. Linear Algebra and its Applications, 108(??):213–230, September 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901899.

Friedland:1988:RPM

[FH88b] Shmuel Friedland and Daniel Hershkowitz. The rank of powers of matrices in a block triangular form. Linear Algebra and its Applications, 107(??):17–22, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902340. Proceedings

of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Frey:1980:NIC

[FHP80] A. E. Frey, C. A. Hall, and T. A. Porsching. Numerical integration and the constant strain condition. Linear Algebra and its Applications, 29(??):151–159, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580902360.

Fujimoto:1985:SAL

[FHV85] Takao Fujimoto, Carmen Herrero, and Antonio Villar. A sensitivity analysis for linear systems involving Mmatrices and its application to the Leontief model. Linear Algebra and its Applications, 64(??):85–91, January 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958590268X.

Fiala:1983:NEA

[Fia83] Tibor Fiala. Normal equivalent to an arbitrary diagonalizable matrix. Linear Algebra and its Applications, 51 (??):153–162, June 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901568.

Fialkow:1989:LDL

[Fia89] Aaron Fialkow. Linear dependence of linear transformations and images. Linear Algebra and its Applications, 126(??):15–37, December 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379589900037.

Fiedler:1981:GNR

[Fie81a] Miroslav Fiedler. Geometry of the numerical range of matrices. Linear Algebra and its Applications, 37(??): 81–96, April 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581901695.

Fiedler:1981:RSC

[Fie81b] Miroslav Fiedler. Remarks on the Schur complement. Linear Algebra and its Applications, 39(??):189–195, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581903025.

Fiedler:1983:NHP

[Fie83] Miroslav Fiedler. A note on the Hadamard product of matrices. Linear Algebra and its Applications, 49(??): 233–235, February 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901052.

Fiedler:1984:HLM

[Fie84a] Miroslav Fiedler. Hankel and Loewner matrices. Linear Algebra and its Applications, 58(??):75–95, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584902052.

Fiedler:1984:QDH

[Fie84b] Miroslav Fiedler. Quasidirect decompositions of Hankel and Toeplitz matrices. Linear Algebra and its Applications, 61(??):155–174, September 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900284.

Fiedler:1984:M

[Fie84c] Miroslav Fiedler. S-matrices. Linear Algebra and its Applications, 57(??):157–167, February 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901848.

Fiedler:1985:PHM

[Fie85] Miroslav Fiedler. Polynomials and Hankel matrices. Linear Algebra and its Applications, 66(??):235–248, April 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585901351.

Fiedler:1988:CBH

[Fie88a] M. Fiedler. Characterizations of Bézout and Hankel– Bézout matrices. Linear Algebra and its Applications, 105 (??):77–89, July 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588900055.

Fiedler:1988:CM

[Fie88b] Miroslav Fiedler. Characterizations of MMA-matrices. Linear Algebra and its Applications, 106(??):233–244, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588900316.

Fischer:1987:SMN

[Fis87] P. Fischer. Substochastic matrices and von Neumann majorization. Linear Algebra and its Applications, 93(??): 247–254, August 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0024379587903284.

Fiedler:1987:NIM

[FJM87] M. Fiedler, Charles R. Johnson, and T. L. Markham. Notes on inverse M-matrices. Linear Algebra and its Applications, 91(??):75–81, June 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379587900619.

Fiedler:1985:TIM

[FJMN85] Miroslav Fiedler, Charles R. Johnson, Thomas L. Markham, and Michael Neumann. A trace inequality for M-matrices and the symmetrizability of a real matrix by a positive diagonal matrix. Linear Algebra and its Applications, 71(??):81–94, November 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958590237X.

Fujimoto:1985:SES

[FK85] Takao Fujimoto and Ulrich Krause. Strong ergodicity for strictly increasing nonlinear operators. Linear Algebra and its Applications, 71(??):101–112, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902393.

Friedland:1987:MI

[FK87] Shmuel Friedland and Moshe Katz. On a matrix inequality. Linear Algebra and its Applications, 85(??):

185–190, January 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587902151.

Franklin:1989:SMM

[FL89a] Joel Franklin and Jens Lorenz. On the scaling of multidimensional matrices. Linear Algebra and its Applications, 114/115(??):717–735, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Friedland:1989:BSO

[FL89b] Shmuel Friedland and David London. Binyamin Schwarz — on the occasion of his 70th birthday. Linear Algebra and its Applications, 120(??):3–8, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589903650. Proceedings

of the Fourth Haifa Matrix Theory Conference (Haifa, 1988).

Fletcher:1988:DPR

[Fle88] R. Fletcher. Degeneracy in the presence of roundoff errors. Linear Algebra and its Applications, 106(??):149– 183, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588900262.

FloresdeChela:1982:ANG

[Flo82] Delia Flores de Chela. Approximations in the ℓ_{∞} norm and the generalized inverse. Linear Algebra and its Applications, 42(??):3–21, February 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582901343.

Fiedler:1981:IAM

[FM81] Miroslav Fiedler and Russell Merris. Irreducibility of associated matrices. Linear Algebra and its Applications, 37(??):1–10, April 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581901622.

Fiedler:1986:CMW

[FM86a] Miroslav Fiedler and Thomas L. Markham. Completing a matrix when certain entries of its inverse are specified. Linear Algebra and its Applications, 74(??):225– 237, February 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901254.

Funderlic:1986:SSD

[FM86b] R. E. Funderlic and C. D. Meyer, Jr. Sensitivity of the stationary distribution vector for an ergodic Markov chain. Linear Algebra and its Applications, 76(??):1–17, April 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902107.

Fiedler:1987:RPD

[FM87] Miroslav Fiedler and Thomas L. Markham. Rankpreserving diagonal completions of a matrix. Linear Algebra and its Applications, 85(??):49–56, January 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587902072.

Fiedler:1988:IHP

[FM88a] M. Fiedler and Thomas L. Markham. An inequality for the Hadamard product of an M-matrix and an inverse M-matrix. Linear Algebra and its Applications, 101 (??):1–8, April 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588901395.

Fiedler:1988:CCI

[FM88b] Miroslav Fiedler and Thomas L. Markham. A characterization of the closure of inverse M-matrices. Linear Algebra and its Applications, 105(??):209–223, July 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588900146.

Frommer:1989:CRP

[FM89] Andreas Frommer and Günter Mayer. Convergence of relaxed parallel multisplitting methods. Linear Algebra and its Applications, 119(??):141–152, July 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589900748.

Fiedler:1983:CPM

[FMN83] Miroslav Fiedler, Thomas L. Markham, and Michael Neumann. Classes of products of M-matrices and inverse M-matrices. Linear Algebra and its Applications, 52/53 (??):265–287, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Forster:1989:CWN

[FN89] K.-H. Förster and B. Nagy. On the Collatz–Wielandt numbers and the local spectral radius of a nonnegative operator. Linear Algebra and its Applications, 120(??): 193–205, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589903789.

Proceedings of the Fourth Haifa Matrix Theory Conference (Haifa, 1988).

Fong:1986:NER

[Fon86] C. K. Fong. Norm estimates related to self-commutators. Linear Algebra and its Applications, 74(??):151–156, February 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901187.

Foregger:1980:MVP

[For80] Thomas H. Foregger. On the minimum value of the permanent of a nearly decomposable doubly stochastic matrix. Linear Algebra and its Applications, 32(??): 75–85, August 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580900087.

Formanek:1986:RBA

[For86] Edward Formanek. Review of Linear Algebra over Commutative Rings, by Bernard R. McDonald. Linear Algebra and its Applications, 76(??):279–280, April 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902296.

Foregger:1987:MPM

[For87] Thomas H. Foregger. Minimum permanents of multiplexes. Linear Algebra and its Applications, 87(??): 197–211, March 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587901674.

Foster:1986:RNS

[Fos86] Leslie V. Foster. Rank and null space calculations using matrix decomposition without column interchanges. Linear Algebra and its Applications, 74(??):47–71, February 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586901151.

Funderlic:1981:DME

[FP81] R. E. Funderlic and R. J. Plemmons. LU decomposition of M-matrices by elimination without pivoting. Linear Algebra and its Applications, 41(??):99–110, December 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900914.

Fiedler:1987:BIM

[FP87a] Miroslav Fiedler and Vlastimil Pták. Bézoutians and intertwining matrices. Linear Algebra and its Applications, 86(??):43–51, February 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379587902862.

Fiedler:1987:ITM

[FP87b] Miroslav Fiedler and Vlastimil Pták. Intertwining and testing matrices corresponding to a polynomial. Linear Algebra and its Applications, 86(??):53–74, February 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587902874.

Fiedler:1988:LBM

[FP88] Miroslav Fiedler and Vlastimil Pták. Loewner and Bézout matrices. Linear Algebra and its Applications, 101(??): 187–220, April 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588901516.

Friedland:1982:STP

[FR82] S. Friedland and J. W. Robbin. Symmetries of tensor products. Linear Algebra and its Applications, 44(??): 97–123, April 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582900076.

Freund:1987:NTB

[Fre87] R. Freund. A note on two block SOR methods for sparse least squares problems. Linear Algebra and its Applications, 88/89(??):211–221, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Friedland:1981:GMT

[Fri81] Shmuel Friedland. A generalization of the Motzkin– Taussky theorem. Linear Algebra and its Applications, 36(??):103–109, March 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379581902238.

Friedland:1985:MEM

[Fri85a] Shmuel Friedland. The maximal eigenvalue of 0-1 matrices with prescribed number of ones. Linear Algebra and its Applications, 69(??):33–69, August 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900680.

Friedland:1985:WIP

[Fri85b] Shmuel Friedland. Weak interlacing properties of totally positive matrices. Linear Algebra and its Applications, 71(??):95–100, November 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902381.

Friedland:1986:LEN

[Fri86] Shmuel Friedland. Limit eigenvalues of nonnegative matrices. Linear Algebra and its Applications, 74(??):173– 178, February 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901205.

Friedland:1988:BSR

[Fri88] Shmuel Friedland. Bounds on the spectral radius of graphs with e edges. Linear Algebra and its Applications, 101(??):81–86, April 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379588901449.

Faudree:1982:SRG

[FRS82] R. J. Faudree, C. C. Rousseau, and J. Sheehan. Strongly regular graphs and finite Ramsey theory. Linear Algebra and its Applications, 46(??):221–241, August 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582900374.

Frydman:1985:SDS

[Fry85] Halina Frydman. A structure of doubly stochastic Markov chains. Linear Algebra and its Applications, 67(??): 51–66, June 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585901855.

Friedland:1982:CDP

[FS82] Shmuel Friedland and Barry Simon. The codimension of degenerate pencils. Linear Algebra and its Applications, 44(??):41–53, April 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582900039.

Fischer:1985:TDL

[FS85] P. Fischer and E. R. Swart. Three-dimensional line stochastic matrices and extreme points. Linear Algebra and its Applications, 69(??):179–203, August 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585900758.

Fabrykowski:1985:SBI

[FSZ85] J. Fabrykowski, A. Sharma, and H. Zassenhaus. Some Birkhoff interpolation problems on the roots of unity. Linear Algebra and its Applications, 65(??):1–23, February 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585900849.

Friedland:1984:OLW

[FT84] Shmuel Friedland and Eitan Tadmor. Optimality of the Lax–Wendroff condition. Linear Algebra and its Applications, 56(??):121–129, January 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901186.

Franchetti:1988:NEP

[FT88] C. Franchetti and U. Tiberio. A numerical evaluation of projection constants. Linear Algebra and its Applications, 109(??):179–196, October 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902091.

Foias:1989:PSS

[FT89a] Ciprian Foiaş and Allen Tannenbaum. On the parametrization of the suboptimal solutions in generalized interpolation. Linear Algebra and its Applications, 122/123/124 (??):145–164, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Frank:1989:ASF

[FT89b] András Frank and Éva Tardos. An application of submodular flows. Linear Algebra and its Applications, 114/115 (??):329–348, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fuhrmann:1983:SRT

[Fuh83] Paul A. Fuhrmann. On symmetric rational transfer functions. Linear Algebra and its Applications, 50(??): 167–250, April 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583900575.

Fuhrmann:1984:HRT

[Fuh84] Paul A. Fuhrmann. On Hamiltonian rational transfer functions. Linear Algebra and its Applications, 63(??): 1–93, December 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901368.

Fuhrmann:1986:RIH

[Fuh86] Paul A. Fuhrmann. Remarks on the inversion of Hankel matrices. Linear Algebra and its Applications, 81(??): 89–104, September 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958690251X.

Fujishige:1984:SLI

[Fuj84] Satoru Fujishige. A system of linear inequalities with a submodular function on {0, ±1} vectors. Linear Algebra and its Applications, 63(??):253–266, December 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901472.

Furuta:1988:OE

[Fur88] Takayuki Furuta. The operator equation T (H^{1/n}T)^{n} = K. Linear Algebra and its Applications, 109(??):149– 152, October 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588902054.

Dedicated to Professor Shōo Koshi on the occasion of his sixtieth birthday.

Furuta:1989:POM

[Fur89] Takayuki Furuta. A proof via operator means of an order preserving inequality. Linear Algebra and its Applications, 113(??):129–130, February 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589902917.

Feig:1984:DSC

[FW84] Ephraim Feig and Shmuel Winograd. On the direct sum conjecture. Linear Algebra and its Applications, 63(??): 193–219, December 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901435.

Friedland:1984:ASD

[FZ84] S. Friedland and C. Zenger. All spectral dominant norms are stable. Linear Algebra and its Applications, 58(??): 97–107, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584902064.

Gunn:1981:CIL

[GA81] E. A. Gunn and G. J. Anders. A comparison of interval linear programming with the simplex method. Linear Algebra and its Applications, 38(??):149–159, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900161.

Guichal:1985:RBG

[GA85] Edgardo N. Güichal and Bahía Blanca Argentina. A relation between Gram and Wronsky determinants. Linear Algebra and its Applications, 72(??):59–72, December 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585901429.

Gabriel:1987:NMC

[Gab87] Richard Gabriel. The normal ΔH-matrices with connection to some Jacobi-like methods. Linear Algebra and its Applications, 91(??):181–194, June 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958790070X.

Gader:1988:TFF

[Gad88] Paul D. Gader. Tridiagonal factorizations of Fourier matrices and applications to parallel computations of discrete Fourier transforms. Linear Algebra and its Applications, 102(??):169–209, April 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588903254.

Gaines:1989:FDC

[Gai89] Fergus Gaines. Fourth Dublin Conference on Matrix Theory and Its Applications. Linear Algebra and its Applications, 117(??):133–183, May 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589905557.

Galkowski:1981:TWC

[Gał81] Krzysztof Gałkowski. Three-way companion matrices of three-variable polynomials. Linear Algebra and its Applications, 37(??):55–75, April 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581901671.

Gamas:1988:CSD

[Gam88a] Carlos Gamas. Conditions for a symmetrized decomposable tensor to be zero. Linear Algebra and its Applications, 108(??):83–119, September 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901802.

Gambini:1988:ASS

[Gam88b] Alessandra Gambini. Algorithms for soluble subalgebras of finite-dimensional Lie algebras. Linear Algebra and its Applications, 104(??):29–38, June 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588903047.

Gardiner:1982:AGD

[Gar82a] A. Gardiner. Antipodal graphs of diameter three. Linear Algebra and its Applications, 46(??):215–219, August 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582900362.

Garloff:1982:CSR

[Gar82b] Jürgen Garloff. Criteria for sign regularity of sets of matrices. Linear Algebra and its Applications, 44(??): 153–160, April 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582900106.

Garloff:1982:MBD

[Gar82c] Jürgen Garloff. Majorization between the diagonal elements and the eigenvalues of an oscillating matrix. Linear Algebra and its Applications, 47(??):181–184, October 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582902361.

Gautschi:1983:CVL

[Gau83a] Walter Gautschi. The condition of Vandermonde-like matrices involving orthogonal polynomials. Linear Algebra and its Applications, 52/53(??):293–300, 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gautschi:1983:AMO

[Gau83b] Walter Gautschi. To Alexander M. Ostrowski on his ninetieth birthday. Linear Algebra and its Applications, 52/53(??):xi–xiv, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gover:1984:GPM

[GB84] M. J. C. Gover and S. Barnett. Generating polynomials for matrices with Toeplitz or conjugate-toeplitz inverses. Linear Algebra and its Applications, 61(??):253– 275, September 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584900351.

Gover:1987:DRQ

[GB87] M. J. C. Gover and S. Barnett. Displacement rank and quasitriangular decomposition for r-toeplitz matrices. Linear Algebra and its Applications, 88/89(??):329– 348, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Greenfield:1984:DT

[GD84] Gary R. Greenfield and Daniel Drucker. On the discriminant of a trinomial. Linear Algebra and its Applications, 62(??):105–112, November 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900892.

Gracia:1989:PLC

[GdHZ89] Juan-M. Gracia, Inmaculada de Hoyos, and Ion Zaballa. Perturbation of linear control systems. Linear Algebra and its Applications, 121(??):353–383, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589907106. Linear

algebra and applications (Valencia, 1987).

Golub:1987:D

[GDM87] Gene H. Golub, Iain Duff, and Cleve Moler. Dedication. Linear Algebra and its Applications, 88/89(??):1– 12, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Geerts:1989:AOC

[Gee89a] Ton Geerts. All optimal controls for the singular linearquadratic problem without stability and a new interpretation of the optimal cost. Linear Algebra and its Applications, 116(??):135–181, April 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589904035.

Geerts:1989:CPC

[Gee89b] Ton Geerts. Continuity properties of the cheap-control problem without stability. Linear Algebra and its Applications, 122/123/124(??):65–104, 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gekeler:1981:PMP

[Gek81] Eckart Gekeler. On the pointwise matrix product and the mean value theorem. Linear Algebra and its Applications, 35(??):183–191, February 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581902731.

Gendreau:1986:LED

[Gen86] Michel Gendreau. On the location of eigenvalues of off-diagonal constant matrices. Linear Algebra and its Applications, 79(??):99–102, July 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902946.

Gerth:1984:FFT

[Ger84] Frank Gerth, III. Finite Fourier transform for functions with restricted support. Linear Algebra and its Applications, 60(??):145–158, August 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900764.

Gerards:1989:SPT

[Ger89] A. M. H. Gerards. A short proof of Tutte’s characterization of totally unimodular matrices. Linear Algebra and its Applications, 114/115(??):207–212, 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gohberg:1982:FDW

[GG82] Israel Gohberg and Seymour Goldberg. Finite-dimensional Wiener–Hopf equations and factorizations of matrices. Linear Algebra and its Applications, 48(??):219– 236, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582901094.

Gander:1989:CEP

[GGvM89] Walter Gander, Gene H. Golub, and Urs von Matt. A constrained eigenvalue problem. Linear Algebra and its Applications, 114/115(??):815–839, 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

George:1980:SSL

[GH80] Alan George and Michael T. Heath. Solution of sparse linear least squares problems using Givens rotations. Linear Algebra and its Applications, 34(??):69–83, December 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580901597.

Goffin:1983:RBH

[GH83] Jean-Louis Goffin and Alan J. Hoffman. On the relationship between the Hausdorff distance and matrix distances of ellipsoids. Linear Algebra and its Applications, 52/53 (??):301–313, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gabriel:1984:DIG

[GH84] Richard Gabriel and Robert E. Hartwig. The Drazin inverse as a gradient. Linear Algebra and its Applications, 63(??):237–252, December 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379584901460.

Gregorac:1987:NAO

[GH87] R. J. Gregorac and Irvin Roy Hentzel. A note on the analogue of Oppenheim’s inequality for permanents. Linear Algebra and its Applications, 94(??):109–112, September 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587900826.

Gamst:1989:IPH

[GH89] J. Gamst and K. Hoechsmann. An invariance property of Hankel forms. Linear Algebra and its Applications, 122/123/124(??):447–456, 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ghatage:1987:SBH

[Gha87] Pratibha G. Ghatage. On the spectrum of the Bergman–Hilbert matrix. Linear Algebra and its Applications, 97(??):57–63, December 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958790139X.

George:1986:PCF

[GHL86] Alan George, Michael T. Heath, and Joseph Liu. Parallel Cholesky factorization on a shared-memory multiprocessor. Linear Algebra and its Applications, 77(??): 165–187, May 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901679.

Golub:1987:GEY

[GHS87] G. H. Golub, Alan Hoffman, and G. W. Stewart. A generalization of the Eckart–Young–Mirsky matrix approximation theorem. Linear Algebra and its Applications, 88/89 (??):317–327, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Grone:1982:PMG

[GHW82] Robert D. Grone, D. G. Hoffman, and J. R. Wall. Permutation-matrix groups with positive sum. Linear Algebra and its Applications, 45(??):29–34, June 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582902087.

Gibson:1980:GDS

[Gib80a] Peter M. Gibson. Generalized doubly stochastic and permutation matrices over a ring. Linear Algebra and its Applications, 30(??):101–107, April 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901858.

Gibson:1980:PPD

[Gib80b] Peter M. Gibson. Permanental Polytopes of Doubly Stochastic Matrices. Linear Algebra and its Applications, 32(??):87–111, August 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379580900099.

Givens:1982:MLF

[Giv82] Clark R. Givens. On the modified Leverrier–Faddeev algorithm. Linear Algebra and its Applications, 44(??): 161–167, April 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582900118.

Givens:1984:SOE

[Giv84] Clark R. Givens. Some observations on eigenvectors of Hadamard matrices of order 2^{n}. Linear Algebra and its Applications, 56(??):245–250, January 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901307.

Grone:1984:PDC

[GJMW84] Robert Grone, Charles R. Johnson, Eduardo Marques de Sá, and Henry Wolkowicz. Positive definite completions of partial Hermitian matrices. Linear Algebra and its Applications, 58(??):109–124, April 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584902076.

Grone:1987:NM

[GJMW87] Robert Grone, Charles R. Johnson, Eduardo Marques de Sá, and Henry Wolkowicz. Normal matrices. Linear Algebra and its Applications, 87(??):213–225, March 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587901686.

Gunawardena:1989:PGN

[GJS89] Ananda D. Gunawardena, S. K. Jain, and Larry Snyder. On the periodicity of the graph of nonnegative matrices. Linear Algebra and its Applications, 120(??): 181–192, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589903777.

Proceedings of the Fourth Haifa Matrix Theory Conference (Haifa, 1988).

Gohberg:1986:ESL

[GKK86] I. Gohberg, T. Kailath, and I. Koltracht. Efficient solution of linear systems of equations with recursive structure. Linear Algebra and its Applications, 80(??):81– 113, August 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958690279X.

Gohberg:1987:LCP

[GKKL87] I. Gohberg, T. Kailath, I. Koltracht, and P. Lancaster. Linear complexity parallel algorithms for linear systems of equations with recursive structure. Linear Algebra and its Applications, 88/89(??):271–315, 1987. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guralnick:1982:PCM

[GL82] Robert Guralnick and Charles Lanski. Pseudosimilarity and cancellation of modules. Linear Algebra and its Applications, 47(??):111–115, October 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582902282.

Gragg:1983:PRP

[GL83] William B. Gragg and Anders Lindquist. On the partial realization problem. Linear Algebra and its Applications, 50(??):277–319, April 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379583900599.

Gow:1984:PAF

[GL84] R. Gow and T. J. Laffey. Pairs of alternating forms and products of two skew-symmetric matrices. Linear Algebra and its Applications, 63(??):119–132, December 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901393.

Gaines:1985:RDM

[GL85a] F. J. Gaines and T. J. Laffey. Report on the Dublin Matrix Theory Conference, March 1984: Matrices over finite fields, Graciano de Oliveira, Departamento de Matemática, Universidad de Coimbra, Coimbra, Portugal. Linear Algebra and its Applications, 68(??): 243–247, July 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585902150.

Gaines:1985:MWW

[GL85b] Fergus J. Gaines and Thomas J. Laffey. The mathematical work of Walter Ledermann. Linear Algebra and its Applications, 69(??):1–8, August 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900667.

George:1987:HRV

[GL87] Alan George and Joseph W. H. Liu. Householder reflections versus Givens rotations in sparse orthogonal decomposition. Linear Algebra and its Applications, 88/89 (??):223–238, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Goberna:1988:TLI

[GL88] M. A. Goberna and M. A. López. A theory of linear inequality systems. Linear Algebra and its Applications, 106(??):77–115, August 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588900249.

George:1984:ROS

[GLN84] Alan George, Joseph Liu, and Esmond Ng. Rowordering schemes for sparse Givens transformations. I. bipartite graph model. Linear Algebra and its Applications, 61(??):55–81, September 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900223.

George:1986:ROSa

[GLN86a] Alan George, Joseph Liu, and Esmond Ng. Row-ordering schemes for sparse Givens transformations. II. implicit graph model. Linear Algebra and its Applications, 75(??):

203–223, March 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901904.

George:1986:ROSb

[GLN86b] Alan George, Joseph Liu, and Esmond Ng. Rowordering schemes for sparse Givens transformations. III. Analyses for a model problem. Linear Algebra and its Applications, 75(??):225–240, March 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586901916.

Gohberg:1983:WHF

[GLR83] I. Gohberg, L. Lerer, and L. Rodman. Wiener–Hopf factorization of piecewise matrix polynomials. Linear Algebra and its Applications, 52/53(??):315–350, 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gaines:1983:PMQ

[GLS83] Fergus J. Gaines, Thomas J. Laffey, and Helene M. Shapiro. Pairs of matrices with quadratic minimal polynomials. Linear Algebra and its Applications, 52/53(??): 289–292, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Godsil:1980:FCE

[GM80] C. D. Godsil and B. D. McKay. Feasibility conditions for the existence of walk-regular graphs. Linear Algebra and its Applications, 30(??):51–61, April 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580901809.

Gleser:1985:PDM

[GM85] Leon J. Gleser and David S. Moore. Positive dependence in Markov chains. Linear Algebra and its Applications, 70(??):131–146, October 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900485.

Grone:1987:FIS

[GM87] Robert Grone and Russell Merris. A Fischer inequality for the second immanant. Linear Algebra and its Applications, 87(??):77–83, March 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587901595.

Godsil:1988:WGF

[GM88a] C. D. Godsil and B. Mohar. Walk generating functions and spectral measures of infinite graphs. Linear Algebra and its Applications, 107(??):191–206, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902455. Proceedings

of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Goodman:1988:CCA

[GM88b] T. N. T. Goodman and Charles A. Micchelli. Corner cutting algorithms for the Bézier representation of free form curves. Linear Algebra and its Applications, 99(??): 225–252, February 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588901346.

Gasca:1989:CDA

[GM89] M. Gasca and J. J. Martínez. On the computation of determinants arising in some bivariate rational interpolation problems. Linear Algebra and its Applications, 121(??):87–93, August 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589906939. Linear alge-

bra and applications (Valencia, 1987).

Gill:1987:MFG

[GMSW87] Philip E. Gill, Walter Murray, Michael A. Saunders, and Margaret H. Wright. Maintaining LU factors of a general sparse matrix. Linear Algebra and its Applications, 88/89 (??):239–270, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gunzburger:1982:SGE

[GN82] Max D. Gunzburger and R. A. Nicolaides. Stability of Gaussian elimination without pivoting on tridiagonal Toeplitz matrices. Linear Algebra and its Applications, 45(??):21–28, June 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582902075.

Gunzburger:1985:ENP

[GN85] M. D. Gunzburger and R. A. Nicolaides. Elimination with Noninvertible Pivots. Linear Algebra and its Applications, 64(??):183–189, January 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902757.

Godsil:1982:EGD

[God82] C. D. Godsil. Eigenvalues of graphs and digraphs. Linear Algebra and its Applications, 46(??):43–50, August 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582900246.

Goff:1981:HFM

[Gof81] Stuart Goff. Hermitian function matrices which commute with their derivative. Linear Algebra and its Applications, 36(??):33–40, March 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581902172.

Goldberg:1984:MNM

[Gol84] Moshe Goldberg. Multiplicativity of ℓ_{p} norms for matrices. II. Linear Algebra and its Applications, 62(??): 1–10, November 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584900843.

Goldberg:1985:EGS

[Gol85] Moshe Goldberg. Ernst G. Straus (1922–1983). Linear Algebra and its Applications, 64(??):1–19, January 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902630.

Goldberg:1986:MMN

[Gol86a] Moshe Goldberg. Mixed multiplicativity and ℓ_{p} norms for matrices. Linear Algebra and its Applications, 73(??): 123–131, January 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958690234X.

Goller:1986:SOR

[Gol86b] Hans Goller. Shorted operators and rank decomposition matrices. Linear Algebra and its Applications, 81(??): 207–236, September 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379586902570.

Goldberg:1987:MFM

[Gol87] Moshe Goldberg. Multiplicativity factors and mixed multiplicativity. Linear Algebra and its Applications, 97(??): 45–56, December 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587901388.

Gorodetsky:1989:BTM

[Gor89] Michael Gorodetsky. On block Toeplitz matrices with analytic symbol. Linear Algebra and its Applications, 116 (??):41–51, April 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589903972.

Gover:1988:RBL

[Gov88a] M. J. C. Gover. Review of Applied Linear Algebra by Riaz A. Usmani. Linear Algebra and its Applications, 106(??):53–55, August 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588900225.

Gover:1988:RBM

[Gov88b] M. J. C. Gover. A ring of Brownian matrices. Linear Algebra and its Applications, 103(??):87–102, May 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588902224.

Gover:1989:CRB

[Gov89a] M. J. C. Gover. Comparative review of Applied Linear Algebra, 3rd ed., by B. Noble and J. W. Daniel and Linear Algebra and Its Applications, 3rd ed., by G. Strang. Linear Algebra and its Applications, 118(??):159–161, June 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958990579X.

Gover:1989:DCM

[Gov89b] M. J. C. Gover. The determination of companion matrices characterizing Toeplitz and r-toeplitz matrices. Linear Algebra and its Applications, 117(??):81–92, May 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589905491.

Gower:1980:MLF

[Gow80] J. C. Gower. A modified Leverrier–Faddeev algorithm for matrices with multiple eigenvalues. Linear Algebra and its Applications, 31(??):61–70, June 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580902062.

Gow:1981:NEC

[Gow81] R. Gow. The number of equivalence classes of nondegenerate bilinear and sesquilinear forms over a finite field. Linear Algebra and its Applications, 41(??):175– 181, December 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581900975.

Gowda:1985:CCP

[Gow85a] M. Seetharama Gowda. Cone characterizations of positive semidefinite operators on a Hilbert space. Linear Algebra and its Applications, 64(??):77–83, January 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902678.

Gower:1985:PEN

[Gow85b] J. C. Gower. Properties of Euclidean and non-Euclidean distance matrices. Linear Algebra and its Applications, 67(??):81–97, June 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585901879.

Gowda:1989:PCS

[Gow89] M. Seetharama Gowda. Pseudomonotone and copositive star matrices. Linear Algebra and its Applications, 113(??):107–118, February 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589902899.

Golub:1980:LSG

[GP80] Gene H. Golub and Robert J. Plemmons. Largescale geodetic least-squares adjustment by dissection and orthogonal decomposition. Linear Algebra and its Applications, 34(??):3–28, December 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901561.

Govaerts:1989:SVI

[GP89] W. Govaerts and J. D. Pryce. A singular value inequality for block matrices. Linear Algebra and its Applications, 125(??):141–148, December 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589900360.

George:1980:SFP

[GR80] Alan George and Hamza Rashwan. On symbolic factorization of partitioned sparse symmetric matrices. Linear Algebra and its Applications, 34(??):145–157, December 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580901640.

Gohberg:1986:DBL

[GR86] I. Gohberg and L. Rodman. On the distance between lattices of invariant subspaces of matrices. Linear Algebra and its Applications, 76(??):85–120, April 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902168.

Gragg:1987:AOP

[GR87] W. B. Gragg and L. Reichel. On the application of orthogonal polynomials to the iterative solution of linear systems of equations with indefinite or non-Hermitian matrices. Linear Algebra and its Applications, 88/89(??):349–371, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gauchman:1989:SPF

[GR89a] Hillel Gauchman and Lee A. Rubel. Sums of products of functions of x times functions of y. Linear Algebra and its Applications, 125(??):19–63, December 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589900311.

Gragg:1989:SVH

[GR89b] W. B. Gragg and L. Reichel. On singular values of Hankel operators of finite rank. Linear Algebra and its Applications, 121(??):53–70, August 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589906915. Linear alge-

bra and applications (Valencia, 1987).

Green:1983:RBI

[Gre83a] William L. Green. Review of Finite and Infinite Dimensional Linear Spaces by Richard D. Järvinen. Linear Algebra and its Applications, 54(??):179–182, October 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583902148.

Greville:1983:TMT

[Gre83b] T. N. E. Greville. Toeplitz matrices with Toeplitz inverses revisited. Linear Algebra and its Applications, 55(??): 87–92, December 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901684.

Green:1988:BSR

[Gre88] Michael Green. Balanced stochastic realizations. Linear Algebra and its Applications, 98(??):211–247, January 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901668.

Greenbaum:1989:BSP

[Gre89] A. Greenbaum. Behavior of slightly perturbed Lanczos and conjugate-gradient recurrences. Linear Algebra and its Applications, 113(??):7–63, February 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589902851.

Grone:1980:CIR

[Gro80a] Robert Grone. Certain isometries of rectangular complex matrices. Linear Algebra and its Applications, 29(??): 161–171, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580902372.

Grosse:1980:TSA

[Gro80b] Eric H. Grosse. Tensor spline approximation. Linear Algebra and its Applications, 34(??):29–41, December 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580901573.

Groh:1982:SOS

[Gro82] Ulrich Groh. Some observations on the spectra of positive operators on finite-dimensional C^{∗}-algebras. Linear Algebra and its Applications, 42(??):213–222, February 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582901501.

Grone:1985:RDM

[Gro85] Bob Grone. Report on the Dublin matrix theory conference, March 1984: Computation of immanants. Linear Algebra and its Applications, 68(??):252–254, July 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902186.

Grunbaum:1981:TMC

[Grü81] F. Alberto Grünbaum. Toeplitz matrices commuting with tridiagonal matrices. Linear Algebra and its Applications, 40(??):25–36, October 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581901385.

Grunbaum:1982:RHM

[Grü82] F. Alberto Grünbaum. A remark on Hilbert’s matrix. Linear Algebra and its Applications, 43(??):119– 124, March 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582902476.

Garola:1981:CVS

[GS81] C. Garola and L. Solombrino. Commutation in vector spaces over division rings with a conjugation. Linear Algebra and its Applications, 36(??):41–56, March 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581902184.

Goldberg:1982:ONM

[GS82] Moshe Goldberg and E. G. Straus. Operator norms, multiplicativity factors, and C-numerical radii. Linear Algebra and its Applications, 43(??):137–159, March 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958290249X.

Goldberg:1983:MFN

[GS83a] Moshe Goldberg and E. G. Straus. Multiplicativity factors for C-numerical radii. Linear Algebra and its Applications, 54(??):1–16, October 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583902021.

Goldberg:1983:MNM

[GS83b] Moshe Goldberg and E. G. Straus. Multiplicativity of ℓ_{p} norms for matrices. Linear Algebra and its Applications, 52/53(??):351–360, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Graham:1984:AHM

[GS84] R. L. Graham and N. J. A. Sloane. Anti-Hadamard matrices. Linear Algebra and its Applications, 62(??):113– 137, November 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584900909.

Geldenhuys:1987:BRE

[GS87] G. Geldenhuys and C. Sippel. Bounds for the real eigenvalues of a cascade matrix. Linear Algebra and its Applications, 93(??):127–130, August 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/S0024379587903181.

Gatto:1989:BDF

[GS89] R. Gatto and G. Sartori. Biunitarily diagonalizable families of compact linear operators. Linear Algebra and its Applications, 119(??):87–96, July 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589900700.

George:1986:P

[GSV86a] Alan George, G. W. Stewart, and Robert Voigt. Preface. Linear Algebra and its Applications, 77(??):1–2, May 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586901588.

George:1986:PCS

[GSV86b] Alan George, G. W. (Gilbert W.) Stewart, and Robert G. Voigt, editors. Parallel computing: special volume, volume 77 of Linear Algebra and its Applications. Elsevier Science Publishers, Amsterdam, The Netherlands, May 1986. ISSN 0024-3795 (print), 1873-1856 (electronic).

Goldberg:1982:NRA

[GT82] Moshe Goldberg and Eitan Tadmor. On the numerical radius and its applications. Linear Algebra and its Applications, 42(??):263–284, February 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582901550.

Gill:1989:MCE

[GT89a] Doron Gill and Eitan Tadmor. An O(N^{2}) method for computing the eigensystem of N × N symmetric tridiagonal matrices by the divide-and-conquer approach. Linear Algebra and its Applications, 120(??):257–258, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589903856.

Grotschel:1989:MPC

[GT89b] M. Grötschel and K. Truemper. Master polytopes for cycles of binary matroids. Linear Algebra and its Applications, 114/115(??):523–540, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guralnick:1980:NLG

[Gur80] Robert M. Guralnick. A note on the local-global principle for similarity of matrices. Linear Algebra and its Applications, 30(??):241–245, April 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901974.

Guralnick:1981:RTD

[Gur81a] Robert M. Guralnick. Roth’s theorems and decomposition of modules. Linear Algebra and its Applications, 39(??):155–165, August 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581903001.

Guralnick:1981:SML

[Gur81b] Robert M. Guralnick. Similarity of matrices over local rings. Linear Algebra and its Applications, 41(??):161– 174, December 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581900963.

Guralnick:1982:MEI

[Gur82] Robert M. Guralnick. Matrix equivalence and isomorphism of modules. Linear Algebra and its Applications, 43(??):125–136, March 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379582902488.

Guralnick:1985:IFL

[Gur85a] Robert M. Guralnick. Invariants of finite linear groups on relatively free algebras. Linear Algebra and its Applications, 72(??):85–92, December 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585901442.

Guralnick:1985:RTS

[Gur85b] Robert M. Guralnick. Roth’s theorems for sets of matrices. Linear Algebra and its Applications, 71(??):113– 117, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958590240X.

Guralnick:1988:SHM

[Gur88] Robert M. Guralnick. Similarity of holomorphic matrices. Linear Algebra and its Applications, 99(??):85– 96, February 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588901267.

Gustafson:1984:NMI

[Gus84] William H. Gustafson. A note on matrix inversion. Linear Algebra and its Applications, 57(??):71–73, February 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901770.

Genin:1983:LTF

[GVK^{+}83] Yves Genin, Paul Van Dooren, Tom Kailath, Jean-Marc Delosme, and Martin Morf. On Σ-lossless transfer functions and related questions. Linear Algebra and its Applications, 50(??):251–275, April 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583900587.

Galperin:1980:POP

[GW80a] A. M. Galperin and Z. Waksman. On pseudoinverses of operator products. Linear Algebra and its Applications, 33(??):123–131, October 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901019.

Gray:1980:NFP

[GW80b] L. J. Gray and D. G. Wilson. Nonnegative factorization of positive semidefinite nonnegative matrices. Linear Algebra and its Applications, 31(??):119–127, June 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580902128.

Giovagnoli:1985:MAM

[GW85] A. Giovagnoli and H. P. Wynn. G-majorization with applications to matrix orderings. Linear Algebra and its Applications, 67(??):111–135, June 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585901909.

Haack:1984:WDT

[Haa84] Joel K. Haack. When do the transpose and dual agree? Linear Algebra and its Applications, 56(??):

63–72, January 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901149.

Hackbusch:1984:MCS

[Hac84] Wolfgang Hackbusch. Multigrid convergence for a singular perturbation problem. Linear Algebra and its Applications, 58(??):125–145, April 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584902088.

Hadeler:1983:CM

[Had83] K. P. Hadeler. On copositive matrices. Linear Algebra and its Applications, 49(??):79–89, February 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583900952.

Hadjidimos:1984:OSP

[Had84] A. Hadjidimos. The optimal solution to the problem of complex extrapolation of a first-order scheme. Linear Algebra and its Applications, 62(??):241–261, November 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901009.

Hadjisavvas:1986:MSS

[Had86] Nicolas Hadjisavvas. Metrics on the set of states of a W^{∗}-algebra. Linear Algebra and its Applications, 84(??): 281–287, December 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586903204.

Proceedings of the symposium on operator theory (Athens, 1985).

Haley:1980:SBM

[Hal80] Stephen B. Haley. Solution of band matrix equations by projection-recurrence. Linear Algebra and its Applications, 32(??):33–48, August 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580900051.

Hall:1983:LRC

[Hal83] Jonathan I. Hall. Linear representations of cotriangular spaces. Linear Algebra and its Applications, 49(??): 257–273, February 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901088.

Hall:1986:CMK

[Hal86a] Alastair Hall. Conditions for a matrix Kronecker lemma. Linear Algebra and its Applications, 76(??): 271–277, April 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586902284.

Halpin:1986:SI

[Hal86b] Patrick Halpin. Symmetric T -ideals. Linear Algebra and its Applications, 79(??):53–59, July 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902922.

Han:1986:REH

[Han86] S.-P. Han. Relative eigenvalues of Hermitian matrices. Linear Algebra and its Applications, 81(??):75– 88, September 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586902508.

Hartfiel:1980:CIS

[Har80] D. J. Hartfiel. Concerning the interior of the D-stable matrices. Linear Algebra and its Applications, 30(??): 201–207, April 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580901950.

Hartfiel:1981:GTM

[Har81a] D. J. Hartfiel. A general theory of measures for nonnegative matrices. Linear Algebra and its Applications, 35(??): 21–35, February 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902639.

Hartwig:1981:DIC

[Har81b] Robert E. Hartwig. Drazin inverses and canonical forms in M_{n}(Z/h). [A projection method for the solution of a class of singular integral equations whose symbol has zeros of noninteger order]. Linear Algebra and its Applications, 37(??):205–233, April 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581901816.

Hartwig:1982:AWG

[Har82] Robert E. Hartwig. Applications of the Wrońskian and Gram matrices of {t^{i}e^{λ_{k}t}}. Linear Algebra and its Applications, 43(??):229–241, March 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582902567.

Hartwig:1983:RRR

[Har83] Robert E. Hartwig. Roth’s removal rule revisited. Linear Algebra and its Applications, 49(??):91–115, February 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583900964.

Hartfiel:1984:CTS

[Har84a] D. J. Hartfiel. A control theory for stochastic eigenvectors of stochastic matrices through variations of entries. Linear Algebra and its Applications, 56(??):139–160, January 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901204.

Hartwig:1984:GST

[Har84b] R. E. Hartwig. Generalizations of the spectral theorem for matrices. II. matrix polynomials over arbitrary fields. Linear Algebra and its Applications, 61(??):37– 53, September 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584900211.

Hari:1986:CCJ

[Har86a] V. Hari. On the convergence of cyclic Jacobi-like processes. Linear Algebra and its Applications, 81(??):105– 127, September 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586902521.

Hartwig:1986:ROL

[Har86b] R. E. Hartwig. The reverse order law revisited. Linear Algebra and its Applications, 76(??):241–246, April 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902260.

Hartwig:1987:NLM

[Har87a] Robert E. Hartwig. A note on light matrices. Linear Algebra and its Applications, 97(??):153–169, December 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587901479.

Hartwig:1987:RCM

[Har87b] Robert E. Hartwig. Resultants, companion matrices, and Jacobson chains. Linear Algebra and its Applications, 94(??):1–34, September 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587900759.

Hartwig:1988:RBT

[Har88] Robert Hartwig. Review of The Theory of Matrices, second edition, by Peter Lancaster and Miron Tismenetsky. Linear Algebra and its Applications, 108(??):173– 175, September 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588901863.

Hartfiel:1989:CCN

[Har89] D. J. Hartfiel. Cone control on nonnegative control equations. Linear Algebra and its Applications, 117(??): 115–126, May 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589905533.

Hautus:1983:SDO

[Hau83] M. L. J. Hautus. Strong detectability and observers. Linear Algebra and its Applications, 50(??):353–368, April 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583900617.

Haynsworth:1982:RBM

[Hay82] Emilie Haynsworth. Review of Circulant Matrices by Philip Davis. Linear Algebra and its Applications, 43(??): 243–244, March 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582902579.

Hershkowitz:1983:LSM

[HB83] Daniel Hershkowitz and Abraham Berman. Localization of the spectra of P - and P_{0}-matrices. Linear Algebra and its Applications, 52/53(??):383–397, 1983. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hershkowitz:1984:NM

[HB84] Daniel Hershkowitz and Abraham Berman. Notes on ωand τ-matrices. Linear Algebra and its Applications, 58 (??):169–183, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584902106.

Heuvers:1988:CPF

[HCR88] Konrad J. Heuvers, L. J. Cummings, and K. P. S. Bhaskara Rao. A characterization of the permanent function by the Binet–Cauchy theorem. Linear Algebra and its Applications, 101(??):49–72, April 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901425.

Heading:1983:IFE

[Hea83] John Heading. Invariant functions of the elements of the powers of a square matrix. Linear Algebra and its Applications, 55(??):133–145, December 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583901726.

Hellegouarch:1987:PGA

[Hel87] Yves Hellegouarch. Propriétés géométrico-arithmétiques des algèbres à involution. (French) [Geometric and arithmetical properties of algebras with involution]. Linear Algebra and its Applications, 97(??):29–43, December 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587901376.

Helmke:1989:RFB

[Hel89] Uwe Helmke. Rational functions and Bézout forms: a functorial correspondence. Linear Algebra and its Applications, 122/123/124(??):623–640, 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Henrici:1983:GTO

[Hen83] Peter Henrici. A general theory of osculation algorithms for conformal mapping. Linear Algebra and its Applications, 52/53(??):361–382, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hershkowitz:1983:SMH

[Her83] Daniel Hershkowitz. On the spectra of matrices having nonnegative sums of principal minors. Linear Algebra and its Applications, 55(??):81–86, December 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583901672.

Hershkowitz:1986:SAM

[Her86] Daniel Hershkowitz. Stability of acyclic matrices. Linear Algebra and its Applications, 73(??):157–169, January 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902387.

Hershkowitz:1989:MRB

[Her89] Daniel Hershkowitz. A majorization relation between the height and the level characteristics. Linear Algebra and its Applications, 125(??):97–101, December 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589900347.

Hess:1986:SEO

[Hes86] Peter Hess. On the spectrum of elliptic operators with respect to indefinite weights. Linear Algebra and its Applications, 84(??):99–109, December 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586903095. Proceedings

of the symposium on operator theory (Athens, 1985).

Han:1985:ITS

[HF85] S.-P. Han and O. Fujiwara. An inertia theorem for symmetric matrices and its application to nonlinear programming. Linear Algebra and its Applications, 72(??):47– 58, December 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585901417.

Helmke:1989:B

[HF89] U. Helmke and P. A. Fuhrmann. Bézoutians. Linear Algebra and its Applications, 122/123/124(??):1039–1097, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ha:1980:NDF

[HG80] T. T. Ha and J. A. Gibson. A note on the determinant of a functional confluent Vandermonde matrix and controllability. Linear Algebra and its Applications, 30 (??):69–75, April 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580901822.

Hochstadt:1987:HMS

[HG87] Harry Hochstadt and Wallace Goldberg. Hill’s matrices with some vanishing instability intervals. Linear Algebra and its Applications, 97(??):121–126, December 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587901443.

Habibi:1988:RSA

[HG88] Javad Faghih Habibi and William H. Gustafson. Reflexive serial algebras. Linear Algebra and its Applications, 99 (??):217–223, February 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379588901334.

Hernandez:1989:ESM

[HG89] Vicente Hernández and Maite Gassó. Explicit solution of the matrix equation AXB − CXD = E. Linear Algebra and its Applications, 121(??):333–344, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589907088. Linear

algebra and applications (Valencia, 1987).

Hammer:1983:FLS

[HH83] Jacob Hammer and Michael Heymann. Factorization of linear systems: a generalized framework. Linear Algebra and its Applications, 50(??):321–352, April 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583900605.

Herzer:1985:SSV

[HH85] Armin Herzer and Bertram Huppert. Ein Satz von I. Schur über vertauschbare Matrizen. (German) [A theorem of I. Schur on commutative matrices]. Linear Algebra and its Applications, 71(??):151–158, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902423.

Hong:1988:CFM

[HH88] Yoo Pyo Hong and Roger A. Horn. A canonical form for matrices under consimilarity. Linear Algebra and its Applications, 102(??):143–168, April 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588903242.

Haemers:1989:SRG

[HH89] W. H. Haemers and D. G. Higman. Strongly regular graphs with strongly regular decomposition. Linear Algebra and its Applications, 114/115(??):379–398, 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hong:1986:RPH

[HHJ86] Yoo Pyo Hong, Roger A. Horn, and Charles R. Johnson. On the reduction of pairs of Hermitian or symmetric matrices to diagonal form by congruence. Linear Algebra and its Applications, 73(??):213–226, January 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902417.

Helmke:1989:CDS

[HHM89] U. Helmke, D. Hinrichsen, and W. Manthey. A cell decomposition of the space of real Hankel matrices of rank ≤ n and some applications. Linear Algebra and its Applications, 122/123/124(??):331–355, 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hernandez:1986:BBF

[HI86] Vicente G. Hernández and Fernando C. Incertis. A block bidiagonal form for block companion matrices. Linear Algebra and its Applications, 75(??):241–256, March 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586901928.

Hiai:1987:SPL

[Hia87] Fumio Hiai. Similarity preserving linear maps on matrices. Linear Algebra and its Applications, 97(??):127– 139, December 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587901455.

Higham:1987:CRS

[Hig87a] Nicholas J. Higham. Computing real square roots of a real matrix. Linear Algebra and its Applications, 88/89 (??):405–430, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Higman:1987:CA

[Hig87b] D. G. Higman. Coherent algebras. Linear Algebra and its Applications, 93(??):209–239, August 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/S0024379587903260.

Higham:1988:CNS

[Hig88] Nicholas J. Higham. Computing a nearest symmetric positive semidefinite matrix. Linear Algebra and its Applications, 103(??):103–118, May 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902236.

Hilgers:1982:TOR

[Hil82] John W. Hilgers. A theory for optimal regularization in the finite-dimensional case. Linear Algebra and its Applications, 48(??):359–379, December 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582901215.

Heinig:1986:HMG

[HJ86a] Georg Heinig and Uwe Jungnickel. Hankel matrices generated by the Markov parameters of rational functions. Linear Algebra and its Applications, 76(??):121–135, April 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958690217X.

Heinig:1986:LEC

[HJ86b] Georg Heinig and Uwe Jungnickel. Lyapunov equations for companion matrices. Linear Algebra and its Applications, 76(??):137–147, April 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902181.

Hershkowitz:1986:LTM

[HJ86c] Daniel Hershkowitz and Charles R. Johnson. Linear transformations that map the P -matrices into themselves. Linear Algebra and its Applications, 74(??):23– 38, February 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901138.

Hershkowitz:1986:SMM

[HJ86d] Daniel Hershkowitz and Charles R. Johnson. Spectra of matrices with P -matrix powers. Linear Algebra and its Applications, 80(??):159–171, August 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902831.

Hall:1981:NIG

[HK81] Frank J. Hall and Irving J. Katz. Nonnegative integral generalized inverses. Linear Algebra and its Applications, 39(??):23–39, August 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581902871.

Hinrichsen:1983:URC

[HK83] D. Hinrichsen and U. Krause. Unique representation in convex sets by extraction of marked components. Linear Algebra and its Applications, 51(??):73–96, June 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583901507.

Hannah:1983:NFC

[HL83] John Hannah and Thomas J. Laffey. Nonnegative factorization of completely positive matrices. Linear Algebra and its Applications, 55(??):1–9, December 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583901623.

Hartfiel:1984:MHE

[HL84] D. J. Hartfiel and R. Loewy. On matrices having equal corresponding principal minors. Linear Algebra and its Applications, 58(??):147–167, April 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958490209X.

Hsuan:1985:IAS

[HLG85] Francis Hsuan, Patricia Langenberg, and Albert Getson. The {2}-inverse with applications in statistics. Linear Algebra and its Applications, 70(??):241–248, October 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585900552.

Hadeler:1981:RCP

[HM81] K. P. Hadeler and G. Meinardus. On the roots of Cauchy polynomials. Linear Algebra and its Applications, 38(??): 81–102, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581900100.

Han:1984:CCC

[HM84a] S.-P. Han and O. L. Mangasarian. Conjugate cone characterization of positive definite and semidefinite matrices. Linear Algebra and its Applications, 56(??):89– 103, January 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901162.

Hudson:1984:ABM

[HM84b] Richard H. Hudson and Thomas L. Markham. Alfred T. Brauer as a mathematician and teacher. Linear Algebra and its Applications, 59(??):1–17, June 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958490154X.

Hershkowitz:1986:LTW

[HM86] Daniel Hershkowitz and Volker Mehrmann. Linear transformations which map the classes of ω-matrices and τ-matrices into or onto themselves. Linear Algebra and its Applications, 78(??):79–106, June 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586900170.

Hassin:1989:OSP

[HM89a] Refael Hassin and Nimrod Megiddo. On orientations and shortest paths. Linear Algebra and its Applications, 114/115(??):589–602, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Horiguchi:1989:NSM

[HM89b] Susumu Horiguchi and Willard L. Miranker. Noisy sort, a memory-intensive sorting algorithm. Linear Algebra and its Applications, 114/115(??):641–658, 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hershkowitz:1987:EIC

[HMS87] Daniel Hershkowitz, Volker Mehrmann, and Hans Schneider. Eigenvalue interlacing for certain classes of matrices with real principal minors. Linear Algebra and its Applications, 88/89(??):373–403, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hadjidimos:1988:NSC

[HN88] A. Hadjidimos and M. Neumann. A note on the SSOR convergence domain due to Neumaier and Varga. Linear Algebra and its Applications, 107(??):207–217, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588902467. Pro-

ceedings of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Hochstadt:1984:IST

[Hoc84] Harry Hochstadt. An inverse spectral theorem for a Hill’s matrix. Linear Algebra and its Applications, 57(??): 21–30, February 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901733.

Hochwald:1987:UAO

[Hoc87] Scott H. Hochwald. The uniqueness of the adjoint operation. II. Linear Algebra and its Applications, 91(??):

107–119, June 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587900656.

Hoffman:1985:NCK

[Hof85] Alan J. Hoffman. Note on a conjecture of Kojima and Saigal. Linear Algebra and its Applications, 71(??):159– 160, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585902435.

Holbrook:1982:PMC

[Hol82] John A. R. Holbrook. Polynomials in a matrix and its commutant. Linear Algebra and its Applications, 48(??): 293–301, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958290115X.

Holland:1985:RDM

[Hol85] Finbarr Holland. Report on the Dublin Matrix Theory Conference, March 1984: Counterparts of Toeplitz and Hankel operators on C^{n}. Linear Algebra and its Applications, 68(??):263–271, July 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902228.

Holgate:1987:IDG

[Hol87] P. Holgate. The interpretation of derivations in genetic algebras. Linear Algebra and its Applications, 85(??): 75–79, January 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587902096.

Host:1986:TDU

[Hos86] Linda H. Host. Tactical decompositions in uniform normal designs. Linear Algebra and its Applications, 75(??): 105–116, March 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901837.

Howard:1980:LMP

[How80] Ralph Howard. Linear maps that preserve matrices annihilated by a polynomial. Linear Algebra and its Applications, 30(??):167–176, April 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901925.

Howland:1983:SMS

[How83] James Lucien Howland. The sign matrix and the separation of matrix eigenvalues. Linear Algebra and its Applications, 49(??):221–232, February 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583901040.

Hartwig:1981:SMD

[HP81] Robert E. Hartwig and Mohan S. Putcha. Semisimilarity for matrices over a division ring. Linear Algebra and its Applications, 39(??):125–132, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581902962.

Hager:1982:PE

[HP82] William W. Hager and Roger N. Pederson. Perturbation in eigenvalues. Linear Algebra and its Applications, 42(??):39–55, February 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379582901379.

Huylebrouck:1986:GIS

[HP86] D. Huylebrouck and R. Puystjens. Generalized inverses of a sum with a radical element. Linear Algebra and its Applications, 84(??):289–300, December 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586903216. Proceedings

of the symposium on operator theory (Athens, 1985).

Hadjidimos:1983:OBI

[HPS83] Apostolos Hadjidimos, Theodore S. Papatheodorou, and Yiannis G. Saridikis. Optimal block iterative schemes for certain large, sparse, and nonsymmetric linear systems. Linear Algebra and its Applications, 110(??):285– 318, November 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901428.

Hadjidimos:1988:OBI

[HPS88] Apostolos Hadjidimos, Theodore S. Papatheodorou, and Yiannis G. Saridakis. Optimal block iterative schemes for certain large, sparse, and nonsymmetric linear systems. Linear Algebra and its Applications, 110(??):285– 318, 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hinrichsen:1987:WQL

[HPW87] D. Hinrichsen and D. Prätzel-Wolters. A wild quiver in linear systems theory. Linear Algebra and its Applications, 91(??):143–175, June 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379587900681.

Hinrichsen:1989:JCF

[HPW89] D. Hinrichsen and D. Prätzel-Wolters. A Jordan canonical form for reachable linear systems. Linear Algebra and its Applications, 122/123/124(??):489–524, 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hadjidimos:1986:CSG

[HPY86] A. Hadjidimos, A. Psimarni, and A. Yeyios. On the convergence of some generalized iterative methods. Linear Algebra and its Applications, 75(??):117–132, March 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586901849.

Haviv:1984:BDB

[HR84] Moshe Haviv and Uriel G. Rothblum. Bounds on distances between eigenvalues. Linear Algebra and its Applications, 63(??):101–118, December 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901381.

Heinig:1988:ITP

[HR88] Georg Heinig and Karla Rost. On the inverses of Toeplitz-plus-Hankel matrices. Linear Algebra and its Applications, 106(??):39–52, August 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588900213.

Heinig:1989:MRT

[HR89] Georg Heinig and Karla Rost. Matrix representations of Toeplitz-plus-Hankel matrix inverses. Linear Algebra and its Applications, 113(??):65–78, February 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589902863.

Haviv:1987:IMA

[HRR87] Moshe Haviv, Yakov Ritov, and Uriel G. Rothblum. Iterative methods for approximating the subdominant modulus of an eigenvalue of a nonnegative matrix. Linear Algebra and its Applications, 87(??):61–75, March 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587901583.

Hershkowitz:1988:CCM

[HRS88] Daniel Hershkowitz, Uriel G. Rothblum, and Hans Schneider. Characterizations and classifications of Mmatrices using generalized nullspaces. Linear Algebra and its Applications, 109(??):59–69, October 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958890198X.

Hershkowitz:1989:CSG

[HRS89] Daniel Hershkowitz, Uriel G. Rothblum, and Hans Schneider. The combinatorial structure of the generalized nullspace of a block triangular matrix. Linear Algebra and its Applications, 116(??):9–26, April 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589903947.

Hautus:1983:SSS

[HS83] M. L. J. Hautus and L. M. Silverman. System structure and singular control. Linear Algebra and its Applications, 50(??):369–402, April 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583900629.

Hoekstra:1984:LTM

[HS84] A. H. Hoekstra and F. W. Steutel. Limit theorems for Markov chains of finite rank. Linear Algebra and its Applications, 60(??):65–77, August 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900715.

Hershkowitz:1985:LDS

[HS85] Daniel Hershkowitz and Hans Schneider. Lyapunov diagonal semistability of real H-matrices. Linear Algebra and its Applications, 71(??):119–149, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902411.

Hartwig:1986:SCS

[HS86a] Robert E. Hartwig and George P. H. Styan. On some characterizations of the “star” partial ordering for matrices and rank subtractivity. Linear Algebra and its Applications, 82(??):145–161, October 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586901485.

Heath:1986:PGM

[HS86b] M. T. Heath and D. C. Sorensen. A pipelined Givens method for computing the QR factorization of a sparse matrix. Linear Algebra and its Applications, 77(??): 189–203, May 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901680.

Helmke:1987:BMF

[HS87] Uwe Helmke and Mark A. Shayman. The biflag manifold and the fixed points of a unipotent transformation on the flag manifold. Linear Algebra and its Applications, 92(??):125–159, July 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587902540.

Hershkowitz:1988:GNM

[HS88a] Daniel Hershkowitz and Hans Schneider. On the generalized nullspace of M-matrices and Z-matrices. Linear Algebra and its Applications, 106(??):5–23, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588900195.

Hershkowitz:1988:SME

[HS88b] Daniel Hershkowitz and Hans Schneider. Solutions of Zmatrix equations. Linear Algebra and its Applications, 106(??):25–38, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379588900201.

Hara:1989:CSC

[HS89] Shinji Hara and Hak-Kyung Sung. Constraints on sensitivity characteristics in linear multivariable discretetime control systems. Linear Algebra and its Applications, 122/123/124(??):889–919, 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hoskins:1984:CIA

[HT84] W. D. Hoskins and R. S. D. Thomas. Conditions for isonemal arrays on a Cartesian grid. Linear Algebra and its Applications, 57(??):87–103, February 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901794.

Hsu:1988:VAP

[HTRG88] I.-S. Hsu, T. K. Truong, I. S. Reed, and N. Glover. A VLSI architecture for performing finite field arithmetic with reduced table lookup. Linear Algebra and its Applications, 98(??):249–262, January 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958890167X.

Hu:1987:ARM

[Hu87] Hui Hu. An algorithm for rescaling a matrix positive definite. Linear Algebra and its Applications, 96(??):131– 147, November 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587903405.

Hughes-Hallett:1985:TWA

[Hug85] A. J. Hughes-Hallett. Techniques which accelerate the convergence of first order iterations automatically. Linear Algebra and its Applications, 68(??):115–130, July 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902101.

Hunter:1982:GIT

[Hun82] Jeffrey J. Hunter. Generalized inverses and their application to applied probability problems. Linear Algebra and its Applications, 45(??):157–198, June 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958290218X.

Hunter:1986:SDP

[Hun86] Jeffrey J. Hunter. Stationary distributions of perturbed Markov chains. Linear Algebra and its Applications, 82 (??):201–214, October 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379586901527.

Hunter:1988:CGI

[Hun88] Jeffrey J. Hunter. Characterizations of generalized inverses associated with Markovian kernels. Linear Algebra and its Applications, 102(??):121–142, April 1988.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588903230.

Hurkens:1989:EIP

[Hur89] C. A. J. Hurkens. On the existence of an integral potential in a weighted bidirected graph. Linear Algebra and its Applications, 114/115(??):541–553, 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Haynsworth:1981:GIC

[HW81] Emilie Haynsworth and J. R. Wall. Group inverses of certain positive operators. Linear Algebra and its Applications, 40(??):143–159, October 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581901464.

Hill:1987:CPS

[HW87] Richard D. Hill and Steven R. Waters. On the cone of positive semidefinite matrices. Linear Algebra and its Applications, 90(??):81–88, May 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587903077.

Hayden:1988:AMP

[HW88] T. L. Hayden and Jim Wells. Approximation by matrices positive semidefinite on a subspace. Linear Algebra and its Applications, 109(??):115–130, October 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588902029.

Hu:1989:ASI

[HW89] Hui Hu and Qing Wang. On approximate solutions of infinite systems of linear inequalities. Linear Algebra and its Applications, 114/115(??):429–438, 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hwang:1986:MDC

[Hwa86a] Suk Geun Hwang. The monotonicity of and the Doković conjectures on permanents of doubly stochastic matrices.

Linear Algebra and its Applications, 79(??):127–151, July 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958690296X.

Hwang:1986:NCP

[Hwa86b] Suk Geun Hwang. A note on a conjecture on permanents. Linear Algebra and its Applications, 76(??):31–44, April 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902120.

Hwang:1987:CD

[Hwa87] Suk Geun Hwang. On a Conjecture of E. Dittert. Linear Algebra and its Applications, 95(??):161–169, October 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587900322.

Hadjidimos:1980:PEC

[HY80] A. Hadjidimos and A. Yeyios. The principle of extrapolation in connection with the accelerated overrelaxation method. Linear Algebra and its Applications, 30(??): 115–128, April 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580901871.

Imam:1983:TUT

[Ima83] I. N. Imam. Tridiagonal and upper triangular inverse M-matrices. Linear Algebra and its Applications, 55(??): 93–104, December 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901696.

Imam:1984:SCI

[Ima84] I. N. Imam. The Schur complement and the inverse M-matrix problem. Linear Algebra and its Applications, 62(??):235–240, November 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900995.

Isaacson:1982:ECI

[IS82] Dean Isaacson and E. Seneta. Ergodicity for countable inhomogeneous Markov chains. Linear Algebra and its Applications, 48(??):37–44, December 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582900970.

Ivanov:1989:ASD

[IS89] A. A. Ivanov and S. V. Shpectorov. The association schemes of dual polar spaces of type ^{2}A_{2d−1}(p^{f}) are characterized by their parameters if d ≥ 3. Linear Algebra and its Applications, 114/115(??):133–139, 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Isaacs:1980:AMA

[Isa80] I. M. Isaacs. Automorphisms of matrix algebras over commutative rings. Linear Algebra and its Applications, 31 (??):215–231, June 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580902219.

Ishibashi:1983:GCG

[Ish83] Hiroyuki Ishibashi. Generation of classical groups over finite fields. Linear Algebra and its Applications, 51 (??):9–15, June 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901465.

Ishihara:1987:SWD

[Ish87] Kazuo Ishihara. Strong and weak discrete maximum principles for matrices associated with elliptic problems. Linear Algebra and its Applications, 88/89(??):431–448, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ipsen:1986:CDL

[ISS86] Ilse C. F. Ipsen, Youcef Saad, and Martin H. Schultz. Complexity of dense-linear-system solution on a multiprocessor ring. Linear Algebra and its Applications, 77 (??):205–239, May 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901692.

Special volume on parallel computing.

Ito:1982:BDR

[Ito82] Tatsuro Ito. Bipartite distance-regular graphs of valency three. Linear Algebra and its Applications, 46(??): 195–213, August 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582900350.

Jain:1984:OLS

[Jai84] S. K. Jain. Overdetermined linear systems satisfying nonnegativity constraints. Linear Algebra and its Applications, 57(??):205–213, February 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901885.

Jodar:1989:KPC

[JAK89] L. Jódar and H. Abou-Kandil. Kronecker products and coupled matrix Riccati differential systems. Linear Algebra and its Applications, 121(??):39–51, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589906903. Linear

algebra and applications (Valencia, 1987).

Jamison:1981:FLH

[Jam81] Robert E. Jamison. Finding little hyperplanes in bigger ones. Linear Algebra and its Applications, 35(??): 11–19, February 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902627.

James:1983:RAU

[Jam83] Donald G. James. Representations of Anisotropic unitary groups. II. Linear Algebra and its Applications, 110(??):271–284, November 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583901416.

James:1988:RAU

[Jam88] Donald G. James. Representations of anisotropic unitary groups. II. Linear Algebra and its Applications, 110 (??):271–284, 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:1985:STE

[JB85] Charles R. Johnson and Wayne W. Barrett. Spanningtree extensions of the Hadamard–Fischer inequalities. Linear Algebra and its Applications, 66(??):177–193, April 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585901314.

Johnson:1987:RBH

[JE87] Charles R. Johnson and Ludwig Elsner. The relationship between Hadamard and conventional multiplication for positive definite matrices. Linear Algebra and its Applications, 92(??):231–240, July 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587902606.

Jeffries:1986:QSC

[Jef86] Clark Jeffries. Qualitative stability of certain nonlinear systems. Linear Algebra and its Applications, 75(??): 133–144, March 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901850.

Jeffries:1988:ESC

[Jef88] Clark Jeffries. Eigenvalues, stability, and color tests. Linear Algebra and its Applications, 107(??):65–76, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588902388. Pro-

ceedings of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Jain:1980:NMH

[JG80] S. K. Jain and V. K. Goel. Nonnegative matrices having nonnegative Drazin pseudoinverses. Linear Algebra and its Applications, 29(??):173–183, February 1980.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580902384.

Jiang:1982:SVN

[Jia82] Er Xiong Jiang. On spectral variation of a nonnormal matrix. Linear Algebra and its Applications, 42(??):223– 241, February 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582901513.

Jinadasa:1988:AMO

[Jin88] K. G. Jinadasa. Applications of the matrix operators vech and vec. Linear Algebra and its Applications, 101 (??):73–79, April 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588901437.

Jonckheere:1989:STL

[JJS89] Edmond A. Jonckheere, Jyh-Ching Juang, and Leonard M. Silverman. Spectral theory of the linear-quadratic and H^{∞} problems. Linear Algebra and its Applications, 122/123/124(??):273–300, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jeffries:1987:QSL

[JKvdD87] Clark Jeffries, Victor Klee, and Pauline van den Driessche. Qualitative stability of linear systems. Linear Algebra and its Applications, 87(??):1–48, March 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958790156X.

Johnson:1985:LBS

[JKW85] Charles R. Johnson, Ravinder Kumar, and Henry Wolkowicz. Lower bounds for the spread of a matrix. Linear Algebra and its Applications, 71(??):161– 173, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585902447.

James:1987:PPM

[JL87] N. S. James and R. Lidl. Permutation polynomials on matrices. Linear Algebra and its Applications, 96(??):181–190, November 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587903430.

Jonckheere:1989:RPR

[JM89a] Edmond Jonckheere and Chingwo Ma. Recursive partial realization from the combined sequence of Markov parameters and moments. Linear Algebra and its Applications, 122/123/124(??):565–590, 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jonckheere:1989:SHI

[JM89b] Edmond Jonckheere and Chingwo Ma. A simple Hankel interpretation of the Berlekamp–Massey algorithm. Linear Algebra and its Applications, 125(??):65–76, December 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589900323.

Jongen:1987:ISC

[JMRT87] H. Th. Jongen, T. Möbert, J. Rückmann, and K. Tammer. On inertia and Schur complement in optimization. Linear Algebra and its Applications, 95(??):97– 109, October 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587900280.

Jodar:1986:BVP

[Jód86] Lucas Jódar. Boundary value problems for secondorder differential operator equations. Linear Algebra and its Applications, 83(??):29–38, November 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902636.

Jodar:1987:BVP

[Jód87] Lucas Jódar. Boundary value problems and Cauchy problem for the second-order Euler operator differential equation. Linear Algebra and its Applications, 91 (??):1–12, June 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587900565.

Jodar:1988:ADO

[Jód88a] Lucas Jódar. Algebraic and differential operator equations. Linear Algebra and its Applications, 102(??): 35–53, April 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588903199.

Jodar:1988:ESS

[Jód88b] Lucas Jódar. Explicit solutions for second order operator differential equations with two boundary value conditions. Linear Algebra and its Applications, 103(??):73–86, May 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588902212.

Joe:1985:ODC

[Joe85] Harry Joe. An ordering of dependence for contingency tables. Linear Algebra and its Applications, 70(??):89– 103, October 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958590045X.

Johnsen:1980:SMS

[Joh80] E. C. Johnsen. Stochastic matrices. I. similarity via stochastic transforms. Linear Algebra and its Applications, 29(??):185–193, February 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580902396.

Johnson:1981:NRP

[Joh81] Charles R. Johnson. Numerical ranges of principal submatrices. Linear Algebra and its Applications, 37(??): 23–34, April 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581901646.

Johnson:1982:IM

[Joh82a] Charles R. Johnson. Inverse M-matrices. Linear Algebra and its Applications, 47(??):195–216, October 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582902385.

Johnson:1982:GM

[Joh82b] Gregery A. Johnson. A generalization of N-matrices. Linear Algebra and its Applications, 48(??):201–217, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582901082.

Johnson:1983:SPI

[Joh83] Charles R. Johnson. Sign patterns of inverse nonnegative matrices. Linear Algebra and its Applications, 55(??): 69–80, December 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901660.

Johnson:1984:PMS

[Joh84] Gregery A. Johnson. Partitioned matrices satisfying certain null space properties. Linear Algebra and its Applications, 57(??):75–85, February 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901782.

Johnson:1985:IM

[Joh85] Gregery A. Johnson. Inverse N_{0}-matrices. Linear Algebra and its Applications, 64(??):215–222, January 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902782.

Johnson:1987:CPC

[Joh87] Charles R. Johnson. Closure properties of certain positivity classes of matrices under various algebraic operations. Linear Algebra and its Applications, 97(??):243– 247, December 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587901534.

Johnson:1988:CMA

[Joh88] Charles R. Johnson. Combinatorial matrix analysis: An overview. Linear Algebra and its Applications, 107(??): 3–15, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588902339.

Proceedings of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Johnson:1989:GTL

[Joh89] Charles R. Johnson. A Gershgorin-type lower bound for the smallest singular value. Linear Algebra and its Applications, 112(??):1–7, January 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589905831.

Jonckheere:1984:NBS

[Jon84] Edmond A. Jonckheere. New bound on the sensitivity of the solution of the Lyapunov equation. Linear Algebra and its Applications, 60(??):57–64, August 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584900703.

Johnson:1987:SDM

[JORvdD87] Charles R. Johnson, D. D. Olesky, Brian Robertson, and P. van den Driessche. Sign determinacy of M-matrix minors. Linear Algebra and its Applications, 91(??): 133–141, June 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958790067X.

Jeter:1981:NNR

[JP81] M. W. Jeter and W. C. Pye. A note on nonnegative rank factorizations. Linear Algebra and its Applications, 38(??):171–173, June 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581900185.

Jeter:1982:NPM

[JP82] Melvyn W. Jeter and Wallace C. Pye. Nonnegative (s, t)-potent matrices. Linear Algebra and its Applications, 45(??):109–121, June 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582902130.

Jeter:1984:SPM

[JP84] Melvyn W. Jeter and Wallace C. Pye. Some properties of Q-matrices. Linear Algebra and its Applications, 57(??): 169–180, February 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958490185X.

Jeter:1987:LCP

[JP87] Melvyn W. Jeter and Wallace C. Pye. The linear complementarity problem and a subclass of fully semimonotone matrices. Linear Algebra and its Applications, 87(??): 243–256, March 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587901704.

Jeter:1988:SPM

[JP88a] Melvyn W. Jeter and Wallace C. Pye. Structure properties of W matrices. Linear Algebra and its Applications, 111(??):219–229, December 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588900614.

Johnson:1988:ISH

[JP88b] Charles R. Johnson and Stephen Pierce. Inequalities for single-hook immanants. Linear Algebra and its Applications, 102(??):55–79, April 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588903205.

Johnson:1981:EIP

[JR81] Charles R. Johnson and Herbert A. Robinson. Eigenvalue inequalities for principal submatrices. Linear Algebra and its Applications, 37(??):11–22, April 1981.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581901634.

Johnson:1989:PST

[JS89] Charles R. Johnson and Tracy A. Summers. The potentially stable tree sign patterns for dimensions less than five. Linear Algebra and its Applications, 126(??): 1–13, December 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589900025.

Jungnickel:1989:PS

[Jun89] Dieter Jungnickel. Partial spreads over Z_{q}. Linear Algebra and its Applications, 114/115(??):95–102, 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jury:1982:CSC

[Jur82] E. I. Jury. Comments on the stability criterion of Liénard and Chipart. Linear Algebra and its Applications, 47(??):169–171, October 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582902348.

Johnson:1982:SPN

[JUW82] Charles R. Johnson, Frank Uhlig, and Dan Warner. Sign patterns, nonsingularity, and the solvability of Ax = b. Linear Algebra and its Applications, 47(??):1–9, October 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958290221X.

Jeffries:1988:EMT

[JvdD88] Clark Jeffries and P. van den Driessche. Eigenvalues of matrices with tree graphs. Linear Algebra and its Applications, 101(??):109–120, April 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901462.

Jea:1983:SGC

[JY83] Kang C. Jea and David M. Young. On the simplification of generalized conjugate- gradient methods for nonsymmetrizable linear systems. Linear Algebra and its Applications, 52/53(??):399–417, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Joubert:1986:NSC

[JY86] Wayne D. Joubert and David M. Young. Necessary and sufficient conditions for the simplification of generalized conjugate gradient algorithms. Linear Algebra and its Applications, 86/87(??):449–485, 1986.

Joubert:1987:NSC

[JY87] Wayne D. Joubert and David M. Young. Necessary and sufficient conditions for the simplification of generalized conjugate-gradient algorithms. Linear Algebra and its Applications, 88/89(??):449–485, 1987. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jiang:1985:NSA

[JZ85] Er Xiong Jiang and Zhen Yue Zhang. A new shift of the QL algorithm for irreducible symmetric tridiagonal matrices. Linear Algebra and its Applications, 65(??):261–272, February 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585901016.

Kabe:1985:MNP

[Kab85] D. G. Kabe. On a multivariate normal probability minimization problem. Linear Algebra and its Applications, 67(??):103–109, June 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585901892.

Kamen:1988:PZL

[Kam88] Edward W. Kamen. The poles and zeros of a linear time-varying system. Linear Algebra and its Applications, 98(??):263–289, January 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901681.

Kamps:1989:CPL

[Kam89] U. Kamps. Chebyshev polynomials and least squares estimation based on one-dependent random variables. Linear Algebra and its Applications, 112(??):217–230, January 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589905971.

Karzanov:1989:PRU

[Kar89] A. V. Karzanov. Polyhedra related to undirected multicommodity flows. Linear Algebra and its Applications, 114/115(??):293–328, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kass:1989:SCF

[Kas89] Seymour Kass. Spaces of closest fit. Linear Algebra and its Applications, 117(??):93–97, May 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589905508.

Kaufman:1987:GHT

[Kau87] Linda Kaufman. The generalized Householder transformation and sparse matrices. Linear Algebra and its Applications, 90(??):221–234, May 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587903144.

Kaykobad:1985:PSC

[Kay85a] M. Kaykobad. Positive solutions of a class of linear systems. Linear Algebra and its Applications, 72(??):97– 105, December 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585901466.

Kaykobad:1985:PSP

[Kay85b] M. Kaykobad. Positive solutions of positive linear systems. Linear Algebra and its Applications, 64(??):133– 140, January 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958590271X.

Kaykobad:1987:NFM

[Kay87] M. Kaykobad. On nonnegative factorization of matrices. Linear Algebra and its Applications, 96(??):27– 33, November 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958790334X.

Kessler:1985:MTG

[KB85] Ofra Kessler and Abraham Berman. Matrices with a transitive graph and inverse M-matrices. Linear Algebra and its Applications, 71(??):175–185, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902459.

Kaszkurewicz:1989:QSD

[KB89a] Eugenius Kaszkurewicz and Amit Bhaya. Qualitative stability of discrete-time systems. Linear Algebra and its Applications, 117(??):65–71, May 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589905478.

Kiers:1989:OCT

[KB89b] Henk A. L. Kiers and Jos M. F. Ten Berge. Optimality conditions for the trace of certain matrix products. Linear Algebra and its Applications, 126(??):125–134, December 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589900104.

Kailath:1986:FMF

[KBM86] T. Kailath, A. Bruckstein, and D. Morgan. Fast matrix factorizations via discrete transmission lines. Linear Algebra and its Applications, 75(??):1–25, March 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586901783.

Kemeny:1981:GFM

[Kem81] John G. Kemeny. Generalization of a fundamental matrix. Linear Algebra and its Applications, 38(??): 193–206, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581900203.

Kershaw:1983:AAE

[Ker83] D. Kershaw. QD algorithms and algebraic eigenvalue problems. Linear Algebra and its Applications, 54(??): 53–75, October 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583902057.

Kershaw:1987:AMF

[Ker87] D. Kershaw. An analysis of the method of L. Fox and L. Hayes for the factorization of a polynomial. Linear Algebra and its Applications, 86(??):179–187, February 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958790293X.

Kestenband:1982:FPG

[Kes82] Barbu C. Kestenband. Finite projective geometries that are incidence structures of caps. Linear Algebra and its Applications, 48(??):303–313, December 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582901161.

Kautsky:1983:CJM

[KG83] J. Kautský and G. H. Golub. On the calculation of Jacobi matrices. Linear Algebra and its Applications, 52/53 (??):439–455, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Karcanias:1989:NSC

[KG89] N. Karcanias and C. Giannakopoulos. Necessary and sufficient conditions for zero assignment by constant squaring down. Linear Algebra and its Applications, 122/123/124(??):415–446, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Khargonekar:1983:SPP

[KGÖ83] Pramod P. Khargonekar, Tryphon T. Georgiou, and A. Bülent Özgüler. Skew-prime polynomial matrices: the polynomial-model approach. Linear Algebra and its Applications, 50(??):403–435, April 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583900630.

Kaszkurewicz:1984:TCM

[KH84] E. Kaszkurewicz and L. Hsu. On two classes of matrices with positive diagonal solutions to the Lyapunov equation. Linear Algebra and its Applications, 59(??): 19–27, June 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901551.

Khatri:1980:PMI

[Kha80] C. G. Khatri. Powers of matrices and idempotency. Linear Algebra and its Applications, 33(??):57–65, October 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958090097X.

Khatri:1983:GLI

[Kha83] C. G. Khatri. A generalization of Lavoie’s inequality concerning the sum of idempotent matrices. Linear Algebra and its Applications, 54(??):97–108, October 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583902082.

Khatri:1985:NIM

[Kha85] C. G. Khatri. A note on idempotent matrices. Linear Algebra and its Applications, 70(??):185–195, October 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585900515.

Kielbasinski:1987:NRE

[Kie87] Andrzej Kiełbasiński. A note on rounding error analysis of Cholesky factorization. Linear Algebra and its Applications, 88/89(??):487–494, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kittappa:1987:GRM

[Kit87] R. K. Kittappa. A generalization of the rotation matrix and related results. Linear Algebra and its Applications, 92(??):251–258, July 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958790262X.

Kittaneh:1989:CAV

[Kit89] Fuad Kittaneh. On the continuity of the absolute value map in the Schatten classes. Linear Algebra and its Applications, 118(??):61–68, June 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589905715.

Kailath:1986:MBT

[KK86] T. Kailath and I. Koltracht. Matrices with block Toeplitz inverses. Linear Algebra and its Applications, 75(??): 145–153, March 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901862.

Karcanias:1989:PEH

[KK89] N. Karcanias and G. Kalogeropoulos. Projective equivalence of homogeneous binary polynomials. Linear Algebra and its Applications, 121(??):433–453, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589907155. Linear

algebra and applications (Valencia, 1987).

Klingler:1986:SSM

[KL86] Lee Klingler and Lawrence S. Levy. Sweeping-similarity of matrices. Linear Algebra and its Applications, 75(??): 67–104, March 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901825.

Kuhn:1987:PAB

[KL87] Dieter Kuhn and Rainer Löwen. Piecewise affine bijections of R^{n}, and the equation Sx^{+} − T x^{−} = y. Linear Algebra and its Applications, 96(??):109–129, November 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587903399.

Kasami:1988:BWD

[KL88a] Tadao Kasami and Shu Lin. The binary weight distribution of the extended (2^{m}, 2^{m} − 4) code of the Reed–Solomon code over GF(2^{m}) with generator polynomial (x − α)(x − α^{2})(x − α^{3}). Linear Algebra and its Applications, 98(??):291–307, January 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901693.

Koltracht:1988:GSP

[KL88b] I. Koltracht and P. Lancaster. Generalized Schur parameters and the effects of perturbations. Linear Algebra and its Applications, 105(??):109–129, July 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588900079.

Klein:1982:TMT

[Kle82] D. J. Klein. Treediagonal matrices and their inverses. Linear Algebra and its Applications, 42(??):109–117, February 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582901422.

Klee:1987:RSM

[Kle87] Victor Klee. Recursive structure of S-matrices and an O(m^{2}) algorithm for recognizing strong sign solvability. Linear Algebra and its Applications, 96(??):233– 247, November 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587903478.

Kleiner:1988:PPO

[Kle88] Mark Kleiner. Pairs of partially ordered sets of tame representation type. Linear Algebra and its Applications, 104 (??):103–115, June 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588903114.

Klee:1984:SR

[KLM84] Victor Klee, Richard Ladner, and Rachel Manber. Signsolvability revisited. Linear Algebra and its Applications, 59(??):131–157, June 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379584901642.

Kravvaritis:1986:CR

[KLM86] D. Kravvaritis, P. Lancaster, and J. Maroulas. Conference report. Linear Algebra and its Applications, 84(??): 373–420, December 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586903289.

Keller-McNulty:1985:EFC

[KMK85] S. Keller-McNulty and W. J. Kennedy. Error-free computation of a reflexive generalized inverse. Linear Algebra and its Applications, 67(??):157–167, June 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585901934.

Kankaanpaa:1987:IIS

[KMV87] Hilkka Kankaanpää, Jorma Kaarlo Merikoski, and Ari Virtanen. Improving an inequality for sums of elements of matrix powers. Linear Algebra and its Applications, 92(??):225–229, July 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958790259X.

Knuppel:1987:PTI

[KN87a] Frieder Knüppel and Klaus Nielsen. On products of two involutions in the orthogonal group of a vector space. Linear Algebra and its Applications, 94(??):209– 216, September 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587900917.

Knuppel:1987:PI

[KN87b] Frieder Knüppel and Klaus Nielsen. Products of involutions in O^{+}(V). Linear Algebra and its Applications, 94(??):217–222, September 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587900929.

Kautsky:1989:RPA

[KNC89] J. Kautský, N. K. Nichols, and E. K.-W. Chu. Robust pole assignment in singular control systems. Linear Algebra and its Applications, 121(??):9–37, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589906897. Linear

algebra and applications (Valencia, 1987).

Karlin:1985:SMP

[KO85a] Samuel Karlin and Friedemann Ost. Some monotonicity properties of Schur powers of matrices and related inequalities. Linear Algebra and its Applications, 68 (??):47–65, July 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585902071.

Khargonekar:1985:STA

[KÖ85b] Pramod P. Khargonekar and A. Bülent Özgüler. Systemtheoretic and algebraic aspects of the rings of stable and proper stable rational functions. Linear Algebra and its Applications, 66(??):123–167, April 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585901296.

Kobilinsky:1985:CRD

[Kob85] André Kobilinsky. Confounding in relation to duality of finite Abelian groups. Linear Algebra and its Applications, 70(??):321–347, October 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900631.

Koecher:1985:GII

[Koe85] Max Koecher. The generalized inverse of integral matrices. Linear Algebra and its Applications, 71(??):187– 198, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585902460.

Koh:1984:RBQ

[Koh84] Kwangil Koh. Review of Ternary Quadratic Forms and Norms, edited by Olga Taussky. Linear Algebra and its Applications, 57(??):291–293, February 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901952.

Koh:1985:MRF

[Koh85] Kwangil Koh. On the matrix ring over a finite field. Linear Algebra and its Applications, 66(??):195–197, April 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585901326.

Komiya:1983:NSC

[Kom83] Hidetoshi Komiya. Necessary and sufficient conditions for multivariate majorization. Linear Algebra and its Applications, 55(??):147–154, December 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583901738.

Konig:1986:ECO

[Kön86] Hermann König. Eigenvalues of compact operators with applications to integral operators. Linear Algebra and its Applications, 84(??):111–122, December 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586903101. Proceedings

of the symposium on operator theory (Athens, 1985).

Kosir:1988:CJC

[Koš88] Tomaž Košir. Common Jordan chains of matrices. Linear Algebra and its Applications, 99(??):253–258, February 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901358.

Kovarik:1980:MFP

[Kov80] Zdislav V. Kovarik. Manifolds of frames of projectors. Linear Algebra and its Applications, 31(??):151–158, June 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580902153.

Kovacs:1987:SEP

[Kov87] Amos Kovacs. Some enumeration problems for matrices over a finite field. Linear Algebra and its Applications, 94(??):223–236, September 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587900930.

Kaneko:1980:SLC

[KP80] Ikuyo Kaneko and Jong Shi Pang. Some n by dn linear complementarity problems. Linear Algebra and its Applications, 34(??):297–319, December 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901706.

Khargonekar:1986:PMF

[KP86] Pramod P. Khargonekar and Kameshwar R. Poolla. Polynomial matrix-fraction representations for linear timevarying systems. Linear Algebra and its Applications, 80 (??):1–37, August 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586902752.

Kociecki:1989:NCL

[KP89] M. Kociȩcki and K. M. Przyłuski. On the number of controllable linear systems over a finite field. Linear Algebra and its Applications, 122/123/124(??):115–122, 1989.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:1981:ECM

[KR81a] K. H. Kim and F. W. Roush. Expressions for certain minors and permanents. Linear Algebra and its Applications, 41(??):93–97, December 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581900902.

Kim:1981:GAL

[KR81b] K. H. Kim and F. W. Roush. Generating all linear transformations. Linear Algebra and its Applications, 37(??): 97–101, April 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581901701.

Karlin:1983:MCM

[KR83] Samuel Karlin and Yosef Rinott. M-matrices as covariance matrices of multinormal distributions. Linear Algebra and its Applications, 52/53(??):419–438, 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kramarz:1981:APM

[Kra81] Luis Kramarz. Algebraic perturbation methods for the solution of singular linear systems. Linear Algebra and its Applications, 36(??):79–88, March 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581902202.

Krafft:1983:MOP

[Kra83] Olaf Krafft. A matrix optimization problem. Linear Algebra and its Applications, 51(??):137–142, June 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583901544.

Krauter:1987:GDB

[Krä87] Arnold R. Kräuter. On the greatest distance between two permanental roots of a matrix. Linear Algebra and its Applications, 93(??):39–55, August 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/S0024379587903107.

Krishna:1984:CSS

[Kri84] Lala B. Krishna. On the convergence of the symmetric successive overrelaxation method. Linear Algebra and its Applications, 56(??):185–194, January 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901241.

Kumar:1982:GEL

[KRS82] Vijaya Kumar, S. B. Rao, and N. M. Singhi. Graphs with eigenvalues at least −2. Linear Algebra and its Applications, 46(??):27–42, August 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582900234.

Kojima:1980:PMP

[KS80] M. Kojima and R. Saigal. A property of matrices with positive determinants. Linear Algebra and its Applications, 30(??):1–8, April 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901779.

Kovarik:1984:CSB

[KS84] Zdislav V. Kovarik and Nagwa Sherif. Characterization of similarities between two n-frames of projectors. Linear Algebra and its Applications, 57(??):57–69, February 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901769.

Kovarik:1985:PIS

[KS85] Zdislav V. Kovarik and Nagwa Sherif. Perturbation of invariant subspaces. Linear Algebra and its Applications, 64(??):93–113, January 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902691.

Khatri:1986:OMC

[KS86a] C. G. Khatri and K. R. Shah. Orthogonality in multiway classifications. Linear Algebra and its Applications, 82(??):215–224, October 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379586901539.

Kravvaritis:1986:PMM

[KS86b] Dimitrios Kravvaritis and Nicolaos Stavrakakis. Perturbations of maximal monotone random operators. Linear Algebra and its Applications, 84(??):301–310, December 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586903228. Pro-

ceedings of the symposium on operator theory (Athens, 1985).

Kovarik:1988:GNG

[KS88] Zdislav V. Kovarik and Nagwa Sherif. Geodesics and near-geodesics in the manifolds of projector frames. Linear Algebra and its Applications, 99(??):259–277, February 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958890136X.

Kung:1981:CSL

[Kun81] Joseph P. S. Kung. The cycle structure of a linear transformation over a finite field. Linear Algebra and its Applications, 36(??):141–155, March 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581902275.

Karcanias:1989:DDD

[KW89a] N. Karcanias and D. R. Wilson. Decentralized diagonal dynamic stabilization of linear multivariable systems. Linear Algebra and its Applications, 121(??):455– 474, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589907167.

Linear algebra and applications (Valencia, 1987).

Kravitsky:1989:SRI

[KW89b] N. Kravitsky and Z. Waksman. On some resultant identities. Linear Algebra and its Applications, 122/123/124 (??):3–21, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kwong:1988:DSC

[Kwo88] Man Kam Kwong. On the definiteness of the solutions of certain matrix equations. Linear Algebra and its Applications, 108(??):177–197, September 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901875.

Kwong:1989:SRM

[Kwo89] Man Kam Kwong. Some results on matrix monotone functions. Linear Algebra and its Applications, 118(??): 129–153, June 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589905776.

Limebeer:1988:ITA

[LA88] D. J. N. Limebeer and B. D. O. Anderson. An interpolation theory approach to H^{∞} controller degree bounds. Linear Algebra and its Applications, 98(??): 347–386, January 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588901723.

Livsic:1989:SSC

[LA89] M. S. Livšic and Y. Avishai. A study of solitonic combinations based on the theory of commuting nonselfadjoint operators. Linear Algebra and its Applications, 122/123/124 (??):357–414, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:1968:LAA

[LAA68] Linear algebra and its applications, 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

North-Holland Publishing Co., Amsterdam, The Netherlands.

Laffey:1980:STP

[Laf80] Thomas J. Laffey. Simultaneous triangularization of a pair of matrices whose commutator has rank two. Linear Algebra and its Applications, 29(??):195–203, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580902402.

Laffey:1981:AGT

[Laf81a] Thomas J. Laffey. Algebras generated by two idempotents. Linear Algebra and its Applications, 37(??): 45–53, April 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958190166X.

Laffey:1981:CGF

[Laf81b] Thomas J. Laffey. A criterion for generating full matrix algebras. Linear Algebra and its Applications, 37(??): 191–198, April 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581901798.

Laffey:1981:NMS

[Laf81c] Thomas J. Laffey. A note on matrix subalgebras containing a full Jordan block. Linear Algebra and its Applications, 37(??):187–189, April 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581901786.

Laffey:1983:CKC

[Laf83a] Thomas J. Laffey. A counterexample to Kippenhahn’s conjecture on Hermitian pencils. Linear Algebra and its Applications, 51(??):179–182, June 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583901593.

Laffey:1983:IGS

[Laf83b] Thomas J. Laffey. Irredundant generating sets for matrix algebras. Linear Algebra and its Applications, 52/53 (??):457–478, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Laffey:1985:MDM

[Laf85a] Thomas J. Laffey. The minimal dimension of maximal commutative subalgebras of full matrix algebras. Linear Algebra and its Applications, 71(??):199–212, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902472.

Laffey:1985:RDM

[Laf85b] Thomas J. Laffey. Report on the Dublin matrix theory conference, March 1984: Products of skew-symmetric matrices. Linear Algebra and its Applications, 68(??): 249–251, July 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585902174.

Laffey:1986:SRS

[Laf86] Thomas J. Laffey. Simultaneous reduction of sets of matrices under similarity. Linear Algebra and its Applications, 84(??):123–138, December 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586903113. Proceedings

of the symposium on operator theory (Athens, 1985).

Laffey:1989:FIM

[Laf89] Thomas J. Laffey. Factorizations of integer matrices as products of idempotents and nilpotents. Linear Algebra and its Applications, 120(??):81–93, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589903716. Proceedings

of the Fourth Haifa Matrix Theory Conference (Haifa,

Lancaster:1981:RBF

[Lan81] Peter Lancaster. Review of Minimal Factorization of Matrix and Operator Functions, by H. Bart, I. Gohberg, and M. A. Kaashoek. Linear Algebra and its Applications, 39(??):295–296, August 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581903128.

Lancaster:1986:CED

[Lan86a] Peter Lancaster. Common eigenvalues, divisors, and multiples of matrix polynomials: a review. Linear Algebra and its Applications, 84(??):139–160, December 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586903125. Proceedings

of the symposium on operator theory (Athens, 1985).

Lancaster:1986:P

[Lan86b] Peter Lancaster. Preface. Linear Algebra and its Applications, 84(??):1–2, December 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379586903046.

Laszlo:1980:BID

[Lás80]  Lajos László.  A bandstructure-invariant deflation

for real symmetric matrices. Linear Algebra and its Applications, 33(??):9–13, October 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580900944.

Laszlo:1981:MMP

[Lás81] L. László. Matrix methods for polynomials. Linear Algebra and its Applications, 36(??):129–131, March 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581902251.

Laszlo:1982:IPB

[Lás82] Lajos László. Imaginary part bounds on polynomial zeros. Linear Algebra and its Applications, 44(??):173– 180, April 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582900131.

Lavoie:1980:DII

[Lav80] J. L. Lavoie. A determinantal inequality involving the Moore–Penrose inverse. Linear Algebra and its Applications, 31(??):77–80, June 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379580902086.

Lawrence:1982:OMM

[Law82] Jim Lawrence. Oriented matroids and multiply ordered sets. Linear Algebra and its Applications, 48(??):1– 12, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582900945.

Lee:1980:CHS

[Lee80] Anna Lee. Centro-Hermitian and skew-centro-Hermitian matrices. Linear Algebra and its Applications, 29(??): 205–210, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580902414.

Lee:1989:SWS

[Lee89] Jon Lee. Subspaces with well-scaled frames. Linear Algebra and its Applications, 114/115(??):21–56, 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Leitman:1981:SFP

[Lei81] M. J. Leitman. On the spectra of a family of positive matrices. Linear Algebra and its Applications, 37(??): 251–258, April 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958190183X.

Lei:1989:CAI

[Lei89a] Li Lei. Convergence of asynchronous iteration with arbitrary splitting form. Linear Algebra and its Applications, 113(??):119–127, February 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589902905.

Leite:1989:PGC

[Lei89b] F. Silva Leite. Pairs of generators for compact real forms of the classical Lie algebras. Linear Algebra and its Applications, 121(??):123–133, August 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589906964. Linear alge-

bra and applications (Valencia, 1987).

Lempel:1988:CSS

[Lem88] Abraham Lempel. Characterization and synthesis of self-complementary normal bases in finite fields. Linear Algebra and its Applications, 98(??):331–346, January 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901711.

Leong:1988:NUB

[Leó88] Tuck Sang Leóng. A note on upper bounds on the maximum modulus of subdominant eigenvalues of nonnegative matrices. Linear Algebra and its Applications, 106 (??):1–4, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588900183.

Leron:1980:MMD

[Ler80] Uri Leron. Matrix methods in decompositions of modules. Linear Algebra and its Applications, 31(??):159–171, June 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580902165.

Leuze:1986:PTS

[Leu86] Michael R. Leuze. Parallel triangularization of substructured finite element problems. Linear Algebra and its Applications, 77(??):241–258, May 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586901709.

Lewis:1982:ICH

[Lew82] D. W. Lewis. The isometry classification of Hermitian forms over division algebras. Linear Algebra and its Applications, 43(??):245–272, March 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582902580.

Lewis:1985:DIS

[Lew85a] D. W. Lewis. A determinantal identity for skewfields. Linear Algebra and its Applications, 71(??):213– 217, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585902484.

Lewis:1985:RDM

[Lew85b] D. W. Lewis. Report on the Dublin matrix theory conference, March 1984: Hermitian forms, full rings, and von Neumann regular matrices. Linear Algebra and its Applications, 68(??):258–260, July 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902204.

Lewin:1989:IM

[Lew89a] Mordechai Lewin. On inverse M-matrices. Linear Algebra and its Applications, 118(??):83–94, June 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589905740.

Lewkowicz:1989:BSV

[Lew89b] Izchak Lewkowicz. Bounds for the singular values of a matrix with nonnegative eigenvalues. Linear Algebra and its Applications, 112(??):29–37, January 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589905867.

Lewkowicz:1989:REJ

[Lew89c] Izchak Lewkowicz. Remarks on equality in Johnson’s lower bounds for the smallest singular value of a matrix. Linear Algebra and its Applications, 120(??):39– 46, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589903686.

Proceedings of the Fourth Haifa Matrix Theory Conference (Haifa, 1988).

Li:1986:DNR

[Li86a] Chi Kwong Li. The decomposable numerical radius and numerical radius of a compound matrix. Linear Algebra and its Applications, 76(??):45–58, April 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902132.

Li:1986:HNR

[Li86b] Chi Kwong Li. On the higher numerical radius and spectral norm. Linear Algebra and its Applications, 80(??): 55–70, August 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586902776.

Li:1987:GSR

[Li87] Chi Kwong Li. A generalization of spectral radius, numerical radius, and spectral norm. Linear Algebra and its Applications, 90(??):105–118, May 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587903090.

Li:1988:MSE

[Li88a] Chi-Kwong Li. Matrices with some extremal properties. Linear Algebra and its Applications, 101(??): 255–267, April 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958890153X.

Li:1988:CBF

[Li88b] Ren Cang Li. A converse to the Bauer–Fike type theorem. Linear Algebra and its Applications, 109(??):167– 178, October 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958890208X.

Li:1989:RSM

[Li89] Wen Li. On regular splittings of an M-matrix. Linear Algebra and its Applications, 113(??):159–172, February 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589902942.

Lickteig:1985:TTR

[Lic85] Thomas Lickteig. Typical tensorial rank. Linear Algebra and its Applications, 69(??):95–120, August 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585900709.

Lin:1986:RIE

[Lin86] Wen-Wei Lin. On reducing infinite eigenvalues of regular pencils by a nonequivalence transformation. Linear Algebra and its Applications, 78(??):207–231, June 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958690025X.

Lin:1987:ACE

[Lin87a] Wen Wei Lin. An algorithm for computing the eigenstructure of a regular matrix polynomial. Linear Algebra and its Applications, 91(??):195–211, June 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587900711.

Lin:1987:NMC

[Lin87b] Wen Wei Lin. A new method for computing the closedloop eigenvalues of a discrete-time algebraic Riccati equation. Linear Algebra and its Applications, 96(??):157– 180, November 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587903429.

Lindqvist:1987:NMP

[Lin87c] Bo Henry Lindqvist. Nonnegative multiplicative processes reaching stationarity in finite time. Linear Algebra and its Applications, 86(??):75–90, February 1987.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587902886.

Lin:1988:CKC

[Lin88] Wen Wei Lin. The computation of the Kronecker canonical form of an arbitrary symmetric pencil. Linear Algebra and its Applications, 103(??):41–71, May 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588902200.

Lindqvist:1989:APP

[Lin89] Bo Henry Lindqvist. Asymptotic properties of powers of nonnegative matrices, with applications. Linear Algebra and its Applications, 114/115(??):555–588, 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lipman:1980:LIG

[Lip80] Marc J. Lipman. The linear index of a group representation. Linear Algebra and its Applications, 31(??): 43–54, June 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580902049.

Lipitakis:1986:ARF

[Lip86] Elias A. Lipitakis. Approximate root-free factorization techniques for solving elliptic difference equations in three space variables. Linear Algebra and its Applications, 76(??):247–269, April 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902272.

Liu:1988:DMT

[Liu88] Kang Man Liu. Decomposition of matrices into three involutions. Linear Algebra and its Applications, 111(??): 1–24, December 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958890047X.

Lancaster:1984:FFS

[LLT84] P. Lancaster, L. Lerer, and M. Tismenetsky. Factored forms for solutions of AX − XB = C and X − AXB = C in companion matrices. Linear Algebra and its Applications, 62(??):19–49, November 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900867.

Longsine:1980:SRQ

[LM80] D. E. Longsine and S. F. McCormick. Simultaneous Rayleigh-quotient minimization methods for Ax = λBx. Linear Algebra and its Applications, 34(??):195– 234, December 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580901664.

Layton:1987:CDA

[LM87] W. Layton and T. D. Morley. On central difference approximations to general second order elliptic equations. Linear Algebra and its Applications, 97(??):65– 75, December 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587901406.

Lancaster:1988:SPS

[LM88] P. Lancaster and J. Maroulas. Selective perturbation of spectral properties of vibrating systems using feedback. Linear Algebra and its Applications, 98(??):309– 330, January 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958890170X.

Lewin:1980:IMP

[LN80] Mordechai Lewin and Michael Neumann. On the inverse M-matrix problem for (0, 1)-matrices. Linear Algebra and its Applications, 30(??):41–50, April 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580901792.

Langer:1985:SIR

[LN85] Heinz Langer and Branko Najman. Some interlacing results for indefinite Hermitian matrices. Linear Algebra and its Applications, 69(??):131–154, August 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585900722.

Luesink:1989:SBS

[LN89] Rob Luesink and Henk Nijmeijer. On the stabilization of bilinear systems via constant feedback. Linear Algebra and its Applications, 122/123/124(??):457–474, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Landau:1981:BEC

[LO81] H. J. Landau and A. M. Odlyzko. Bounds for eigenvalues of certain stochastic matrices. Linear Algebra and its Applications, 38(??):5–15, June 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581900033.

Loewy:1980:QFC

[Loe80] Raphael Loewy. Quadratic forms on completely symmetric spaces. Linear Algebra and its Applications, 29(??): 211–216, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580902426.

Loewy:1983:DNS

[Loe83] Raphael Loewy. Determinants of nonprincipal submatrices of normal matrices. Linear Algebra and its Applications, 52/53(??):497–507, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Loewy:1984:DNS

[Loe84] R. Loewy. Determinants of nonprincipal submatrices of positive semidefinite matrices. Linear Algebra and its Applications, 56(??):1–16, January 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901095.

Loewy:1986:PMD

[Loe86] Raphael Loewy. Principal minors and diagonal similarity of matrices. Linear Algebra and its Applications, 78 (??):23–64, June 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586900157.

Loewy:1988:DFG

[Loe88] Raphael Loewy. Dimension of faces generated by certain positive linear operators. Linear Algebra and its Applications, 105(??):199–207, July 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588900134.

Loewy:1989:LTW

[Loe89a] Raphael Loewy. Linear transformations which preserve or decrease rank. Linear Algebra and its Applications, 121(??):151–161, August 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589906988. Linear alge-

bra and applications (Valencia, 1987).

Loewy:1989:FGC

[Loe89b] Raphael Loewy. On faces generated by certain positive linear transformations. Linear Algebra and its Applications, 118(??):95–106, June 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589905752.

London:1981:MPC

[Lon81] David London. Monotonicity of permanents of certain doubly stochastic circulant matrices. Linear Algebra and its Applications, 37(??):235–249, April 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581901828.

London:1982:VWC

[Lon82] David London. On the van der Waerden conjecture and zeros of polynomials. Linear Algebra and its Applications, 45(??):35–41, June 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582902099.

Longstaff:1984:PLI

[Lon84] W. E. Longstaff. Picturing the lattice of invariant subspaces of a nilpotent complex matrix. Linear Algebra and its Applications, 56(??):161–168, January 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901216.

Longstaff:1988:TM

[Lon88] W. E. Longstaff. On tridiagonalization of matrices. Linear Algebra and its Applications, 109(??):153–163, October 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588902066.

Lorenz:1981:RED

[Lor81] Jens Lorenz. Resonance in equations with a diagonal field. Linear Algebra and its Applications, 38(??):103–107, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900112.

Luk:1980:QPM

[LP80] Franklin T. Luk and Marcello Pagano. Quadratic programming with M-matrices. Linear Algebra and its Applications, 33(??):15–40, October 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580900956.

Loebl:1981:SRC

[LP81] Richard I. Loebl and Vern I. Paulsen. Some remarks on C^{∗}-convexity. Linear Algebra and its Applications, 35(??): 63–78, February 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902664.

Lancaster:1982:LDJ

[LP82a] P. Lancaster and M. V. Pattabhiraman. The local determination of Jordan bases for H-selfadjoint operators. Linear Algebra and its Applications, 48(??):191– 199, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582901070.

Lefevre-Percsy:1982:GDA

[LP82b] Christiane Lefèvre-Percsy. Geometries with dual affine planes and symplectic quadrics. Linear Algebra and its Applications, 42(??):31–37, February 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582901367.

Longhi:1989:GAS

[LPC89] S. Longhi, A. Perdon, and G. Conte. Geometric and algebraic structure at infinity of discrete time linear periodic systems. Linear Algebra and its Applications, 122/123/124(??):245–271, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Luk:1987:FUO

[LQ87] Franklin T. Luk and San Zheng Qiao. A fast but unstable orthogonal triangularization technique for Toeplitz matrices. Linear Algebra and its Applications, 88/89(??):495– 506, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lau:1981:CBA

[LR81] K. K. Lau and W. O. J. Riha. Characterization of best approximations in normed linear spaces of matrices by elements of finite-dimensional linear subspaces. Linear Algebra and its Applications, 35(??):109–120, February 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581902688.

Lutz:1985:ISF

[LS85] D. A. Lutz and R. Schäfke. On the identification and stability of formal invariants for singular differential equations. Linear Algebra and its Applications, 72(??):1– 46, December 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585901405.

Li:1987:HMG

[LS87] T. Y. Li and T. Sauer. Homotopy method for generalized eigenvalue problems Ax = λBx. Linear Algebra and its Applications, 91(??):65–74, June 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587900607.

Lovasz:1989:ORC

[LSS89] L. Lovász, M. Saks, and A. Schrijver. Orthogonal representations and connectivity of graphs. Linear Algebra and its Applications, 114/115(??):439–454, 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lypchuk:1988:WSM

[LST88] Tanya A. Lypchuk, Malcolm C. Smith, and Allen Tannenbaum. Weighted sensitivity minimization: general plants in H^{∞} and rational weights. Linear Algebra and its Applications, 109(??):71–90, October 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901991.

Lancaster:1983:ICS

[LT83a] P. Lancaster and M. Tismenetsky. Inertia characteristics of selfadjoint matrix polynomials. Linear Algebra and its Applications, 52/53(??):479–496, 1983. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:1983:DBS

[LT83b] Chi-Kwong Li and Nam-Kiu Tsing. Duality between some linear preserver problems. II. Isometries with respect to c-special norms and matrices with fixed singular values. Linear Algebra and its Applications, 110(??):181–212, November 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583901362.

Loewy:1986:CFL

[LT86] Raphael Loewy and Bit Shun Tam. Complementation in the face lattice of a proper cone. Linear Algebra and its Applications, 79(??):195–207, July 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902995.

Lerer:1988:GBM

[LT88a] L. Lerer and M. Tismenetsky. Generalized Bézoutian and matrix equations. Linear Algebra and its Applications, 99 (??):123–160, February 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379588901292.

Li:1988:DBS

[LT88b] Chi-Kwong Li and Nam-Kiu Tsing. Duality between some linear preserver problems. II. isometries with respect to c-spectral norms and matrices with fixed singular values. Linear Algebra and its Applications, 110(??):181– 212, 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:1989:NRD

[LT89] Chi-Kwong Li and Nam-Kiu Tsing. The numerical range of derivations. Linear Algebra and its Applications, 119 (??):97–119, July 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589900712.

Li:1983:LOP

[LTT83] Chi-Kwong Li, Bit-Shun Tam, and Nam-Kiu Tsing. Linear operators preserving the (p, q)-numerical range. Linear Algebra and its Applications, 110(??):75–89, November 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901337.

Li:1988:LOP

[LTT88] Chi-Kwong Li, Bit Shun Tam, and Nam Kiu Tsing. Linear operators preserving the (p, q)-numerical range. Linear Algebra and its Applications, 110(??):75–89, 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lubachevsky:1987:SIS

[Lub87] Boris D. Lubachevsky. The structure of the inverse to the Sylvester resultant matrix. Linear Algebra and its Applications, 85(??):191–202, January 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587902163.

Luk:1981:CPS

[Luk81] Franklin T. Luk. The community problem for Stieltjes matrices. Linear Algebra and its Applications, 38(??): 161–169, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581900173.

Luk:1986:TPA

[Luk86] Franklin T. Luk. A triangular processor array for computing singular values. Linear Algebra and its Applications, 77(??):259–274, May 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379586901710.

Lutwak:1984:PMP

[Lut84] Erwin Lutwak. On power means of positive quadratic forms. Linear Algebra and its Applications, 57(??):13– 19, February 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901721.

Lou:1988:CSB

[LVWC88] Sheldon X.-C. Lou, George C. Verghese, Alan S. Willsky, and Pamela G. Coxson. Conditions for scale-based decompositions in singularly perturbed systems. Linear Algebra and its Applications, 98(??):387–413, January 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901735.

Lawrence:1981:UOM

[LW81] Jim Lawrence and Louis Weinberg. Unions of oriented matroids. Linear Algebra and its Applications, 41(??): 183–200, December 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581900987.

Lancaster:1988:ISP

[LY88] Peter Lancaster and Qiang Ye. Inverse spectral problems for linear and quadratic matrix pencils. Linear Algebra and its Applications, 107(??):293–309, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902522. Proceedings

of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Ma:1984:RCS

[Ma84] S. L. Ma. On rational circulants satisfying A^{m} = dl + λJ. Linear Algebra and its Applications, 62(??):155– 161, November 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584900922.

Ma:1985:MPA

[Ma85a] S. L. Ma. On multipliers of polynomial addition sets. Linear Algebra and its Applications, 69(??):231–238, August 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585900783.

Ma:1985:PAS

[Ma85b] S. L. Ma. Polynomial addition sets and polynomial digraphs. Linear Algebra and its Applications, 69(??): 213–230, August 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585900771.

MacLeod:1988:FDR

[Mac88] Allan J. MacLeod. Finite-dimensional regularization with nonidentity smoothing matrices. Linear Algebra and its Applications, 111(??):191–207, December 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588900596.

MacRae:1989:TDA

[Mac89] R. E. MacRae. On the theory of division algebras. Linear Algebra and its Applications, 119(??):121–128, July 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589900724.

Madsen:1983:COE

[Mad83] Niel K. Madsen. Cyclic odd-even reduction for symmetric circulant matrices. Linear Algebra and its Applications, 51(??):17–35, June 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901477.

Maddocks:1988:RQF

[Mad88] J. H. Maddocks. Restricted quadratic forms, inertia theorems, and the Schur complement. Linear Algebra and its Applications, 108(??):1–36, September 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901772.

Magnus:1987:RT

[Mag87] Jan R. Magnus. A representation theorem for (trA^{p})^{1/p}. Linear Algebra and its Applications, 95(??):127–134, October 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587900309.

Malabre:1989:GLS

[Mal89] Michel Malabre. Generalized linear systems: geometric and structural approaches. Linear Algebra and its Applications, 122/123/124(??):591–621, 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mangasarian:1981:STA

[Man81] O. L. Mangasarian. A stable theorem of the alternative: an extension of the Gordan theorem. Linear Algebra and its Applications, 41(??):209–223, December 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581901002.

Manber:1982:GTA

[Man82] Rachel Manber. Graph-theoretical approach to qualitative solvability of linear systems. Linear Algebra and its Applications, 48(??):457–470, December 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582901276.

Mandel:1984:CAI

[Man84] Jan Mandel. A convergence analysis of the iterative aggregation method with one parameter. Linear Algebra and its Applications, 59(??):159–169, June 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901654.

Markham:1981:CUP

[Mar81a] Thomas L. Markham. Class G, and a unification problem. Linear Algebra and its Applications, 39(??):197– 204, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581903037.

MarquesdeSa:1981:ISH

[Mar81b] E. Marques de Sá. On the inertia of sums of Hermitian matrices. Linear Algebra and its Applications, 37(??): 143–159, April 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581901749.

Martin:1981:FCC

[Mar81c] D. H. Martin. Finite criteria for conditional definiteness of quadratic forms. Linear Algebra and its Applications, 39(??):9–21, August 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958190286X.

Marek:1984:PRC

[Mar84a] Ivo Marek. Perron roots of a convex combination of a cone preserving map and its adjoint. Linear Algebra and its Applications, 58(??):185–200, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584902118.

Markham:1984:RBA

[Mar84b] Thomas Lowell Markham. Review of Recent Applications of Generalized Inverses: edited by S. L. Campbell. Linear Algebra and its Applications, 57(??):295–296, February 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901964.

Martignon:1984:DSM

[Mar84c] L. F. Martignon. Doubly stochastic matrices with prescribed positive spectrum. Linear Algebra and its Applications, 61(??):11–13, September 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900181.

Maroulas:1985:FMP

[Mar85] J. Maroulas. Factorization of matrix polynomials with multiple roots. Linear Algebra and its Applications, 69 (??):9–32, August 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585900679.

Maroulas:1986:TFM

[Mar86a] J. Maroulas. A theorem on the factorization of matrix polynomials. Linear Algebra and its Applications, 84(??): 311–322, December 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958690323X.

Proceedings of the symposium on operator theory (Athens, 1985).

Martins:1986:CMO

[Mar86b] M. Madalena Martins. On the convergence of the modified overrelaxation method. Linear Algebra and its Applications, 81(??):55–73, September 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902491.

MarquesdeSa:1988:CLI

[Mar88] E. Marques de Sá. A convexity lemma on the interlacing inequalities for invariant factors. Linear Algebra and its Applications, 109(??):107–113, October 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588902017.

Mathar:1985:BEF

[Mat85] Rudolf Mathar. The best Euclidian fit to a given distance matrix in prescribed dimensions. Linear Algebra and its Applications, 67(??):1–6, June 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585901818.

Mathew:1986:HNP

[Mat86] Thomas Mathew. Hermitian and nonnegativity preserving subspaces. Linear Algebra and its Applications, 82 (??):225–235, October 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379586901540.

Mayer:1984:CPI

[May84] Günter Mayer. On the convergence of powers of interval matrices. Linear Algebra and its Applications, 58(??): 201–216, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958490212X.

Mayer:1985:CNS

[May85] Günter Mayer. On the convergence of the Neumann series in interval analysis. Linear Algebra and its Applications, 65(??):63–70, February 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900874.

Mayer:1987:ACF

[May87] Günter Mayer. On the asymptotic convergence factor of the total step method in interval computation. Linear Algebra and its Applications, 85(??):153–164, January 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587902138.

Maybee:1988:SPN

[May88] John S. Maybee. Some possible new directions for combinatorial matrix analysis. Linear Algebra and its Applications, 107(??):23–40, August 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902352. Proceedings

of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Miranker:1981:DCG

[MC81] W. L. Miranker and I.-L. Chern. Dichotomy and conjugate gradients in the stiff initial value problem. Linear Algebra and its Applications, 36(??):57–77, March 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581902196.

Micali:1985:DAG

[MCCF85] Artibano Micali, Tânia M. M. Campos, Maria C. Costa e Silva, and Silvia M. M. Ferreira. Dérivations dans les algèbres gamétiques. II. (French) [Derivations in gametic algebras. II]. Linear Algebra and its Applications, 64(??):175–181, January 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902745.

Micali:1989:DAG

[MCCF89] Artibano Micali, Tânia M. M. Campos, Maria C. Costa e Silva, and Silvia M. M. Ferreira. Dérivations dans les algèbres gamétiques. III. (French) [derivations in gametic algebras. III]. Linear Algebra and its Applications, 113(??):79–99, February 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589902875.

McDonald:1980:SZC

[McD80] John N. McDonald. On some zero configurations associated with the van der Waerden conjecture. Linear Algebra and its Applications, 32(??):137–155, August 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580900130.

McFall:1980:EDT

[McF80] J. D. McFall. On elementary divisors of the tensor product of two matrices. Linear Algebra and its Applications, 33(??):67–86, October 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580900981.

McKay:1981:EED

[McK81] Brendan D. McKay. The expected eigenvalue distribution of a large regular graph. Linear Algebra and its Applications, 40(??):203–216, October 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581901506.

Meek:1983:ITB

[Mee83] D. S. Meek. The inverses of Toeplitz band matrices. Linear Algebra and its Applications, 49(??):117– 129, February 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583900976.

Meek:1987:CIN

[Mee87] D. S. Meek. Comparison of infinity norms of matrix inverses. Linear Algebra and its Applications, 96(??):149– 156, November 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587903417.

Mehrmann:1984:GFI

[Meh84a] Volker Mehrmann. On a generalized Fan inequality. Linear Algebra and its Applications, 58(??):235–245, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584902143.

Mehrmann:1984:CMC

[Meh84b] Volker Mehrmann. On classes of matrices containing M-matrices and Hermitian positive semidefinite matrices. Linear Algebra and its Applications, 58(??):217– 234, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584902131.

Mehrmann:1984:DM

[Meh84c] Volker Mehrmann. On the LU decomposition of V - matrices. Linear Algebra and its Applications, 61(??): 175–186, September 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379584900296.

Mehrmann:1989:EUS

[Meh89] Volker Mehrmann. Existence, uniqueness, and stability of solutions to singular linear quadratic optimal control problems. Linear Algebra and its Applications, 121(??): 291–331, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589907076.

Linear algebra and applications (Valencia, 1987).

Mena:1986:DLP

[Men86] R. A. Mena. On discriminant-like polynomials. Linear Algebra and its Applications, 75(??):193–201, March 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586901898.

Merris:1980:MPC

[Mer80a] Russell Merris. Manifestations of Polya’s Counting Theorem. Linear Algebra and its Applications, 32(??):209– 232, August 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958090018X.

Merris:1980:PIA

[Mer80b] Russell Merris. Pattern inventories associated with symmetry classes of tensors. Linear Algebra and its Applications, 29(??):225–230, February 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958090244X.

Merikoski:1981:ONS

[Mer81a] Jorma Kaarlo Merikoski. On operator norms of submatrices. Linear Algebra and its Applications, 36(??): 173–183, March 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902305.

Merris:1981:IBS

[Mer81b] Russell Merris. Induced bases of symmetry classes of tensors. Linear Algebra and its Applications, 39(??): 103–110, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902949.

Merikoski:1984:TSE

[Mer84] Jorma Kaarlo Merikoski. On the trace and the sum of elements of a matrix. Linear Algebra and its Applications, 60(??):177–185, August 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900788.

Meshulam:1989:TEM

[Mes89] Roy Meshulam. On two extremal matrix problems. Linear Algebra and its Applications, 114/115(??):261–271, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Meyer:1984:RBN

[Mey84] Carl Meyer. Review of Non-Negative Matrices and Markov Chains, by E. Seneta. Linear Algebra and its Applications, 61(??):287–288, September 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900375.

Meyer:1989:UPE

[Mey89] Carl D. Meyer. Uncoupling the Perron eigenvector problem. Linear Algebra and its Applications, 114/115(??): 69–94, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Muller-Funk:1985:DBL

[MFPW85] U. Müller-Funk, F. Pukelsheim, and H. Witting. On the duality between locally optimal tests and optimal experimental designs. Linear Algebra and its Applications, 67 (??):19–34, June 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585901831.

Muhlbach:1985:GSI

[MG85] G. Mühlbach and M. Gasca. A generalization of Sylvester’s identity on determinants and some applications. Linear Algebra and its Applications, 66(??): 221–234, April 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958590134X.

Mattis:1981:CBM

[MH81] Michael P. Mattis and Harry Hochstadt. On the construction of band matrices from spectral data. Linear Algebra and its Applications, 38(??):109–119, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900124.

Milaszewicz:1983:MJL

[Mil83] J. P. Milaszewicz. On modified Jacobi linear operators. Linear Algebra and its Applications, 51(??):127–136, June 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583901532.

Milaszewicz:1987:IJG

[Mil87] J. P. Milaszewicz. Improving Jacobi and Gauss–Seidel Iterations. Linear Algebra and its Applications, 93(??): 161–170, August 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0024379587903211.

Miller:1988:ISM

[Mil88] D. F. Miller. The iterative solution of the matrix equation XA + BX + C = 0. Linear Algebra and its Applications, 105(??):131–137, July 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588900080.

Minc:1985:RFP

[Min85] Henryk Minc. Recurrence formulas for permanents of (0, 1)-circulants. Linear Algebra and its Applications, 71 (??):241–265, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379585902502.

Minc:1987:CRB

[Min87a] Henryk Minc. On a conjecture of R. A. Brualdi. Linear Algebra and its Applications, 94(??):61–66, September 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587900784.

Minc:1987:PCP

[Min87b] Henryk Minc. Permanental compounds and permanents of (0, 1)-circulants. Linear Algebra and its Applications, 86(??):11–42, February 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587902850.

Mirman:1983:AME

[Mir83] Boris Mirman. Antieigenvalues: method of estimation and calculation. Linear Algebra and its Applications, 49(??):247–255, February 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583901076.

Missirlis:1985:PIS

[Mis85] Nikolaos M. Missirlis. A parallel iterative system solver. Linear Algebra and its Applications, 65(??):25– 44, February 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585900850.

Mitra:1982:SDR

[Mit82] Sujit Kumar Mitra. Simultaneous diagonalization of rectangular matrices. Linear Algebra and its Applications, 47(??):139–150, October 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582902312.

Mitra:1984:ME

[Mit84] Sujit Kumar Mitra. The matrix equations AX = C, XB = D. Linear Algebra and its Applications, 59(??): 171–181, June 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901666.

Mitra:1986:MPO

[Mit86] Sujit Kumar Mitra. The minus partial order and the shorted matrix. Linear Algebra and its Applications, 83 (??):1–27, November 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379586902624.

Mitra:1987:GIS

[Mit87] Sujit Kumar Mitra. On group inverses and the sharp order. Linear Algebra and its Applications, 92(??): 17–37, July 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587902485.

Mitra:1988:IPM

[Mit88] Sujit Kumar Mitra. Infimum of a pair of matrices. Linear Algebra and its Applications, 105(??):163–182, July 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588900109.

Martin:1981:CMD

[MJ81] D. H. Martin and D. H. Jacobson. Copositive matrices and definiteness of quadratic forms subject to homogeneous linear inequality constraints. Linear Algebra and its Applications, 35(??):227–258, February 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581902767.

Martins:1988:SNR

[MK88] M. Madalena Martins and Lala B. Krishna. Some new results on the convergence of the SSOR and USSOR methods. Linear Algebra and its Applications, 106(??): 185–193, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588900274.

Mehrmann:1989:LTW

[MK89] Volker Mehrmann and Gerd M. Krause. Linear transformations which leave controllable multi-input descriptor systems controllable. Linear Algebra and its Applications, 120(??):47–64, August 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589903698. Proceedings

of the Fourth Haifa Matrix Theory Conference (Haifa, 1988).

Mucha:1983:PSC

[ML83] M. Mucha and T. Lulek. Permutational symmetry of Clebsch–Gordan coefficients for the metacyclic groups. Linear Algebra and its Applications, 51(??):49–71, June 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583901490.

Muirhead:1985:EFC

[ML85] Robb J. Muirhead and Pui Lam Leung. Estimating functions of canonical correlation coefficients. Linear Algebra and its Applications, 70(??):173–183, October 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585900503.

Martinez-Legaz:1983:SH

[MLS83] Juan-Enrique Martinez-Legaz and Ivan Singer. The structure of hemispaces in R^{n}. Linear Algebra and its Applications, 110(??):117–179, November 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583901350.

Martinez-Legaz:1987:LS

[MLS87] Juan-Enrique Martínez-Legaz and Ivan Singer. Lexicographical separation in r^{n}. Linear Algebra and its Applications, 90(??):147–163, May 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587903120.

Martinez-Legaz:1988:SH

[MLS88] Juan-Enrique Martínez-Legaz and Ivan Singer. The structure of hemispaces in r^{n}. Linear Algebra and its Applications, 110(??):117–179, 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Marcus:1980:SIN

[MM80] Marvin Marcus and Kenneth Moore. A subdeterminant inequality for normal matrices. Linear Algebra and its Applications, 31(??):129–143, June 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958090213X.

Marcus:1981:DFC

[MM81] Marvin Marcus and Kenneth Moore. A determinant formulation of the Cauchy–Schwarz inequality. Linear Algebra and its Applications, 36(??):111–127, March 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958190224X.

McAsey:1982:PEI

[MM82] Michael J. McAsey and Paul S. Muhly. On projective equivalence of invariant subspace lattices. Linear Algebra and its Applications, 43(??):167–179, March 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582902518.

Mukherjee:1988:SPP

[MM88] Bishwa Nath Mukherjee and Sadhan Samar Maiti. On some properties of positive definite Toeplitz matrices and their possible applications. Linear Algebra and its Applications, 102(??):211–240, April 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588903266.

Mallol:1989:ABI

[MM89] Cristian Mallol and Artibano Micali. Sur les algèbres de Bernstein. II. (French) [On Bernstein algebras. II]. Linear Algebra and its Applications, 117(??):11–19, May 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589905430.

Miller:1987:SOM

[MN87] Valerie A. Miller and Michael Neumann. Successive overrelaxation methods for solving the rank deficient linear least squares problem. Linear Algebra and its Applications, 88/89(??):533–557, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mneimne:1989:FTP

[Mne89] R. Mneimné. Formule de Taylor pour le déterminant et deux applications. (French) [Taylor’s formula for the determinant, and two applications]. Linear Algebra and its Applications, 112(??):39–47, January 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589905879.

Markham:1985:CDI

[MNP85] T. L. Markham, M. Neumann, and R. J. Plemmons. Convergence of a direct-iterative method for large-scale least-squares problems. Linear Algebra and its Applications, 69(??):155–167, August 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900734.

Mohan:1984:NPM

[MNR84] S. R. Mohan, M. Neumann, and K. G. Ramamurthy. Nonnegativity of principal minors of generalized inverses of M-matrices. Linear Algebra and its Applications, 58(??): 247–259, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584902155.

Mitra:1986:PSM

[MO86] Sujit Kumar Mitra and Patrick L. Odell. On parallel summability of matrices. Linear Algebra and its Applications, 74(??):239–255, February 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586901266.

Mohar:1987:DSI

[MO87] Bojan Mohar and Matjaž Omladič. Divisors and the spectrum of infinite graphs. Linear Algebra and its Applications, 91(??):99–106, June 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587900644.

Mohar:1982:SIG

[Moh82] Bojan Mohar. The spectrum of an infinite graph. Linear Algebra and its Applications, 48(??):245–256, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582901112.

Mohar:1988:IIG

[Moh88]  Bojan Mohar.  Isoperimetric inequalities, growth,

and the spectrum of graphs. Linear Algebra and its Applications, 103(??):119–131, May 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902248.

Monahan:1984:RBL

[Mon84] John F. Monahan. Review of Fitting Linear Models: an Application of Conjugate Gradient Algorithms, by Allen McIntosh. Linear Algebra and its Applications, 61(??): 293, September 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584900405.

Moore:1982:DBN

[Moo82] Kenneth Moore. Decomposable bilinear numerical radii. Linear Algebra and its Applications, 42(??):253–260, February 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582901537.

Morgan:1985:EAA

[Mor85] Frank Morgan. The exterior algebra Λ^{k}R^{n} and area minimization. Linear Algebra and its Applications, 66 (??):1–28, April 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585901235.

Morris:1988:CMC

[Mor88] Walter D. Morris, Jr. Counterexamples to Q-matrix conjectures. Linear Algebra and its Applications, 111(??): 135–145, December 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588900559.

Miranker:1980:MA

[MP80] W. L. Miranker and V. Ya. Pan. Methods of aggregation. Linear Algebra and its Applications, 29(??):231– 257, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580902451.

Mitra:1982:SME

[MP82] Sujit Kumar Mitra and Madan L. Puri. Shorted matrices—an extended concept and some applications. Linear Algebra and its Applications, 42(??):57–79, February 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582901380.

Markus:1983:CJS

[MP83] A. S. Markus and E. È. Parilis. The change of the Jordan structure of a matrix under small perturbations. Linear Algebra and its Applications, 54(??):139–152, October 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583902100. Trans-

lated from the Russian by Chandler Davis.

Mendelsohn:1989:SIP

[MP89a] N. S. Mendelsohn and R. Padmanabhan. Self-inscribed polygons with vertices on nonsingular cubic curves. Linear Algebra and its Applications, 114/115(??):603–611, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Micchelli:1989:URC

[MP89b] Charles A. Micchelli and Hartmut Prautzsch. Uniform refinement of curves. Linear Algebra and its Applications, 114/115(??):841–870, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Martin:1981:DCP

[MPJ81] D. H. Martin, M. J. D. Powell, and D. H. Jacobson. On a decomposition of conditionally positive-semidefinite matrices. Linear Algebra and its Applications, 39(??): 51–59, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902895.

Miao:1983:GMP

[MR83] Jian-Ming Miao and Donald W. Robinson. Group and Moore–Penrose inverses of regular morphisms with kernel and cokernel. Linear Algebra and its Applications, 110(??):263–270, November 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583901404.

Melter:1984:LEI

[MR84] Robert A. Melter and Sergiu Rudeanu. Linear equations and interpolation in Boolean algebra. Linear Algebra and its Applications, 57(??):31–40, February 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901745.

Maybee:1988:SPG

[MR88a] John S. Maybee and Daniel J. Richman. Some properties of GM-matrices and their inverses. Linear Algebra and its Applications, 107(??):219–236, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902479. Proceedings

of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Miao:1988:GMP

[MR88b] Jian Ming Miao and Donald W. Robinson. Group and Moore–Penrose inverses of regular morphisms with kernel and cokernel. Linear Algebra and its Applications, 110 (??):263–270, 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Merris:1981:PPG

[MRW81] Russell Merris, Kenneth R. Rebman, and William Watkins. Permanental polynomials of graphs. Linear Algebra and its Applications, 38(??):273–288, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900264.

Malek-Shahmirzadi:1980:ANL

[MS80a] Massoud Malek-Shahmirzadi. Application of Neumann’s lemma to the minimum and characteristic polynomials. Linear Algebra and its Applications, 29(??):217– 224, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580902438.

Malek-Shahmirzadi:1980:CDZ

[MS80b] Massoud Malek-Shahmirzadi. On a conjecture by D. Ž. Djoković. Linear Algebra and its Applications, 30(??): 177–182, April 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580901937.

Mandel:1983:LCP

[MS83] Jan Mandel and Bohuslav Sekerka. A local convergence proof for the iterative aggregation method. Linear Algebra and its Applications, 51(??):163–172, June 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958390157X.

Malek-Shahmirzadi:1984:MOM

[MS84] Massoud Malek-Shahmirzadi. On monotonic and orthant monotonic norms. Linear Algebra and its Applications, 56(??):169–175, January 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901228.

Marcus:1985:CGN

[MS85] Marvin Marcus and Markus Sandy. Conditions for the generalized numerical range to be real. Linear Algebra and its Applications, 71(??):219–239, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902496.

Mattheij:1986:EIT

[MS86]  R. M. M. Mattheij and M. D. Smooke.  Esti-

mates for the inverse of tridiagonal matrices arising in boundary value problems. Linear Algebra and its Applications, 73(??):33–57, January 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902326.

Mattheij:1987:IEP

[MS87] R. M. M. Mattheij and G. Söderlind. On inhomogeneous eigenvalue problems. I. Linear Algebra and its Applications, 88/89(??):507–531, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Marcus:1988:SPH

[MS88] Marvin Marcus and Markus Sandy. Symmetry properties of higher numerical ranges. Linear Algebra and its Applications, 104(??):141–164, June 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958890314X.

McAllister:1984:RRR

[MSS84] D. F. McAllister, G. W. Stewart, and W. J. Stewart. On a Rayleigh–Ritz refinement technique for nearly uncoupled stochastic matrices. Linear Algebra and its Applications, 60(??):1–25, August 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900673.

Marusic:1989:CPS

[MSS89] D. Marušič, R. Scapellato, and N. Zagaglia Salvi. A characterization of particular symmetric (0, 1) matrices. Linear Algebra and its Applications, 119(??):153–162, July 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958990075X.

Merikoski:1983:BRE

[MSW83] Jorma Kaarlo Merikoski, George P. H. Styan, and Henry Wolkowicz. Bounds for ratios of eigenvalues using traces. Linear Algebra and its Applications, 55(??):105– 124, December 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901702.

Marsaglia:1985:MSR

[MT85] George Marsaglia and Liang-Huei Tsay. Matrices and the structure of random number sequences. Linear Algebra and its Applications, 67(??):147–156, June 1985.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585901922.

Muller:1987:SIG

[Mül87] Vladimír Müller. On the spectrum of an infinite graph. Linear Algebra and its Applications, 93(??):187– 189, August 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0024379587903247.

Murphy:1982:NPC

[Mur82] Ian S. Murphy. A note on the product of complementary principal minors of a positive definite matrix. Linear Algebra and its Applications, 44(??):169–172, April 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958290012X.

Murota:1983:DME

[Mur83] Kazuo Murota. LU-decomposition of a matrix with entries of different kinds. Linear Algebra and its Applications, 49(??):275–283, February 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958390109X.

Murota:1989:SRR

[Mur89] Kazuo Murota. Some recent results in combinatorial approaches to dynamical systems. Linear Algebra and its Applications, 122/123/124(??):725–759, 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mertens:1988:SCS

[MV88a] L. Mertens and H. Van de Vel. A special class of structured matrices constructed with the Kronecker product and its use for difference equations. Linear Algebra and its Applications, 106(??):117–147, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588900250.

Mullen:1988:CLP

[MV88b] Gary L. Mullen and Theresa P. Vaughan. Cycles of linear permutations over a finite field. Linear Algebra and its Applications, 108(??):63–82, September 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901796.

Merikoski:1989:SNO

[MV89] Jorma Kaarlo Merikoski and Ari Virtanen. Some notes on de Oliveira’s determinantal conjecture. Linear Algebra and its Applications, 121(??):345–352, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958990709X. Linear

algebra and applications (Valencia, 1987).

Merris:1981:EDI

[MW81] Russell Merris and William Watkins. Elementary divisors of induced transformations on symmetry classes of tensors. Linear Algebra and its Applications, 38(??): 17–26, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581900045.

Merris:1984:ITA

[MW84] Russell Merris and William Watkins. An invariant theory approach to graph enumeration. Linear Algebra and its Applications, 61(??):277–285, September 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584900363.

Merikoski:1985:IEB

[MW85a] Jorma Kaarlo Merikoski and Henry Wolkowicz. Improving eigenvalue bounds using extra bounds. Linear Algebra and its Applications, 68(??):93–113, July 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902095.

Merris:1985:IIG

[MW85b] Russell Merris and William Watkins. Inequalities and identities for generalized matrix functions. Linear Algebra and its Applications, 64(??):223–242, January 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902794.

Ma:1987:CS

[MW87a] S. L. Ma and William C. Waterhouse. The gcirculant solutions of A^{m} = λJ. Linear Algebra and its Applications, 85(??):211–220, January 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587902187.

McKay:1987:TBN

[MW87b] James H. McKay and Stuart Sui Sheng Wang. On a theorem of Brualdi and Newman. Linear Algebra and its Applications, 92(??):39–43, July 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587902497.

Marek:1983:TSS

[MŽ83] Ivo Marek and Karel Žitný. Two-sided spectral bounds of two irreducible linear maps. Linear Algebra and its Applications, 52/53(??):509–523, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Makai:1989:PCB

[MZ89] Endre Makai, Jr. and Jaroslav Zemánek. On polynomial connections between projections. Linear Algebra and its Applications, 126(??):91–94, December 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589900074.

Nakamura:1984:HML

[Nak84] Yoshihiro Nakamura. Any Hermitian matrix is a linear combination of four projections. Linear Algebra and its Applications, 61(??):133–139, September 1984.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584900260.

Narayanan:1986:DVS

[Nar86] H. Narayanan. On the decomposition of vector spaces. Linear Algebra and its Applications, 79(??):61–98, July 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902934.

Naude:1987:RBS

[Nau87] Gert Naudé. Review of Linear Systems over Commutative Rings, by J. W. Brewer, J. W. Bunce, and F. S. Van Vleck. Linear Algebra and its Applications, 90(??): 241–244, May 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958790317X.

Navot:1989:OFH

[Nav89] I. Navot. Optimization of functionals of Hurwitz polynomials. Linear Algebra and its Applications, 120(??): 258–260, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589903868.

Nazareth:1984:NBA

[Naz84] Larry Nazareth. Numerical behavior of LP algorithms based upon the decomposition principle. Linear Algebra and its Applications, 57(??):181–189, February 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901861.

Neumann:1989:PVC

[NdOH89] Michael Neumann, Graciano N. de. Oliveira, and Vincente Hernandez, editors. Proceedings of the Valencia Conference on Linear Algebra and Applications: September 28–30, 1987, volume 121 of Linear Algebra and its Applications. North-Holland Publishing Co., Amsterdam, The Netherlands, 1989. ISSN 0024-3795 (print), 1873-1856 (electronic).

Niethammer:1984:CBI

[NdPV84] W. Niethammer, J. de Pillis, and R. S. Varga. Convergence of block iterative methods applied to sparse least-squares problems. Linear Algebra and its Applications, 58(??):327–341, April 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584902180.

Nel:1985:MDA

[Nel85] D. G. Nel. A matrix derivation of the asymptotic covariance matrix of sample correlation coefficients. Linear Algebra and its Applications, 67(??):137–145, June 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585901910.

Neumann:1980:CDI

[Neu80] M. Neumann. A combined direct-iterative approach for solving large scale singular and rectangular consistent systems of linear equations. Linear Algebra and its Applications, 34(??):85–101, December 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901603.

Neuberger:1981:FDP

[Neu81a] J. W. Neuberger. Finite-dimensional potential theory applied to numerical analysis of linear systems. Linear Algebra and its Applications, 35(??):193–202, February 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581902743.

Neumann:1981:KRC

[Neu81b] M. Neumann. The Kahan S.O.R. convergence bound for nonsingular and irreducible M-matrices. Linear Algebra and its Applications, 39(??):205–222, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581903049.

Neumaier:1982:SLE

[Neu82a] A. Neumaier. The second largest eigenvalue of a tree. Linear Algebra and its Applications, 46(??):9–25, August 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582900222.

Neumann:1982:USC

[Neu82b] Michael Neumann. The University of South Carolina mini-conference on matrix theory and linear algebra. Linear Algebra and its Applications, 43(??):273–284, March 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582902592.

Neumaier:1984:NTA

[Neu84] Arnold Neumaier. New techniques for the analysis of linear interval equations. Linear Algebra and its Applications, 58(??):273–325, April 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584902179.

Neudecker:1985:DMM

[Neu85] H. Neudecker. On the dispersion matrix of a matrix quadratic form connected with the noncentral Wishart distribution. Linear Algebra and its Applications, 70(??): 257–262, October 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585900576.

Neumaier:1986:CMM

[Neu86] A. Neumaier. On the comparison of H-matrices with M-matrices. Linear Algebra and its Applications, 83(??): 135–141, November 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586902703.

Neumaier:1987:FRL

[Neu87] A. Neumaier. Further results on linear interval equations. Linear Algebra and its Applications, 87(??):155– 179, March 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587901649.

Neumaier:1989:GRT

[Neu89] A. Neumaier. Graph representations, two-distance sets, and equiangular lines. Linear Algebra and its Applications, 114/115(??):141–156, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Newman:1980:GR

[New80] Morris Newman. Geršgorin revisited. Linear Algebra and its Applications, 30(??):247–249, April 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901986.

Newman:1983:LR

[New83] Morris Newman. Lyapunov revisited. Linear Algebra and its Applications, 52/53(??):525–528, 1983. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ng:1987:EHB

[Ng87] C. T. Ng. The equation F (x) + M(x)G(1/x) = 0 and homogeneous biadditive forms. Linear Algebra and its Applications, 93(??):255–279, August 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/S0024379587903296.

Nickel:1982:ALK

[Nic82] Karl Nickel. Die Auflösbarkeit linearer Kreisscheibenund Intervall-Gleichungssysteme. (German) [The solvability of linear systems of disc and interval equations]. Linear Algebra and its Applications, 44(??):19–40, April 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582900027.

Nielsen:1987:BTO

[Nie87] Klaus Nielsen. On bireflectionality and trireflectionality of orthogonal groups. Linear Algebra and its Applications, 94(??):197–208, September 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587900905.

Nieuwenhuis:1989:MLS

[Nie89] J. W. Nieuwenhuis. Modern linear systems theory. Linear Algebra and its Applications, 122/123/124(??):655–680, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nishio:1985:SRL

[Nis85] Katsuyoshi Nishio. The structure of a real linear combination of two projections. Linear Algebra and its Applications, 66(??):169–176, April 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585901302.

Nishi:1988:GBM

[Nis88] Akihiro Nishi. A generalization of the Binet–Minc formula for the evaluation of permanents. Linear Algebra and its Applications, 111(??):209–218, December 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588900602.

Nehorai:1985:UDF

[NM85] Arye Nehorai and Martin Morf. A unified derivation for fast estimation algorithms by the conjugate direction method. Linear Algebra and its Applications, 72(??):119– 143, December 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958590148X.

Muoneke:1987:SPN

[nM87]  n’Ekwunife Muoneke.  On the stochastic powers

of nonnegative reducible matrices. Linear Algebra and its Applications, 90(??):57–63, May 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587903041.

Nomakuchi:1980:CGI

[Nom80] Kentaro Nomakuchi. On the characterization of generalized inverses by bordered matrices. Linear Algebra and its Applications, 33(??):1–8, October 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580900932.

Norris:1984:RBT

[Nor84] L. K. Norris. Review of Statistical Theory and Random Matrices by Moshe Carmeli. Linear Algebra and its Applications, 61(??):291–292, September 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900399.

Norman:1989:RLM

[Nor89] C. W. Norman. Resultants and Lyapunov matrix equations. Linear Algebra and its Applications, 116(??): 109–120, April 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589904011.

Neumann:1987:CPM

[NP87] M. Neumann and R. J. Plemmons. Convergence of parallel multisplitting iterative methods for M-matrices. Linear Algebra and its Applications, 88/89(??):559–573, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Neumann:1982:MGM

[NPW82] Michael Neumann, George D. Poole, and Hans-Joachim Werner. More on generalizations of matrix monotonicity. Linear Algebra and its Applications, 48(??):413– 435, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582901240.

Nguyen:1985:GCF

[NS85] Truc T. Nguyen and Allan R. Sampson. The geometry of certain fixed marginal probability distributions. Linear Algebra and its Applications, 70(??):73–87, October 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585900448.

Neumann:1989:DAR

[NS89] Michael Neumann and Ronald J. Stern. Discrete approximations to reachability cones of linear differential systems. Linear Algebra and its Applications, 120(??): 65–79, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589903704.

Proceedings of the Fourth Haifa Matrix Theory Conference (Haifa, 1988).

Nowosad:1980:SIF

[NT80] Pedro Nowosad and Raul Tovar. Spectral inequalities and G-functions. Linear Algebra and its Applications, 31(??): 179–197, June 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580902189.

Nussbaum:1986:CLC

[Nus86] Roger D. Nussbaum. Convexity and log convexity for the spectral radius. Linear Algebra and its Applications, 73(??):59–122, January 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902338.

Neumaier:1984:ECD

[NV84] A. Neumaier and R. S. Varga. Exact convergence and divergence domains for the symmetric successive overrelaxation iterative (SSOR) method applied to Hmatrices. Linear Algebra and its Applications, 58(??): 261–272, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584902167.

Najfeld:1980:MIS

[NVD80] Igor Najfeld, Richard A. Vitale, and Philip J. Davis. Minkowski iteration of sets. Linear Algebra and its Applications, 29(??):259–291, February 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580902463.

Narcowich:1984:THS

[NW84] F. J. Narcowich and J. D. Ward. Toeplitz–Hausdorff systems. Linear Algebra and its Applications, 62(??):183– 193, November 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584900946.

Neudecker:1987:FOP

[NW87] Heinz Neudecker and Tom Wansbeek. Fourth-order properties of normally distributed random matrices. Linear Algebra and its Applications, 97(??):13–21, December 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587901352.

OConnor:1985:RDM

[O’C85] Derek O’Connor. Report on the Dublin matrix theory conference, March 1984: An introduction to sparse matrices. Linear Algebra and its Applications, 68(??): 271–272, July 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958590223X.

Ogryczak:1988:SMA

[Ogr88] Włodzimierz Ogryczak. The simplex method is not always well behaved. Linear Algebra and its Applications, 109(??):41–57, October 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379588901978.

Oxenrider:1985:MR

[OH85] Clinton J. Oxenrider and Richard D. Hill. On the matrix reorderings Γ and Ψ. Linear Algebra and its Applications, 69(??):205–212, August 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958590076X.

Okoh:1986:ALFa

[Oko86a] Frank Okoh. Applications of linear functionals to Kronecker modules. I. Linear Algebra and its Applications, 76 (??):165–188, April 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586902211.

Okoh:1986:ALFb

[Oko86b] Frank Okoh. Applications of linear functionals to Kronecker modules. II. Linear Algebra and its Applications, 76(??):189–204, April 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379586902223.

Okubo:1987:HTN

[Oku87] K. Okubo. Hölder-type norm inequalities for Schur products of matrices. Linear Algebra and its Applications, 91 (??):13–28, June 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587900577.

OLeary:1980:BCG

[O’L80a] Dianne P. O’Leary. The block conjugate gradient algorithm and related methods. Linear Algebra and its Applications, 29(??):293–322, February 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580902475.

OLeary:1980:GCG

[O’L80b] Dianne Prost O’Leary. A generalized conjugate gradient algorithm for solving a class of quadratic programming problems. Linear Algebra and its Applications, 34(??): 371–399, December 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580901731.

OLeary:1990:BSP

[O’L90] Dianne P. O’Leary. On bounds for scaled projections and pseudo-inverses. Linear Algebra and its Applications, 132:115–117, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See [Ste89].

Olkin:1983:ISF

[Olk83] Ingram Olkin. An inequality for a sum of forms. Linear Algebra and its Applications, 52/53(??):529–532, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Olkin:1985:ECD

[Olk85] Ingram Olkin. Estimating a Cholesky decomposition. Linear Algebra and its Applications, 67(??):201–205, June 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958590196X.

Ober:1989:BCF

[OM89] Raimund Ober and Duncan McFarlane. Balanced canonical forms for minimal systems: a normalized coprime factor approach. Linear Algebra and its Applications, 122/123/124(??):23–64, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ong:1982:RIS

[Ong82] Sing Cheong Ong. Remarks on invariant subspaces for finite-dimensional operators. Linear Algebra and its Applications, 42(??):99–101, February 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582901409.

Ong:1984:SMN

[Ong84] Sing-Cheong Ong. On the Schur multiplier norm of matrices. Linear Algebra and its Applications, 56(??): 45–55, January 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901125.

Ong:1987:WKO

[Ong87] Sing-Cheong Ong. What Kind of operators have few invariant subspaces? Linear Algebra and its Applications, 95(??):181–185, October 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587900346.

Olkin:1982:DBT

[OP82] I. Olkin and F. Pukelsheim. The distance between two random vectors with given dispersion matrices. Linear Algebra and its Applications, 48(??):257–263, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582901124.

Olsder:1988:CCH

[OR88] G. J. Olsder and C. Roos. Cramér and Cayley– Hamilton in the max algebra. Linear Algebra and its Applications, 101(??):87–108, April 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901450.

Olkin:1985:P

[ORS85a] Ingram Olkin, C. Radhakrishna Rao, and George P. H. Styan. Preface. Linear Algebra and its Applications, 67 (??):vii, June 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585901806.

Olkin:1985:SIS

[ORS85b] Ingram Olkin, C. Radhakrishna (Calyampudi Radhakrishna) Rao, and George P. H. Styan, editors. Special issue on statistics, volume 67 of Linear Algebra and its Applications. Elsevier, Amsterdam, The Netherlands, 1985. ISSN 0024-3795 (print), 1873-1856 (electronic).

Olivier:1988:NPS

[ORV88] P. Olivier, Q. I. Rahman, and R. S. Varga. On a new proof and sharpenings of a result of Fejér on bounded partial sums. Linear Algebra and its Applications, 107(??):237–251, August 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902480. Proceedings

of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Opfer:1984:RIN

[OS84] Gerhard Opfer and Glenn Schober. Richardson’s iteration for nonsymmetric matrices. Linear Algebra and its Applications, 58(??):343–361, April 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584902192.

OLeary:1986:ASP

[OS86] Dianne P. O’Leary and G. W. Stewart. Assignment and scheduling in parallel matrix factorization. Linear Algebra and its Applications, 77(??):275–299, May 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586901722. Special

volume on parallel computing.

Olesky:1989:ROA

[OS89] D. D. Olesky and T. A. Slater. Recognition and ordering algorithms for global inheritance in LU factorizations. Linear Algebra and its Applications, 120(??): 260–262, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958990387X.

Ouellette:1981:SCS

[Oue81] Diane Valérie Ouellette. Schur complements and statistics. Linear Algebra and its Applications, 36(??):187– 295, March 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902329.

Olesky:1982:MPS

[OvdD82] D. D. Olesky and P. van den Driessche. Monotone positive stable matrices. Linear Algebra and its Applications, 48(??):381–401, December 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582901227.

Olesky:1988:CMA

[OvdD88] D. D. Olesky and P. van den Driessche. Combinatorial Matrix Analysis Conference, 20–23 May 1987, University of Victoria, Victoria, BC, Canada. Linear Algebra and its Applications, 107(??):311–357, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588902534.

Ozguler:1986:PCI

[Özg86] A. Bülent Özgüler. Polynomial characterizations of (H, F)-invariant subspaces with applications. Linear Algebra and its Applications, 73(??):1–31, January 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902314.

Paige:1980:AEL

[Pai80] C. C. Paige. Accuracy and effectiveness of the Lanczos algorithm for the symmetric eigenproblem. Linear Algebra and its Applications, 34(??):235–258, December 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580901676.

Paige:1985:GLM

[Pai85] C. C. Paige. The general linear model and the generalized singular value decomposition. Linear Algebra and its Applications, 70(??):269–284, October 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958590059X.

Pan:1989:VMM

[Pan89a] C.-T. Pan. A vector majorization method for solving a nonlinear programming problem. Linear Algebra and its Applications, 119(??):129–139, July 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589900736.

Pandolfi:1989:CPP

[Pan89b] L. Pandolfi. Controllability properties of perturbed distributed parameter systems. Linear Algebra and its Applications, 122/123/124(??):525–538, 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pang:1989:CPN

[Pan89c] Ming Xian Pang. On constructions of partially nonpositive matrices. Linear Algebra and its Applications, 112 (??):49–55, January 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379589905880.

Parlett:1980:NLL

[Par80] B. N. Parlett. A new look at the Lanczos algorithm for solving symmetric systems of linear equations. Linear Algebra and its Applications, 29(??):323–346, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580902487.

Parlett:1982:WFT

[Par82] B. N. Parlett. Winograd’s Fourier transform via circulants. Linear Algebra and its Applications, 45(??): 137–155, June 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582902178.

Parter:1986:DSV

[Par86] Seymour V. Parter. On the distribution of the singular values of Toeplitz matrices. Linear Algebra and its Applications, 80(??):115–130, August 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902806.

Pate:1982:POI

[Pat82] T. H. Pate. Permutation-operator inequalities on certain symmetry classes of tensors. Linear Algebra and its Applications, 45(??):1–12, June 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582902051.

Pate:1984:IIP

[Pat84] Thomas H. Pate. An inequality involving permanents of certain direct products. Linear Algebra and its Applications, 57(??):147–155, February 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901836.

Pate:1985:DSM

[Pat85] Thomas H. Pate. Decomposability of symmetric multilinear functions. Linear Algebra and its Applications, 65(??): 71–81, February 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585900886.

Pate:1987:GFI

[Pat87] Thomas H. Pate. Generalizing the Fischer inequality. Linear Algebra and its Applications, 92(??):1–15, July 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587902473.

Pate:1988:GI

[Pat88a] Thomas H. Pate. Generalizing the inequality per(J_{2} ⊗ M) ≥ 2^{n}[per(M)]^{2}. Linear Algebra and its Applications, 102(??):81–93, April 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379588903217.

Pate:1988:LBP

[Pat88b] Thomas H. Pate. Lower bounds for permanents of Gram matrices having a rank one principal submatrix. Linear Algebra and its Applications, 111(??):249–259, December 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588900638.

Peresi:1986:BAP

[Per86] Luiz A. Peresi. On baric algebras with prescribed automorphisms. Linear Algebra and its Applications, 78(??): 163–185, June 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586900224.

Peresi:1988:NDA

[Per88a] Luiz A. Peresi. A note on duplication of algebras. Linear Algebra and its Applications, 104(??):65–69, June 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588903072.

Peresi:1988:DBA

[Per88b] Luiz A. Peresi. On derivations of baric algebras with prescribed automorphisms. Linear Algebra and its Applications, 104(??):71–74, June 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588903084.

Pettet:1982:NFL

[Pet82] Martin R. Pettet. Near-fields and linear transformations of finite fields. Linear Algebra and its Applications, 48(??):443–456, December 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582901264.

Poluikis:1981:CPH

[PH81] John A. Poluikis and Richard D. Hill. Completely positive and Hermitian-preserving linear transformations. Linear Algebra and its Applications, 35(??):1–10, February 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581902615.

Piff:1987:IBH

[Pif87] M. J. Piff. Inverses of banded and k-Hessenberg matrices. Linear Algebra and its Applications, 85(??):9–15, January 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587902047.

Pinkus:1985:SEP

[Pin85] A. Pinkus. Some extremal problems for strictly totally positive matrices. Linear Algebra and its Applications, 64(??):141–156, January 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902721.

Pizarro:1983:SCM

[Piz83] Antonio Pizarro. Similarity classes of 3 × 3 matrices over a discrete valuation ring. Linear Algebra and its Applications, 54(??):29–51, October 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583902045.

Pommer:1987:SNM

[PK87] Christian Pommer and Wolfhard Kliem. Simultaneously normalizable matrices. Linear Algebra and its Applications, 94(??):113–125, September 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587900838.

Parlett:1982:MSD

[PL82] B. N. Parlett and T. L. Landis. Methods for scaling to double stochastic form. Linear Algebra and its Applications, 48(??):53–79, December 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582900994.

Pavur:1986:RPM

[PL86] R. J. Pavur and T. O. Lewis. The role of powers of matrices in determining the distribution of quadratic forms involving the complex normal distribution. Linear Algebra and its Applications, 76(??):71–84, April 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902156.

Poole:1988:PEC

[PL88] George D. Poole and Michael Laidacker. Projectionally exposed cones in R^{3}. Linear Algebra and its Applications, 111(??):183–190, December 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588900584.

Plemmons:1981:RBM

[Ple81] R. J. Plemmons. Review of Sparse matrix proceedings 1978, edited by Iain S. Duff and G. W. Stewart. Linear Algebra and its Applications, 40(??):275–278, October 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581901580.

Plemmons:1983:RBM

[Ple83] Robert J. Plemmons. Review of Boolean Matrix Theory and Applications, by Ki Hang Kim. Linear Algebra and its Applications, 54(??):177–178, October 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583902136.

Porter:1980:ESR

[PM80] A. Duane Porter and Nick Mousouris. Exponential sums and rectangular partitions. Linear Algebra and its Applications, 29(??):347–355, February 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580902499.

Puschmann:1985:ZUV

[PN85] Heinrich Puschmann and Marcelo Nordio. Zwei unzulässige Verstärkungen der Vermutung von Wilkinson. (German) [Two inadmissible strengthenings of Wilkinson’s conjecture]. Linear Algebra and its Applications, 72 (??):167–176, December 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379585901521.

Parlett:1985:URE

[PNO85] B. N. Parlett and B. Nour-Omid. The use of a refined error bound when updating eigenvalues of tridiagonals. Linear Algebra and its Applications, 68(??):179–219, July 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902137.

Pohlmeyer:1989:SCT

[Poh89] K. Pohlmeyer. Solution of the constraints for tensors generated by iterated integrals. Linear Algebra and its Applications, 118(??):11–60, June 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589905703.

Pokrzywa:1985:CSN

[Pok85] Andrzej Pokrzywa. On continuity of spectra in norm ideals. Linear Algebra and its Applications, 69(??):121– 130, August 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585900710.

Pokrzywa:1986:PEO

[Pok86] Andrzej Pokrzywa. On perturbations and the equivalence orbit of a matrix pencil. Linear Algebra and its Applications, 82(??):99–121, October 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586901448.

Pollock:1985:TPM

[Pol85] D. S. G. Pollock. Tensor products and matrix differential calculus. Linear Algebra and its Applications, 67(??): 169–193, June 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585901946.

Polyrakis:1986:CLI

[Pol86] Ioannis A. Polyrakis. Cones locally isomorphic to the positive cone of l_{1}(Γ). Linear Algebra and its Applications, 84(??):323–334, December 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586903241. Proceedings

of the symposium on operator theory (Athens, 1985).

Polman:1987:IBF

[Pol87] Ben Polman. Incomplete blockwise factorizations of (block) H-matrices. Linear Algebra and its Applications, 90(??):119–132, May 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379587903107.

Polderman:1989:ALC

[Pol89] J. W. Polderman. Adaptive LQ control: conflict between identification and control. Linear Algebra and its Applications, 122/123/124(??):219–244, 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ponstein:1987:MMP

[Pon87] J. Ponstein. A matrix minimization problem involving ranks. Linear Algebra and its Applications, 92(??): 81–106, July 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587902527.

Poole:1981:RBI

[Poo81] George D. Poole. Review of Generalized Inverses of Linear Transformations, by Stephen L. Campbell and Carl D. Meyer, Jr. Linear Algebra and its Applications, 37(??):293–295, April 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581901889.

Power:1982:FRM

[Pow82] S. C. Power. Finite rank multivariable Hankel forms. Linear Algebra and its Applications, 48(??):237–244, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582901100.

Powell:1987:EGB

[Pow87] M. J. D. Powell. On error growth in the Bartels–Golub and Fletcher–Matthews algorithms for updating matrix factorizations. Linear Algebra and its Applications, 88/89 (??):597–621, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Powers:1988:GPE

[Pow88] David L. Powers. Graph partitioning by Eigenvectors. Linear Algebra and its Applications, 101(??):

121–133, April 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588901474.

Powers:1989:BGE

[Pow89] David L. Powers. Bounds on graph eigenvalues. Linear Algebra and its Applications, 117(??):1–6, May 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589905417.

Puystjens:1981:MPI

[PR81] Roland Puystjens and Donald W. Robinson. The Moore– Penrose inverse of a morphism with factorization. Linear Algebra and its Applications, 40(??):129–141, October 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581901452.

Pless:1988:TC

[PR88] Vera Pless and Joseph J. Rushanan. Triadic Codes. Linear Algebra and its Applications, 98(??):415–433, January 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901747.

Przyluski:1983:ASP

[Prz83] K. Maciej Przyłuski. About some problems of spectral synthesis. Linear Algebra and its Applications, 50(??): 475–485, April 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583900654.

Powers:1982:WPC

[PS82] David L. Powers and Mohammad M. Sulaiman. The walk partition and colorations of a graph. Linear Algebra and its Applications, 48(??):145–159, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582901045.

Parlett:1987:CSI

[PS87] Beresford N. Parlett and Youcef Saad. Complex shift and invert strategies for real matrices. Linear Algebra and its Applications, 88/89(??):575–595, 1987. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pullman:1988:SMC

[PS88] Norman J. Pullman and Miriam Stanford. Singular (0, 1) matrices with constant row and column sums. Linear Algebra and its Applications, 106(??):195–208, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588900286.

Peled:1989:PDS

[PS89] Uri N. Peled and Murali K. Srinivasan. The polytope of degree sequences. Linear Algebra and its Applications, 114/115(??):349–377, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ptak:1981:ELT

[Ptá81] Vlastimil Pták. An equation of Lyapunov type. Linear Algebra and its Applications, 39(??):73–82, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581902913.

Ptak:1983:BSI

[Ptá83a] Vlastimil Pták. Biorthogonal systems and the infinite companion matrix. Linear Algebra and its Applications, 49(??):57–78, February 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379583900940.

Ptak:1983:LEG

[Ptá83b] Vlastimil Pták. Lyapunov equations and Gram matrices. Linear Algebra and its Applications, 49(??):33– 55, February 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583900939.

Ptak:1984:BP

[Ptá84a] Vlastimil Pták. Bézoutians and projections. Linear Algebra and its Applications, 59(??):29–42, June 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901563.

Ptak:1984:EEB

[Ptá84b] Vlastimil Pták. Explicit expressions for Bézoutians. Linear Algebra and its Applications, 59(??):43–54, June 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901575.

Ptak:1984:LBH

[Ptá84c] Vlastimil Pták. Lyapunov, Bézout, and Hankel. Linear Algebra and its Applications, 58(??):363–390, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584902209.

Ptak:1986:EPO

[Ptá86a] Vlastimil Pták. An extremal problem for operators. Linear Algebra and its Applications, 84(??):213–226, December 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586903150.

Proceedings of the symposium on operator theory (Athens, 1985).

Ptak:1986:HMI

[Ptá86b] Vlastimil Pták. Hankel matrices and the infinite companion. Linear Algebra and its Applications, 81(??):199– 206, September 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586902569.

Pullman:1987:RBA

[Pul87] Norman J. Pullman. Review of Incline Algebra and Applications, by Z-Q Cao, K. H. Kim, and F. W. Roush. Linear Algebra and its Applications, 90(??):239–240, May 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587903168.

Paige:1981:SDH

[PV81] Chris C. Paige and Charles Van Loan. A Schur decomposition for Hamiltonian matrices. Linear Algebra and its Applications, 41(??):11–32, December 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900860.

Puystjens:1982:DMG

[PV82] R. Puystjens and J. Van Geel. Diagonalization of matrices over graded principal ideal domains. Linear Algebra and its Applications, 48(??):265–281, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582901136.

Paige:1986:QCK

[PV86] C. C. Paige and P. Van Dooren. On the quadratic convergence of Kogbetliantz’s algorithm for computing the singular value decomposition. Linear Algebra and its Applications, 77(??):301–313, May 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586901734. Special vol-

ume on parallel computing.

Pratzel-Wolters:1983:CFL

[PW83] D. Prätzel-Wolters. Canonical forms for linear systems. Linear Algebra and its Applications, 50(??):437– 473, April 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583900642.

Ptak:1985:BPM

[PW85] Vlastimil Pták and Harald K. Wimmer. On the Bézoutian for polynomial matrices. Linear Algebra and its Applications, 71(??):267–272, November 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902514.

Plemmons:1989:EPS

[PW89] Robert J. Plemmons and Stephen J. Wright. An efficient parallel scheme for minimizing a sum of Euclidean norms. Linear Algebra and its Applications, 121(??): 71–85, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589906927.

Linear algebra and applications (Valencia, 1987).

Ptak:1980:FOS

[PY80] Vlastimil Pták and N. J. Young. Functions of operators and the spectral radius. Linear Algebra and its Applications, 29(??):357–392, February 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580902505.

Ptak:1982:ZLH

[PY82] Vlastimil Pták and N. J. Young. Zero location by Hermitian forms: the singular case. Linear Algebra and its Applications, 43(??):181–196, March 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958290252X.

Pye:1980:NDI

[Pye80] Wallace C. Pye. Nonnegative Drazin inverses. Linear Algebra and its Applications, 30(??):149–153, April 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580901901.

Petz:1988:CT

[PZ88] Dénes Petz and Jaroslav Zemánek. Characterizations of the trace. Linear Algebra and its Applications, 111(??): 43–52, December 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958890050X.

Patera:1989:LG

[PZ89] J. Patera and H. Zassenhaus. On Lie gradings. I. Linear Algebra and its Applications, 112(??):87–159, January 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589905910.

Qi:1984:SSE

[Qi84] Li Qun Qi. Some simple estimates for singular values of a matrix. Linear Algebra and its Applications, 56(??): 105–119, January 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901174.

Queiro:1983:IFA

[Que83] João Filipe Queiró. Invariant factors as approximation numbers. Linear Algebra and its Applications, 49(??): 131–136, February 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583900988.

Quebbemann:1987:RTA

[Que87a] H.-G. Quebbemann. A representation theorem for algebras with involution. Linear Algebra and its Applications, 94(??):193–195, September 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587900899.

Queiro:1987:IPS

[Que87b] João Filipe Queiró. On the interlacing property for singular values and eigenvalues. Linear Algebra and its Applications, 97(??):23–28, December 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587901364.

Radjavi:1984:CPO

[Rad84]  Heydar Radjavi.  Commutativity-preserving opera-

tors on symmetric matrices. Linear Algebra and its Applications, 61(??):219–224, September 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900326.

Raghavan:1984:PMM

[Rag84] T. E. S. Raghavan. On pairs of multidimensional matrices. Linear Algebra and its Applications, 62(??):263– 268, November 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901010.

Rall:1986:RBI

[Ral86] L. B. Rall. Review of Introduction to Interval Computations, by Götz Alefeld and Jürgen Herzberger. Linear Algebra and its Applications, 78(??):233–234, June 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586900261.

Ramamurthy:1986:PAT

[Ram86] K. G. Ramamurthy. A polynomial algorithm for testing the nonnegativity of principal minors of Z-matrices. Linear Algebra and its Applications, 83(??):39–47, November 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902648.

Rantzer:1989:EBS

[Ran89] Anders Rantzer. Equivalence between stability of partial realizations and feedback stabilization—applications to reduced order stabilization. Linear Algebra and its Applications, 122/123/124(??):641–653, 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rao:1983:GIM

[Rao83] K. P. S. Bhaskara Rao. On generalized inverses of matrices over integral domains. Linear Algebra and its Applications, 49(??):179–189, February 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583901027.

Rao:1985:ILS

[Rao85] C. Radhakrishna Rao. The inefficiency of least squares: extensions of the Kantorovich inequality. Linear Algebra and its Applications, 70(??):249–255, October 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585900564.

Raykhshteyn:1983:RIS

[Ray83] Boris Raykhshteyn. On rational invariants of a set of symmetric matrices. Linear Algebra and its Applications, 54(??):17–28, October 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379583902033.

Robinson:1989:JSM

[RB89] Donald W. Robinson and Wayne W. Barrett. The Jordan 1-structure of a matrix of Redheffer. Linear Algebra and its Applications, 112(??):57–73, January 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589905892.

Rathore:1981:SAI

[RC81] R. K. S. Rathore and C. S. K. Chetty. Some angularity and inertia theorems related to normal matrices. Linear Algebra and its Applications, 40(??):69–77, October 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581901403.

Ruiz-Claeyssen:1988:DSD

[RCA88] Julio Cesar Ruiz-Claeyssen and Liara Aparecida dos Santos Leal. Diagonalization and spectral decomposition of factor block circulant matrices. Linear Algebra and its Applications, 99(??):41–61, February 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901243.

Ray-Chaudhuri:1989:ADT

[RCS89] D. K. Ray-Chaudhuri and N. M. Singhi. q-analogues of t-designs and their existence. Linear Algebra and its Applications, 114/115(??):57–68, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Reichstein:1986:SOH

[Rei86] Boris Reichstein. On symmetric operators of higher degree and their applications. Linear Algebra and its Applications, 75(??):155–172, March 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586901874.

Reichstein:1987:ECF

[Rei87] Boris Reichstein. On expressing a cubic form as a sum of cubes of linear forms. Linear Algebra and its Applications, 86(??):91–122, February 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587902898.

Rhodius:1989:ASS

[Rho89] Adolf Rhodius. On almost scrambling stochastic matrices. Linear Algebra and its Applications, 126(??):79– 86, December 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589900050.

Rice:1983:PAM

[Ric83] John R. Rice. Performance analysis of 13 methods to solve the Galerkin method equations. Linear Algebra and its Applications, 52/53(??):533–546, 1983. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rizvi:1984:IQM

[Riz84] S. A. H. Rizvi. Inverses of quasitridiagonal matrices. Linear Algebra and its Applications, 56(??):177–184, January 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958490123X.

Ramamurthy:1985:NPM

[RM85] K. G. Ramamurthy and S. R. Mohan. Nonnegativity of principal minors of generalized inverses of P_{0}- matrices. Linear Algebra and its Applications, 65(??):

125–131, February 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585900916.

Rogers:1989:APS

[RO89] E. Rogers and D. H. Owens. Axis positivity and the stability of linear multipass processes. Linear Algebra and its Applications, 122/123/124(??):779–796, 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Robert:1980:IEF

[Rob80a] F. Robert. Iterations sur des ensembles finis et automates cellulaires contractants. (French) [Iterations of finite sets and cellular automata contractors]. Linear Algebra and its Applications, 29(??):393–412, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580902517.

Robinson:1980:FRE

[Rob80b] Donald W. Robinson. The Frobenius rank equality for morphisms. Linear Algebra and its Applications, 29(??): 413–421, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580902529.

Robert:1982:RIF

[Rob82] Yves Robert. Regular incomplete factorizations of real positive definite matrices. Linear Algebra and its Applications, 48(??):105–117, December 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958290101X.

Robert:1983:DDC

[Rob83] F. Robert. Dérivée discrète et comportement local d’une itération discrète. (French) [discrete derivative and local behavior of a discrete iteration]. Linear Algebra and its Applications, 52/53(??):547–589, 1983. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Robinson:1984:CMP

[Rob84] Donald W. Robinson. On the covariance of the Moore– Penrose inverse. Linear Algebra and its Applications, 61 (??):91–99, September 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379584900247.

Robinson:1985:CMP

[Rob85] Donald W. Robinson. Covariance of Moore–Penrose inverses with respect to an invertible matrix. Linear Algebra and its Applications, 71(??):275–281, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902538.

Robinson:1986:GRS

[Rob86] Cecil E. Robinson, Jr. Green’s relations for substochastic matrices. Linear Algebra and its Applications, 80(??): 39–53, August 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586902764.

Robinson:1987:NSM

[Rob87a] Donald W. Robinson. Nullities of submatrices of the Moore–Penrose inverse. Linear Algebra and its Applications, 94(??):127–132, September 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958790084X.

Robinson:1987:GIS

[Rob87b] Cecil E. Robinson, Jr. Generalized inverses of substochastic matrices. Linear Algebra and its Applications, 91 (??):89–98, June 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587900632.

Rodman:1980:CCS

[Rod80] Leiba Rodman. On connected components in the set of divisors of a monic matrix polynomial. Linear Algebra and its Applications, 31(??):33–42, June 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580902037.

Roesler:1986:RHE

[Roe86] Friedrich Roesler. Riemann’s hypothesis as an eigenvalue problem. Linear Algebra and its Applications, 81(??):153– 198, September 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586902557.

Roesler:1987:RHE

[Roe87] Friedrich Roesler. Riemann’s hypothesis as an eigenvalue problem. II. Linear Algebra and its Applications, 92(??):45–73, July 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587902503.

Rohn:1981:ILS

[Roh81] Jiří Rohn. Interval linear systems with prescribed column sums. Linear Algebra and its Applications, 39(??): 143–148, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902986.

Rohn:1989:SLI

[Roh89] Jiří Rohn. Systems of linear interval equations. Linear Algebra and its Applications, 126(??):39–78, December 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589900049.

Romani:1986:ASI

[Rom86] Francesco Romani. On the additive structure of the inverses of banded matrices. Linear Algebra and its Applications, 80(??):131–140, August 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902818.

Rong:1982:EBP

[Ron82] Guang Hao Rong. The error bound of the perturbation of the Drazin inverse. Linear Algebra and its Applications, 47(??):159–168, October 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582902336.

Rose:1980:MIF

[Ros80] Donald J. Rose. Matrix identities of the Fast Fourier Transform. Linear Algebra and its Applications, 29(??): 423–443, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580902530.

Rose:1984:RBS

[Ros84a] N. J. Rose. Review of Singular Systems of Differential Equations. Linear Algebra and its Applications, 57(??): 283–286, February 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901939.

Rosenthal:1984:RBT

[Ros84b] Erik Rosenthal. A remark on Burnside’s theorem on matrix algebras. Linear Algebra and its Applications, 63(??): 175–177, December 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901411.

Rothblum:1980:BIS

[Rot80] Uriel G. Rothblum. Bounds on the indices of the spectral-circle eigenvalues of a nonnegative matrix. Linear Algebra and its Applications, 29(??):445–450, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580902542.

Rothblum:1981:REM

[Rot81] Uriel G. Rothblum. Resolvent expansions of matrices and applications. Linear Algebra and its Applications, 38 (??):33–49, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581900069.

Roth:1989:EBM

[Rot89a] Robert Roth. On the eigenvectors belonging to the minimum eigenvalue of an essentially nonnegative symmetric matrix with bipartite graph. Linear Algebra and its Applications, 118(??):1–10, June 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589905697.

Rothblum:1989:P

[Rot89b] Uriel Rothblum. Preface. Linear Algebra and its Applications, 114/115(??):1–16, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rothblum:1989:GSS

[Rot89c] Uriel G. Rothblum. Generalized scalings satisfying linear equations. Linear Algebra and its Applications, 114/115 (??):765–783, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rowlinson:1983:MIG

[Row83] Peter Rowlinson. On the maximal index of graphs with a prescribed number of edges. Linear Algebra and its Applications, 110(??):43–53, November 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583901313.

Rowlinson:1988:MIG

[Row88] Peter Rowlinson. On the maximal index of graphs with a prescribed number of edges. Linear Algebra and its Applications, 110(??):43–53, 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Runckel:1983:PCM

[RP83] Hans-J. Runckel and Uwe Pittelkow. Practical computation of matrix functions. Linear Algebra and its Applications, 49(??):161–178, February 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583901015.

Rao:1985:CHK

[RP85a] P. S. S. N. V. P. Rao and M. Precht. On a conjecture of Hoerl and Kennard on a property of least squares estimators of regression coefficients. Linear Algebra and its Applications, 67(??):99–101, June 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585901880.

Robinson:1985:EM

[RP85b] Donald W. Robinson and Roland Puystjens. EP morphisms. Linear Algebra and its Applications, 64(??): 157–174, January 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585902733.

Robinson:1987:GIM

[RP87] Donald W. Robinson and Roland Puystjens. Generalized inverses of morphisms with kernels. Linear Algebra and its Applications, 96(??):65–85, November 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587903363.

Robinson:1987:CMO

[RPV87] Donald W. Robinson, Roland Puystjens, and Jan Van Geel. Categories of matrices with only obvious Moore–Penrose inverses. Linear Algebra and its Applications, 97(??):93–102, December 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958790142X.

Rao:1982:GIB

[RR82] K. P. S. Bhaskara Rao and P. S. S. N. V. Prasada Rao. On generalized inverses of Boolean matrices — II. Linear Algebra and its Applications, 42(??):133–144, February 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582901446.

Ran:1984:SIMa

[RR84a] A. C. M. Ran and L. Rodman. Stability of invariant maximal semidefinite subspaces. I. Linear Algebra and its Applications, 62(??):51–86, November 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584900879.

Ran:1984:SIMb

[RR84b] A. C. M. Ran and L. Rodman. Stability of invariant maximal semidefinite subspaces. II. applications: selfadjoint rational matrix functions, algebraic Riccati equations. Linear Algebra and its Applications, 63(??):133–173, December 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958490140X.

Rothblum:1982:CEN

[RS82] Uriel G. Rothblum and Hans Schneider. Characterizations of extreme normalized circulations satisfying linear constraints. Linear Algebra and its Applications, 46(??): 61–72, August 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958290026X.

Rao:1986:SPO

[RS86a] M. L. Narayana Rao and K. Satyanarayana. On Sherk’s plane of order 27. Linear Algebra and its Applications, 74 (??):1–9, February 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901114.

Ruess:1986:AAP

[RS86b] W. M. Ruess and W. H. Summers. Asymptotic almost periodicity and motions of semigroups of operators. Linear Algebra and its Applications, 84(??):335–351, December 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586903253. Pro-

ceedings of the symposium on operator theory (Athens,

Rothblum:1989:SMW

[RS89] Uriel G. Rothblum and Hans Schneider. Scalings of matrices which have prespecified row sums and column sums via optimization. Linear Algebra and its Applications, 114/115(??):737–764, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rao:1989:CDT

[RSR89] M. L. Narayana Rao, K. Satyanarayana, and G. Vithal Rao. On a class of derived translation planes of square order. Linear Algebra and its Applications, 116(??): 81–99, April 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589903996.

Rothblum:1983:MPR

[RT83] Uriel G. Rothblum and Choon Peng Tan. Multiplicative processes reaching stationarity in finite time. Linear Algebra and its Applications, 55(??):155–167, December 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958390174X.

Rothblum:1985:UBM

[RT85] Uriel G. Rothblum and Choon Peng Tan. Upper bounds on the maximum modulus of subdominant eigenvalues of nonnegative matrices. Linear Algebra and its Applications, 66(??):45–86, April 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585901259.

Ries:1983:NMM

[RTW83] Manfred Ries, Ulrich Trottenberg, and Gerd Winter. A note on MGR methods. Linear Algebra and its Applications, 49(??):1–26, February 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583900915.

Ran:1985:NND

[RU85] André Ran and Frank Uhlig. A note on a new description of invariant maximal nonnegative subspaces in an indefinite inner product space. Linear Algebra and its Applications, 71(??):273–274, November 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902526.

Russell:1987:CPS

[RU87] A. M. Russell and C. J. F. Upton. A class of positive semidefinite matrices. Linear Algebra and its Applications, 93(??):121–126, August 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/S002437958790317X.

Ruhe:1983:NAG

[Ruh83] Axel Ruhe. Numerical aspects of Gram–Schmidt orthogonalization of vectors. Linear Algebra and its Applications, 52/53(??):591–601, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ruhe:1984:RKS

[Ruh84] Axel Ruhe. Rational Krylov sequence methods for eigenvalue computation. Linear Algebra and its Applications, 58(??):391–405, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379584902210.

Ran:1988:ECT

[RV88] A. C. M. Ran and R. Vreugdenhil. Existence and comparison theorems for algebraic Riccati equations for continuous- and discrete-time systems. Linear Algebra and its Applications, 99(??):63–83, February 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901255.

Redheffer:1989:PLF

[RV89] Ray Redheffer and Peter Volkmann. Positive linear functionals and the order cone. Linear Algebra and its Applications, 118(??):77–82, June 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589905739.

Rocha:1989:SDS

[RW89] P. Rocha and J. C. Willems. State for 2-D systems. Linear Algebra and its Applications, 122/123/124(??):1003–1038, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rand:1988:ILA

[RWZ88] D. Rand, P. Winternitz, and H. Zassenhaus. On the identification of a Lie algebra given by its structure constants. I. direct decompositions, Levi decompositions, and nilradicals. Linear Algebra and its Applications, 109(??):197–246, October 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902108.

Rao:1985:GIL

[RY85a] C. Radhakrishna Rao and Haruo Yanai. Generalized inverse of linear transformations: a geometric approach. Linear Algebra and its Applications, 66(??):87–98, April 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585901260.

Rao:1985:GIP

[RY85b] C. Radhakrishna Rao and Haruo Yanai. Generalized inverses of partitioned matrices useful in statistical applications. Linear Algebra and its Applications, 70(??): 105–113, October 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585900461.

Ryser:1980:TCM

[Rys80] H. J. Ryser. A theorem in combinatorial matrix theory. Linear Algebra and its Applications, 29(??):451– 458, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580902554.

Ryser:1981:MSI

[Rys81] H. J. Ryser. Matrices and set intersections. Linear Algebra and its Applications, 37(??):267–275, April 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581901865.

Ryser:1985:NLO

[Rys85] H. J. Ryser. A new look at the optimal assignment problem. Linear Algebra and its Applications, 66(??): 113–121, April 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585901284.

Sansigre:1989:BRV

[SA89] Gabriela Sansigre and Manuel Alvarez. On Bézoutian reduction with the Vandermonde matrix. Linear Algebra and its Applications, 121(??):401–408, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958990712X. Linear

algebra and applications (Valencia, 1987).

Saad:1980:VAM

[Saa80] Y. Saad. Variations on Arnoldi’s method for computing eigenelements of large unsymmetric matrices. Linear Algebra and its Applications, 34(??):269–295, December 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958090169X.

Saad:1986:CCG

[Saa86] Youcef Saad. Communication complexity of the Gaussian elimination algorithm on multiprocessors. Linear Algebra and its Applications, 77(??):315–340, May 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586901746. Special

volume on parallel computing.

Saitoh:1982:PDH

[Sai82] Saburou Saitoh. Positive definite Hermitian matrices and reproducing kernels. Linear Algebra and its Applications, 48(??):119–130, December 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582901021.

Saitoh:1987:QID

[Sai87] Saburou Saitoh. Quadratic inequalities deduced from the theory of reproducing kernels. Linear Algebra and its Applications, 93(??):171–178, August 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/S0024379587903223.

Saitoh:1988:QIA

[Sai88] Saburou Saitoh. Quadratic inequalities associated with integrals of reproducing kernels. Linear Algebra and its Applications, 101(??):269–280, April 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901541.

Sakai:1985:CVI

[Sak85] Yūji Sakai. Continuous versions of an inequality due to Hoffman and Wielandt. Linear Algebra and its Applications, 71(??):283–287, November 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958590254X.

Salzberg:1981:MGS

[Sal81] Pablo M. Salzberg. The minimal generating space of a polynomial. Linear Algebra and its Applications, 36(??): 89–96, March 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902214.

Santos:1987:CMC

[San87] Maria Célia Santos. On a class of methods for the computation of the spectral radius and the Perron vector of a nonnegative, irreducible matrix. Linear Algebra and its Applications, 92(??):209–223, July 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587902588.

Shasha:1987:ULS

[SB87] Dafna Shasha and Abraham Berman. On the uniqueness of the Lyapunov scaling factors. Linear Algebra and its Applications, 91(??):53–63, June 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587900590.

Shasha:1988:MUL

[SB88] Dafna Shasha and Abraham Berman. More on the uniqueness of the Lyapunov scaling factors. Linear Algebra and its Applications, 107(??):253–273, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902492. Proceedings

of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Schneider:1981:GCE

[Sch81a] Hans Schneider. Geometric conditions for the existence of positive eigenvalues of matrices. Linear Algebra and its Applications, 38(??):253–271, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900252.

Schrijver:1981:TDI

[Sch81b] A. Schrijver. On total dual integrality. Linear Algebra and its Applications, 38(??):27–32, June 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581900057.

Schaefer:1982:MVF

[Sch82] H. H. Schaefer. A metric variant of Frobenius’s theorem and some other remarks on positive matrices. Linear Algebra and its Applications, 42(??):175–182, February 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582901471.

Schoenberg:1983:ISL

[Sch83a] I. J. Schoenberg. Interpolating splines as limits of polynomials. Linear Algebra and its Applications, 52/53(??): 617–628, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Schumacher:1983:ARP

[Sch83b] J. M. Schumacher. The algebraic regulator problem from the state-space point of view. Linear Algebra and its Applications, 50(??):487–520, April 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583900666.

Schwerdtfeger:1983:CMP

[Sch83c] Hans Schwerdtfeger. On the covariance of the Moore– Penrose inverse. Linear Algebra and its Applications, 52/53(??):629–643, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Schneider:1984:TSS

[Sch84a] Hans Schneider. Theorems on M-splittings of a singular M-matrix which depend on graph structure. Linear Algebra and its Applications, 58(??):407–424, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584902222.

Schreiber:1984:GIM

[Sch84b] Robert Schreiber. Generalized iterative methods for semidefinite linear systems. Linear Algebra and its Applications, 62(??):219–230, November 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900971.

Schrijver:1984:RBA

[Sch84c]  Alexander Schrijver.  Review of Modern Applied

Mathematics: Optimization and Operations Research, edited by Bernhard Korte. Linear Algebra and its Applications, 62(??):289–292, November 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901058.

Schmid:1985:RMG

[Sch85] Peter Schmid. Rational matrix groups of a special type. Linear Algebra and its Applications, 71(??):289– 293, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585902551.

Schneider:1986:IMR

[Sch86a] Hans Schneider. The influence of the marked reduced graph of a nonnegative matrix on the Jordan form and on related properties: a survey. Linear Algebra and its Applications, 84(??):161–189, December 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586903137. Proceedings

of the symposium on operator theory (Athens, 1985).

Schock:1986:WPC

[Sch86b] Eberhard Schock. What are the proper condition numbers of discretized ill-posed problems. Linear Algebra and its Applications, 81(??):129–136, September 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902533.

Schwenk:1986:TBS

[Sch86c] Allen J. Schwenk. Tight bounds on the spectral radius of asymmetric nonnegative matrices. Linear Algebra and its Applications, 75(??):257–265, March 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958690193X.

Schumacher:1988:TLS

[Sch88] J. M. Schumacher. Transformations of linear systems under external equivalence. Linear Algebra and its Applications, 102(??):1–33, April 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588903187.

Schneider:1989:MSE

[Sch89a] Michael H. Schneider. Matrix scaling, entropy minimization, and conjugate duality. I. existence conditions. Linear Algebra and its Applications, 114/115(??):785–813, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Schrijver:1989:HCS

[Sch89b] A. Schrijver. Homotopy and crossings of systems of curves on a surface. Linear Algebra and its Applications, 114/115 (??):157–167, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Scott:1985:AGC

[Sco85] David S. Scott. On the accuracy of the Gershgorin circle theorem for bounding the spread of a real symmetric matrix. Linear Algebra and its Applications, 65(??): 147–155, February 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958590093X.

Scott:1986:ASR

[Sco86] David S. Scott. Avoiding the square-root bottleneck in the Choleski factorization of a matrix on a parallel computer. Linear Algebra and its Applications, 77(??):341–344, May 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586901758.

S-Dehesa:1980:EDR

[SD80] Jesús S.-Dehesa. The eigenvalue density of rational Jacobi matrices. II. Linear Algebra and its Applications, 33(??): 41–55, October 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580900968.

Schneider:1981:HAM

[SDGT81] Hans Schneider, Chandler Davis, Walter Gautschi, and Olga Taussky Todd. Honoring Alexander M. Ostrowski. Linear Algebra and its Applications, 38(??):3, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900021.

Seneta:1980:CSD

[Sen80] E. Seneta. Computing the stationary distribution for infinite Markov chains. Linear Algebra and its Applications, 34(??):259–267, December 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580901688.

Seneta:1984:EFE

[Sen84] E. Seneta. Explicit forms for ergodicity coefficients and spectrum localization. Linear Algebra and its Applications, 60(??):187–197, August 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958490079X.

Seneta:1988:SPS

[Sen88] E. Seneta. Sensitivity to perturbation of the stationary distribution: some refinements. Linear Algebra and its Applications, 108(??):121–126, September 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901814.

Shamir:1985:MFR

[SF85] T. Shamir and P. A. Fuhrmann. Minimal factorizations of rational matrix functions in terms of polynomial models. Linear Algebra and its Applications, 68(??):67–91, July 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902083.

Shasha:1988:MLS

[SH88] Dafna Shasha and Daniel Hershkowitz. Maximal Lyapunov scaling factors and their applications in the study of Lyapunov diagonal semistability of block triangular matrices. Linear Algebra and its Applications, 103(??): 21–39, May 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588902194.

Shapiro:1982:CKAb

[Sha82a] Helene Shapiro. A conjecture of Kippenhahn about the characteristic polynomial of a pencil generated by two Hermitian matrices. II. Linear Algebra and its Applications, 45(??):97–108, June 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582902129.

Shapiro:1982:HPC

[Sha82b] Helene Shapiro. Hermitian pencils with a cubic minimal polynomial. Linear Algebra and its Applications, 48(??): 81–103, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582901008.

Shapiro:1982:CKAa

[Sha82c] Helene Shapiro. On a conjecture of Kippenhahn about the characteristic polynomial of a pencil generated by two Hermitian matrices. I. Linear Algebra and its Applications, 43(??):201–221, March 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582902543.

Shapiro:1983:UIE

[Sha83] Alexander Shapiro. On the unsolvability of inverse eigenvalues problems almost everywhere. Linear Algebra and its Applications, 49(??):27–31, February 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583900927.

Shao:1985:MPE

[Sha85a] Jia Yu Shao. Matrices permutation equivalent to primitive matrices. Linear Algebra and its Applications, 65(??): 225–247, February 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585900990.

Shao:1985:CAE

[Sha85b] Jia Yu Shao. On a conjecture about the exponent set of primitive matrices. Linear Algebra and its Applications, 65(??):91–123, February 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900904.

Shao:1985:EPD

[Sha85c] Jia Yu Shao. On the exponent of a primitive digraph. Linear Algebra and its Applications, 64(??):21–31, January 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902642.

Shao:1985:PIM

[Sha85d] Jia Yu Shao. Products of irreducible matrices. Linear Algebra and its Applications, 68(??):131–143, July 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902113.

Shapiro:1985:EPS

[Sha85e] Alexander Shapiro. Extremal problems on the set of nonnegative definite matrices. Linear Algebra and its Applications, 67(??):7–18, June 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958590182X.

Shapiro:1985:IFA

[Sha85f] Alexander Shapiro. Identifiability of factor analysis: some results and open problems. Linear Algebra and its Applications, 70(??):1–7, October 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900382. See erratum

[Sha89].

Sha:1986:EE

[Sha86a] Hu Yun Sha. Estimation of the eigenvalues of AB for A > 0, B > 0. Linear Algebra and its Applications, 73(??):147–150, January 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902363.

Shamir:1986:AAP

[Sha86b]  T. Shamir.  An algebraic approach to the par-

tial realization problem. Linear Algebra and its Applications, 73(??):171–196, January 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902399.

Shalom:1987:ATM

[Sha87] Tamir Shalom. On algebras of Toeplitz matrices. Linear Algebra and its Applications, 96(??):211–226, November 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587903454.

Shamir:1988:PRS

[Sha88] T. Shamir. Partial realization of symmetric matrix rational functions. Linear Algebra and its Applications, 105 (??):139–161, July 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588900092.

Shapiro:1989:EIF

[Sha89] Alexander Shapiro. Erratum: “Identifiability of factor analysis: some results and open problems” [Linear Algebra Appl. 70 (1985), 1–7, MR 86m:62114]. Linear Algebra and its Applications, 125(??):149, December 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589900384. See [?].

Shemesh:1984:CET

[She84] Dan Shemesh. Common eigenvectors of two matrices. Linear Algebra and its Applications, 62(??):11– 18, November 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584900855.

Shekhtman:1986:SPM

[She86] Boris Shekhtman. On some problems of M. Z. Nashed on outer inverses. Linear Algebra and its Applications, 76 (??):149–152, April 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586902193.

Shemesh:1987:USC

[She87a] Dan Shemesh. Unobservable subspaces and common roots of several polynomials. Linear Algebra and its Applications, 95(??):187–210, October 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587900358.

Sheng:1987:PGW

[She87b] Rong Qin Sheng. Properties of the Golub–Wilkinson algorithm. Linear Algebra and its Applications, 92(??): 75–80, July 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587902515.

Shearer:1989:DME

[She89] James B. Shearer. On the distribution of the maximum eigenvalue of graphs. Linear Algebra and its Applications, 114/115(??):17–20, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shrikhande:1986:SSP

[Shr86] Mohan Shrikhande. A survey of some problems in combinatorial designs—a matrix approach. Linear Algebra and its Applications, 79(??):215–247, July 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586903010.

Scott:1986:PBJ

[SHW86] David S. Scott, Michael T. Heath, and Robert C. Ward. Parallel block Jacobi eigenvalue algorithms using systolic arrays. Linear Algebra and its Applications, 77(??): 345–355, May 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958690176X.

Sidi:1989:EPM

[Sid89] Avram Sidi. On extensions of the power method for normal operators. Linear Algebra and its Applications, 120 (??):207–224, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379589903790.

Proceedings of the Fourth Haifa Matrix Theory Conference (Haifa, 1988).

Silva:1986:NIP

[Sil86] F. Conceição Silva. On the number of invariant polynomials of the matrix XAX^{−1} + B. Linear Algebra and its Applications, 79(??):1–21, July 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902880.

Silva:1987:MPE

[Sil87a] F. Conceição Silva. Matrices with prescribed eigenvalues and principal submatrices. Linear Algebra and its Applications, 92(??):241–250, July 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587902618.

Silva:1987:IFM

[Sil87b] F. Conceição Silva. On the invariant factors of the matrix XAY + B. Linear Algebra and its Applications, 96(??): 1–16, November 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587903326.

Silva:1988:SCP

[Sil88] F. Conceição Silva. Spectrally complete pairs of matrices. Linear Algebra and its Applications, 108(??):239– 262, September 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588901917.

Simon:1984:ASL

[Sim84] Horst D. Simon. Analysis of the symmetric Lanczos algorithm with reorthogonalization methods. Linear Algebra and its Applications, 61(??):101–131, September 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584900259.

Sinkhorn:1981:PSS

[Sin81] Richard Sinkhorn. Power symmetric stochastic matrices. Linear Algebra and its Applications, 40(??):225– 228, October 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958190152X.

Singh:1988:NLR

[Sin88] Surjeet Singh. A note on linear recurring sequences. Linear Algebra and its Applications, 104(??):97–101, June 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588903102.

Siskakis:1986:WCS

[Sis86] Aristomenis G. Siskakis. Weighted composition semigroups on Hardy spaces. Linear Algebra and its Applications, 84(??):359–371, December 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586903277. Proceedings

of the symposium on operator theory (Athens, 1985).

Sterna-Karwat:1982:WMC

[SK82] Alicja Sterna-Karwat. On weak monomorphisms in the category of convex processes. Linear Algebra and its Applications, 47(??):173–179, October 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958290235X.

Sterna-Karwat:1984:NAC

[SK84] Alicja Sterna-Karwat. A note on additive convex processes. Linear Algebra and its Applications, 57(??):1– 12, February 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958490171X.

Seberry:1982:OMC

[SL82] Jennifer Seberry and Clement W. H. Lam. On orthogonal matrices with constant diagonal. Linear Algebra and its Applications, 46(??):117–129, August 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582900313.

Shao:1987:ICI

[SL87] Jia Yu Shao and Qiao Li. On the index of convergence of an irreducible Boolean matrix. Linear Algebra and its Applications, 97(??):185–210, December 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587901492.

Sweet:1988:AOT

[SM88] L. G. Sweet and J. A. MacDougall. On automorphisms of order three of division algebras. Linear Algebra and its Applications, 108(??):231–237, September 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901905.

Smeds:1981:LDM

[Sme81] Per A. Smeds. Line digraphs and the Moore–Penrose inverse. Linear Algebra and its Applications, 36(??): 165–172, March 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902299.

Smith:1986:SM

[Smi86] Ronald L. Smith. On the spectrum of N_{0}-matrices. Linear Algebra and its Applications, 83(??):129–134, November 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902697.

Smith:1987:MMP

[Smi87] Ronald L. Smith. On Markham’s M-matrix properties. Linear Algebra and its Applications, 87(??):189– 195, March 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587901662.

Smith:1988:SNM

[Smi88] Ronald L. Smith. Some notes on Z-matrices. Linear Algebra and its Applications, 106(??):219–231, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588900304.

Sugie:1989:DCF

[SO89] Toshiharu Sugie and Toshiro Ono. On doubly coprime factorizations. Linear Algebra and its Applications, 122/123/124(??):681–696, 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sodupe:1987:IIG

[Sod87] María-José Sodupe. Interlacing inequalities for generalized singular values. Linear Algebra and its Applications, 87(??):85–92, March 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379587901601.

Sontag:1987:RBS

[Son87] Eduardo D. Sontag. Review of Multidimensional Systems Theory, edited by N. K. Bose. Linear Algebra and its Applications, 87(??):273–277, March 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587901741.

Song:1988:MPC

[Son88] Seok-Zun Song. Minimum permanents on certain faces of matrices containing an identity submatrix. Linear Algebra and its Applications, 108(??):263–280, September 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901929.

Sourour:1986:AIS

[Sou86a] A. R. Sourour. Anti-invariant subspaces of maximum dimension. Linear Algebra and its Applications, 74(??):39– 45, 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sourour:1986:ASM

[Sou86b] A. R. Sourour. Antiinvariant subspaces of maximum dimension. Linear Algebra and its Applications, 74(??): 38–45, February 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958690114X.

Spigler:1982:AGT

[Spi82] Renato Spigler. An application of group theory to matrices and to ordinary differential equations. Linear Algebra and its Applications, 44(??):143–151, April 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958290009X.

Spingarn:1985:PDP

[Spi85] Jonathan E. Spingarn. A primal-dual projection method for solving systems of linear inequalities. Linear Algebra and its Applications, 65(??):45–62, February 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585900862.

Spingarn:1987:PML

[Spi87] Jonathan E. Spingarn. A projection method for leastsquares solutions to overdetermined systems of linear inequalities. Linear Algebra and its Applications, 86(??): 211–236, February 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587902965.

Spreen:1981:SPC

[Spr81] Dieter Spreen. On some properties of the Cesàro limit of a stochastic matrix. Linear Algebra and its Applications, 41(??):81–91, December 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581900896.

Simoes-Pereira:1982:SNT

[SPZ82] J. M. S. Simões-Pereira and Christina M. Zamfirescu. Submatrices of non-tree-realizable distance matrices. Linear Algebra and its Applications, 44(??):1–17, April 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582900015.

Stadelmaier:1982:NMP

[SRPM82] Michael W. Stadelmaier, Nicholas J. Rose, George D. Poole, and Carl D. Meyer, Jr. Nonnegative matrices with power invariant zero patterns. Linear Algebra and its Applications, 42(??):23–29, February 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582901355.

Seneta:1981:SEN

[SS81] E. Seneta and Suzanne Sheridan. Strong ergodicity of nonnegative matrix products. Linear Algebra and its Applications, 37(??):277–292, April 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581901877.

Scott:1985:STG

[SS85] Alastair J. Scott and George P. H. Styan. On a separation theorem for generalized eigenvalues and a problem in the analysis of sample surveys. Linear Algebra and its Applications, 70(??):209–224, October 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585900539.

Sierksma:1986:SMM

[SS86] Gerard Sierksma and Elmer Sterken. The structure matrix of (0, 1)-matrices: its rank, trace, and eigenvalues. an application to econometric models. Linear Algebra and its Applications, 83(??):151–166, November 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902727.

Saad:1987:PDM

[SS87a] Y. Saad and M. Schultz. Parallel direct methods for solving banded linear systems. Linear Algebra and its Applications, 88/89(??):623–650, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sierksma:1987:TMS

[SS87b] Gerard Sierksma and Albert E. Steenge. Two-minor stable matrices and the open Leontief model. Linear Algebra and its Applications, 95(??):1–10, October 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587900231.

Schneider:1988:ASP

[SS88] Hans Schneider and Jeffrey Stuart. Allowable spectral perturbations of ZME-matrices. Linear Algebra and its Applications, 111(??):63–118, December 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588900523.

Sexton:1982:RDR

[SSS82] H. Sexton, M. Shensa, and J. Speiser. Remarks on a displacement-rank inversion method for Toeplitz systems. Linear Algebra and its Applications, 45(??):127–130, June 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582902154.

Saxena:1983:HBI

[SSZ83] R. B. Saxena, A. Sharma, and Z. Ziegler. Hermite– Birkhoff interpolation on roots of unity and Walsh equiconvergence. Linear Algebra and its Applications, 52/53(??):603–615, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stembridge:1981:TS

[ST81] John R. Stembridge and John Todd. On a trigonometrical sum. Linear Algebra and its Applications, 35(??): 287–291, February 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902809.

Stern:1987:EMS

[ST87a] Ronald J. Stern and Michael Tsatsomeros. Extended M-matrices and subtangentiality. Linear Algebra and its Applications, 97(??):1–11, December 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587901340.

Sung:1987:CFD

[ST87b] C. H. Sung and B. S. Tam. On the cone of a finitedimensional compact convex set at a point. Linear Algebra and its Applications, 90(??):47–55, May 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958790303X.

Schreck:1988:NCS

[ST88a] H. Schreck and G. Tinhofer. A note on certain subpolytopes of the assignment polytope associated with circulant graphs. Linear Algebra and its Applications, 111(??):125–134, December 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588900547.

Stahlecker:1988:FPM

[ST88b] P. Stahlecker and G. Trenkler. Full and partial minimax estimation in regression analysis with additional linear constraints. Linear Algebra and its Applications, 111(??):279–292, December 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588900651.

Stafney:1981:CFM

[Sta81] James D. Stafney. Correction to: “Functions of a matrix and their norms” [Linear Algebra Appl. 20 (1978), no. 1, 87–94, MR 57 #3162]. Linear Algebra and its Applications, 39(??):259–260, August 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581903086.

Stanley:1987:BSR

[Sta87] Richard P. Stanley. A bound on the spectral radius of graphs with e edges. Linear Algebra and its Applications, 87(??):267–269, March 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587901728.

Steeb:1981:RBK

[Ste81a] W.-H. Steeb. The relation between the Kronecker product, the trace calculation, and the one-dimensional Ising model. Linear Algebra and its Applications, 37(??): 261–265, April 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581901853.

Stern:1981:GM

[Ste81b] Ronald J. Stern. Generalized M-matrices. Linear Algebra and its Applications, 41(??):201–208, December 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900999.

Stern:1982:SPI

[Ste82] Ronald J. Stern. On strictly positively invariant cones. Linear Algebra and its Applications, 48(??):13–24, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582900957.

Stewart:1984:SOP

[Ste84] G. W. Stewart. A second order perturbation expansion for small singular values. Linear Algebra and its Applications, 56(??):231–235, January 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901289.

Stepniak:1985:OND

[Ste85] Czesław Stepniak. Ordering of nonnegative definite matrices with application to comparison of linear models. Linear Algebra and its Applications, 70(??):67–71, October 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585900436.

Stegall:1986:ODB

[Ste86] Charles Stegall. Optimization and differentiation in Banach spaces. Linear Algebra and its Applications, 84(??): 191–211, December 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586903149.

Proceedings of the symposium on operator theory (Athens, 1985).

Stepniak:1987:TOC

[Ste87a] Czesław Stepniak. Two orderings on a convex cone of nonnegative definite matrices. Linear Algebra and its Applications, 94(??):263–272, September 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587900954.

Stewart:1987:RBC

[Ste87b] G. W. Stewart. Review of Matrix Computations, by Gene H. Golub and Charles F. Van Loan. Linear Algebra and its Applications, 95(??):211–215, October 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958790036X.

Stewart:1989:SPP

[Ste89] G. W. Stewart. On scaled projections and pseudoinverses. Linear Algebra and its Applications, 112(??): 189–193, January 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589905946.

See answer [O’L90] to a question raised in this paper.

Stickel:1987:FDM

[Sti87] Eberhard Stickel. On the Fréchet derivative of matrix functions. Linear Algebra and its Applications, 91(??): 83–88, June 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587900620.

Stoyan:1986:MPM

[Sto86] Gisbert Stoyan. On maximum principles for monotone matrices. Linear Algebra and its Applications, 78(??): 147–161, June 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586900212.

Strassen:1983:ROC

[Str83] V. Strassen. Rank and optimal computation of generic tensors. Linear Algebra and its Applications, 52/53(??): 645–685, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Streit:1985:SPO

[Str85] Roy L. Streit. Saddle points and overdetermined complex equations. Linear Algebra and its Applications, 64(??): 57–76, January 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585902666.

Strauss:1987:JFB

[Str87] Nicholas Strauss. Jordan form of a binomial coefficient matrix over Z_{p}. Linear Algebra and its Applications, 90 (??):65–72, May 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587903053.

Stuart:1987:DIA

[Stu87] Jeffrey L. Stuart. The decomposition of idempotents associated with inflators. Linear Algebra and its Applications, 97(??):171–184, December 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587901480.

Stuart:1988:SIF

[Stu88a] Jeffrey L. Stuart. The solutions to an infinite family of matrix inequalities involving ZME-matrices. Linear Algebra and its Applications, 108(??):141–156, September 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958890184X.

Sturmfels:1988:TPM

[Stu88b] Bernd Sturmfels. Totally positive matrices and cyclic polytopes. Linear Algebra and its Applications, 107(??): 275–281, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588902509.

Proceedings of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Sturmfels:1988:TCM

[Stu88c] Bernd Sturmfels. Tridiagonalization of complex matrices and a problem of Longstaff. Linear Algebra and its Applications, 109(??):165–166, October 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588902078.

Styan:1981:LDI

[Sty81] George P. H. Styan. On Lavoie’s determinantal inequality. Linear Algebra and its Applications, 37(??):77–80, April 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581901683.

Sun:1982:NST

[Sun82a] Ji Guang Sun. A note on Stewart’s theorem for definite matrix pairs. Linear Algebra and its Applications, 48(??): 331–339, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582901197.

Sunder:1982:NA

[Sun82b] V. S. Sunder. A noncommutative analogue of |D(X^{k})| = |kX^{k−1}|. Linear Algebra and its Applications, 44(??): 87–95, April 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582900064.

Sunder:1982:PCH

[Sun82c] V. S. Sunder. On permutations, convex hulls, and normal operators. Linear Algebra and its Applications, 48(??):

403–411, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582901239.

Smith:1984:DBS

[SW84] P. W. Smith and Henry Wolkowicz. Dimensionality of biinfinite systems. Linear Algebra and its Applications, 57(??):115–130, February 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901812.

Smyth:1985:RDM

[SW85] M. R. F. Smyth and T. T. West. Report on the Dublin matrix theory conference, March 1984: Barnes and support idempotents. Linear Algebra and its Applications, 68(??):247–249, July 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902162.

Shivakumar:1987:EIM

[SWR87] P. N. Shivakumar, J. J. Williams, and N. Rudraiah. Eigenvalues for infinite matrices. Linear Algebra and its Applications, 96(??):35–63, November 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587903351.

Sugimoto:1989:SPA

[SY89] Kenji Sugimoto and Yutaka Yamamoto. On successive pole assignment by linear-quadratic optimal feedbacks. Linear Algebra and its Applications, 122/123/124(??): 697–723, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sylvester:1986:DSL

[Syl86] John Sylvester. On the dimension of spaces of linear transformations satisfying rank conditions. Linear Algebra and its Applications, 78(??):1–10, June 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586900133.

Schwarz:1986:CMU

[SZ86] Binyamin Schwarz and Abraham Zaks. Contraction of the matrix unit disk. Linear Algebra and its Applications, 84(??):353–357, December 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586903265. Proceedings

of the symposium on operator theory (Athens, 1985).

Schwarz:1989:GCM

[SZ89] Binyamin Schwarz and Abraham Zaks. Generalized contractions of the matrix unit disk. Linear Algebra and its Applications, 120(??):177–180, August 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589903765. Proceedings

of the Fourth Haifa Matrix Theory Conference (Haifa, 1988).

Szatrowski:1985:ADT

[Sza85] Ted H. Szatrowski. Asymptotic distributions in the testing and estimation of the missing-data multivariate normal linear patterned mean and correlation matrix. Linear Algebra and its Applications, 67(??):215–231, June 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585901983.

Szulc:1988:LBP

[Szu88] Tomasz Szulc. A lower bound for the Perron root of a nonnegative matrix. Linear Algebra and its Applications, 101 (??):181–186, April 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588901504.

Szulc:1989:CTM

[Szu89a] Tomasz Szulc. A contribution to the theory of Mmatrices. Linear Algebra and its Applications, 126(??): 87–89, December 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589900062.

Szulc:1989:LBP

[Szu89b] Tomasz Szulc. A lower bound for the Perron root of a nonnegative matrix. II. Linear Algebra and its Applications, 112(??):19–27, January 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589905855.

Tadmor:1981:ESR

[Tad81] Eitan Tadmor. The equivalence of L_{2}-stability, the resolvent condition, and strict H-stability. Linear Algebra and its Applications, 41(??):151–159, December 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900951.

Tadmor:1986:CSM

[Tad86] Eitan Tadmor. Complex symmetric matrices with strongly stable iterates. Linear Algebra and its Applications, 78(??):65–77, June 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379586900169.

Takacs:1982:RWG

[Tak82] Lajos Takács. Random walks on groups. Linear Algebra and its Applications, 43(??):49–67, March 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582902439.

Takahashi:1983:RPO

[Tak83] Katsutoshi Takahashi. On relative primeness of operator polynomials. Linear Algebra and its Applications, 50(??): 521–526, April 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583900678.

Takemura:1985:EPC

[Tak85] Akimichi Takemura. On the equivalence of proportional cell frequencies and orthogonality of interaction spaces in n-way ANOVA. Linear Algebra and its Applications, 67(??):35–49, June 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585901843.

Tam:1981:GIC

[Tam81a] Bit Shun Tam. Generalized inverses of cone preserving maps. Linear Algebra and its Applications, 40(??): 189–202, October 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958190149X.

Tam:1981:GTG

[Tam81b] Bit Shun Tam. A geometric treatment of generalized inverses and semigroups of nonnegative matrices. Linear Algebra and its Applications, 41(??):225–272, December 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581901014.

Tam:1981:SCP

[Tam81c] Bit Shun Tam. On the semiring of cone preserving maps. Linear Algebra and its Applications, 35(??):79– 108, February 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902676.

Tam:1985:DOC

[Tam85] Bit Shun Tam. On the duality operator of a convex cone. Linear Algebra and its Applications, 64(??):33– 56, January 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585902654.

Tam:1987:GSM

[Tam87a] Tin Yau Tam. The generalized spectral matrices and radial matrices on symmetry classes of tensors. Linear Algebra and its Applications, 85(??):29–48, January 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587902060.

Tam:1987:LOMb

[Tam87b] Tin Yau Tam. Linear operators on matrices: the invariance of the decomposable numerical radius. Linear Algebra and its Applications, 87(??):147–153, March 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587901637.

Tam:1987:LOMa

[Tam87c] Tin Yau Tam. Linear operators on matrices: the invariance of the decomposable numerical range. Linear Algebra and its Applications, 85(??):1–7, January 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587902035.

Tam:1987:LOMc

[Tam87d] Tin Yau Tam. Linear operators on matrices: the invariance of the decomposable numerical range. II. Linear Algebra and its Applications, 92(??):197–207, July 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587902576.

Tannenbaum:1983:SSP

[Tan83] Allen Tannenbaum. On the stabilizer subgroup of a pair of matrices. Linear Algebra and its Applications, 50(??): 527–544, April 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958390068X.

Tang:1986:PLM

[Tan86] Wai-Shing Tang. On positive linear maps between matrix algebras. Linear Algebra and its Applications, 79 (??):33–44, July 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586902909.

Tao:1984:CSM

[Tao84] Pham Dinh Tao. Convergence of a subgradient method for computing the bound norm of matrices. Linear Algebra and its Applications, 62(??):163–182, November 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584900934.

Taussky:1980:MNA

[Tau80a] Olga Taussky. More on norms from algebraic number fields, commutators and matrices which transform a rational matrix into its transpose. Linear Algebra and its Applications, 29(??):459–464, February 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580902566.

Taussky:1980:SFC

[Tau80b] Olga Taussky. Some facts concerning integral representations of the ideals in an algebraic number field. Linear Algebra and its Applications, 31(??):245–248, June 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580902232.

Taussky:1982:MAP

[Tau82] Olga Taussky. The many aspects of the Pythagorean triangles. Linear Algebra and its Applications, 43(??): 285–295, March 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582902609.

Taussky:1983:CTP

[Tau83] Olga Taussky. On the congruence transformation of a pencil of real symmetric matrices to a pencil with identical characteristic polynomial. Linear Algebra and its Applications, 52/53(??):687–691, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Taussky:1985:RCP

[Tau85] Olga Taussky. Remarks concerning possible connections between Fermat’s last theorem and integral p × p circulants. Linear Algebra and its Applications, 71(??):295–297, November 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902563.

Troskie:1986:DRC

[TC86] C. G. Troskie and W. J. Conradie. The distribution of the ratios of characteristic roots (condition numbers) and their applications in principal component or ridge regression. Linear Algebra and its Applications, 82(??):255–279, October 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586901564.

Tchon:1989:SIN

[Tch89] Krzysztof Tchoń. On structural instability of normal forms of affine control systems subject to static state feedback. Linear Algebra and its Applications, 121(??): 95–104, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589906940.

Linear algebra and applications (Valencia, 1987).

Terwilliger:1987:RSG

[Ter87] Paul Terwilliger. Root system graphs. Linear Algebra and its Applications, 94(??):157–163, September 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587900863.

Taub:1987:RCP

[TG87] Hedi Taub and Shaul Gutman. Roots of composite polynomials—an application to root clustering. Linear Algebra and its Applications, 87(??):181–188, March 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587901650.

Thas:1989:IPG

[Tha89] J. A. Thas. Interesting pointsets in generalized quadrangles and partial geometries. Linear Algebra and its Applications, 114/115(??):103–131, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Thompson:1982:JGF

[Tho82a]  Robert C. Thompson.  The Jacobi–Gundelfinger–

Frobenius–Iohvidov rule and the Hasse symbol. Linear Algebra and its Applications, 44(??):189–196, April 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582900167.

Thompson:1982:SFI

[Tho82b] Robert C. Thompson. The Smith form, the inversion rule for 2 × 2 matrices, and the uniqueness of the invariant factors for finitely generated modules. Linear Algebra and its Applications, 44(??):197–201, April 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582900179.

Thompson:1983:RBT

[Tho83a] R. C. Thompson. Review of Inequalities: Theory of Majorization and Its Applications, by Albert W. Marshall and Ingram Olkin. Linear Algebra and its Applications, 54(??):173–175, October 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583902124.

Thompson:1983:COT

[Tho83b] Robert C. Thompson. A converse to the Ostrowski– Taussky determinantal inequality. Linear Algebra and its Applications, 52/53(??):693–699, 1983. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Thompson:1983:CCD

[Tho83c] Robert C. Thompson. Counterexamples concerning the diagonal elements of normal matrices. Linear Algebra and its Applications, 49(??):285–291, February 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583901106.

Thompson:1985:AMV

[Tho85] R. C. Thompson. p-adic matrix valued inequalities. Linear Algebra and its Applications, 71(??):299–308, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585902575.

Thomassen:1986:SNM

[Tho86] Carsten Thomassen. Sign-nonsingular matrices and even cycles in directed graphs. Linear Algebra and its Applications, 75(??):27–41, March 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586901795.

Thompson:1988:SCM

[Tho88] Robert C. Thompson. Special cases of a matrix exponential formula. Linear Algebra and its Applications, 107(??): 283–292, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588902510.

Proceedings of the Victoria Conference on Combinatorial Matrix Analysis (Victoria, BC, 1987).

Thomassen:1989:WSP

[Tho89a] Carsten Thomassen. When the sign pattern of a square matrix determines uniquely the sign pattern of its inverse. Linear Algebra and its Applications, 119(??):27–34, July 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589900669.

Thompson:1989:CPG

[Tho89b] Robert C. Thompson. Convergence proof for Goldberg’s exponential series. Linear Algebra and its Applications, 121(??):3–7, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379589906885.

Linear algebra and applications (Valencia, 1987).

Thompson:1989:CSE

[Tho89c] Robert C. Thompson. Cyclic shifts and an exponential formula. Linear Algebra and its Applications, 120(??): 9–21, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589903662.

Proceedings of the Fourth Haifa Matrix Theory Conference (Haifa, 1988).

Timoney:1985:RDM

[Tim85] Richard M. Timoney. Report on the Dublin matrix theory conference, March 1984: Reinhardt decompositions of operator matrix spaces. Linear Algebra and its Applications, 68(??):255–258, July 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902198.

Tismenetsky:1987:DBT

[Tis87] M. Tismenetsky. Determinant of block-Toeplitz band matrices. Linear Algebra and its Applications, 85(??): 165–184, January 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958790214X.

Togawa:1980:GSS

[Tog80] Yoshio Togawa. A geometric study of the D-stability problem. Linear Algebra and its Applications, 33(??): 133–151, October 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580901020.

Tomiyama:1985:GPM

[Tom85] Jun Tomiyama. On the geometry of positive maps in matrix algebras. II. Linear Algebra and its Applications, 69 (??):169–177, August 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379585900746.

Torgasev:1986:IGL

[Tor86a] Aleksandar Torgašev. Infinite graphs with the least limiting eigenvalue greater than −2. Linear Algebra and its Applications, 82(??):133–141, October 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586901461.

Torsney:1986:MIO

[Tor86b] B. Torsney. Moment inequalities via optimal design theory. Linear Algebra and its Applications, 82(??):237– 253, October 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901552.

Towers:1980:LAA

[Tow80] David Towers. Lie algebras all of whose proper subalgebras are nilpotent. Linear Algebra and its Applications, 32 (??):61–73, August 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580900075.

Tremon:1985:PDM

[Tre85a] Michel Tremon. Polynômes de degré minimum connectant deux projections dans une algèbre de Banach. (French) [Minimum degree polynomials connecting two projections in a Banach algebra]. Linear Algebra and its Applications, 64(??):115–132, January 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902708.

Trench:1985:EPT

[Tre85b]  William F. Trench.  On the eigenvalue problem

for Toeplitz band matrices. Linear Algebra and its Applications, 64(??):199–214, January 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902770.

Trench:1986:SST

[Tre86] William F. Trench. Solution of systems with Toeplitz matrices generated by rational functions. Linear Algebra and its Applications, 74(??):191–211, February 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586901229.

Trench:1987:NSN

[Tre87] William F. Trench. A note on solving nearly triangular Toeplitz systems. Linear Algebra and its Applications, 93 (??):57–65, August 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0024379587903119.

Truemper:1980:CTU

[Tru80] K. Truemper. Complement total unimodularity. Linear Algebra and its Applications, 30(??):77–92, April 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580901834.

Tsing:1984:CBF

[Tsi84]  Nam Kiu Tsing.  The constrained bilinear form

and the C-numerical range. Linear Algebra and its Applications, 56(??):195–206, January 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901253.

Tsay:1986:FMC

[TST86] Yih T. Tsay, Leang S. Shieh, and Jason S. H. Tsai. A fast method for computing the principal nth roots of complex matrices. Linear Algebra and its Applications, 76(??):205–221, April 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902235.

Turisco:1985:FQF

[Tur85] JoAnn S. Turisco. A family of quadratic forms associated to quadratic mappings of spheres. Linear Algebra and its Applications, 65(??):249–260, February 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585901004.

Trentelman:1988:AIN

[TvdW88] H. L. Trentelman and J. W. van der Woude. Almost invariance and noninteracting control: a frequency-domain analysis. Linear Algebra and its Applications, 101(??): 221–254, April 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588901528.

Trottenberg:1980:FIO

[TW80] Ulrich Trottenberg and Gerd Winter. On faces and W - irreducible operators in finite-dimensional linear spaces.

Linear Algebra and its Applications, 33(??):87–109, October 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580900993.

Tam:1989:CWS

[TW89] Bit-Shun Tam and Shiow Fang Wu. On the Collatz– Wielandt sets associated with a cone-preserving map. Linear Algebra and its Applications, 125(??):77–95, December 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589900335.

Tanny:1987:SEU

[TZ87] Stephen M. Tanny and Michael Zuker. The sensitivity of eigenvalues under elementary matrix perturbations. Linear Algebra and its Applications, 86(??):123– 143, February 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587902904.

Uchimura:1986:PFT

[Uch86] Keisuke Uchimura. Perron–Frobenius theory of some algebraic equations associated with context free grammars. Linear Algebra and its Applications, 78(??):11–22, June 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586900145.

Uhlig:1981:EPD

[Uhl81a] Frank Uhlig. Explicit polar decomposition and a nearcharacteristic polynomial: the 2 × 2 case. Linear Algebra and its Applications, 38(??):239–249, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900239.

Uhlig:1981:IER

[Uhl81b] Frank Uhlig. Inertia and eigenvalue relations between symmetrized and symmetrizing matrices for the real and the general field case. Linear Algebra and its Applications, 35(??):203–226, February 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581902755.

Uhlig:1983:RBF

[Uhl83] Frank Uhlig. Relations between the fields of values of a matrix and of its polar factors: the 2×2 real and complex case. Linear Algebra and its Applications, 52/53(??):701– 715, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Uhlig:1987:MEA

[Uhl87] Frank Uhlig. On the matrix equation AX = B with applications to the generators of a controllability matrix. Linear Algebra and its Applications, 85(??):203–209, January 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587902175.

Ursic:1982:IMR

[Urs82] Silvio Ursic. Inverse matrix representation with one triangular array. Linear Algebra and its Applications, 47(??): 151–157, October 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582902324.

Usmani:1987:EIB

[Usm87] Riaz A. Usmani. Explicit inverse of a band matrix with application to computing eigenvalues of a Sturm–Liouville system. Linear Algebra and its Applications, 86(??):189–198, February 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587902941.

Valiaho:1986:CCM

[Väl86] Hannu Väliaho. Criteria for copositive matrices. Linear Algebra and its Applications, 81(??):19–34, September 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902466.

Valenzuela:1988:MCL

[Val88a] Oscar Adolfo Sánchez Valenzuela. Matrix computations in linear superalgebra. Linear Algebra and its Applications, 111(??):151–181, December 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588900572.

Valiaho:1988:DIM

[Väl88b] H. Väliaho. Determining the inertia of a matrix pencil as a function of the parameter. Linear Algebra and its Applications, 106(??):245–258, August 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588900328.

Valiaho:1988:TDM

[Väl88c] H. Väliaho. Testing the definiteness of matrices on polyhedral cones. Linear Algebra and its Applications, 101 (??):135–165, April 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588901486.

Valiaho:1989:ACM

[Väl89a] H. Väliaho. Almost copositive matrices. Linear Algebra and its Applications, 116(??):121–134, April 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589904023.

Valiaho:1989:QPC

[Väl89b] H. Väliaho. Quadratic-programming criteria for copositive matrices. Linear Algebra and its Applications, 119 (??):163–182, July 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589900761.

VanLoan:1984:SMA

[Van84] C. F. Van Loan. A symplectic method for approximating all the eigenvalues of a Hamiltonian matrix. Linear Algebra and its Applications, 61(??):233–251, September 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958490034X.

VanLoan:1987:ECE

[Van87] Charles Van Loan. On estimating the condition of eigenvalues and eigenvectors. Linear Algebra and its Applications, 88/89(??):715–732, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vaserstein:1985:ASS

[Vas85] L. N. Vaserstein. On arithmetic subgroups of simple algebraic groups. Linear Algebra and its Applications, 72(??): 93–96, December 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585901454.

Vavrin:1984:ILM

[Vav84] Zdeněk Vavřín. Inverses of Löwner matrices. Linear Algebra and its Applications, 63(??):227–236, December 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901459.

Vavrin:1988:MDT

[Vav88] Zdeněk Vavřín. Multiplication of diagonal transforms of Loewner matrices. Linear Algebra and its Applications, 99(??):1–40, February 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901231.

Vavrin:1989:RCP

[Vav89] Zdeněk Vavřín. Remarks on complexity of polynomial and special matrix computations. Linear Algebra and its Applications, 122/123/124(??):539–564, 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vonBelow:1985:CEA

[vB85] Joachim von Below. A characteristic equation associated to an eigenvalue problem on c^{2}-networks. Linear Algebra and its Applications, 71(??):309–325, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902587.

VanBarel:1988:AMS

[VB88] M. Van Barel and A. Bultheel. An algebraic method to solve the minimal partial realization problem for scalar sequences. Linear Algebra and its Applications, 104(??): 117–129, June 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588903126.

VanBarel:1989:CMC

[VB89] Marc Van Barel and Adhemar Bultheel. A canonical matrix continued fraction solution of the minimal (partial) realization problem. Linear Algebra and its Applications, 122/123/124(??):973–1002, 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

VanDooren:1983:EAP

[VD83] Paul Van Dooren and Patrick Dewilde. The eigenstructure of an arbitrary polynomial matrix: computational aspects. Linear Algebra and its Applications, 50(??): 545–579, April 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583900691.

Vein:1982:SSS

[Vei82] P. R. Vein. A short survey of some recent applications of determinants. Linear Algebra and its Applications, 42(??):287–297, February 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582901574.

Vein:1986:IAC

[Vei86] P. R. Vein. Identities among certain triangular matrices. Linear Algebra and its Applications, 82(??):27– 79, October 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901424.

Veinott:1989:CDC

[Vei89a] Arthur F. Veinott, Jr. Conjugate duality for convex programs: a geometric development. Linear Algebra and its Applications, 114/115(??):663–667, 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Veinott:1989:RGP

[Vei89b] Arthur F. Veinott, Jr. Representation of general and polyhedral subsemilattices and sublattices of product spaces. Linear Algebra and its Applications, 114/115(??):681– 704, 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Verghese:1982:CRL

[Ver82] George C. Verghese. A “Cramer rule” for the least-norm, least-squared-error solution of inconsistent linear equations. Linear Algebra and its Applications, 48(??):315– 316, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582901173.

Verfurth:1984:CNM

[Ver84] R. Verfürth. The contraction number of a multigrid method with mesh ratio 2 for solving poisson’s equation. Linear Algebra and its Applications, 60(??):113– 128, August 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584900740.

Vinnikov:1989:CDD

[Vin89] Victor Vinnikov. Complete description of determinantal representations of smooth irreducible curves. Linear Algebra and its Applications, 125(??):103–140, December 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589900359.

Vitoria:1982:BEB

[Vit82] José Vitória. Block Eigenvalues of block compound matrices. Linear Algebra and its Applications, 47(??):23– 34, October 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582902233.

Vere-Jones:1984:IIP

[VJ84] D. Vere-Jones. An identity involving permanents. Linear Algebra and its Applications, 63(??):267–270, December 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584901484.

Vere-Jones:1988:GPD

[VJ88] D. Vere-Jones. A generalization of permanents and determinants. Linear Algebra and its Applications, 111(??): 119–124, December 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588900535.

vanLint:1981:NEP

[vL81] J. H. van Lint. Notes on Egoritsjev’s proof of the van der Waerden conjecture. Linear Algebra and its Applications, 39(??):1–8, August 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379581902858.

Vaserstein:1987:PBA

[VM87] L. N. Vaserstein and B. A. Magurnt. Prestabilization for K_{1} of Banach algebras. Linear Algebra and its Applications, 95(??):69–96, October 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587900279.

Varga:1984:CMS

[VNC84] R. S. Varga, W. Niethammer, and D.-Y. Cai. p-cyclic matrices and the symmetric successive overrelaxation method. Linear Algebra and its Applications, 58(??): 425–439, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584902234.

Vidyasagar:1986:CIP

[VV86] M. Vidyasagar and N. Viswanadham. Construction of inverses with prescribed zero minors and applications to decentralized stabilization. Linear Algebra and its Applications, 83(??):103–115, November 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379586902673.

vanderSluis:1987:CBR

[vV87a] A. van der Sluis and H. A. Van der Vorst. The convergence behavior of Ritz values in the presence of close eigenvalues. Linear Algebra and its Applications, 88/89 (??):651–694, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

VanHuffel:1987:ARB

[VV87b] Sabine Van Huffel and Joos Vandewalle. Algebraic relationships between classical regression and total leastsquares estimation. Linear Algebra and its Applications, 93(??):149–160, August 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S002437958790320X.

VanHuffel:1987:SSU

[VV87c] Sabine Van Huffel and Joos Vandewalle. Subset selection using the total least squares approach in collinearity problems with errors in the variables. Linear Algebra and its Applications, 88/89(??):695–714, 1987. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vonzurGathen:1987:PD

[vzG87] Joachim von zur Gathen. Permanent and determinant. Linear Algebra and its Applications, 96(??):87– 100, November 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587903375.

Wagowski:1989:TGW

[Wag89] Marc Wagowski. The Tutte group of a weakly orientable matroid. Linear Algebra and its Applications, 117(??): 21–24, May 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589905442.

Walter:1986:NSP

[Wal86] Martin E. Walter. On the norm of a Schur product. Linear Algebra and its Applications, 79(??):209–213, July 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586903009.

Wang:1980:RGM

[Wan80] Kai Wang. Resultants and group matrices. Linear Algebra and its Applications, 33(??):111–122, October 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580901007.

Wang:1981:GR

[Wan81a] Kai Wang. On the generalization of a retrocirculant. Linear Algebra and its Applications, 37(??):35–43, April 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581901658.

Wang:1981:GMG

[Wan81b] Kai Wang. On the group matrices for a generalized dihedral group. Linear Algebra and its Applications, 39(??): 83–89, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902925.

Wang:1983:CPS

[Wan83] Kai Wang. Characteristic polynomials of symmetric graphs. Linear Algebra and its Applications, 51(??): 121–125, June 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901520.

Wang:1986:CRM

[Wan86] Guo Rong Wang. A Cramer rule for minimum-norm (T) least-squares (S) solution of inconsistent linear equations. Linear Algebra and its Applications, 74(??):213– 218, February 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901230.

Wang:1989:CRF

[Wan89] Guo Rong Wang. A Cramer rule for finding the solution of a class of singular equations. Linear Algebra and its Applications, 116(??):27–34, April 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589903959.

Wasan:1984:NPK

[Was84] Siri Krishan Wasan. A new proof of Korzhik’s result on burst distributions. Linear Algebra and its Applications, 59(??):115–119, June 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584901629.

Watters:1980:BTA

[Wat80] J. F. Watters. Block Triangularization of Algebras of Matrices. Linear Algebra and its Applications, 32(??): 3–7, August 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580900038.

Waterhouse:1982:LMP

[Wat82a] William C. Waterhouse. Linear maps preserving reduced norms. Linear Algebra and its Applications, 43(??): 197–200, March 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582902531.

Waterhouse:1982:RCC

[Wat82b] William C. Waterhouse. Real classification of complex quadrics. Linear Algebra and its Applications, 48(??): 45–52, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379582900982.

Waterhouse:1983:CPH

[Wat83] William C. Waterhouse. A conjectured property of Hermitian pencils. Linear Algebra and its Applications, 51 (??):173–177, June 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901581.

Waterhouse:1984:CSM

[Wat84] William C. Waterhouse. The codimension of singular matrix pairs. Linear Algebra and its Applications, 57(??): 227–245, February 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901903.

Waterhouse:1986:LTS

[Wat86] William C. Waterhouse. Linear transformations on selfadjoint matrices: the preservation of rank-one-plus-scalar. Linear Algebra and its Applications, 74(??):73– 85, February 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586901163.

Waterhouse:1987:EDS

[Wat87a] William C. Waterhouse. Elementary divisors of ssymmetric matrices. Linear Algebra and its Applications, 93(??):241–245, August 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0024379587903272.

Waterhouse:1987:NFI

[Wat87b] William C. Waterhouse. A note on Fan’s identity for symmetric bilinear forms. Linear Algebra and its Applications, 90(??):235–237, May 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587903156.

Watson:1987:MDK

[Wat87c] Geoffrey S. Watson. A method for discovering Kantorovichtype inequalities and a probabilistic interpretation. Linear Algebra and its Applications, 97(??):211–217, December 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587901509.

Waterhouse:1988:IPI

[Wat88a] William C. Waterhouse. Inner product identities on the unit circle. Linear Algebra and its Applications, 108(??):127–134, September 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588901826.

Waterhouse:1988:MBB

[Wat88b] William C. Waterhouse. The map behind a binomial coefficient matrix over z/pz. Linear Algebra and its Applications, 105(??):195–198, July 1988. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379588900122.

Watkins:1988:DIC

[Wat88c] William Watkins. A determinantal inequality for correlation matrices. Linear Algebra and its Applications, 104 (??):59–63, June 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588903060.

Worz-Busekros:1981:RBG

[WB81] Angelika Wörz-Busekros. Relationship between genetic algebras and semicommutative matrices. Linear Algebra and its Applications, 39(??):111–123, August 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581902950.

Watkins:1983:SSF

[WE83] D. S. Watkins and L. Elsner. Self-similar flows. Linear Algebra and its Applications, 110(??):213–242, November 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379583901374.

Watkins:1988:SSF

[WE88] D. S. Watkins and L. Elsner. Self-similar flows. Linear Algebra and its Applications, 110(??):213–242, 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Watkins:1989:SEF

[WE89] D. S. Watkins and L. Elsner. Self-equivalent flows associated with the generalized eigenvalue problem. Linear Algebra and its Applications, 118(??):107–127, June 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589905764.

Weaver:1983:REN

[Wea83] James R. Weaver. Real eigenvalues of nonnegative matrices which commute with a symmetric matrix involution. Linear Algebra and its Applications, 110(??):243– 253, November 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901386.

Weaver:1988:REN

[Wea88] James R. Weaver. Real eigenvalues of nonnegative matrices which commute with a symmetric matrix involution. Linear Algebra and its Applications, 110(??):243– 253, 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Webb:1986:TDP

[Web86] J. R. L. Webb. Topological degree and A-proper operators. Linear Algebra and its Applications, 84(??):227– 242, December 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586903162.

Proceedings of the symposium on operator theory (Athens, 1985).

Chu:1987:SVG

[wEC87a] King wah Eric Chu. Singular value and generalized singular value decompositions and the solution of linear matrix equations. Linear Algebra and its Applications, 88/89(??): 83–98, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chu:1987:SME

[wEC87b]  King wah Eric Chu.  The solution of the ma-

trix equations AXB − CXD = E and (Y A − DZ, Y C − BZ) = (E, F). Linear Algebra and its Applications, 93(??):93–105, August 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/S0024379587903144.

Chu:1989:SSL

[wEC89] King wah Eric Chu. Symmetric solutions of linear matrix equations by matrix decompositions. Linear Algebra and its Applications, 119(??):35–50, July 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589900670.

Weinberger:1983:ONA

[Wei83] H. F. Weinberger. Optimal numerical approximation of a linear operator. Linear Algebra and its Applications, 52/53(??):717–737, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Weinberg:1984:CFS

[Wei84] David A. Weinberg. Canonical forms for symmetric tensors. Linear Algebra and its Applications, 57(??):271– 282, February 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901927.

Weichsel:1987:PG

[Wei87] Paul M. Weichsel. Polynomials on graphs. Linear Algebra and its Applications, 93(??):179–186, August 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0024379587903235.

Wei:1989:PCL

[Wei89a] Mu Sheng Wei. The perturbation of consistent least squares problems. Linear Algebra and its Applications, 112(??):231–245, January 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0024379589905983.

Weichsel:1989:DRG

[Wei89b] Paul M. Weichsel. Distance-regular graphs in block form. Linear Algebra and its Applications, 126(??):135– 148, December 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589900116.

Werner:1983:GCM

[Wer83] Wilhelm Werner. A generalized companion matrix of a polynomial and some applications. Linear Algebra and its Applications, 55(??):19–36, December 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583901647.

Werner:1984:CMM

[Wer84] Hans Joachim Werner. Charakterisierungen von monotonen matrizen. (German) [Characterizations of monotone matrices]. Linear Algebra and its Applications, 60(??): 79–90, August 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584900727.

Werner:1985:DMC

[Wer85a] Hans Joachim Werner. Drazin-monotonicity characterizations for property-n matrices. Linear Algebra and its Applications, 71(??):327–337, November 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902599.

Werner:1985:MBE

[Wer85b] Hans Joachim Werner. More on BLU estimation in regression models with possibly singular covariances. Linear Algebra and its Applications, 67(??):207–214, June 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585901971.

Werner:1987:WM

[Wer87] Hans Joachim Werner. On weak r-monotonicity. Linear Algebra and its Applications, 86(??):199–209, February 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587902953.

Wermuth:1989:TRM

[Wer89] Edgar M. E. Wermuth. Two remarks on matrix exponentials. Linear Algebra and its Applications, 117(??):

127–132, May 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589905545.

Waters:1985:GCI

[WH85a] Steven R. Waters and Richard D. Hill. Generalized controllability and inertia theory. II. Linear Algebra and its Applications, 72(??):161–166, December 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958590151X.

Waters:1985:IDR

[WH85b] Steven R. Waters and Richard D. Hill. An inertial decomposition and related range results. Linear Algebra and its Applications, 72(??):145–153, December 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585901491.

Waters:1985:IRT

[WH85c] Steven R. Waters and Richard D. Hill. On the inertia and range of transformations of positive definite matrices. Linear Algebra and its Applications, 72(??):155– 160, December 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585901508.

White:1980:UEP

[Whi80] Neil L. White. A unique exchange property for bases. Linear Algebra and its Applications, 31(??):81–91, June 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580902098.

Wiener:1982:SSR

[Wie82] Gerry Wiener. Some structural results concerning certain classes of qualitative matrices and their inverses. Linear Algebra and its Applications, 48(??):161–175, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582901057.

Wiener:1984:SMS

[Wie84] Gerry Wiener. Spectral multiplicity and splitting results for a class of qualitative matrices. Linear Algebra and its Applications, 61(??):15–29, September 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584900193.

Wiener:1985:SRC

[Wie85a] Gerry Wiener. Some results concerning invertible Morishima and anti-Morishima matrices whose associated digraphs are trees. Linear Algebra and its Applications, 66(??):199–219, April 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585901338.

Wiener:1985:TCC

[Wie85b] Gerry Wiener. A theorem concerning certain sign symmetric matrices whose inverses are Morishima. Linear Algebra and its Applications, 68(??):31–45, July 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958590206X.

Wiens:1985:SPR

[Wie85c] Douglas P. Wiens. On some pattern-reduction matrices which appear in statistics. Linear Algebra and its Applications, 67(??):233–258, June 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585901995.

Wilker:1982:AOI

[Wil82a] J. B. Wilker. The Apollonian octets and an inversive form of Krause’s theorem. Linear Algebra and its Applications, 48(??):131–144, December 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582901033.

Wilson:1982:IMD

[Wil82b] Richard M. Wilson. Incidence matrices of t-designs. Linear Algebra and its Applications, 46(??):73–82, August 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582900271.

Willems:1983:IOS

[Wil83] Jan C. Willems. Input-output and state-space representations of finite-dimensional linear time-invariant systems. Linear Algebra and its Applications, 50(??): 581–608, April 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583900708.

Wilkinson:1987:TF

[Wil87] J. H. Wilkinson. On a theorem of Feingold. Linear Algebra and its Applications, 88/89(??):13–30, 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wimmer:1980:TLE

[Wim80] H. K. Wimmer. Two linear equations in formal power series. Linear Algebra and its Applications, 29(??):465– 469, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580902578.

Wimmer:1983:IIP

[Wim83a] Harald K. Wimmer. Indefinite inner-product spaces associated with Hermitian polynomial matrices. Linear Algebra and its Applications, 50(??):609–619, April 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958390071X.

Wimmer:1983:OGS

[Wim83b] Harald K. Wimmer. On Ostrowski’s generalization of Sylvester’s law of inertia. Linear Algebra and its Applications, 52/53(??):739–741, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wimmer:1984:ARE

[Wim84] Harald K. Wimmer. The algebraic Riccati equation: conditions for the existence and uniqueness of solutions. Linear Algebra and its Applications, 58(??):441–452, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379584902246.

Wimmer:1985:RDM

[Wim85a] H. K. Wimmer. Report on the Dublin matrix theory conference, March 1984: Unmixed solutions of the algebraic Riccati equation. Linear Algebra and its Applications, 68(??):272–274, July 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585902241.

Wimmer:1985:PMF

[Wim85b] Hårald K. Wimmer. Powers of a matrix and a formula for the moduli of its eigenvalues. Linear Algebra and its Applications, 71(??):339–340, November 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585902605.

Wimmer:1987:DFA

[Wim87] Harald K. Wimmer. Duality of finite Abelian groups and a Jordan normal form of integer matrices. Linear Algebra and its Applications, 90(??):89–104, May 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587903089.

Wimmer:1988:MEA

[Wim88] Harald K. Wimmer. The matrix equation X − AXB = C and an analogue of Roth’s theorem. Linear Algebra and its Applications, 109(??):145–147, October 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588902042.

Wimmer:1989:BPM

[Wim89a] Harald K. Wimmer. Bézoutians of polynomial matrices and their generalized inverses. Linear Algebra and its Applications, 122/123/124(??):475–487, 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wimmer:1989:GBB

[Wim89b] Harald K. Wimmer. Generalized Bézoutians and block Hankel matrices. Linear Algebra and its Applications, 117 (??):105–113, May 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589905521.

Wimmer:1989:LME

[Wim89c] Harald K. Wimmer. Linear matrix equations: the module theoretic approach. Linear Algebra and its Applications, 120(??):149–164, August 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589903741. Proceedings

of the Fourth Haifa Matrix Theory Conference (Haifa, 1988).

Wimmer:1989:RVJ

[Wim89d] Harald K. Wimmer. Root vectors and Jordan chains of integer matrices. Linear Algebra and its Applications, 112 (??):161–168, January 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379589905922.

Wild:1987:ULS

[WN87] P. Wild and W. Niethammer. Over- and underrelaxation for linear systems with weakly cyclic Jacobi matrices of index p. Linear Algebra and its Applications, 91 (??):29–52, June 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587900589.

Woerdeman:1989:MRC

[Woe89] H. J. Woerdeman. Minimal rank completions for block matrices. Linear Algebra and its Applications, 121(??):

105–122, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589906952.

Linear algebra and applications (Valencia, 1987).

Wolkowicz:1981:SAO

[Wol81] Henry Wolkowicz. Some applications of optimization in matrix theory. Linear Algebra and its Applications, 40 (??):101–118, October 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379581901439.

Wolsson:1989:LDF

[Wol89a] K. Wolsson. Linear dependence of a function set of m variables with vanishing generalized Wronskians. Linear Algebra and its Applications, 117(??):73–80, May 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/002437958990548X.

Wolsson:1989:CEL

[Wol89b] Kenneth Wolsson. A condition equivalent to linear dependence for functions with vanishing Wronskian. Linear Algebra and its Applications, 116(??):1–8, April 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589903935.

Wong:1982:CPS

[Won82a] Chi Song Wong. Characterizations of products of symmetric matrices. Linear Algebra and its Applications, 42 (??):243–251, February 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379582901525.

Wong:1982:IFM

[Won82b] Edward T. Wong. Involutory functions and Moore– Penrose inverses of matrices in an arbitrary field. Linear Algebra and its Applications, 48(??):283–291, December 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379582901148.

Wong:1985:UDS

[Won85] Chi Song Wong. On the use of differentials in statistics. Linear Algebra and its Applications, 70(??):285– 299, October 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585900606.

Wozniakowski:1980:REA

[Woź80] H. Woźniakowski. Roundoff error analysis of a new class of conjugate gradient algorithms. Linear Algebra and its Applications, 29(??):507–529, February 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580902591.

Wolkowicz:1980:BEU

[WS80a] Henry Wolkowicz and George P. H. Styan. Bounds for eigenvalues using traces. Linear Algebra and its Applications, 29(??):471–506, February 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958090258X.

Wolkowicz:1980:MBE

[WS80b] Henry Wolkowicz and George P. H. Styan. More bounds for eigenvalues using traces. Linear Algebra and its Applications, 31(??):1–17, June 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379580902013.

Wyman:1983:ZMR

[WS83] Bostwick F. Wyman and Michael K. Sain. On the zeros of a minimal realization. Linear Algebra and its Applications, 50(??):621–637, April 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379583900721.

Wyman:1989:ZPT

[WSCP89] Bostwick F. Wyman, Michael K. Sain, Giuseppe Conte, and Anna-Maria Perdon. On the zeros and poles of a transfer function. Linear Algebra and its Applications, 122/123/124(??):123–144, 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wright:1985:OSM

[WT85] Tommy Wright and How Tsao. On an optimal solution for maximizing the probability of retention in PPS sampling. Linear Algebra and its Applications, 67(??):67–80, June 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585901867.

Wu:1980:SOP

[Wu80] Chien-Fu Wu. On some ordering properties of the generalized inverses of nonnegative definite matrices. Linear Algebra and its Applications, 32(??):49–60, August 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379580900063.

Wu:1987:PNM

[Wu87] Pei Yuan Wu. Products of nilpotent matrices. Linear Algebra and its Applications, 96(??):227–232, November 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587903466.

Wu:1988:PPS

[Wu88] Pei Yuan Wu. Products of positive semidefinite matrices. Linear Algebra and its Applications, 111(??):53– 61, December 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588900511.

Wu:1989:OFP

[Wu89] Pei Yuan Wu. The operator factorization problems. Linear Algebra and its Applications, 117(??):35–63, May 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379589905466.

Weiss:1986:ACN

[WWWS86] N. Weiss, G. W. Wasilkowski, H. Woźniakowski, and M. Shub. Average condition number for solving linear equations. Linear Algebra and its Applications, 83(??): 79–102, November 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586902661.

Yang:1984:SCS

[Yan84] Shang Jun Yang. Structure of ⟨-classes in the semigroup of nonnegative matrices. Linear Algebra and its Applications, 60(??):91–111, August 1984. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379584900739.

Yip:1981:CCM

[YB81] E. L. Yip and C. S. Ballantine. Congruence and conjunctivity of matrices to their adjoints. Linear Algebra and its Applications, 41(??):33–72, December 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379581900872.

Yip:1984:NSM

[YB84] E. L. Yip and C. S. Ballantine. A note on stable matrices. Linear Algebra and its Applications, 56(??):17– 28, January 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901101.

Yang:1988:EPM

[YB88] Shang Jun Yang and George P. Barker. On the exponent of a primitive, minimally strong digraph. Linear Algebra and its Applications, 99(??):177–198, February 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901310.

Ycart:1982:EDC

[Yca82] Bernard Ycart. Extrémales du cône des matrices de type non négatif, à coefficients positifs ou nuls. (French)

[Extremals of the cone of nonnegative matrices with positive or zero coefficients]. Linear Algebra and its Applications, 48(??):317–330, December 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582901185.

Yeyios:1984:OEM

[Yey84] A. Yeyios. On the optimization of an extrapolation method. Linear Algebra and its Applications, 57(??): 191–203, February 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901873.

Yarlagadda:1982:NEH

[YH82] R. Yarlagadda and J. Hershey. A note on the eigenvectors of Hadamard matrices of order 2^{n}. Linear Algebra and its Applications, 45(??):43–53, June 1982. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379582902105.

Young:1980:GCG

[YJ80] David M. Young and Kang C. Jea. Generalized conjugate gradient acceleration of non- symmetrizable iterative methods. Linear Algebra and its Applications, 34(??): 159–194, December 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379580901652.

Yamada:1985:GPC

[YL85] Takeo Yamada and David G. Luenberger. Generic properties of column-structured matrices. Linear Algebra and its Applications, 65(??):189–206, February 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379585900977.

Young:1981:BNF

[You81a] N. J. Young. A bound for norms of functions of matrices. Linear Algebra and its Applications, 37(??): 181–186, April 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581901774.

Young:1981:RCM

[You81b] N. J. Young. The rate of convergence of a matrix power series. Linear Algebra and its Applications, 35(??):261– 278, February 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379581902780.

Young:1983:SVD

[You83] N. J. Young. The singular-value decomposition of an infinite Hankel matrix. Linear Algebra and its Applications, 50(??):639–656, April 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379583900733.

Young:1984:LFT

[You84] N. J. Young. Linear fractional transformations in rings and modules. Linear Algebra and its Applications, 56(??): 251–290, January 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901319.

Yuan:1986:LET

[Yua86] Hong Yuan. The Kth largest eigenvalue of a tree. Linear Algebra and its Applications, 73(??):151–155, January 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379586902375.

Yuan:1988:BSR

[Yua88] Hong Yuan. A bound on the spectral radius of graphs. Linear Algebra and its Applications, 108(??):135–139, September 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588901838.

Yuan:1989:SLB

[Yua89] Hong Yuan. Sharp lower bounds on the eigenvalues of trees. Linear Algebra and its Applications, 113(??):101– 105, February 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589902887.

Zaballa:1986:EMP

[Zab86] Ion Zaballa. Existence of matrices with prescribed entries. Linear Algebra and its Applications, 73(??):227– 280, January 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379586902429.

Zaballa:1987:MPR

[Zab87] Ion Zaballa. Matrices with prescribed rows and invariant factors. Linear Algebra and its Applications, 87(??): 113–146, March 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587901625.

Zaballa:1988:IIC

[Zab88] Ion Zaballa. Interlacing inequalities and control theory. Linear Algebra and its Applications, 101(??):9–31, April 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588901401.

Zaballa:1989:IMI

[Zab89] Ion Zaballa. Interlacing and majorization in invariant factor assignment problems. Linear Algebra and its Applications, 121(??):409–421, August 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589907131. Linear alge-

bra and applications (Valencia, 1987).

Zalinescu:1985:CSG

[Zăl85] Constantin Zălinescu. On convex sets in general position. Linear Algebra and its Applications, 64(??):191– 198, January 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379585902769.

Zassenhaus:1983:CSL

[Zas83] Hans Zassenhaus. On the Cartan subalgebra of a Lie algebra. Linear Algebra and its Applications, 52/53(??):743– 761, 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zenger:1984:DON

[Zen84] Christoph Zenger. Dual operator norms and the spectra of matrices. Linear Algebra and its Applications, 58(??): 453–460, April 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584902258.

Zenios:1989:NNM

[Zen89] Stavros A. Zenios. Nonlinear network models for matrix balancing and the estimation of social accounting matrices. Linear Algebra and its Applications, 120(??): 262–269, August 1989. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379589903881.

Zhang:1987:LVC

[Zha87a] Ke Min Zhang. On Lewin and Vitek’s conjecture about the exponent set of primitive matrices. Linear Algebra and its Applications, 96(??):101–108, November 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587903387.

Zhang:1987:GCS

[Zha87b] Mou Cheng Zhang. On graph compatible Splittings of M-matrices. Linear Algebra and its Applications, 91(??): 177–179, June 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587900693.

Zhang:1989:CRA

[Zha89] Guo Dong Zhang. On the convergence rate of the QL algorithm with Wilkinson’s shift. Linear Algebra and its Applications, 113(??):131–137, February 1989. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379589902929.

Zheng:1987:PDS

[Zhe87] Yu Fan Zheng. The properties of decentralized systems under local feedbacks via polynomial models. Linear Algebra and its Applications, 92(??):161–185, July 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379587902552.

Zhong:1988:IGI

[Zho88] Xu Zhong. On inverse and generalized inverses of Hessenberg matrices. Linear Algebra and its Applications, 101 (??):167–180, April 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379588901498.

Zielke:1983:SRM

[Zie83] Gerhard Zielke. Some remarks on matrix norms, condition numbers, and error estimates for linear equations. Linear Algebra and its Applications, 110(??):29– 41, November 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379583901301.

Zietak:1985:PCI

[Zie85] K. Zietak. On a particular case of the inconsistent linear matrix equation AX + Y B = C. Linear Algebra and its Applications, 66(??):249–258, April 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379585901363.

Zietak:1987:PAG

[Zie87] K. Zietak. Properties of the approximate generalized inverses of a class of matrices. Linear Algebra and its Applications, 92(??):259–272, July 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/0024379587902631.

Zielke:1988:SRM

[Zie88a] Gerhard Zielke. Some remarks on matrix norms, condition numbers, and error estimates for linear equations.

Linear Algebra and its Applications, 110(??):29–41, 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zietak:1988:CEP

[Zie88b] K. Zietak. On the characterization of the extremal points of the unit sphere of matrices. Linear Algebra and its Applications, 106(??):57–75, August 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0024379588900237.

Zijm:1984:GES

[Zij84] W. H. M. Zijm. Generalized eigenvectors and sets of nonnegative matrices. Linear Algebra and its Applications, 59 (??):91–113, June 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379584901617.

Zimmermann:1980:DAL

[Zim80] Uwe Zimmermann. Duality for Algebraic Linear Programming. Linear Algebra and its Applications, 32(??): 9–31, August 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437958090004X.

Zellini:1981:STC

[ZM81] Paolo Zellini and Angela Mack. On some theorems on circulant matrices. Linear Algebra and its Applications, 41(??):137–149, December 1981. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958190094X.